A Complete Bibliography of Publications in Biometrics: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

20 September 2024 Version 2.01

Title word cross-reference

(θ_{1} + dθ_{2})/θ_{3} [1442]. 2 × 2 [1224, 1400, 1669, 1730, 2, 232, 1235, 1650, 1236, 1599, 1204, 960, 1225, 1532, 145]. 2 × 2 × 2 [633]. 2 × C [2]. 2 × J [1722]. 2 × J × K [1099]. 2 × K [1419, 1037]. A [642]. B [1605]. Δ [1596]. K [1722]. κ [1733]. M [1592]. N = 80 [387]. P [174, 1650, 1532]. P r(X < Y) [1799, 1609, 1320, 1801]. Pr(X_{1} < X_{2}) [1363]. R [961]. R^{2} [1362]. t [689].

-Circular [642]. -Distribution [1596]. -Regression [1592]. -Series [174]. -Value [1605]. -Values [1650, 1532].

1 [1841, 1044]. 1876 [63]. 1900 [1511, 1511]. 1950 [937, 939]. 1950-59 [939]. 1979 [374, 485, 203, 487]. 1981 [390, 432]. 1982 [554, 1129, 1126]. 1983 [768]. 1984 [999, 1172, 1172]. 1984/1985 [1172]. 1985 [1172, 1350, 1350, 1749]. 1986 [1300, 1525, 1525]. 1987 [1530, 1706, 1706]. 1988 [1895]. 1989 [1899]. 1990 [1434].

2 [1346, 1392, 833, 298]. 21 [1384]. 2nd [1127, 1221, 1217, 1219, 1394, 1473, 1460, 1253, 1513, 1122].

3 [289]. 3/4 [1895]. 37 [776]. 38 [1000].

40 [1225]. 41 [1178, 1260, 1179, 1132]. 42 [1354, 1528, 1625, 1801, 1581, 1529]. 43 [1847]. 44 [1904, 1671, 1846, 1905, 1755]. 45 [1902].

59 [939].

79 [174].

83m [1000, 776]. 86d [1225]. 87g [1581]. 88c [1801]. 89d [1905]. 89h [1904].

900 [1220]. 90k [1902].

Âge [483, 483].

Abbott [1577]. Abnormalities [272]. ABO [950, 1066]. ABO-Like [1066, 950]. Aborigines [1198]. Absence [636]. Absorption [1850]. Abstract [281]. Abstracts [28, 65, 124, 164, 234, 370, 417, 466, 527, 694, 748, 817, 928, 988, 1039, 1080, 1118, 1164]. Abundance [886, 733, 729, 587, 1024, 1229, 1139]. Abuse [1687]. Accelerating [1026]. Acceleration [965]. Access [1714]. Accessions [155]. Accident [721]. Accidents [1597]. Accompany [1174, 1175, 1176, 1180, 1181, 1182, 1183]. According [265]. Accrual [1019, 1495]. Accumulated [1193]. Accumulating [1862]. Accuracy [1238]. Acquisitions [995]. Acquistin [1809]. Across [1543, 1018]. Act [63]. Action [355, 217, 1241, 1607, 367, 113, 1082]. Activity [309, 70, 1602, 970]. Activity-Dependent [1602]. Actuaries [485, 487]. Adaptive [689, 1592, 1516]. Adaptively [1773]. Adaptively-Ordered [1773]. Addendum [1184]. Additif [717]. Addition [309, 805]. Additional [1019, 1238]. Additive [787, 571]. Additive-Hazards [571]. Adelman [548]. Adena [828]. Adequacy [1429, 1481, 1843, 1847]. Adipose [1267]. Adjusted [1776, 1377, 859, 1073]. Adjusting [967, 1074]. Adjustment [1140, 445, 592, 348, 1761, 444]. Adjustments [1356]. Administration [63]. Advanced [876, 1563, 1793]. Advances [701, 1518, 1518]. Advisory [63]. Aerial [1761]. Affected [399]. Affects [444]. after [1852, 1193]. against [1177, 409]. Age [1133, 1452, 703, 1377, 744, 617, 866, 405, 1408, 1358, 153, 956, 223, 108, 725, 859, 1073, 1447, 1038]. Age- [1358]. Age-Adjusted [1377, 859, 1073]. Age-Dependent [866, 223]. Age-Specific [153, 108]. Age-Standardized [1452]. Age-Structured [703]. Age-Year [405]. Aggregate [723]. Aggregated [1406]. Aggregation [1586, 730]. Agreement [1601, 1730, 517, 1319, 1416, 1820, 958]. Agresti [1120].

Agricultural [377, 1043, 923, 1025, 818, 419]. Agriculture [820, 1523, 101]. Agronomic [1444]. Agronomie [101]. Agronomie/Statistics [101]. AI [1616]. Aickin [994]. Aid [1277]. AIDS [1735, 1708]. Air [274, 121]. Aitchison [1748]. Akin [376]. Albert [1886]. Albumin [970]. Alfred [1786]. Algebra [762]. Algebras [245]. Algorithm [1608, 1535, 1604, 1197, 1829, 1264]. Algorithmes [60]. Algorithms [836, 450, 1167]. All-or-None [1055, 854]. Allele [312, 112, 220, 313]. Alleles [950, 1485]. Allgemeinen [934]. Allied [1255]. Allocation [15, 857, 511, 746, 1640]. Allocations [636]. Allometric [796]. Allometry [905]. Allow [582]. Allowance [1353, 622, 1266]. Allowing [223]. Almeder [1704]. along [1764]. Alpha [1888, 413, 1522]. Alpha-Emitters-BEIR [1888]. Alpha-Fetoprotein [413]. Alström [546]. Alternative [1177, 1204, 409, 1283, 1320, 1801, 1799]. Altman [646]. Altri [386]. Altruism [490]. Alvey [697]. Amaral [812]. America [485, 487]. American [117]. among [346, 795, 1198, 669, 1692, 1770]. amongst [272]. Análisis [1517]. analogue [961]. Analogy [1159]. Analyse [1474, 938, 600, 60, 1514, 821, 429, 473, 1424]. Analysen [1515]. Analyses [1353, 1439, 1340, 267, 382, 1375, 443, 1811, 197, 1266, 788, 274, 353, 1230, 859, 1661]. Analysing [410, 184, 1098, 1316, 272]. Analysis [239, 994, 403, 306, 314, 556, 1090, 1095, 1174, 1259, 1391, 1397, 1523, 1469, 1400, 1570, 1569, 1573, 1577, 1669, 1753, 1800, 698, 1512, 931, 522, 421, 229, 574, 1227, 452, 448, 1338, 563, 1451, 1062, 1609, 2, 72, 1631, 627, 1292, 1497, 1107, 1148, 1239, 1027, 449, 655, 1251, 1462, 438, 1494, 1071, 544, 1263, 1708, 1653, 651, 1678, 9, 228, 1783, 1218, 1743, 1513, 516, 872, 1697, 598, 325, 1145, 1725, 1010, 671, 40, 20, 1068, 1244, 1625, 1559, 572, 1744, 122, 475, 758, 1851, 1149, 608, 1544, 673, 654, 114]. Analysis [1742, 745, 339, 580, 149, 450, 442, 45, 446, 405, 669, 287, 447, 905, 1720, 1615, 1640, 156, 304, 1205, 876, 1210, 1423, 1009, 800, 891, 357, 1686, 439, 1882, 1441, 225, 78, 510, 1212, 1445, 1859, 1072, 1599, 1159, 1826, 441, 861, 899, 1884, 591, 620, 1407, 1286, 46, 560, 186, 826, 1211, 540, 79, 89, 1197, 241, 592, 1367, 623, 720, 1280, 1214, 1289, 687, 168, 1143, 1748, 519, 130, 399, 537, 1438, 1715, 1890, 1366, 380, 827, 763, 440, 1484, 1691, 1812, 909, 1660, 727, 742, 564, 626, 300, 1775, 451, 1677, 862, 1813, 453]. Analysis [1731, 1048, 471, 84, 29, 1818, 518, 257, 1551, 190, 1014, 1807, 1675, 1275, 113, 148, 514, 4, 1277, 1262, 501, 1125, 677, 1384, 680, 1195, 1311, 1646, 1139, 1260, 1362, 1584, 310, 1001, 1624, 602, 1127, 1128, 1217, 1395, 1663, 1795, 1839, 1894, 1833, 237, 1124, 695, 1428, 1215, 1743, 707, 881, 1420, 1249, 653, 207, 759, 1213, 1655, 165, 1083, 644, 238, 645, 478, 1336, 419, 696, 708, 296, 756, 1048, 294, 1654, 167, 1084, 1386, 1250, 378, 381, 1120, 383, 892]. Analysis- [1210]. Analysis-Methods [1212]. Analytic [901]. Analytical [1461]. Analyzing [969, 1824, 1682, 1504, 856, 1759]. Anamorphic [514]. Anatomy [1052]. Ancienne [883]. Andal [1474]. Andersen [129]. Anderson [199, 1789, 752, 1890, 480, 1127, 64, 24]. Andersson [1384, 1620]. Angeles [117]. Angelika [245]. Angewandte [936]. Angewandten [1467, 830]. Angina [325]. Angler [1714]. Animal [1800, 647, 1410, 1497, 64, 886, 1590, 733, 1377, 197, 1075, 130, 1647, 734, 1102, 1073, 785, 1229, 169, 1181, 753]. Animals [1330, 1680, 63, 1715, 285]. Annette [935]. Annotated [937, 939]. Announcements [31, 67, 103, 135, 178, 211, 251, 308, 391, 434, 495, 555, 615, 657, 712, 769, 843]. Annual [1172, 1350, 1525, 1706, 1895]. ANOVA [1508, 1314, 1417, 1011, 1813]. ANOVA-like [1813]. Ant [1482]. Anthelmintic [1331]. Anthropology [1047]. Antibiotic [1809, 265]. Antibiotic-Resistant [1809]. Aphasia [955]. APL. [247]. APLSTAT [247]. Apparent [970]. Appear [1555]. Appeared [324]. Applicability [1307]. Applicable [1247, 1597, 547]. Application [1144, 310, 770, 1664, 1795, 1796, 1735, 266, 264, 974, 267, 1708, 789, 110, 903, 795, 405, 1067, 1358, 197, 411, 1872, 1197, 1629, 1688, 1194, 1723, 1758, 963, 871, 1267, 1153, 719, 4, 1709, 196, 1646, 776]. Applications [1705, 1461, 1892, 606, 836, 265, 959, 1213, 666, 321, 1211, 121, 1858, 412, 250, 1757, 396, 731, 1045, 307, 1064, 492, 1798, 248, 429, 129, 1742, 1212, 1659, 1051, 379, 383, 471, 29, 696, 760]. Applied [841, 1221, 1518, 673, 1508, 1852, 877, 195, 1270, 471, 1342, 1091, 1394, 1570, 1663, 1702, 1460, 1252, 707, 1167, 532, 1661, 644, 1089, 645, 294, 832, 428]. Applies [1055, 854]. Appreciation [980]. Approach [388, 1754, 1630, 506, 1549, 907, 328, 1642, 1678, 1776, 1783, 1504, 653, 122, 1712, 1085, 156, 510, 186, 1405, 353, 1152, 725, 17, 1855, 145, 1641, 1807, 1685, 1683, 1755, 1435, 288, 544, 996, 885, 1046]. Approaches [327, 1779, 825]. Appropriateness [687]. Approximate [403, 865, 232, 686, 1583, 1200, 54, 524]. Approximately [1236, 1372]. Approximating [681, 875, 1869]. Approximation [1371, 53, 1744, 809, 1629, 1455, 871, 305]. Approximations [10]. Arbitrarily [110]. Arch [626]. Architecture [1388]. Area [1575, 1619, 189, 624, 1356]. Areas [1642, 126]. Argomenti [386]. Arithmetical [1701]. ARMA [1724]. Armitage [1871]. Armitage-Doll [1871]. Arnold [296]. Aroian [493]. Art [546]. Arteriovenous [1237]. Arthanari [535]. Artificial [1616, 1617, 1616, 1616, 1617]. Ascertainment [853, 10]. Ashton [375]. Aspects [1658, 1465, 438, 327, 336]. Assad [1297]. Assay [275, 1008, 1805]. Assays [1318, 1107, 184, 567, 1272, 77]. Assess [1774]. Assessing [309, 972, 1717, 573, 1331, 795, 1820, 620, 1536, 70]. Assessment [764, 332, 577, 270, 193, 1075, 182, 930, 242, 1429, 276, 925, 333, 1398, 1704, 889]. Assisted [371]. Associated [1305, 1304, 1735, 325, 1306]. Association [1730, 910, 1235, 800, 1236, 1099, 485, 487]. Associations [453]. Assuming [1035]. Assumptions [1726, 678]. Assurance [1794]. Asymptotic [1498, 961, 1068, 1587, 1681, 406, 1335, 1346]. Asymptotically [1110]. Asynchronous [1402]. Asynchronously [1713]. Atkinson [1285, 1420]. Atlantic [871]. Atomic [1124]. Attacks [16]. Attributable [575, 1585, 188]. Attribute [1820]. Auditor [554, 768, 1300, 1530, 1899, 390]. Auditors [999].

Aufl [1740]. Augmented [660]. Auquier [480]. ausführlichen [542]. Australian [1198]. Auswertung [934]. Auswertungs [1837]. Auswertungs- [1837]. Autocorrelated [1143]. Automation [1703]. Automatique [60, 298]. Autonomic [925]. Autoregression [1309]. Autoregressive [137, 21, 1035]. aux [382, 751]. Auxiliary [1562, 1142, 849, 1716]. Avian [1782, 1026].

B [174, 100, 552, 1221, 1348, 1345, 1297, 1518, 1523, 1472, 1436, 1840, 1796, 1839, 936, 1246, 612, 700, 606, 833, 1040, 1653, 247, 703, 204, 881, 470, 129, 1121, 207, 608, 1742, 1615, 381, 547, 1168, 1210, 1247, 1737, 1787, 384, 1747, 891, 1883, 1509, 642, 1614, 1385, 824, 1336, 1337, 1042, 1654, 1567, 830, 291]. Background [114]. Bacon [1464]. Bailey [595, 887, 885]. Balakrishnan [1613]. Balance [739]. Balanced [1543, 1640, 88, 867]. Baldessari [606]. Ball [1216]. Bancroft [530]. Band [472, 1062, 911, 1637]. Band-Recovery [1062, 911, 1637]. Bands [1635, 1449]. Banks [155]. Barigozzi [374]. Barnett [1165, 767, 474, 596]. Barrie [538]. Barry [249]. Basal [36, 402]. Base [50]. Based [1182, 1131, 1670, 563, 350, 1734, 1364, 1496, 454, 966, 1166, 40, 1711, 1068, 1851, 1406, 1104, 924, 1550, 1407, 1700, 1028, 977, 362, 1890, 730, 909, 1268, 897, 1807, 1671, 1132]. Baseline [1371]. Baselines [1494]. Bases [52]. BASIC [1431, 1620, 930, 113, 1790, 539, 656, 946]. Basics [1508, 696]. Basis [1092, 1714, 975]. Batschelet [477]. Bauer [1741]. Bayes [1452, 1766, 565]. Bayesian [776, 1260, 770, 1259, 522, 506, 1538, 264, 261, 1149, 156, 848, 1405, 1025, 1152, 564, 17, 145, 971, 1696, 1641, 1014, 1392, 533]. BBB [351]. Be [1546, 32, 68, 104, 142, 180, 212, 252, 315, 392, 496, 557, 435]. Bean [421]. Bean-San [421]. Becker [1797, 1169]. Becoming [1433]. been [1018]. Before [1511, 683]. Behavioral [1397]. Behaviour [1535]. beim [542]. Being [343]. BEIR [1888]. Belfort [1258]. Bell [612]. Bellman [1519, 944]. Belsley [293]. Belts [1310]. Benard [1860]. Benard-van [1860]. ’Bending [219]. Bennett [989]. Berchtold [98]. Berenson [1702]. Bergman [424]. Bernard [1380]. Bernardo [1392, 533]. Bernoulli [1553, 1151]. Best [1752, 109, 521, 1870, 1307, 1857, 18, 1900]. Beta [1906, 1486, 1270, 1561]. Beta-Binomial [1906, 1486, 1561]. Beta-Geometric [1270]. Between [1601, 1821, 800, 583, 138, 1402, 1498, 1680, 1717, 910, 55, 1058, 686, 1072, 1075, 79, 984, 1241, 348, 1097]. Beutel [894]. Beyond [1508]. Bhat [1221, 1348]. Bias [577, 263, 502, 1357, 670, 858, 1235, 1359, 1017, 44, 480, 1561, 139]. Biased [788, 1230]. Biases [1480]. Bibby [766, 1521, 1473, 165]. Bibliography [937, 939, 1248, 100]. Bicentennial [1434]. Bidirectional [915]. Bifurcating [1309]. Binaires [855]. Binary [1175, 1903, 1276, 1499, 398, 1638, 224, 966, 1207, 1157, 673, 1016, 1684, 1598, 1679, 957, 1904, 1179]. Binomial [779, 1579, 1906, 1498, 185, 1502, 1332, 1595, 811, 266, 1733, 267, 674, 1364, 1583, 868, 809, 685, 1610, 1405, 732, 1284, 1726, 926, 1532, 1486, 1561, 793, 780]. Binomially [629]. Bio [831, 206]. Bio- [831]. Bio-Sciences [206]. Bioassay [772, 1580, 590, 907, 218, 1076, 1688, 566, 1757, 1487, 1196, 37, 356, 1644, 1581, 1835]. Bioassays [1580, 1317, 1242, 1196, 1581]. Bioeconomic [1294]. Bioequivalence [231, 182, 1489, 807, 145, 971, 1078, 1328]. Biological [173, 421, 649, 372, 652, 199, 475, 282, 954, 845, 397, 993, 331, 1432, 172]. Biologie [751, 704]. Biologischen [1740]. Biologist [1793, 1433]. Biologists [1468, 171, 235]. Biology [374, 1129, 175, 1513, 203, 1577, 246, 375, 595, 1293, 1892, 30, 607, 820, 1288, 477, 991, 1337, 1092]. Biomass [795]. Biomathematical [1052]. Biomathematics [887, 1211, 468, 1425]. Biomedical [174, 1222, 1380, 959, 761]. Biometric [117, 134, 177, 210, 390, 554, 433, 494, 614, 768, 842, 999, 1130, 1300, 1530, 1672, 1899, 102, 66, 551]. Biometrical [288, 613]. Biometrics [1000, 1304, 776, 1904, 1053, 1354, 1528, 1178, 1625, 1260, 1671, 70, 1179, 1225, 1132, 1801, 1581, 1846, 1847, 1529, 1902, 1905, 1755, 715, 755, 32, 68, 104, 142, 180, 212, 252, 315, 392, 496, 557, 438]. Biometrie [1170, 1170, 1837, 1044, 1514]. Biometrische [1514]. Biometry [1136, 420, 97, 922, 686, 923, 870, 869, 467]. Biomonitoring [242]. Biopsies [846]. Biostastistics [204]. Biostatistical [715, 1390]. Biostatistics [1114, 1086, 997, 1518, 946, 1564, 1569, 1380, 1656, 826, 711]. Biostatistik [472, 550]. Bird [153, 154]. Birds [612, 754]. Birks [1839]. Birth [1778]. Bisby [1577]. Bithell [545]. Bivariate [1321, 1325, 1597]. Blackith [1122]. Blanks [1645]. Blinding [927]. Bloch [1461]. Block [1361, 1244, 1202, 88, 1884, 77, 1624, 1625]. Blocked [455]. Blocking [804, 1877, 1413]. Blocks [20, 446]. Blood [789, 1197, 1237, 47, 485]. BLUP [542]. BLUP-Metode [542]. BMDP [1661, 174, 432]. BMDP-79 [174]. Bock [166]. Body [36, 402]. Boen [594]. Boitard [603]. Boker [1623]. Bollinger [831]. Bologna [386]. Bologna-Istituto [386]. Bomb [1124]. Bonferroni [1477, 1476, 1414]. Book [1430, 602, 239, 994, 174, 173, 697, 945, 1342, 100, 305, 176, 249, 306, 432, 551, 387, 552, 388, 492, 493, 553, 764, 765, 766, 840, 841, 1091, 1092, 1170, 1127, 1093, 1090, 1094, 1128, 1171, 1095, 1172, 1346, 1298, 1347, 1255, 1221, 1254, 1256, 1344, 1216, 1295, 1299, 1348, 1217, 1219, 1220, 1345, 1296, 1257, 1258, 1297, 1349, 1350, 1391, 1394, 1517, 1518, 1522, 1392, 1432, 1433, 1434, 1519, 1520, 1395, 1396, 1397, 1467, 1470, 1393, 1398, 1399, 1523, 1521, 1471, 1472, 1473, 1435, 1469]. Book [1474, 1524, 1475, 1436, 1525, 1570, 1663, 1702, 1703, 1569, 1664, 1667, 1618, 1571, 1574, 1665, 1572, 1622, 1705, 1666, 1704, 1575, 1576, 1619, 1620, 1623, 1573, 1577, 1621, 1706, 1837, 1749, 1838, 1840, 1795, 1796, 1750, 1751, 1797, 1839, 1893, 1894, 1798, 1842, 1752, 1753, 1895, 484, 698, 1380, 1512, 290, 536, 702, 1612, 931, 1611, 1658, 377, 755, 1833, 749, 420, 937, 939, 374, 531, 95, 934, 421, 649, 705, 468, 472, 550, 936, 1246, 1740, 1741, 1464, 97, 821, 1794, 246, 375, 1052, 595, 1738, 98, 1338]. Book [647, 1084, 243, 236, 535, 611, 1656, 612, 244, 1169, 237, 700, 288, 389, 757, 1124, 1292, 1461, 289, 1129, 938, 1087, 1343, 64, 600, 1468, 372, 597, 879, 1088, 1745, 883, 992, 1340, 828, 1293, 1465, 1892, 539, 942, 1388, 62, 248, 606, 701, 1460, 655, 609, 886, 823, 1251, 1462, 1565, 303, 423, 30, 285, 1086, 530, 767, 60, 297, 422, 170, 544, 943, 1253, 1888, 61, 695, 1165, 479, 833, 1040, 889, 941, 1252, 1653, 483, 1428, 1424, 247, 651, 932, 175, 652, 476, 607, 835, 1887, 543, 1386, 299, 172, 541, 604, 829, 429]. Book [382, 245, 301, 1041, 1218, 1215, 1743, 1513, 425, 707, 598, 117, 703, 63, 204, 593, 989, 820, 292, 881, 1420, 1746, 1044, 1514, 1515, 1788, 706, 656, 946, 1123, 101, 295, 470, 610, 699, 1043, 1166, 1250, 1249, 1510, 199, 1885, 653, 129, 1121, 207, 201, 474, 940, 1049, 1744, 761, 1290, 475, 758, 836, 1422, 426, 750, 762, 1390, 608, 893, 894, 654, 1789, 1742, 378, 548, 481, 759, 825, 1047, 1122, 613, 996, 128, 376, 1463, 282, 1085, 709, 538, 1891, 287, 1288, 1889, 1508, 753, 1213, 489, 752, 477, 990, 882, 1615, 995, 304, 286]. Book [381, 547, 876, 834, 1167, 1168, 1285, 1248, 1210, 1247, 1341, 1381, 1511, 1423, 1431, 1655, 1563, 1701, 1613, 1832, 1737, 1787, 1793, 469, 171, 384, 529, 997, 1747, 546, 891, 1883, 1882, 532, 1212, 165, 1083, 1509, 1661, 642, 877, 1884, 751, 240, 528, 822, 1286, 371, 998, 545, 473, 644, 826, 1089, 1050, 1211, 486, 887, 540, 1659, 206, 385, 648, 241, 467, 125, 238, 645, 1421, 1051, 200, 754, 991, 1662, 1614, 603, 1214, 838, 133, 1385, 646, 824, 1289, 1700, 1291, 1466, 605, 478, 168, 485, 839, 1336, 1389, 1748, 386, 424, 819]. Book [130, 293, 537, 1082, 1836, 1739, 302, 1383, 1287, 131, 373, 933, 1120, 482, 710, 59, 930, 1890, 205, 419, 380, 827, 711, 763, 1886, 696, 166, 1881, 885, 250, 379, 534, 760, 643, 1126, 1425, 1337, 1660, 427, 1834, 96, 704, 708, 599, 126, 944, 888, 208, 430, 1516, 242, 300, 1339, 831, 296, 1294, 99, 1564, 132, 818, 549, 837, 431, 533, 488, 596, 756, 1042, 1429, 1387, 1048, 203, 383, 1616, 1617, 294, 471, 1046, 1654, 1567, 29, 650, 1427, 1791, 832, 480, 1426, 487, 890, 169, 601, 892, 880, 830, 291, 298]. Book [1045, 1566, 594, 542, 127, 94, 283, 284, 202, 490, 929, 993, 1568, 307, 1657, 209, 428, 491, 878, 935, 1835, 235, 1125, 1790, 1792, 884, 1384, 1081, 167, 418, 1054, 249, 1575]. Bookstein [1337]. Boorman [490]. Bootstrap [700, 1759]. Bootstrapping [113]. Boswell [1248]. Both [1232, 514]. Bothe [1840]. Boundaries [1549, 1158, 1372]. Bounding [1301, 1206]. Bounds [1600, 984]. Bouroche [297]. Box [1090]. Bozler [654]. Braak [1621]. bracteatum [1482]. Bradley [581]. Brain [611]. Branching [1105]. Breast [1852]. Breeding [1800, 1497, 96, 647, 1426, 169]. Breiman [940]. Breslow [300]. Brillinger [1254, 306]. Brister [885]. Britain [531]. British [325, 326]. Britton [1892]. Broken [358]. Bross [385, 1836]. Brown [1393, 204, 174]. Brownie [1890]. Bruni [128]. Bruno [1040]. Build [487]. Building [198]. Bulmer [1399, 283]. Buncher [418]. Bunke [1421, 1421]. Bureau [1703]. Burman [1210]. Burn [160]. Burnham [64, 199, 752, 1890, 24]. Busekros [245]. Butler [1752, 1752]. Butz [819]. Buyse [1123].

C [552, 388, 1093, 1090, 1095, 1347, 1221, 1345, 1434, 1571, 1619, 1621, 749, 420, 374, 705, 1246, 1052, 1129, 64, 1468, 883, 539, 606, 1251, 1086, 422, 833, 1040, 541, 598, 1420, 1510, 1885, 940, 608, 1789, 759, 128, 282, 489, 752, 286, 1285, 1210, 1381, 1382, 1655, 1737, 1787, 997, 642, 877, 1050, 1659, 206, 1051, 754, 424, 1739, 131, 482, 1890, 205, 696, 1881, 885, 534, 643, 1660, 704, 599, 208, 1516, 1294, 1564, 818, 1042, 892, 284, 418]. C.-F [1090]. Cairns [242].

Calculated [1773]. Calculating [53, 1455, 354]. Calculation [345, 785, 1328]. Calculations [1694, 1537]. Calculus [1705, 282]. Calendar [1155]. Calibration [1133, 123, 956, 735, 280, 1032]. Calibrations [567]. California [117]. Cam [941]. Campbell [237, 1122]. Can [1546, 1555]. Cancer [640, 1223, 1451, 974, 943, 175, 1368, 150, 504, 405, 1316, 1205, 1852, 1441, 336, 186, 827, 300, 1814, 1014, 941, 1123]. Cannings [541]. Cannot [691]. Canonical [627, 1859, 1646, 1211]. Cantillon [101]. Capability [1711]. Capillary [564]. Capture [1062, 1271, 72, 1483, 1762, 64, 1154, 1760, 1060, 23, 912, 256, 223, 734, 849, 1061, 1696, 195, 1765, 752]. Carbohydrate [424]. Carcinogenesis [40, 785]. Carcinogenic [311, 41, 193, 785]. Carcinogenicity [332, 1591, 266, 1590, 1725, 1377, 1073, 333]. Carcinogens [314, 1075, 148]. Cardiac [925]. Care [160]. CARICOM [923]. Carl [1393]. Carlo [605, 920, 218, 1721, 871, 862]. Carolina [186]. Carryover [740, 1274]. Carter [827, 841]. Casagrande [232]. Case [1648, 865, 828, 1098, 575, 1269, 1586, 1694, 670, 1606, 1327, 1851, 1587, 1279, 1639, 446, 1416, 1585, 591, 809, 1693, 788, 1012, 962, 1322, 901, 860, 1231, 271, 1230, 1854, 300, 1649, 257, 1323, 961, 708]. Case-Control [1648, 828, 1098, 575, 1586, 1694, 670, 1327, 1851, 1639, 446, 1416, 1585, 591, 1693, 788, 1012, 962, 1322, 901, 1231, 271, 1230, 1854, 300, 1649, 257, 1323, 708]. Casebook [204, 1836]. Casella [1510]. Catch [733]. Catch-Effort [733]. Catchability [1181, 1483, 1102]. Categorical [635, 1140, 1352, 1150, 1112, 544, 856, 115, 439, 78, 510, 186, 540, 921, 1120, 18, 1273, 383, 1551]. Categories [572, 1768, 1391]. Categorized [1799, 1609, 1320, 1801]. Category [18, 271, 43]. Category-Matching [271]. Caulcutt [699]. Causal [1739]. Causation [1856]. Cause [483, 978]. Causes [1555]. Causton [420, 1793]. Cautionary [969]. ce [600]. Celebration [1285]. Cell [1309, 175, 624, 186, 455]. Cell-Area [624]. Cell-Cycle [455]. Cells [789, 1197]. Censored [138, 775, 1180, 1800, 1595, 965, 865, 1497, 569, 515, 1708, 1371, 1145, 1315, 55, 110, 339, 672, 259, 1370, 1200, 121, 1767, 1268, 1534, 1267, 897, 1101, 677, 777, 1184]. Censored-Data [965, 672]. Censoring [773, 1896, 1901, 1735, 340, 1768, 570, 1030, 407, 1545, 1819, 1902, 1905]. Census [1434, 1186]. Censuses [430]. Center [1616, 1617]. Central [1703]. Cerebral [548]. Certain [683]. Certainty [1036]. Chain [39, 185, 461]. Chains [1376, 792]. Chambers [1797, 881]. Chandon [429]. Change [770, 1181, 1177, 264, 966, 1763, 409, 22, 1102, 154, 1554, 776, 168]. Change-in-Ratio [1181, 1763, 22, 1102]. Change-Over [966]. Change-Point [770, 1177, 264, 409, 1554, 776]. Changeover [415, 145, 518]. Changes [1896, 1901, 1053, 1450, 1545, 1819, 1902, 1905]. Changing [1826, 1806]. Chaos [1298]. Chapman [609]. Character [1137]. Characterisations [606]. Characteristic [1642]. Characterization [1430, 265]. Characterizations [1884]. Characterize [1774]. Characterizing [1602]. Charles [610, 1880]. Charles-Edwards [610]. Charlesworth [703]. Chart [1787]. Chatfield [1095, 877]. Chaudhary [1573]. Chemical [1794, 789]. Chemisch [1837, 1044, 1514]. Chemisch-Pharmazeutischen [1837, 1044, 1514]. Chemistry [1221]. Chemoprevention [1441]. Chemotheraphy [175]. Chemotherapy [827]. Chernoff [1337]. Cheuk [176]. Chevalier [938]. Chi [369, 1776, 513, 1115, 512, 862]. Chi-Square [1776, 1115]. Chiang [1051]. Chien [530]. Chien-Pai [530]. Child [870]. Children [360, 158]. Childress [423]. Choice [523, 1419, 88, 1160, 629]. Choosing [636, 1272, 1770]. Christine [483]. Chromosome [625]. Chronic [783, 1887, 1317, 51]. Chronological [937, 939]. Chudnovsky [1842, 1842]. Circulant [562]. Circular [1410, 642, 477]. Citro [1434]. Clark [1570, 1294, 132]. Clarke [1885, 656, 209]. Class [1774, 1811, 1720, 1767]. Classes [74, 632, 1700]. Classification [1049, 265, 1050, 298, 975, 60, 60, 940, 758]. Classifications [572]. Classified [100, 1391, 1248]. Cleveland [881, 1749, 1566]. Cliff [1658, 379]. Clinical [583, 675, 310, 771, 1259, 1471, 1670, 1057, 1512, 1111, 1412, 15, 857, 222, 1549, 1862, 14, 1656, 52, 1461, 1148, 1491, 1734, 1388, 1494, 322, 917, 117, 966, 352, 896, 1123, 120, 1375, 122, 415, 1149, 1492, 118, 401, 1490, 279, 152, 1550, 157, 998, 578, 648, 334, 1495, 1291, 1778, 119, 482, 353, 1438, 763, 1886, 1283, 1193, 354, 724, 1077, 262, 1818, 1641, 1770, 739, 457, 585, 4, 579, 1059, 1262, 1274, 1548, 1260, 1671, 778, 1388, 299, 825, 996]. Clinician [118]. Clinimetrics [1657]. Clinton [97, 99]. Clonal [970]. Clone [625]. Clone-Tissue [625]. Closed [1501, 64, 752, 734, 790]. Cluster [1795, 201, 1851, 1407, 1424, 207, 1214, 1048]. Cluster-analyse [1424]. Clustered [1822, 1264]. Clustering [398, 1011, 1533, 12, 986, 1691, 914, 147, 576, 784, 1335, 970, 1208, 884]. Clutter [885]. Cobelli [424]. Cochran [1081, 107, 931, 389, 593, 115, 716, 1323]. Code [536]. Cody [1342]. Coefficient [1351, 1526, 1898, 681, 684, 1240, 875, 1319, 1593, 1712, 35, 1729, 1448, 1354, 1528]. Coefficients [1730, 1538, 1009, 1413]. Coggeshall [764]. Cohen [1434, 1878, 1466]. Cohort [744, 617, 618, 788, 1230, 621, 1018]. Cohorts [1480, 1854]. Coles [206]. Collected [1393, 1749]. Collection [1224, 1292, 960, 1225]. Collinearity [293]. Colton [157]. Column [1419]. Combination [1580, 335, 369, 1196, 356, 1581]. Combinations [84]. Combining [339, 1863, 1242, 1828]. Commentés [938]. Comments [328, 326, 1306]. Commerce [1575]. Committee [1222, 63]. Common [684, 1586, 686, 1013, 973, 1204, 1883]. Commonly [1235, 412]. Communautés [821]. Communications [366, 983]. Communities [1462]. Community [1251, 268, 349]. Companion [1290]. Comparative [774, 586, 3, 40, 973, 123, 480, 1270, 1832, 596]. Compared [793]. Comparing [1539, 1680, 904, 1642, 53, 1808, 1034, 809, 411, 968, 1532, 690, 1193, 54, 1201, 258]. Comparison [1351, 1526, 1896, 1901, 360, 511, 726, 523, 263, 1240, 111, 746, 966, 673, 116, 460, 959, 191, 1159, 581, 620, 182, 1872, 1228, 1030, 687, 1359, 1881, 397, 897, 862, 459, 1545, 1354, 1528, 1819, 1902, 1905, 1221]. Comparisons [1174, 688, 689, 1496, 1010, 1692, 188, 57, 1860, 1311]. Compartment [866, 1803]. Compartmental [667, 1630, 277, 847, 1866, 1513, 608]. Competing [1205, 1600, 861, 1012, 1699, 1831]. Competing-Risks [1205, 861]. Competition [795, 909, 991]. Complete [404]. Complex [1670, 303, 1327, 1550, 1671]. Complexity [51]. Component [664, 1215, 1445, 1744]. Componentes [1517]. Components [905, 1443, 1883, 867, 1035, 186, 1021, 636, 169]. Composite [152, 262]. Composition [1408, 735]. Compositional [1748]. Compounds [789]. Computation [1341]. Computational [1091, 247]. Computations [774, 3]. Computer [174, 536, 101, 945, 371, 837]. Computer-Assisted [371]. Computers [484]. Computing [1411, 19, 1324, 696, 451, 656, 761, 998, 133]. Comtex [1616, 1617]. Conception [402]. Concepts [1620, 630, 890, 1700]. Conceptual [1435]. Concerning [162, 684]. Conclusion [1262]. Concomitant [339, 620]. Concordance [850, 1729, 407]. Concordant [1693]. Condition [1246]. Conditional [1901, 973, 12, 1147, 1819, 1902]. Conditioning [1558]. Conduct [322, 708]. Conference [1666]. Conferences [525, 693]. Confidence [403, 773, 1351, 1526, 1111, 1373, 1498, 1821, 1502, 1539, 341, 920, 974, 1728, 1240, 1634, 1364, 340, 1418, 1556, 1363, 1768, 1020, 618, 1865, 1490, 868, 1859, 1766, 188, 898, 1442, 912, 984, 1688, 1284, 1209, 1635, 730, 897, 682, 1449, 918, 1151, 1644, 1354, 1528]. Configural [781]. Configurations [1721]. Confounding [1874, 1481, 1847, 1843, 1843]. Congeneric [123]. Congress [1345]. Conic [626]. Conjunction [1860]. Connectivity [819]. Conolly [547, 384]. Conover [236]. Consent [1493]. Consideration [846]. Considerations [50, 187, 902, 457]. Consistent [1232, 1710]. Constant [1177, 1271, 394, 409]. Constant-Parameter [1271]. Constrained [676, 682, 739]. Constraints [1688]. Construction [531, 1634, 219]. Consulting [594]. Consumption [1429]. Contact [76]. Contagion [76, 721]. Contagious [381]. Contain [1504]. Contained [1552]. Containing [121]. Contaminants [274, 121]. Contemporary [243]. Content [742]. Context [1059]. Contingency [61, 1419, 347, 791, 149, 1686, 406, 1532]. Continuation [771, 579, 778]. Continuous [1140, 556, 1248, 357, 78, 510, 271, 1195]. Continuum [1519]. Contraception [486]. Contrasts [919]. Contributions [1574, 1086, 322, 323, 320, 324, 1664, 593, 834]. Control [1893, 1648, 335, 828, 1098, 575, 1586, 1694, 670, 1327, 1851, 1639, 446, 1416, 1585, 591, 1693, 788, 1012, 962, 1238, 1322, 901, 1231, 271, 1230, 1854, 300, 1649, 1234, 359, 408, 1770, 257, 1481, 1201, 1323, 1843, 1847, 708]. Controlled [567]. Controlling [1187]. Controls [161, 788, 1230, 961]. Convergence [1430]. Cook [702]. Cooke [612, 656, 649]. Cooperative [120]. Cope [840]. Coppi [545]. Cores [742]. Cormack [173, 241]. Cornell [380]. Cornfield [331, 58, 324, 316, 322, 323, 1459, 321, 320]. Cornwell [493]. Coronary [330]. Corporation [1616, 1617]. Corps [1827]. Correct [452, 1034]. Corrected [504]. Correcting [670]. Correction [1000, 139, 136, 137, 138, 140, 310, 314, 312, 770, 779, 771, 773, 775, 1133, 1131, 1259, 1301, 1526, 1476, 1400, 1624, 1799, 1843, 1896, 1897, 1898, 1754, 1800, 1844, 1845, 1903, 776, 1904, 780, 1354, 1528, 1831, 1178, 1625, 313, 1260, 141, 1477, 1671, 1179, 1225, 1132, 1801, 1581, 1846, 54, 778, 1847, 1529, 1902, 1905, 1755, 1352]. Corrections [139, 136, 137, 138, 140, 311, 556, 772, 774, 1001, 1002, 1003, 1004, 1005, 1054, 1173, 1055, 1174, 1175, 1176, 1180, 1181, 1182, 1183, 1177, 1351, 1352, 1223, 1353, 1224, 1526, 1527, 1578, 1669, 1579, 1580, 1670, 1900, 1901, 1906, 777, 1583]. Correlated [1631, 622, 86, 1642, 218, 364, 1456, 1560, 1684, 1598, 159]. Correlation [1351, 1717, 681, 9, 684, 1240, 875, 580, 1729, 623, 1359, 348, 1413, 1074, 1354, 1528]. Correlations [363, 851, 1634, 1022, 1280, 1542, 1633]. Correspondence [989, 1744, 1213, 1139]. Cortex [548]. Cost [1236, 1141]. Cost-Effective [1236]. Cost-Efficient [1141]. Costa [812]. Cot [196]. Council [1253, 1888, 1429, 1792]. Count [1697, 1547, 1844, 1846]. Counties [186]. Counting [189, 1157, 741]. Country [1097]. Counts [1199, 1647, 730]. Counts-a [730]. Course [295, 1423, 279, 877, 933, 1091]. Covariables [621]. Covariance [1001, 452, 448, 445, 797, 449, 438, 1494, 663, 20, 1861, 450, 442, 443, 446, 1825, 1312, 439, 441, 1161, 592, 440, 1724, 451, 1551, 444, 677]. Covariances [798, 346]. Covariate [1140, 1175, 447, 861, 1016, 1356, 444, 514, 1179]. Covariates [1784, 1371, 1719, 1009, 898, 1333, 1209, 1684, 84, 565]. Coverage [1834]. Cows [1584]. Cox [1000, 1084, 428, 344, 1144, 1003, 1004, 1571, 874, 447, 898, 46, 863]. Cram [1087]. Craven [656]. Crimes [1836]. Criss [747]. Criss-Cross [747]. Critchley [1833]. Criteria [147]. Criterion [904, 1867, 1533]. Critical [803]. Crop [1885]. Crops [1732, 1832]. Cross [1391, 7, 1321, 572, 339, 1445, 88, 747]. Cross-Classifications [572]. Cross-Classified [1391]. Cross-Ratio [1321]. Cross-Sectional-Censored [339]. Cross-Validation [1445]. Cross-Validatory [88]. Crosses [562, 7]. Crossing [1180, 1101, 1184]. Crossover [967, 310, 1259, 14, 1148, 1457, 228, 806, 1783, 1245, 1149, 1488, 740, 692, 4, 1274, 1548, 1260]. Cruelty [63]. Crustacean [1713]. cryptic [220]. Csörgo [880]. Cullen [1288]. Culling [508]. Culling-Level [508]. Cultivars [1187]. Culture [454]. Cumulative [1869]. Cure [1816]. Curnow [820]. Current [546, 334]. Curve [1174, 907, 797, 688, 1010, 1313, 906, 1681, 197, 57, 718, 1679, 565, 1448]. Curves [1642, 1191, 653, 1310, 1009, 977, 1710, 1311]. Curves-the [653]. Curvilinear [229]. Cut [904]. Cut-Off [904]. Cutting [922]. Cycle [411, 455]. Cyclic [1746]. Cycling [1376]. cygnus [659]. Cytogenetic [1191]. Cytometry [1395]. Cytopathological [745, 501].

D [602, 239, 173, 306, 387, 840, 1221, 1254, 1344, 1295, 1217, 1219, 1296, 1392, 1393, 1523, 1472, 1663, 1702, 1618, 1573, 1577, 1839, 1798, 1842, 1752, 290, 702, 1658, 420, 649, 1084, 288, 1124, 1461, 64, 1745, 701, 609, 1086, 61, 1653, 543, 299, 172, 1514, 1515, 656, 946, 610, 1250, 199, 201, 758, 1390, 481, 759, 489, 752, 995, 304, 1167, 1381, 1431, 1382, 1793, 1747, 1884, 644, 385, 1614, 646, 824, 1289, 1700, 1389, 293, 1836, 482, 710, 1890, 827, 696, 379, 704, 1564, 549, 533, 1046, 480, 169, 880, 291, 594]. D [202, 428, 235, 167]. Daget [821]. Dagnelie [553, 101, 528]. D’Agostino [1567]. Dahlberg [604]. Daley [1221]. Damned [766, 1473]. d’Analyse [717]. d’Analyses [603]. Daniel [208, 1569, 826]. dans [821, 1154, 483]. Darstellung [542]. Darstellungsmethoden [1837]. Darwin [989, 989]. Data [13, 1144, 1110, 138, 306, 311, 775, 1090, 1128, 1391, 1395, 1397, 1575, 1799, 1754, 1903, 698, 568, 1833, 343, 1150, 1862, 965, 1019, 1451, 1062, 1506, 1609, 410, 398, 52, 865, 627, 1124, 955, 335, 1483, 1762, 64, 569, 1638, 267, 911, 655, 1251, 1071, 184, 515, 50, 544, 695, 517, 1708, 1868, 404, 9, 664, 670, 1504, 1371, 1743, 161, 516, 872, 1697, 707, 966, 1145, 1315, 1725, 55, 881, 110, 215, 1250, 1645, 24, 1244, 1378, 1013, 796, 1544, 1406, 1556, 673, 654, 856, 745, 339, 759, 116, 401, 1719, 447, 1692]. Data [723, 1205, 672, 1381, 1382, 1655, 115, 439, 41, 1022, 406, 90, 1533, 1585, 507, 1605, 861, 1805, 1560, 659, 912, 274, 186, 1050, 1604, 1197, 921, 1405, 864, 623, 1853, 1264, 1066, 1370, 1584, 1200, 1336, 121, 519, 293, 815, 1767, 1383, 845, 1241, 901, 827, 696, 1660, 465, 1232, 1758, 963, 1598, 599, 17, 1193, 461, 914, 1775, 1534, 1267, 1320, 897, 756, 1273, 1048, 383, 1617, 1654, 1542, 1101, 625, 1547, 892, 571, 1448, 1810, 1710, 501, 196, 884, 258, 677, 680, 167, 1195, 1685, 1186, 1139, 1624, 1844, 777, 1904, 1625, 1811, 1801, 1184]. data [1846, 1755, 239, 994, 1084, 1292, 1165, 1653, 1218, 706, 474, 475, 758, 1615, 891, 540, 1748, 1120, 380, 1886, 208, 1566, 1125, 1180, 1352]. Datalink [1616]. Datenanalyse [893]. Datenbanksystem [1667]. Datenverarbeitung [893]. David [292]. Davis [1660]. Dawson [880]. Day [839, 300]. DBMS [1667]. Dean [1395]. Death [311, 900, 1784, 924, 41, 978, 1555, 196]. Décès [483]. d’Échantillonnage [822]. Decision [1469, 1894, 1549, 1595, 835, 182, 371]. Decision- [1549]. Decision-Making [182, 371]. Decisions [526, 1295, 648]. Declining [1709]. Deconvolution [1584]. Deer [924]. Degli [386]. Degree [1858]. Degrees [814, 873]. DeGroot [1392]. Dehnad [1893]. Deischsel [1424]. Delay [1866]. Delayed [157]. Deleterious [345]. Deletion [1138]. Demets [482]. Demografia [386]. Démographiques [483]. Demography [1465, 1465]. Demonstrations [1700]. Dempfle [542]. Dendroclimatology [1675]. Dennis [548]. Densities [24]. Density [1410, 802, 1308, 199, 952, 977, 362, 629, 1654]. Dental [176]. Department [1575]. Departure [1726]. Dependance [1154]. Dependence [1175, 572, 1016, 1147, 1179]. Dependent [1903, 1499, 1784, 1638, 1606, 866, 1600, 223, 1699, 1555, 84, 1602, 226, 1765, 1904]. Deposited [1888]. DePriest [304]. Derivative [1067]. Derived [679]. Descriptive [1083]. Design [355, 967, 675, 310, 1259, 1893, 1512, 1611, 1148, 1745, 1281, 1374, 655, 502, 1732, 1510, 20, 415, 1149, 270, 1202, 666, 876, 171, 1628, 660, 259, 1493, 1021, 1367, 566, 168, 636, 119, 419, 1231, 763, 1714, 708, 145, 1771, 1818, 1641, 518, 1770, 457, 4, 43, 1260, 209, 598, 1655, 130, 1890]. Designed [1174, 1010, 1250, 1075, 716, 742]. Designing [774, 3, 1059]. Designs [314, 900, 1318, 1491, 228, 806, 1817, 1244, 805, 1369, 88, 1488, 740, 1035, 380, 747, 1854, 262, 148, 77, 1624, 1625, 1573, 1746, 470, 1210, 1884, 819].

Desirability [967]. Detect [1376]. Detectability [1409]. Detecting [1546, 1450, 1455, 36]. Detection [770, 264, 5, 949, 1636, 986, 1691, 784, 1764, 776, 136, 1612]. Determination [1821, 327, 1492, 741, 659, 1723]. Determinations [661]. Determining [1533, 1532]. Deterministes [704]. Deterministic [1465, 488]. Deux [855]. Developing [1097]. Development [654, 1106, 699]. Developments [289]. Devices [691]. Devore [597]. Devroye [1662]. d’Experimentation [553]. Diabetes [789, 1144]. Diagnosis [619, 1500, 226]. Diagnostic [577, 904, 1420, 1500, 26, 1863, 269, 1147]. Diagnostics [293]. Diagonal [792]. Diallel [562, 7]. Diameters [1237]. Dichotomous [519, 1241, 457, 160]. Dictionary [1838, 942, 1248, 1248]. Diday [298]. Diet [908, 330]. Difference [583, 1498, 1821, 1067, 1610, 984, 1284, 1875]. Differences [1374, 851, 1559, 1816]. Different [73, 670, 809, 411, 182, 560]. Different-Sample-Size [809]. Differential [1705, 972, 736, 1378, 282]. Differentials [325]. Differentiation [741]. Difficult [1183, 1116]. Difficulties [859]. Diffusion [1892, 200]. Digby [1462]. Digesta [1803]. Digital [1299, 1622]. Dilla [1395]. Dillon [1212]. Dilution [1689, 952]. Dimension [1744]. Dimensional [909]. Dimensionality [1049]. Diploid [312, 112, 313]. Direct [398]. Direction [905]. Directions [1434]. Directors [485, 487]. Directory [1347]. Diritto [386]. Disadvantages [927]. Discipline [715]. Discordance [1676]. Discovering [683, 1739]. Discrepancy [138, 55]. Discrete [994, 556, 1321, 910, 1774, 1809, 1248, 1454, 357, 1195, 129]. Discrete-Time [1809]. Discriminant [602, 1263, 673, 357, 78, 510, 79, 89, 1731, 556]. Discussant [329]. Discussion [1305, 1304, 335, 438, 328, 326, 331, 715, 333, 126]. Disease [1220, 1223, 577, 1276, 904, 1452, 783, 503, 1887, 150, 1544, 1198, 1316, 800, 954, 342, 986, 902, 860, 1691, 453, 330, 985, 784, 514, 1208]. Diseases [185, 74, 903, 1677, 786]. Disequilibrium [1712]. Disinfection [400]. Diskriminanzanalyse [1424]. Disorders [1771]. Dispersed [629]. Dispersion [1325]. Displacement [1137]. Display [1041, 276]. Dissection [1214]. Dissimilarity [908]. Distance [1402, 1023, 42, 350, 561, 1072, 1767]. Distance-Based [350]. Distances [362]. Distinctive [451]. Distributed [493, 1492, 157]. Distribution [1579, 1800, 1906, 1332, 1595, 811, 1497, 1105, 369, 632, 853, 624, 24, 1711, 201, 221, 637, 1363, 1768, 570, 1563, 1869, 197, 1236, 1072, 345, 1459, 1325, 1199, 1197, 1237, 1589, 351, 1153, 1447, 1596, 1486, 1335, 151, 1698, 731, 786, 1270, 790, 1561, 721, 552, 100]. Distribution-Free [24, 1363, 1768, 197, 1589, 100]. Distributional [1541]. Distributions [1033, 779, 915, 955, 606, 674, 503, 146, 850, 795, 191, 286, 1248, 1454, 1737, 1358, 83, 809, 729, 349, 1647, 1715, 17, 862, 780, 539, 476, 381, 1289, 291]. Distributions-Regression [1737]. Diversity [361, 737, 1199, 1541, 131, 731, 284]. Dixon [174, 432]. DNA [891, 1402]. Do [247, 586, 1463]. Do-it-Yourself [247]. Dobson [935]. Doctorate [764].

Dodge [535]. Doll [1871]. Domain [1254, 1256]. Domestic [285]. Donald [97, 99]. Données [600, 60, 855, 240, 473, 938, 297]. Dose [1527, 332, 1440, 1191, 1192, 183, 722, 192, 1007, 1855, 159, 1779, 1690, 584, 1365, 1278, 1201, 1505, 333, 1529]. Dose-Response [1440, 1191, 1192, 192, 1007, 584, 1365, 1505]. Doses [261, 1036]. Dosis [1515]. Double [1330, 1645, 1066, 1627]. Double- [1330]. Doubling [455]. Doubly [1708, 1197]. Doubly-Censored [1708]. Doubly-Truncated [1197]. Douglas [381]. Dowdy [879]. Draper [294]. Drawing [461]. Dropouts [1868]. Drug [972, 1457, 122, 654, 1872, 1291, 368, 84, 1036]. Drugs [217, 1867, 1283, 367, 113]. Drying [1558]. Dual [383]. Duchene [483]. Dudewicz [536, 1396]. Due [1017]. Duncan [997]. Dunn [1570]. Duo [1695]. Duo-Trio [1695]. Dupac [431]. Duration [1495, 621]. during [654]. Dyck [1160, 1877]. Dyke [1885, 1832]. Dynamical [1840]. Dynamics [649, 1464, 1298, 609, 753, 489]. Dynamiques [704].

Each [74, 1684]. Ear [1198]. Early [1220, 780, 674, 660, 1117, 779]. Eason [206]. Ebeling [1840]. Eberhart [665]. Ecological [1462, 422, 286, 1050, 200, 1383, 730, 29, 1444, 1790, 1064, 1889, 131]. Ecologie [240, 821, 240, 822]. Écologiques [240, 240]. Ecology [243, 1126, 1425, 203, 823, 1251, 1211, 241, 373, 488]. Ecometrics [243]. Econometrics [1705]. Economia [386]. Economic [1523]. Economica [386]. Ecosystem [608]. Ecosystems [287]. ED50 [1736, 1487]. Edges [922]. Edgington [469]. Edilberto [812]. Edition [1127, 1221, 1217, 1219, 1473, 1513, 1394, 1460, 1253, 1122]. Editor [437, 179, 49, 365, 982, 366, 983]. Editorial [1079, 526, 1053, 56]. Edmond [711]. Edmonton [1129]. Educated [840]. Edward [981]. Edwards [249, 610, 1701]. Effect [136, 972, 1203, 1019, 363, 851, 670, 1543, 5, 1013, 1026, 1718, 1778, 1757, 1147]. Effective [261, 1236, 1356, 1503]. Effects [967, 314, 1352, 1903, 1591, 1150, 665, 1239, 1638, 1546, 1058, 1207, 573, 1558, 744, 1851, 617, 1334, 1205, 1443, 1192, 507, 581, 740, 620, 1152, 1772, 1283, 1077, 957, 148, 514, 1904, 172]. Efficacy [1331, 400]. Efficiencies [900]. Efficiency [1361, 675, 515, 1269, 1496, 146, 1377, 1587, 458, 1640, 406, 187, 1379, 668, 1493, 1073, 1596]. Efficient [1110, 1369, 1141, 587, 455, 1448]. Effort [733]. Efron [204, 700, 1453]. Egg [402]. Eggs [96]. Ehrenberg [599, 158]. eimer [542]. Einführung [893, 1667]. Eisen [175]. Elder [1337]. Electric [611]. Elementare [98]. Elementary [1431, 373, 491]. Elements [1785, 1324, 1566, 1221, 1168, 425]. Elimination [903]. Elliott [1347]. Elliptic [977]. Elpelt [830]. Elston [1564]. Elteren [1860]. Embleton [1615]. Embryos [1026]. Emitters [1888]. Emphasis [667, 1492]. Empirical [919, 1452, 1358, 1766, 565]. Encyclopedia [1787, 1246, 833, 1040, 1210, 1737, 1787, 642, 1042]. Endocrinology [484]. Endothelial [624]. Endpoint [1808, 354]. Endpoints [968, 1438]. Endrenyi [655]. Engineering [597]. Engineers [1622]. Entries [1141].

Entropy [268]. Entry [1353, 1266]. Environment [899, 1621]. Environmental [1576, 1840, 1380, 172, 171, 889, 242, 549]. Environments [411]. Enzyme [655, 970]. Epidemic [39, 1735, 683, 461]. Epidemics [1658]. Epidemiologic [1504, 1851, 1874, 1369, 187, 1481, 1843, 1847, 710]. Epidemiological [828, 461]. Epidemiology [1388, 1086, 889, 323, 336, 1427, 1838, 942, 1887, 481, 650]. Episode [1412]. Episodic [1584]. Equal [1021]. Equality [1406, 191]. Equation [1754, 736, 1378, 1778, 1484, 158, 1685, 1755]. Equations [1892, 208, 282]. Equilibré [855]. Equilibria [662]. Equilibrium [111, 345, 25]. Equivalence [1645, 921, 679]. Erkenntnis [166]. Error [137, 1774, 1500, 26, 79, 1359, 1731, 1444, 642]. Errors [1631, 622, 218, 1504, 1034, 21, 1035, 979, 1143]. Espèces [821]. Essai [483]. Essays [834, 643, 1255, 601]. Essentials [1564]. Establish [921]. Estimate [1898, 813, 1593, 1830, 256, 636, 1415, 629]. Estimated [1556, 1644, 144]. Estimates [1173, 445, 619, 1452, 1031, 1313, 509, 219, 639, 1015, 126, 1690, 160]. Estimating [1330, 583, 1754, 919, 1501, 1609, 1410, 1483, 1762, 1281, 1643, 502, 73, 1590, 1689, 733, 853, 624, 24, 509, 737, 559, 1713, 26, 1507, 666, 1594, 1404, 47, 1030, 153, 956, 1597, 1024, 725, 1716, 1781, 17, 661, 1487, 1229, 1077, 1038, 196, 846, 1685, 1186, 1755, 1133, 1897, 220, 754]. Estimation [667, 1221, 1469, 1527, 1799, 915, 189, 222, 1595, 358, 74, 1326, 628, 263, 955, 1733, 1745, 802, 886, 663, 23, 783, 575, 503, 404, 9, 664, 1586, 454, 1191, 1608, 146, 1583, 1552, 1403, 1013, 570, 973, 1, 1736, 617, 618, 1067, 1408, 1782, 728, 723, 924, 1777, 1540, 88, 1488, 1585, 1866, 848, 729, 1161, 952, 1199, 71, 1035, 1864, 1237, 1629, 1021, 416, 1029, 1495, 1238, 1066, 1360, 1065, 976, 1063, 121, 362, 1778, 734, 1715, 978, 1806, 730, 1188, 1724, 1758, 159, 464, 1320, 794, 1447, 1731, 1103, 1654, 1542, 455, 11, 565, 400]. Estimation [584, 1014, 786, 1876, 514, 1710, 1265, 1278, 1561, 793, 1545, 1819, 1902, 1905, 1801, 1189, 1529, 1578, 1896, 1901, 199]. Estimations [1154]. Estimative [327]. Estimator [1005, 1176, 1182, 1906, 505, 1069, 1269, 1100, 1453, 637, 1587, 1825, 1104, 1379, 1099, 1875, 1596, 1486, 524, 1178]. Estimators [1033, 1110, 350, 1308, 109, 1722, 1763, 1307, 1865, 858, 342, 1599, 867, 1204, 668, 183, 1359, 977, 587, 1415, 1232, 1633, 1669]. Ethics [646]. Ethier [1430]. Eugene [469]. Eugenics [989]. Europe [888]. ’Evaluable’ [987]. ’Evaluatable [987]. Evaluate [1729]. Evaluating [150, 1416, 1596, 269, 1223]. Evaluation [1457, 521, 573, 270, 654, 1198, 1369, 156, 1266, 18, 272, 1147, 1353]. Event [1810]. Events [576]. Everitt [291, 1796, 1121, 207]. Evidence [309, 70]. Evoked [1806]. Evolution [1387, 1791, 929, 244, 703]. Evolutionary [1092, 1337, 1568]. Ewens [127]. Exact [266, 633, 1639, 1443, 1485, 1610, 921, 1532, 918, 1644, 1554, 961]. Examination [818]. Example [1133, 969, 1378, 1279, 956, 1336, 160]. Examples [1251, 1214, 428]. Excess [1769]. Exercices [528, 938]. Exercises [170]. Expectation [1197]. Expectation- [1197]. Expected [1546, 19, 1442, 1813, 43]. Experiences [59]. Experiment [14, 261, 454, 1725, 197, 952, 1864, 742, 818, 1765]. Experimental [355, 136, 900, 563, 1468, 1027, 502, 820, 5, 666, 566, 815, 1770, 209]. Experimentation [419, 1885]. Experimenting [808]. Experiments [1174, 1578, 1227, 1762, 1281, 1071, 285, 1233, 1590, 1010, 230, 40, 1377, 1403, 509, 1, 1865, 1832, 342, 1441, 1766, 581, 1075, 8, 716, 519, 130, 399, 1873, 734, 849, 1890, 1366, 1007, 909, 1724, 1073, 785, 1032, 682, 455, 786, 1505, 1512, 1745, 655, 598, 1250, 304, 380]. Expert [1627, 1520]. Explained [1620]. Exploration [1824]. Exploratory [1743, 758, 759, 696]. Exploring [1382, 1607, 1381]. Exponential [1313, 686, 1602]. Exponentially [1492]. Exposure [1098, 575, 573, 274, 121, 1649, 514]. Extended [811, 1371, 1404, 685]. Extension [568, 717, 35, 22]. Extensions [1112]. Extinction [75]. Extra [566]. Extrapolation [722, 192, 1855]. Extreme [1252, 1030, 396]. Eyes [348].

F [1090, 1345, 1392, 1434, 1704, 1623, 1573, 1577, 1621, 1841, 1612, 1833, 1741, 647, 1656, 1892, 609, 886, 30, 170, 943, 695, 706, 1422, 850, 548, 759, 1891, 547, 1381, 1382, 1701, 532, 1509, 545, 467, 125, 603, 1289, 131, 205, 419, 696, 760, 1337, 888, 208, 430, 296, 837, 533, 471, 284, 79]. Faà [1040]. Facteurs [1154, 855]. Factor [1822, 1743, 270, 905, 819, 444]. Factorial [470]. Factorially [1075]. Factors [1241, 860, 1555, 205]. Failure [239, 404, 1145, 1315, 46, 1012, 1855, 1699]. Falconer [543]. Fallible [1627]. False [846]. Familial [1276, 363, 1586, 1022, 1542, 1633]. Families [852, 951, 1454, 1647, 244]. Family [1502, 1156, 9, 853, 1737, 899, 631, 571, 1349]. Family-Size [853]. Farewell [1389]. Farris [1220]. fathers [220]. Faulbaum [1841]. Fears [943]. Features [451]. Fecundability [1507, 1270]. Federer [470]. Feinstein [1388, 1657]. Fenelon [600]. Feoli [1125]. Fertilities [1758]. Fertilization [1109, 346]. Fetal [1133, 413, 956]. Fetoprotein [413]. Fichiers [1340]. Field [1227, 1403, 1832, 1141, 659, 427, 818]. Fields [611]. Fieller [1581, 1580, 1196]. Fienberg [1285]. Fifteenth [945]. Figurale [427]. Filter [719, 1675]. Final [349]. Financial [390, 554, 768, 999, 1300, 1530, 1899]. Fingleton [1336]. Finite [75, 460, 345, 1289, 431, 459, 248, 291]. Finney [235]. First [1091, 62, 1559, 1853, 792]. First-Order [792]. First-Year [1853]. Fischer [1298]. Fish [1027, 1408, 587, 1024, 1890, 930]. Fisher [989, 1615, 232]. Fisheries [1294]. Fishery [1891]. Fit [344, 1003, 1004, 1822, 1871, 863, 1061, 584, 158, 1000, 1567]. Fits [412]. Fitted [1009, 1710]. Fitting [814, 1332, 1606, 1313, 1378, 796, 906, 1604, 845, 718, 461, 1579, 208]. Fixed [1057, 1155, 1269, 443, 793, 220]. Fixed-Sample-Size [793]. Fixed-Strata [1269]. Flamant [299]. Fleiss [1512, 301]. Flow [1803, 790, 1395]. Flowerheads [1482]. Fluctuating [705]. Fluctuation [267]. Flury [936, 1883]. Flushing [977]. Flying [423]. Foetuses [399]. Folks [529]. Follow [1110, 1353, 1057, 1019, 1013, 270, 1266, 1199]. Follow-Up [1110, 1353, 1057, 1013, 270, 1266, 1019]. Followed [1018]. Following [1728, 1490, 1777, 918]. Food [1027, 1429, 756]. Foods [377]. Forecasting [1025, 59]. Foregut [659]. Forest [736, 668, 735, 1384]. Forestry [1384, 1646]. Foreword [317]. Form [1085]. Formula [1040, 232, 634]. Formulae [43]. Formulation [1867]. Forthofer [544]. Fortran [1214]. Fortschritte [1841]. Fortson [885]. Fortunati [386]. Foss [1313]. Foucart [992]. found [272]. Foundation [1569, 826]. Foundations [372, 1618, 481]. Four [1783, 1272]. Four-Parameter [1272]. Four-Sequence [1783]. Fourfold [387]. Fourier [410, 345]. Fowler [489, 1466]. Fox [1045]. Fractiles [1298]. Francis [1345]. Fraumeni [943]. Free [24, 1363, 1768, 197, 1589, 100]. Freedman [1129, 488]. Freedom [814, 873, 1858]. French [1894]. Frequencies [572, 18, 1014, 1189]. Frequency [1254, 454, 1711, 1826, 255, 629, 1018]. Frequently [688]. Freund [167]. Friedman [940, 482, 115, 641]. Front [6, 33, 69, 105, 143, 181, 213, 253, 337, 338, 393, 497, 498, 558, 616, 658, 713, 782, 844, 895, 947, 1006, 1056, 1096, 1134, 1185, 1226, 1261, 1302, 1355, 1401, 1437, 1478, 1531, 1582, 1626, 1673, 1707, 1756, 1802, 1848]. Frontier [822]. Frontiers [1396]. Fujikoshi [1423]. Full [588]. Full-Sib [588]. Function [343, 1371, 1850, 1869, 898, 1030, 1209, 978, 925]. Functional [1557, 653, 1744, 85]. Functions [1539, 342, 1072, 1858, 718, 1409]. Fundamentals [482]. Furberg [482]. Further [805, 724]. Future [18, 214, 715].

G [1430, 173, 93, 764, 1170, 1090, 1258, 1518, 1399, 1472, 1618, 1620, 1840, 1842, 1752, 1380, 536, 931, 531, 472, 550, 243, 288, 389, 64, 1088, 606, 609, 886, 1462, 303, 297, 544, 1428, 1424, 247, 204, 593, 656, 1250, 1510, 1885, 1422, 608, 654, 1742, 548, 1047, 538, 287, 1889, 1508, 752, 882, 286, 876, 1248, 1832, 997, 206, 238, 646, 478, 1336, 131, 710, 59, 1890, 827, 885, 760, 643, 1126, 1425, 1337, 242, 1339, 831, 1426, 283, 209, 1081]. G. [91]. Gains [631]. Galilean [526]. Gallant [1788]. Gallucci [609]. Galwey [697]. Games [1159]. Gani [1255, 601]. Gap [803]. Gardiner [1221]. Gardner [665]. Garki [1627]. Gas [17]. Gaylord [246]. Gelfand [1664]. Gemmell [107, 716]. Gene [454, 345, 899, 255, 1187, 1189]. Gene-Environment [899]. Gene-for-Gene [1187]. Gene-Frequency [255]. Genealogical [541]. Genealogy [1446]. General [1397, 217, 185, 665, 1108, 911, 1504, 1418, 1720, 958, 1287, 625, 257]. Generalised [828]. Generalization [1069, 225, 639, 1099, 628]. Generalized [13, 1754, 663, 950, 1722, 796, 641, 1536, 1803, 183, 1759, 1066, 1858, 192, 1537, 721, 1685, 1755, 878]. Generation [536, 1105, 1662]. Generation-Time [1105]. Genes [1775, 1187]. Genetic [140, 1239, 1086, 1386, 541, 662, 950, 509, 38, 219, 916, 636, 1188, 1758, 1677, 650, 396, 94, 284, 1277, 141]. Genetics [647, 665, 1891, 96, 1426, 490, 1646, 1399, 1338, 1169, 543, 245, 613, 376, 127, 283, 202]. Génétique [425]. Genital [1412]. Genome [1792]. Genotype [800]. Genotypes [312, 112, 1603, 313]. Genotypic [663, 1190]. Genstat [697]. Gentle [945, 133]. Gentleman [837]. Geographic [255]. Geographical [622]. Geology [1660]. Geometric [1557, 85, 1270]. Geometry [1798, 652, 376]. George [246]. Geostatistics [426, 132]. Germination [155]. Gesichtfeld [427]. Gettinby [206]. Getz [203]. Ghosh [643]. Gibbons [1217, 1219]. Gilbert [1576]. Gill [1453, 130]. Giri [598]. Gittins [1211]. Glasshouse [1732]. GLIM [1775]. Global [1321]. Glymour [1739]. GMDH [1535]. Godehardt [1795]. Godron [821]. Goh [421]. Gokhale [61]. Golden [1297]. Goldstein [1212, 168]. Gomez [1043, 1043]. Goodman [1391]. Goodness [344, 1003, 1004, 1822, 1871, 863, 1061, 1000, 1567]. Goodness-of-Fit [1822, 1061, 1567]. Gordon [1839, 758]. Gore [646]. Gorman [208]. Govindarajulu [1469, 1835]. Gradients [186]. Graduate [1423]. Graphical [1420, 276, 881]. Graphics [1749]. Graphing [1566]. Graphs [1795, 282, 1795]. Gräsbeck [546]. Grass [1558]. Grassle [131]. Gray [1347]. Green [171]. Greenacre [1213]. Greenberg [1380]. Grid [1264]. Griffiths [1167]. Groot [533]. Gross [479]. Group [222, 523, 1491, 1728, 1494, 1155, 1817, 1490, 1777, 79, 353, 914, 918, 1372, 778, 579, 1481, 1843, 1847, 771, 486]. Group-Specific [914]. Grouped [1140, 568, 1604, 1197, 1029]. Groups [1680, 1692, 1533, 958, 914]. Growth [1174, 506, 360, 1680, 907, 797, 1010, 736, 653, 796, 953, 1310, 1681, 1009, 197, 411, 668, 205, 1679, 565, 1448, 1311, 420, 1085]. Growth-Curve [907, 797, 565]. Grundbegriffe [1170]. Grundriss [1740]. GSK [1723]. Guessing [840, 582]. Guest [437]. Guide [1467, 247, 193]. Guidelines [1375]. Gulland [930]. Gunst [471]. Günter [427]. Gupta [835, 1884]. Gurney [705].

H [387, 1095, 1347, 1297, 1392, 1393, 1572, 1705, 1840, 1839, 1753, 1841, 937, 939, 934, 936, 375, 1738, 1656, 237, 288, 1292, 1129, 701, 1460, 1424, 1887, 829, 989, 1514, 1515, 656, 295, 940, 761, 1390, 608, 893, 825, 1341, 1431, 171, 1089, 486, 1421, 1214, 838, 168, 710, 1890, 827, 166, 885, 888, 533, 488, 756, 294, 480, 830, 491]. H.-M [1841, 893]. H.-P [387]. Haas [612]. Habitable [910]. Habitat [1763, 560]. Haenszel [1176, 1178, 1005, 505, 1069, 1269, 1100, 1722, 637, 1587, 1825, 1404, 1379, 1099, 1415, 1232, 1875, 524]. Haggett [1658]. Haight [532]. Haith [549]. Hajek [431]. Hall [1834]. Hallam [1425, 1126]. Hampel [1612]. Han [530]. Hand [602, 291]. Handbook [536, 1253, 1661, 832, 1247, 1522, 1423, 552, 1221, 1254, 1256, 1470, 1571, 1049, 1247, 537]. Handbuch [830]. Hannan [1256, 601]. Haplotype [850, 1676]. Hardy [111, 1712, 25, 1485]. Harris [1886]. Harter [937, 939]. Hartl [202]. Hartley [463]. Hartung [830]. Harvested [1732]. Harvesting [108, 871]. Harvey [534]. Hauck [480, 524]. Haux [1520]. Having [1391, 74, 572, 1551]. Hawkins [290]. Hayakawa [1423]. Hazard [1177, 222, 1116, 1371, 898, 409, 1458, 1554, 1183]. Hazard-Rate [1554]. Hazards [1110, 1003, 1180, 568, 865, 1315, 1453, 964, 416, 1029, 1333, 863, 1718, 634, 84, 1101, 571, 1184]. Hazards-Ratio [416]. Hazell [1523]. Health [1172, 1350, 1525, 1569, 1706, 1895, 1380, 997, 870, 826, 302, 1888]. Healy [299]. Heart [330, 925]. Heart-Rate [925]. Heather [224]. Hedayat [470]. Hedges [1249]. Heiden [378]. Height [736, 1681, 158]. Heiser [1833]. Heitfeld [1667]. Helichrysum [1482]. Heligman [1292]. Hemophilia [1735]. Hemophilia-Associated [1735]. Henderson [1880]. Henley [426]. Hepatocyte [1067]. Herbert [1664, 292, 1344]. Heredity [989]. Heritability [1578, 1020, 1, 630]. Herman [463]. Herpes [1412]. Herrmann [831]. Heterogeneity [504, 275, 443, 564]. Heterogeneous [218, 186, 1280]. Hettmansperger [1166]. Hewlett [607]. Hiérarchiques [855]. Higher [442, 1009]. Higher-Level [442]. Higher-Order [1009]. Hill [1167]. Hillion [1475]. Hills [419]. Hiorns [649]. Historical [246, 161, 1234, 359, 408]. History [1521, 1368, 1511, 1117]. HLA [872, 1645, 850, 1676, 902, 453]. HLA-Disease [453]. Hoaglin [759, 1381, 1382, 696]. Hochberg [1881]. Hocking [1286]. Hogg [492]. Holmes [375]. Homer [371]. Homogeneity [635, 1721, 1718, 679, 1218]. Homologous [1402]. Honor [1664]. Honour [834, 643, 601]. Hooke [765]. Hormone [184, 1584]. Hormone-Receptor [184]. Horton [1397]. Hospital [888]. Hospitalization [1072]. Hospitalized [1809]. Host [1187]. Household [349]. Households [349]. Howells [1047]. Humak [1565]. Human [1338, 664, 1386, 1681, 831, 594, 1792, 205]. Human-und [831]. Humber [1704]. Humphries [1337]. Hunt [653]. Huskova [1346]. Hybridization [1775]. Hypergeometric [1459]. Hyperparameter [1629]. Hypotheses [162, 1022, 152, 1283]. Hypothesis [1469, 1040, 510, 962, 1329]. Hypothesis-Testing [510].

IBM [1349]. Icing [1042]. Ideas [529]. Identification [1002, 799, 290]. Identified [1857]. Identify [1773, 860]. Identifying [293]. Ihm [604]. II [1393, 1247, 386, 1111, 1491, 352, 1268, 1641, 1651, 1059, 1572, 649, 472]. III [166, 97, 99, 1189, 1393, 1565]. Illness [311, 900, 1784, 41, 51]. Illness-Death [311, 900, 1784, 41]. im [934, 427]. Image [651]. Immunology [128, 1339]. Immunotherapy [197]. Impact [1192, 1064, 1081]. Impagliazzo [1465]. Implementation [24, 156]. Implications [890]. Improved [1476, 567, 1414, 925, 1477]. Improving [1015]. Imputation [1747]. In-situ [1775]. Inactivation [625]. Inbred [1578, 1]. Inbreeding [852, 951, 588]. Incidence [1590, 1692, 1366, 846, 1139, 943]. Including [1108, 989]. Incompatible [662]. Incomplete [1361, 1411, 1071, 1682, 503, 1218, 1244, 88, 1724, 1017, 77, 1186, 1624, 1625]. Incomplete-Block [1361]. Incorporating [896, 405, 1686, 566, 1716, 1365]. Incorporation [1193, 1234]. Increasing [1587, 1855, 1036]. Increasing-Number-of-Strata [1587]. Independence [1822, 1419, 347, 1243, 1334, 407, 367, 81]. Independent [1400, 1899, 2, 523, 53, 809, 1610, 1241, 1607, 508, 54]. Independently [679]. Index [140, 215, 38, 1325, 1541, 216, 1335, 1710, 1208, 141, 1787]. Indices [219, 730, 690, 731]. Indirect [47]. Individual [958, 1856, 1758, 565, 226].

Individually [1322]. Individuals [1578, 1]. Induced [1591]. Industrial [698]. Industrialisés [483]. Industrie [1837, 1044, 1514]. Industries [1384]. Industry [418]. Inequalities [476, 396]. Inequality [1688, 113]. Inert [17]. Infection [1193]. Infections [1809, 349, 48, 1065]. Infectious [185, 1544]. Inference [1090, 1396, 1538, 64, 911, 530, 261, 1166, 1157, 1639, 1307, 1421, 1268, 1448, 1765, 1171, 1219, 1613, 596]. Inferences [684, 461]. Inferential [606]. Inferring [1187]. Infestation [976]. Influence [702, 865, 1500, 1536, 716]. Influential [293]. Influenza [1658]. Information [456, 106, 254, 499, 714, 948, 1135, 1303, 1432, 1479, 1674, 1849, 7, 339, 1686, 1238, 566, 1716, 1234, 359, 408, 1041, 61, 1125]. Informatique [101, 473]. Informative [1896, 1901, 1545, 1819, 1902, 1905]. Ingelfinger [1656]. Ingram [1461]. Initial [1313]. Innis [287]. Insect [1482, 976, 1106, 1588]. Instances [441]. Institute [1257, 1524, 1668]. Instrumentation [1395]. Insulin [1850]. Insurance [485, 487]. Intact [564]. Integrals [666]. Integration [1313, 838]. Intelligence [1616, 1617]. Intensity [350, 976]. Inter [88]. Inter- [88]. Interaction [1457, 1867, 1632, 633, 405, 899, 1075, 1063, 692]. Interactions [1314, 1482, 1058, 427]. Interactive [414, 314, 148]. Interactively [198]. Intercept [728]. Interchangeability [1867]. Interclass [1022, 1542]. Intercropping [1228, 1603]. Intercrops [808]. Interest [1816]. Interface [945, 1297]. Interim [1375, 1826, 353]. Intermediate [538]. Intermolt [1713]. Internally [1888]. International [1666, 1616, 1617]. Interpretation [607, 810, 1886, 1362, 1050]. Interpreted [1314]. Interpreting [474]. Interrater [1416, 1820]. Intersite [144]. Interval [1469, 1526, 341, 974, 1031, 575, 1240, 1634, 1145, 1315, 1583, 339, 460, 618, 152, 952, 1864, 1267, 1103, 459, 1547, 1278, 1529, 1644, 1844, 1528, 1846, 1173, 1527]. Interval-Censored [1145, 1315, 339, 1267]. Intervals [403, 773, 1351, 1498, 1821, 1502, 920, 1728, 1364, 340, 1418, 1556, 1363, 1768, 1020, 1865, 1490, 868, 1766, 188, 898, 1442, 1284, 1209, 897, 918, 1151, 1354]. Intervened [1153]. Intervention [364, 1503]. Interyear [154]. Intra [88]. Intra-Block [88]. Intraclass [1351, 1526, 9, 684, 1240, 347, 1280, 348, 1542, 1074, 1354, 1528]. Intrasubject [1161]. Introduction [697, 841, 1127, 1095, 436, 1420, 1121, 1425, 29, 935, 1790, 1095, 1093, 892, 176, 388, 1796, 468, 1343, 1251, 543, 382, 751, 1834]. Introductory [1571, 997]. Invariant [729]. Inverse [1501, 1281]. Inversion [232]. Investigating [1070, 914, 51]. Investigation [1721]. Invited [1304, 438, 70, 715]. Involving [660, 519, 1007, 508]. Iosifescu [248]. Iréneusz [589]. Irregular [906]. Irrelevant [1828]. Irreversible [866]. Irwin [385, 232, 638]. Ischemic [16]. Isobel [132]. Isolates [1187]. Issue [435]. Issues [825]. Istituto [386]. Italian [1574]. Iterative [1358]. IV [1888]. Izsak [468].

J [602, 239, 174, 945, 1342, 305, 432, 551, 552, 1170, 1094, 1255, 1221, 1256, 1217, 1219, 1345, 1349, 1518, 1392, 1432, 1433, 1396, 1393, 1521, 1473, 1474, 1570, 1622, 1705, 1704, 1619, 1577, 1621, 1837, 1838, 1750, 1797, 1839, 1798, 1753, 484, 698, 1512, 536, 1612, 1658, 1833, 420, 468, 1464, 1794, 595, 236, 1656, 288, 1124, 1292, 938, 1087, 1343, 597, 1088, 992, 1465, 539, 942, 1460, 1086, 297, 170, 943, 479, 941, 1653, 1424, 247, 651, 1386, 299, 429, 382, 301, 1218, 1513, 989, 881, 1746, 1044, 1514, 706, 1123]. J [295, 1510, 199, 201, 940, 1744, 750, 1390, 608, 654, 759, 613, 996, 282, 753, 1213, 1615, 995, 304, 286, 381, 876, 1168, 1248, 1341, 1381, 1382, 1563, 529, 1882, 165, 1509, 1661, 545, 1089, 887, 385, 648, 241, 467, 754, 603, 133, 1385, 1466, 1748, 130, 1082, 131, 930, 419, 380, 763, 885, 379, 643, 1337, 1660, 96, 704, 781, 708, 126, 208, 242, 837, 431, 533, 29, 601, 880, 830, 291, 1045, 594, 127, 94, 284, 428, 878, 935, 235, 884, 1081, 167, 418]. J. [93]. J.-F [603]. J.-M [1086]. J.-P [938]. J.-Y [992, 418]. Jack [1393, 1393]. Jackknife [700, 813, 1689, 559, 1593, 795, 1898]. Jacquez [1513]. Jambu [60, 1048]. James [645, 94]. Janacek [1087]. Janovy [1433]. January [390, 554, 768, 999, 1300, 1530, 1899]. Japan [1220]. Japanese [1026]. Jean [600, 1085]. Jean-Pierre [600]. Jeffers [750, 29]. Jeffreys [1095]. Jenkins [59]. Jerome [58, 324, 316, 322, 323, 321, 320]. Jerzy [980]. Jinks [613]. John [766, 1749, 373, 1746]. Johnson [1663, 1246, 833, 1040, 1210, 1737, 1787, 642, 644, 1042, 1751, 1250, 1564, 834]. Johnston [205, 539]. Joint [355, 217, 342, 1188, 367, 113]. Jolivet [751]. Jolly [72, 628, 912]. Jolly-Seber [72]. Jones [764]. Joseph [638]. Joshi [1248, 1745]. Josianne [483]. Jr [493, 204, 133, 827, 1337, 94, 1433]. Jr. [756]. Judgment [11]. Juhasz [468]. Juhasz-Nagy [468]. Juhl [648]. June [117]. Juvancz [755, 589].

K.-D [387]. K.-H [830]. Kalbfleisch [239]. Kallianpur [643]. Kallio [1384]. Kalman [719, 1675]. Kaluzny [1889]. Kanal [1049]. Kanzerogenität [1837]. Kapadia [1620]. Kaplan [738, 1767]. Kappa [1730, 1319, 35, 1878, 1117]. Karlin [1791, 1568]. Kelly [125, 1739]. Kempson [1885]. Kempton [1462]. Kendall [1798, 876, 804, 238, 81, 1413, 1563]. Kenkel [1251]. Kennedy [133]. Kent [165]. Kernel [1850, 1690, 602]. Key [1358]. Keyfitz [1293]. Kiefer [1393]. Kimura [929]. Kinetic [655]. Kinetics [1850, 1365]. Kingman [284]. Kinnison [1252]. Kish [1611]. Klagge [1349]. Klassifikation [604, 166, 166]. Kleinbaum [710]. Kleiner [881]. Klösener [830]. Klotz [1390]. Knapp [997]. Knowledge [1092, 636, 531]. Known [57, 1772, 514]. Kohler [1170]. Kolmogorov [110]. Kombinationswirkungen [1514]. Kotz [840, 1246, 833, 1040, 1210, 1737, 1787, 642, 1042]. Krafft [95]. Krishnaiah [1256, 1470, 1049, 643, 1254, 289, 537]. Kroonenberg [1215]. Kruger [387]. Krzanowski [1882]. Kshirsagar [933]. Kuh [293]. Kullback [61]. Kupper [710]. Kurtz [1430]. Kurzanski [1840]. Kutner [1460, 1089]. Kutoyants L [239, 176, 493, 764, 1347, 1221, 1344, 1296, 1297, 1349, 1391, 1517, 1522, 1434, 1397, 1467, 1393, 1472, 1435, 1436, 1702, 1577, 1752, 1512, 1611, 937, 939, 1246, 375, 243, 611, 1656, 1124, 1292, 64, 597, 655, 823, 1251, 422, 695, 833, 1040, 889, 941, 247, 476, 607, 1887, 172, 429, 301, 425, 204, 470, 1249, 199, 940, 1049, 1744, 613, 1889, 752, 995, 1210, 1655, 1613, 1737, 1787, 529, 1083, 642, 240, 540, 1051, 1662, 1466, 819, 130, 1287, 482, 710, 827, 885, 1337, 1042, 471, 832, 169, 542, 202, 1125]. Laake [199]. Laboratory [309, 285, 401, 546, 70, 321, 519, 1886, 99, 1471, 1753, 1616]. Lachin [648]. Lactate [1067]. Laerum [1395]. Lafaye [992]. Lagonegro [1125]. Lai [1344]. Lalouel [1086]. Lam [1345]. Lambert [888]. Lambeth [196]. L’Analyse [297]. Land [1220]. Lane [697]. Language [1342, 1797]. Large [1269, 1068, 1119, 637, 489, 660, 1359, 1232, 524, 43, 884]. Large- [1359]. Large-Sample [1269, 1068, 637, 524]. Large-Strata [1232]. Last [1838, 942]. Late [985]. Late-Onset [985]. Latent [1774, 1121, 1699]. Latent-Failure [1699]. Latin [805]. Lattice [230]. Launer [695, 995]. Laval [1340]. Law [1454]. Lawless [706]. LD50 [63]. Leach [388]. Least [1314, 663, 932, 1119, 1782, 46]. Least-Square [932]. Least-Squares [1119]. Lebart [1083]. Lebeaux [60, 1048]. Lebreton [704]. Lecture [1425, 1665]. Ledermann [1247, 1247]. Leeuw [1833, 1744]. Lefebvre [603, 382]. Lefort [91]. Left [1033, 121]. Left- [121]. Left-Truncated [1033]. Legendre [823, 240, 823, 240]. Lehmacher [387]. Lehmann [427]. Lehnen [544]. Lehr [830]. Lehr- [830]. Lellouch [299]. Lemeshow [302]. Length [358, 1358, 1495, 1447]. Lengths [1131, 1028, 1132]. Leonard [1090, 989]. Leroy [1612]. Leslie [1785, 1324]. Lesquoy [883]. Lessard [1791]. Lethal [1527, 159, 1278, 1529]. Leukocyte [454]. Level [442, 683, 508]. Levels [266, 575, 1267, 970, 1201]. Levin [1126, 1425]. Levine [1702]. Levitt [490]. Levy [302]. Lewi [698]. Lewis [1615, 1165]. Lexian [1002, 799]. Li [892]. Liars [765]. Lienert [472, 550]. Life [344, 403, 775, 1004, 1216, 1539, 569, 606, 520, 707, 1808, 401, 1051, 485, 1387, 487, 1000, 777, 749, 385, 1659]. Life-Table [403]. Lifetime [503, 706, 1681]. Lifetimes [402]. Light [909]. Like [1066, 950, 1813]. Likelihood [74, 263, 797, 1733, 911, 1191, 1608, 1403, 973, 729, 1035, 1115, 1021, 916, 1200, 1063, 735, 725, 1188, 1724, 1758, 1677, 862, 1447, 195, 1547, 1698, 1277, 1633, 1554, 1683, 1844, 1846]. Likelihood-Ratio [1554]. Likelihoods [1446]. Lilienfeld [481, 481]. Limit [1453, 718, 1042]. Limiting [1232]. Limits [1111, 1373, 912, 1610, 1688]. Lindeberg [1246]. Linder [98]. Lindley [1392, 533, 1295]. Lindzey [764]. Line [1033, 1182, 410, 1023, 1107, 1357, 24, 199, 561, 1104, 728, 977, 1449, 77, 1764, 201]. Line-Intercept [728]. Line-Segment [1449]. Line-Transect [561, 977]. Lineage [1309]. Linéaire [992]. Linear [27, 1394, 1397, 1901, 1112, 1538, 828, 1460, 1760, 109, 369, 932, 260, 1418, 521, 1190, 572, 1816, 801, 1020, 45, 1811, 1307, 800, 1773, 861, 1560, 1536, 1286, 1089, 864, 1421, 1759, 1455, 1287, 1858, 933, 735, 1194, 718, 18, 1272, 465, 296, 1267, 1537, 471, 1551, 1282, 878, 1819, 1902, 994, 1750, 1745, 836, 1288, 1045]. Lineare [95]. linearen [934]. Lines [460, 959, 660, 459]. Linhart [1738].

Link [1858, 1617, 1097]. Linkage [916, 1386]. Linked [665, 801, 1282]. Liquid [741]. Lists [815]. Literatur [1467]. Literature [324, 1467, 193]. Litter [1192, 864, 1367, 159]. Litter-Matched [864, 1367]. Litters [272]. Little [1653, 419]. Live [1463]. Liver [564]. Lloyd [834, 1247]. Lobsters [659]. Locally [1546]. Locate [140, 38, 141]. Location [1643, 1636]. Loci [801, 1282, 1277]. Locus [395]. Lod [1239]. Log [27, 1670, 919, 1326, 1760, 572, 801, 45, 1825, 1310, 800, 1550, 861, 1415, 1282, 1671]. Log-Linear [27, 1760, 572, 801, 45, 861, 1282]. Log-Linear-Model [800]. Log-Normal [1310]. Log-Odds-Ratio [1825, 1415]. Log-Rank [1670, 1550, 1671]. Logarithmic [1417]. Logiciels [60, 603]. Logistic [1900, 857, 1276, 1499, 1648, 1327, 1870, 1404, 1635, 412, 1272]. Logit [1108, 1487, 862]. Logits [919]. Lognormal [274, 1197, 121]. Logrank [515, 1155, 151, 258]. Long [516, 1075, 785, 1709]. Long-Term [516, 1075, 785, 1709]. Longitudinal [1754, 360, 215, 339, 1811, 507, 1161, 168, 1547, 1195, 1685, 1844, 1846, 1755]. Look [715, 62]. Lorenz [1170]. Losing [312, 112, 313]. Loss [1540]. Losses [1353, 1266, 1070]. Louis [825]. Low [332, 722, 192, 1855, 333]. Low-Dose [332, 192, 1855, 333]. Low-Dose-Rate [722]. Loynes [1789]. LR [162]. Lumped [461]. Lung [1368, 186, 1026].

M [994, 174, 173, 841, 1346, 1255, 1256, 1216, 1296, 1392, 1434, 1395, 1399, 1702, 1572, 1622, 1704, 1619, 1838, 1840, 1751, 1797, 1841, 290, 1612, 705, 97, 821, 243, 612, 828, 942, 248, 1460, 1565, 30, 1086, 60, 297, 422, 175, 299, 1215, 598, 881, 656, 1123, 1885, 761, 893, 894, 654, 481, 1463, 1288, 1213, 876, 1248, 1511, 1423, 1613, 997, 1212, 165, 1083, 1089, 1659, 648, 241, 238, 754, 603, 1385, 646, 1289, 1287, 933, 482, 59, 419, 1337, 888, 99, 533, 1048, 203, 1567, 880, 94, 283, 929, 209, 1125, 1790, 1384]. M.-O [1048]. MacDonald [993]. MacKenzie [531]. MacNeill [1518]. Made [906]. Magnitudes [636]. Magnus [1705, 1750]. Mahalanobis [139, 44]. Mahalanobis-Metric [139, 44]. Main [1816, 1443]. Maindonald [1341]. Mainland [97, 99, 97, 99]. Make [1555]. Making [1595, 182, 255, 371, 1295]. Makov [1289]. Malaria [732, 1065, 1627, 887]. Mallows [1655]. Mammal [489]. Man [1052]. Management [1383, 119, 1025, 885, 1516, 1384, 1891, 242, 1294, 421]. Mandl [1346]. Manifestation [74]. Manly [1345, 1509]. Mann [1037, 1768]. Mantel [1176, 1178, 1005, 505, 1069, 1269, 1100, 1722, 637, 1587, 1825, 1404, 1379, 1099, 1415, 1232, 1875, 524]. Manton [1428]. Manual [930, 1169, 237, 1885]. Many [914]. Mapping [1452, 1277, 1792]. Maps [255]. Marchuk [303, 1339]. Mardia [165]. Marginal [635, 1825]. Mark [1506, 1060, 1865, 394, 1540, 71, 1864]. Mark-Resighting [1506]. Marked [1034, 1063]. Marker [1223, 150, 954]. Markers [1316, 1277]. Markov [248, 1179, 1412, 1376, 723, 1316, 1016, 226, 1683, 792, 1430, 1175]. Markov-Dependent [226]. Markovian [1002, 799]. Marshall [1852]. Martens [756]. Martin [1787]. Mason [471]. Mass [783]. Matched [445, 1694, 1639, 858, 1585, 591, 1693, 864, 1367, 1322, 901, 257, 43].

Matches [43]. Matching [139, 1898, 1593, 187, 440, 271, 1017, 44, 1018]. Materials [377]. Maternal [665, 413]. Mathematical [305, 1257, 1399, 1524, 1668, 374, 649, 1465, 651, 200, 373, 488, 283, 1054, 1523, 1622, 484, 535, 1461, 175, 1085, 990, 1126, 1425, 944, 1339, 203, 127]. Mathematics [1522, 610, 547, 1247, 1793, 1216, 282, 206, 284]. Mathématiques [1475, 938, 751]. Mather [613]. Maths [249]. Mating [798, 346, 636, 1188]. Matis [608]. Matrices [20, 1312, 1021, 1551, 1246]. Matrix [797, 663, 1785, 1324, 792, 1705, 762]. Matter [6, 33, 69, 105, 143, 181, 213, 253, 337, 338, 393, 497, 498, 558, 616, 658, 713, 782, 844, 895, 947, 1006, 1056, 1096, 1134, 1185, 1226, 1261, 1302, 1355, 1401, 1437, 1478, 1531, 1582, 1626, 1673, 1707, 1756, 1802, 1848]. Matthews [1389]. Maturation [205]. Matusita [1128]. Maximally [513, 512]. Maximization [1197]. Maximizing [631]. Maximum [74, 263, 1733, 911, 1191, 1608, 1403, 201, 973, 729, 659, 1035, 1115, 1021, 1063, 1070, 735, 725, 1188, 1724, 1758, 862, 1447, 195, 151, 1277, 1633, 1548]. McCullagh [878]. McCulloch [1510]. McIntosh [484, 484]. McNemar [1845, 961, 1780]. Mead [820, 913]. Mean [775, 1557, 1821, 1539, 1728, 569, 19, 83, 1459, 1442, 1359, 976, 121, 1077, 1596, 85, 1444, 777]. Means [1173, 1821, 1011, 1031, 1279, 276, 408]. Measure [1236, 1481, 1843, 1847]. Measurement [1071, 1868, 1559, 1511, 168, 121, 1398]. Measurements [273, 310, 1133, 1253, 1682, 1678, 1774, 1861, 906, 1008, 123, 458, 956, 1812, 1275, 4, 462, 214, 1253, 1794]. Measures [1411, 1730, 590, 218, 850, 580, 1823, 1312, 1235, 1035, 1759, 1360, 958, 1724, 1828, 1449]. Measuring [263, 517, 1072, 925, 790]. Meat [96]. Mechanics [423]. Mecke [1798]. Median [773, 1527, 341, 340, 159, 897, 1278, 1529]. Medical [1665, 1136, 97, 757, 619, 170, 50, 1290, 371, 545, 646, 485, 1389, 130, 412, 1153, 99, 487, 755]. Medicine [1461, 546, 1472, 1656, 1513, 711, 944, 1427, 1054]. Meeker [493]. Meeting [117]. Meier [738, 1767]. Melamed [1395]. mellitus [1144]. Membrane [1258, 1087]. Memoranda [1616]. Memoriam [316]. Mende [1463]. Menhaden [871]. Menten [502, 1484, 1781]. Meritorious [1156, 631]. Messy [1250]. Meta [1249]. Meta-Analysis [1249]. Metabolic [903]. Metabolism [1067, 424]. Metastases [1852]. Method [1361, 1182, 574, 1410, 398, 802, 1098, 733, 1313, 24, 903, 561, 405, 1104, 1407, 47, 153, 1063, 22, 1025, 1324, 1723, 1716, 36, 1487, 1187, 925, 846, 1893, 605]. Methode [883]. Methoden [1623, 472, 550, 1565, 98]. Méthodes [60, 528]. Methodologies [205, 1292]. Methodology [1434, 1574, 911, 520, 835, 872, 279, 334, 1291]. Methods [100, 551, 492, 1054, 1221, 1299, 1217, 1296, 1467, 1526, 1576, 1796, 1839, 969, 246, 595, 1326, 1062, 42, 1461, 263, 372, 799, 422, 1682, 1155, 1233, 932, 1689, 1240, 1776, 1887, 301, 1327, 820, 881, 1420, 706, 1123, 1249, 573, 758, 1157, 673, 1047, 752, 171, 413, 439, 1212, 1509, 1012, 182, 1030, 1695, 1287, 1890, 901, 902, 678, 725, 1758, 348, 963, 944, 430, 300, 971, 272, 1696, 1761, 480, 518, 1045, 1074, 1709, 167, 1528, 1002, 1470, 1571, 1620, 389, 1343, 701, 538, 710, 930, 1086]. Metode [542]. Método [1517]. Metric [139, 44]. Metropolitan [1575].

Meulman [1833, 1218]. Meyer [1390]. Michaelis [502, 1484, 1781]. Michener [1383]. Micro [1340, 992, 603]. Micro- [1340]. Micro-Ordinateurs [992, 603]. Microbial [48, 400]. Microcomputer [1347, 1700]. Microcomputer-Based [1700]. Microcomputing [946, 1390]. Microspheres [1237]. Middle [1198]. Miettinen [1427]. Migrant [325, 326]. Migration [847, 1636]. Miké [757]. Milieu [1154]. Milk [96]. Miller [97, 204, 99, 1508, 997, 478, 704]. Millier [883]. Milliken [1250]. Minimal [790]. Minimum [1115, 1723, 862, 1187]. Minton [167]. Misclassification [670, 1504]. Misclassified [406]. Mises [915]. Missing [1669, 1653, 115, 1599]. Misspecified [964]. Misspecifying [1239]. Misused [688]. Mitosis [455]. Mixed [1352, 1150, 1497, 798, 1263, 454, 966, 346, 1068, 856, 1020, 442, 1334, 867, 1021, 1547, 1846, 1800, 1844]. Mixed-Effects [1352, 1150, 1334]. Mixed-Model [1497, 856, 1800]. Mixed-Model-Based [966]. Mixed-Variable [1263]. Mixing [146]. Mixte [855]. Mixture [915, 516, 1697, 146, 1604, 1289, 380, 1879, 1698, 291]. Mixtures [217, 1358, 78, 510, 1804, 380, 147]. ML [1066]. Mode [1215]. Model [1601, 640, 344, 1144, 1140, 311, 1003, 1004, 1055, 1352, 1397, 1800, 1901, 1903, 568, 814, 1412, 39, 185, 857, 1150, 506, 1112, 900, 72, 410, 1538, 865, 665, 1497, 907, 582, 797, 1108, 1638, 1309, 854, 789, 260, 16, 1783, 277, 1109, 1697, 966, 1145, 1315, 1418, 1608, 1809, 736, 1632, 1068, 1403, 856, 1020, 791, 442, 953, 1463, 874, 866, 1334, 803, 447, 1316, 800, 1443, 954, 41, 581, 157, 1805, 867, 898, 46, 8, 912, 691, 1021, 1026, 1333, 1829, 1853, 863, 732, 1065, 1718, 1597, 926, 1589, 722, 678, 192, 1007]. Model [1272, 108, 1532, 1856, 1781, 634, 1699, 195, 84, 1602, 1106, 625, 1187, 154, 1807, 1365, 1554, 1819, 1186, 1139, 1000, 628, 1904, 1902, 1738]. Model-Based [1807]. Model-Fitting [814]. Modele [717]. Modèles [751, 704]. Modeling [1622, 1896, 1192, 1046, 1545, 1905]. Modelisation [717, 483]. Modellbildung [1565]. Modelle [95]. Modelling [414, 484, 1482, 1060, 479, 1861, 1719, 847, 48, 500, 59, 1294, 203, 1588, 935, 705, 303, 750]. Modells [934]. Models [27, 273, 667, 1769, 1175, 1222, 1621, 1795, 217, 1630, 374, 522, 857, 360, 1276, 1326, 1451, 1271, 1506, 1023, 787, 828, 606, 1460, 1760, 1440, 1321, 941, 783, 175, 1831, 1774, 1606, 516, 706, 671, 686, 1645, 1207, 129, 572, 796, 801, 1544, 233, 561, 198, 443, 45, 1811, 1288, 1312, 1307, 394, 990, 723, 1871, 156, 1787, 384, 1852, 1869, 964, 507, 861, 1536, 1866, 1404, 186, 1803, 1604, 991, 1759, 1016, 1695, 979, 256, 80, 1873, 1858, 223, 1061, 380, 412, 1194, 1152, 379, 1232, 1757, 461, 1339, 1637, 296, 1537, 488, 1273, 1679, 957]. Models [1547, 257, 1551, 571, 1045, 1448, 1444, 462, 1282, 1505, 1685, 1683, 1755, 1687, 1844, 1179, 1846, 1221, 1754, 1248, 1840, 1788, 1121, 608, 1248, 1883, 1286, 1089, 1421, 200, 1287, 933, 534, 993, 878, 1790, 1082]. Modern [1396, 492, 1299, 1348, 1572, 695, 1887, 1423, 1431]. Modification [582, 219, 304]. Modified [72, 588, 110, 115, 660, 151]. Mogyoródi [1094]. Mohapatra [761, 761]. Mohsenin [377]. Moisture [742]. Molecular [929]. Molecules [358]. Molt [1782]. Molting [1713]. Moment [1906, 811, 1486].

Moments [1105]. Monitoring [1862, 1734, 1146, 172, 413, 1605, 1826, 719, 155, 1709, 1576]. Monocropping [1228]. Monoculture [909]. Monotone [775, 569, 777]. Monotonic [160]. Monte [605, 920, 218, 1721, 871, 862]. Montmann [831]. Moolgavkar [1887]. Moran [1095, 1786, 601]. Morbidity [725, 985, 1038]. Morgan [1385, 1168, 1384]. Morgenstern [710]. Morgenthaler [781]. Morice [938]. Morineau [1083]. Morphogenesis [990]. Morphology [651]. Morphometrics [1122, 475, 1337]. Morphometry [372]. Mortality [640, 1144, 1769, 1591, 574, 237, 1292, 622, 1428, 325, 405, 1072, 186, 1014]. Morton [1086, 639, 650]. Mosaics [224, 1157]. Moses [204]. Mosquitoes [732, 926]. Moss [753]. Most [1546]. Mosteller [1656, 759, 1381, 1382]. Mouse [519]. Mover [1608, 1829]. Mover-Stayer [1608, 1829]. Moves [1541]. Mown [1558]. MR0675315 [776]. MR0677963 [1000]. MR0786176 [1225]. MR0845630 [1581]. MR0872965 [1801]. MR0931633 [1905]. MR0963914 [1904]. MR1029611 [1902]. Much [1183, 1116]. Multi [819]. Multi-Factor [819]. Multicenter [1148]. Multicompartment [1606]. Multidimensional [791, 679, 1833]. Multidimensionnelles [603, 382]. Multifactor [1607]. Multifactorial [1055, 854]. Multigene [244]. Multigraphs [1795]. Multihit [192]. Multilocus [727]. Multimethod [1246]. Multinomial [517, 632, 1726, 1186]. Multiperiod [1245]. Multiple [1669, 1824, 1501, 266, 1062, 1108, 622, 688, 575, 689, 1496, 1725, 352, 1010, 40, 1319, 509, 114, 198, 669, 1852, 1820, 1599, 240, 1485, 1012, 1872, 1065, 1860, 687, 968, 1438, 902, 678, 724, 1696, 1311, 1646, 1174, 835, 1747, 1881]. Multiple-Comparison [687]. Multiple-Recapture [1501]. Multiple-Set [1646]. Multiple-Testing [724]. Multiple-Trait [509]. Multiplicative [787, 717]. Multiplicity [1405]. Multiprocess [719]. Multisite [82]. Multistage [1111, 619, 1364, 1869, 722, 793]. Multitrait [1246]. Multitrait-Multimethod [1246]. Multivariables [1340]. Multivariado [1517]. Multivariate [1140, 841, 1127, 1133, 1663, 907, 1107, 797, 1071, 476, 20, 474, 475, 758, 796, 570, 116, 1008, 905, 1248, 1210, 1423, 1883, 1882, 225, 1076, 1404, 644, 1242, 1021, 645, 956, 963, 296, 453, 1032, 1196, 1581, 1542, 147, 37, 356, 1275, 1151, 1644, 936, 772, 1580, 1751, 698, 1462, 422, 1743, 1422, 1047, 1122, 1212, 165, 1083, 1509, 238, 1886]. Murphy [1290, 711]. Musgrave [247]. Mutagenicity [267, 1279]. Muttonbirds [953]. Mutual [1334].

N [1430, 697, 1221, 1395, 1472, 1619, 377, 1246, 595, 1293, 1892, 1251, 1462, 1086, 544, 833, 1040, 598, 820, 201, 1049, 761, 750, 1047, 1122, 709, 1891, 1615, 1210, 1613, 1737, 1787, 642, 887, 648, 424, 300, 1042, 294, 29, 650, 880, 94, 993, 307, 850]. Naddeo [1574]. Nagy [468]. Nakache [938]. Nalimov [1387]. Namboodiri [1659]. Natality [1588]. National [1253, 1888, 1429, 1792]. Natural [1221, 1305, 1304, 1368, 1716, 929, 220, 1189, 1306, 989]. Naturally [465]. Naus [201]. Near [1272]. Near-Linear [1272]. Nearest [194]. Nearly [1865]. Neave [491]. Need [586, 231]. Needed [354]. Neff [201]. Negative [1579, 1332, 1595, 811, 1733, 732, 926, 793]. Negatives [846]. Neighbour [1403]. Neighbouring [1227, 592]. Neighbouring-Plot [592]. Neighbours [194]. Nelder [878]. Nelson [707]. Neoplasia [486]. Neoplastic [309, 70]. Nest [1589]. Nested [1143]. Net [1600]. Neter [1460, 1089]. Netherlands [1703]. Networks [76, 993, 790, 378, 125]. Neudecker [1705]. Neural [378]. Neurocommunications [1093]. Nevo [1568]. Newby [282]. Newell [760]. Newer [890]. News [31, 67, 103, 135, 178, 211, 251, 308, 391, 434, 495, 555, 615, 657, 712, 769, 843]. Newton [1753]. Next [435]. Neyman [941, 980, 749]. Nicht [604]. Nicht-Numerische [604]. Nigam [1884]. Nigeria [1627]. Nigerian [158]. Nisbet [705]. Nishisato [383]. No [1645, 633, 1000, 776, 1904, 1354, 1528, 1178, 1625, 1260, 1671, 1179, 1225, 1066, 1132, 1801, 1581, 1846, 1847, 1529, 1902, 1905, 1755]. Nomograms [354]. Non [100, 1400, 236, 2, 547, 87, 1662]. Non-Independent [1400, 2]. Non-Parametric [100, 236]. Non-Physical [547]. Non-Responders [87]. Non-Uniform [1662]. Nonadditivity [1858]. Nonagreements [669]. Noncompliance [1353, 1266]. Nonconstant [1069]. Nondecreasing [1682]. Nondimensional [398]. None [1055, 854]. Nonforced [582]. Nonhomogeneous [442]. Nonindependent [1669, 1599]. Noninteractive [1108]. Nonlethal [1113]. Nonlinear [667, 522, 857, 506, 21, 233, 156, 80, 1152, 108, 1757, 1365, 137, 1788, 1046]. Nonmelanoma [943]. Nonparametric [1769, 388, 1182, 1296, 1470, 1682, 1642, 1233, 503, 404, 261, 1783, 340, 570, 1202, 1823, 1104, 1720, 342, 1859, 1495, 1360, 362, 873, 440, 978, 1812, 794, 518, 584, 786, 1201, 773, 1217, 1219, 426, 1287]. Nonproportional [1453]. Nonrandom [363]. Nonreplicated [1556]. Nonresponders [1203]. Nonresponse [1747]. Nonstratified [1070]. Nonuniform [1353, 1266]. Nonzero [114]. Normal [552, 1173, 1800, 1497, 1728, 1031, 146, 1310, 1236, 152, 678, 147]. Normally [493, 157]. Normally-Distributed [157]. Normals [1879, 1698]. Norman [834]. North [117, 186, 1385]. Norton [1523]. Norwegian [325, 326]. Note [27, 179, 49, 1680, 410, 23, 86, 232, 1037, 1313, 1552, 1825, 1869, 1830, 8, 47, 976, 1032, 1877, 518, 1074, 1201, 1400, 2]. Notes [1665, 97, 1425, 99, 1521]. Nouveaux [883]. Nucleotide [1711]. Null [962, 1698]. Number [1057, 536, 632, 1319, 1587, 741, 1533, 848, 660, 1141, 1541, 47, 354, 1187, 629, 43, 220]. Numbers [754, 272]. Numerical [853, 1313, 666, 838, 1839, 823]. Numérique [240]. Numerische [604, 166, 604]. Nunez [611]. Nutrient [1429]. Nutrition [890].

O [1395, 1471, 1570, 1576, 539, 60, 548, 240, 486, 1421, 1048, 1427]. Oakes [480, 1084]. Obituaries [92, 743]. Obituary [91, 58, 93, 463, 589, 1786, 981, 1880, 812]. Objective [36]. O’Brien [1160, 1877]. Observation [1684]. Observational [931, 1369, 1772]. Observations [1669, 1499, 1822, 503, 364, 221, 1773, 115, 1599, 157, 1805, 720, 957, 975, 226, 1422]. Observed [632, 1645, 1066, 258]. Observer [958]. Observer-Agreement [958]. Observers [1319]. Observing [1409]. Obtaining [1313, 912]. Occupancy [676]. Occupational [573]. Occur [1450]. Occurrence [16, 1711, 1852, 1691, 1427]. Odds [13, 1176, 1326, 1069, 1586, 1100, 572, 637, 973, 149, 1825, 858, 1721, 1204, 1415, 1178]. Odds-Ratio [637]. Off [904]. Offerings [279]. Official [1836]. Offspring [1156, 1634, 1190, 1359, 631, 220]. Offspring-Parent [1190]. Ohta [244]. O’Keefe [1349]. Okubo [200]. Olaf [95]. Oliviera [547]. Oliviera-Pinto [547]. Olkin [1393, 1249, 1852]. Ollason [753]. Ollivier [425]. Olshen [940]. Omitting [1333]. Onchocerca [847]. Onchocerciasis [846]. One [1549, 352, 630, 1652, 1858, 444, 1372, 258, 680]. One-Parameter [1372]. One-Sample [352, 258]. One-Sided [1549]. One-Tailed [1652]. One-Way [680]. O’Neill [287]. Only [1546]. Onore [386]. Onset [503, 1869, 1584, 725, 985, 1038, 786]. Open [1061, 1188, 1854]. Open-Pollinated [1188]. Operating [1642]. Operational [384]. Ophthalmology [1776, 1560, 348, 963]. Opinion [896]. Opportunities [870]. Optima [1824, 682]. Optimal [511, 746, 1408, 1628, 740, 1779, 508, 190, 148, 1372, 1651, 314, 1471]. Optimale [95]. Optimisation [298]. Optimization [1297, 1796, 303, 871, 549]. Optimum [1158, 362, 108, 1827]. Ord [286, 1658, 876, 1563, 241, 379]. Order [969, 937, 939, 1052, 805, 10, 1009, 631, 396, 792, 292]. Ordered [1140, 1391, 1321, 1037, 572, 1419, 1768, 856, 1279, 1773, 1456, 115, 921, 362, 1273, 11]. Ordinal [13, 635, 1601, 1098, 1120, 1649]. Ordinateurs [992, 1340, 603]. Ordination [1050]. Organisms [73, 952]. Organization [1172, 1350, 1525, 1706, 1895, 955, 548, 1172, 1350, 1525, 1706, 1895]. Oriented [1808]. Orloci [1251, 422, 1125]. Ornithologists [1466]. Ornithology [1385]. Osborn [170]. Oscar [638]. Other [314, 1221, 1276, 700, 1888, 963, 148]. Otis [64, 752]. Otolaryngology [1560]. Ott [1386]. Otto [463]. Outcome [1598, 1828, 457]. Outcomes [1195]. Outlier [1612, 260]. Outliers [1773, 290, 1165]. Outline [650]. Output [451]. Overall [79]. Overdispersion [1329]. Overlap [1711]. Ovulation [770, 264, 402, 776]. Oxnard [1052].

P [173, 697, 1342, 387, 553, 764, 1170, 1090, 1095, 1346, 1298, 1254, 1256, 1299, 1395, 1470, 1393, 1523, 1571, 1619, 484, 698, 1380, 1612, 1658, 468, 1741, 1464, 243, 611, 938, 64, 1340, 606, 823, 1462, 932, 607, 604, 1215, 881, 101, 1166, 199, 1049, 894, 1889, 752, 286, 834, 1167, 1248, 1884, 528, 125, 991, 1385, 1291, 537, 1739, 302, 1287, 131, 1890, 696, 643, 1834, 888, 242, 890, 601, 490, 878, 1081, 167]. Pacific [1345]. Packages [1349, 451]. Pai [530]. Pair [591]. Pair-Matched [591]. Paired [1374, 1553, 221, 1719, 581, 1370, 963, 457]. Paired-Comparison [581]. Paired-Data [963]. Paired-Sample [1374]. Pairing [227]. Pairs [802, 910, 1693]. Pairwise [1243]. Paksy [755]. Panchapakesan [835].

Panulirus [659]. Paolo [386]. Paper [1304, 438, 70, 715, 839]. Papers [32, 68, 104, 142, 180, 212, 252, 315, 392, 496, 557, 435, 1393, 117, 525, 693, 126, 601, 1344]. Paradox [1787]. Parallel [1107, 1494, 77]. Parallel-Group [1494]. Parallel-Line [1107, 77]. Parameter [667, 1271, 1733, 1689, 1364, 1013, 1768, 1272, 1806, 1372, 793, 1221]. Parameterization [82]. Parameters [1498, 811, 1239, 502, 328, 853, 325, 1583, 509, 1556, 219, 394, 729, 349, 912, 1688, 1597, 1272, 1781, 565, 886]. Parametric [1769, 100, 236, 339]. Parent [1634, 1190, 1359]. Parent-Offspring [1634, 1359]. Parental [1446]. Parents [1156, 631]. Parsimonious [1060]. Part [1461]. Partial [1314, 562, 1536, 1099]. Partially [1669, 404, 1599, 83]. Partially-Complete [404]. Partially-Truncated [83]. Particles [741, 47]. Particular [454]. Parts [1572]. Paruchuri [289]. Pascal [1390, 1701]. Passage [1237]. Patch [803]. Patch-Gap [803]. Patel [552]. Path [899]. Pathogen [1187]. Pathogens [265]. Pathways [1387]. Patient [1353, 1058, 504, 1266]. Patients [1223, 1057, 1546, 150, 1809, 354]. Patil [173, 243, 606, 286, 1248, 131, 242]. Patrick [1786]. Patten [608]. Pattern [224, 1049, 949, 953, 1085, 1336, 913]. Patterned [797]. Patterns [969, 42, 265, 1406, 163, 1063]. Pays [483]. PC [1349]. PCB [1267]. Pearce [1885, 818]. Pedigree [190, 1338]. Percentage [1698, 155, 1554]. Perfectly [1543]. Performance [1453, 916]. Perfusion [17]. Period [310, 1259, 1148, 1457, 806, 1783, 744, 415, 1149, 617, 518, 4, 1274, 1548, 1260]. Periodically [1368]. Periods [1713]. Permutation [679]. Perpendicular [1023]. Perry [998]. Personal [321, 107]. Personality [1398]. Perspective [1743, 1882]. Perspectives [483, 375, 545]. Pertinent [1467]. Peschel [1463, 1840]. Pest [1025]. Pesticides [612]. Petersburg [1787]. Petrie [1665]. Ph [821]. Pharmaceutical [418]. Pharmacokinetic [1457, 655, 666]. Pharmacokinetics [1629, 654]. Pharmazeutischen [1837, 1044, 1514]. Phase [1408, 1771, 1111, 1491, 352, 1818, 1641, 1651, 1059]. Phenology [1106]. Phenomena [547]. Phenotypic [1823, 1360]. Photosynthesis [610]. Physical [606, 1047, 547, 205, 250]. Physics [1221]. Piecewise [718, 1602]. Piecewise-Linear [718]. Pielou [1050]. Pienaar [885]. Pierce [1786]. Pierre [600]. Pig [564]. Pike [232]. Pimentel [475]. Pinson [429]. Pinto [547]. Pla [1517]. Plackett [1210, 1737, 607, 540]. Plan [855]. Plankton [73, 587]. Planning [1236, 455, 931]. Plans [562, 1189, 700]. Plant [420, 653, 795, 1085, 660, 636, 909, 1187, 1426, 653]. Plant-Mating [636]. Plants [1109, 1087]. Planung [934]. Platek [1619]. Plausibility [331]. Plot [1410, 230, 1559, 592, 1710]. Plots [1227, 1407, 1420]. Pocock [996]. Point [770, 1177, 1469, 969, 189, 904, 264, 974, 1308, 858, 1777, 409, 1063, 1336, 1103, 1409, 1554, 776]. Point-Counting [189]. Points [201, 906, 411, 508, 1698, 1554]. Poisson [1846, 774, 1002, 1844, 799, 671, 3, 1279, 1871, 1325, 1199, 979, 1153, 629, 408, 1547]. Policy [890]. Pollard [1292]. Pollen [1839, 735]. Pollinated [1188]. Pollock [1890]. Pollution [1576, 612]. Polychotomous [1163]. Polymorphic [560]. Polymorphisms [662]. Polynomial [1281, 1009]. Polysampling [781]. Polyspermy [500]. Polytomous [1163]. Pooling [1543, 1018]. Population [1330, 1897, 1821, 1483, 1762, 798, 609, 1251, 73, 1368, 75, 1851, 1763, 376, 1408, 753, 1450, 1594, 1864, 1629, 991, 256, 977, 362, 734, 108, 1758, 488, 455, 1106, 127, 202, 258, 1186, 1752, 431, 1220, 1464, 647, 1129, 1293, 1891]. Population-Based [1851]. Population-Doubling [455]. Populations [1173, 649, 1501, 1031, 1034, 346, 116, 1713, 186, 1238, 519, 734, 1716, 1761, 629, 154, 220, 1189, 173, 1475, 421, 705, 64, 835, 703, 199, 489, 752]. Porter [1511]. Possible [950]. Post [816, 578, 1637]. Post-Release [1637]. Post-Stratification [816, 578]. Potency [785, 1644]. Potential [411, 1806]. Potentials [1063]. pour [938, 1340, 60, 855]. Power [1353, 1643, 1694, 3, 1454, 1266, 960, 1225, 1070, 1537, 1828, 1677, 457, 961, 774, 1224]. Power/Sample [1537]. Powerful [1546]. Practical [372, 1789, 996, 1700, 781, 1046, 236, 1468, 59, 132]. Practice [1472, 646, 698, 1123, 1510, 131, 818]. Practitioners [1375]. Prakasa [1221]. Praxis [604]. Pre [816, 892, 937]. Pre-1950 [937]. Pre-Program [892]. Pre-Stratification [816]. Precision [1331, 509]. Preclinical [974]. Predator [1463]. Predatory [612]. Predicting [1361, 954]. Prediction [456, 770, 264, 441, 18, 242, 1779, 1444, 214, 776, 932]. Predictive [1734, 328, 327, 1863, 82]. Predictor [1635]. Predictors [521]. Preface [437]. Preference [1027]. Preferences [228]. Prentice [1124, 889, 239, 1887, 1370]. Preparation [1340]. Prerandomization [675]. Preschool [360]. Presence [773, 1896, 1901, 1203, 340, 950, 570, 740, 1012, 1280, 1209, 407, 1413, 1545, 1819, 1902, 1905]. Presentation [525, 693]. Presented [117]. Press [645]. Pressure [485]. Pretreatment [1494]. Prevalence [1113, 1015, 683, 1627, 1771]. Prevalences [904]. Prevalent [1480]. Prey [1463]. Price [738]. Prichner [647]. Priest [995]. Priestley [1255]. Primate [375]. Primates [1052]. Primer [1050, 599, 1889, 1509]. Principal [1215, 905, 1883, 1445]. Principales [1517]. Principes [553]. Principles [119, 710, 1427, 428, 1472, 288, 1882, 202]. Prior [456, 1686, 636]. Priority [641]. Probabilistic [1387]. Probabilities [583, 403, 919, 327, 10, 1600, 984, 196, 1328, 1765]. Probability [312, 1095, 1348, 1518, 1111, 343, 1734, 75, 112, 1597, 643, 1532, 1856, 155, 1764, 220, 313, 1094, 1095, 1247, 532, 597, 941, 476]. Probability-of-Being-in-Response [343]. Probable [642]. Probit [1108, 114, 225]. Probits [1804]. Problem [770, 264, 1119, 1628, 1072, 1405, 416, 396, 776]. problemorientierte [893]. Problems [246, 606, 676, 200, 461]. Procedure [1181, 1476, 15, 352, 903, 233, 856, 791, 559, 1414, 1404, 1872, 256, 1860, 1200, 80, 1102, 1489, 730, 724, 1234, 1765, 1477]. Procedures [1799, 288, 1253, 783, 835, 1327, 110, 1043, 191, 864, 687, 968, 1320, 679, 1801, 1396, 1881]. Proceedings [945, 1129, 1126, 203, 1345, 1666, 374]. Process [311, 1896, 1784, 1105, 847, 41, 1545, 1905, 995]. Processes [1430, 1518, 1308, 1717, 62, 248, 379, 1602, 1409, 1568, 1221, 1255, 1258, 1834].

Processing [1432, 1703]. Product [493, 1453]. Production [970]. Productivity [610]. Professionals [302]. Progenies [1188]. Progeny [854, 1142, 18, 1055]. Program [544, 892]. Programing [836]. Programm [894]. Programme [1059]. Programmes [783]. Programming [1342, 1523, 535, 1431]. Programs [174, 764, 1214]. Progress [1097, 30]. Progressive [1768]. Projections [1407]. Promoting [870]. Proneness [721]. Proof [1183, 1116, 1458]. Propertie [140]. Properties [377, 1557, 606, 38, 672, 964, 1541, 1677, 731, 1633, 1139, 141]. Proportion [1330, 146, 1450, 661, 1827]. Proportional [1003, 568, 222, 865, 1315, 964, 898, 1333, 863, 1718, 634, 571]. Proportional- [898, 571]. Proportional-Hazards [568, 964, 634]. Proportions [1301, 1502, 74, 511, 523, 1374, 515, 746, 53, 1456, 868, 1206, 1560, 1610, 1455, 1284, 1329, 399, 1607, 1532, 1234, 359, 54, 301]. Propriety [691]. Prostate [405, 1205]. Protection [1253]. Protective [1555]. Provenance [810]. Puberty [1584]. Public [1380, 544]. Publications [324]. Publish [839]. Published [32, 68, 104, 142, 180, 212, 252, 315, 392, 496, 557, 435]. Pukkila [1666]. Pulse [1810]. Punktprozessen [1623]. Puntanen [1666]. Puri [1884, 1287]. Purity [661]. Purposes [586].

Q [493]. Quadrat [737, 949]. Quadratic [1281]. Quadrats [629]. Quadrature [666]. Quail [1026]. Qualitative [138, 55, 1058, 1515]. Quality [1808, 567, 1709, 1893, 1794]. Quality-Controlled [567]. Quality-of-Life-Oriented [1808]. Quantal [907, 261, 607, 1805, 1688, 1873, 1007, 1487]. Quantal-Response [261]. Quantile [1407]. Quantile-Quantile [1407]. Quantiles [403, 624]. Quantit [720]. Quantitative [1217, 1399, 243, 1169, 543, 1041, 1336, 710, 885, 283, 1277, 425, 1704, 609, 1426, 94, 1790]. Quarterly [1172, 1350, 1525, 1706, 1895]. Quasi [1733, 1683]. Quasi-Likelihood [1733, 1683]. Quaternary [1839]. que [600]. Query [273, 588, 810, 230, 521, 85]. Qu’est [600]. Qu’est-ce [600]. Question [646]. Questionnaires [670]. Queueing [760]. Queues [1221]. Quick [809]. Quotes [766, 1473, 766, 1473]. Quotient [1067].

R [239, 173, 1342, 305, 176, 306, 492, 765, 1092, 1170, 1298, 1254, 1256, 1299, 1348, 1432, 1519, 1520, 1395, 1397, 1470, 1523, 1472, 1663, 1705, 1704, 1576, 1619, 1620, 1750, 1797, 484, 702, 1612, 420, 649, 705, 375, 1084, 288, 1124, 1461, 289, 1087, 64, 372, 883, 828, 1465, 1388, 62, 1462, 30, 422, 544, 943, 695, 889, 1252, 1653, 607, 1887, 299, 1041, 204, 989, 820, 1788, 1123, 295, 699, 1510, 199, 1885, 653, 940, 1049, 761, 475, 750, 762, 1390, 1789, 1122, 1085, 287, 1288, 1508, 753, 752, 990, 995, 1431, 1793]. R [171, 997, 546, 1212, 1286, 371, 998, 545, 473, 644, 1211, 540, 241, 467, 605, 478, 424, 293, 537, 1739, 1890, 885, 643, 1337, 944, 1564, 294, 471, 1567, 29, 169, 594, 490, 1657, 428, 491, 1790, 1384, 1081, 167, 418]. Rabies [1464]. Rade [1522]. Radial [561]. Radial-Distance [561]. Radiation [1253, 573].

Radio [1715]. Radio-Tagged [1715]. Radioactivity [1253]. Radioimmunoassay [280]. Radioligand [1272]. Radon [1888]. Rahmen [934]. Raktoe [470]. Ralley [536]. Ralph [754]. Ramakumar [1465]. Ramifications [1375]. Ramsey [247]. Random [493, 1903, 536, 802, 798, 1638, 1868, 346, 1207, 442, 1443, 507, 581, 1872, 1662, 1152, 1757, 957, 1448, 1904, 62]. Random-Coefficient [1448]. Random-Effect [1757]. Random-Effects [1207, 507, 957]. Random-Pairs [802]. Randomization [1174, 1010, 87, 1493, 1070, 1812, 262, 739, 1311, 384, 469]. Randomized [52, 335, 20, 1375, 1202, 578, 648, 353, 1193]. Randomly [775, 569, 777]. Randomness [163]. Rank [138, 1001, 1400, 1670, 965, 2, 449, 55, 1816, 1450, 672, 1550, 1159, 864, 1267, 258, 677, 1671]. Ranked [11]. Ranking [835, 1141]. Ranks [845, 1166]. Rao [1619, 1086, 422, 643, 880, 1081, 1221, 1256, 874]. Rare [1771, 576]. Rarefaction [361]. Rasch [1296]. Rate [1558, 79, 722, 1855, 925, 1554]. Raters [669]. Rates [640, 1305, 1304, 726, 1680, 787, 1113, 1590, 301, 586, 34, 671, 1034, 1816, 45, 617, 1825, 618, 866, 26, 1782, 1317, 1540, 186, 979, 153, 223, 1366, 859, 1193, 1731, 1106, 1014, 846, 628, 1306]. Ratings [1601]. Ratio [1110, 1002, 1173, 1181, 1373, 1502, 1326, 1318, 797, 1031, 799, 1321, 1586, 1100, 1583, 637, 1020, 973, 1763, 1825, 618, 1008, 858, 1721, 868, 1442, 1204, 1610, 416, 1029, 916, 1284, 22, 1415, 1102, 1677, 1791, 1698, 155, 1554, 1178, 1176]. Ratios [13, 1538, 1069, 149]. Ratkowsky [1046]. Ratnaparkhi [1248]. Reached [683]. Reaction [1892]. Reaction-Diffusion [1892]. Reactions [1187]. Read [1787, 552, 1246, 833, 1040, 1210, 1737, 1787, 642, 1042]. Readable [249]. Readings [1894]. Real [161, 401, 998]. Real-Time [998]. Reality [1082]. Realized [509]. Reappraisal [803]. Reason [1092]. Recapture [1501, 1062, 1271, 1506, 72, 1483, 1762, 1154, 1760, 1060, 23, 1865, 394, 752, 1540, 71, 912, 1864, 256, 223, 734, 849, 1061, 1696, 195, 1765, 1890]. Receiver [1642]. Receptor [184]. Recessive [950]. Recessives [1066]. Recognition [1049]. Recombination [727]. Reconstruction [1446]. Recoveries [153]. Recovery [1062, 911, 1853, 1637]. Recurrent [1223, 1412, 150]. Recursive [1773]. Red [789, 1197]. Reduction [139, 1822, 1049, 44, 1744, 599, 480]. Rees [1618]. Reference [583, 1800, 1497, 402, 546]. Reflect [815]. Reflected [897]. Reflexive [194]. Regards [883]. Regier [761]. Region [117, 923]. Regions [771, 1859, 682, 579, 778]. Regression [1314, 273, 414, 344, 1144, 137, 1004, 1394, 1900, 1612, 1499, 1326, 1648, 865, 883, 622, 701, 86, 1371, 1327, 1145, 1420, 671, 1190, 21, 940, 233, 745, 198, 460, 959, 1870, 1737, 1773, 964, 1404, 411, 1089, 1280, 1718, 80, 1635, 902, 1194, 1684, 465, 859, 1757, 634, 1534, 1077, 1273, 294, 1046, 459, 1547, 462, 501, 1064, 160, 1592, 1683, 1687, 1000, 1844, 1362, 1846, 471, 1570, 702, 295, 1742, 293, 827, 471, 167, 992]. Regressions [442]. Regressionsgerade [829]. Regressive [1276]. Regressograms [1787]. Regressograms-St [1787]. Regulation [932]. Reid [749]. Rejecting [1328]. Rejective [1476, 1414, 1477]. Rejoinder [329]. Relate [1621]. Related [1173, 1578, 243, 951, 886, 1031, 1, 1883, 1766, 1272, 880, 1045]. Relation [686, 721]. Relationale [1667]. Relations [584]. Relationship [1557, 79, 85, 565]. Relative [675, 1144, 1769, 162, 445, 726, 787, 523, 1452, 1269, 1774, 278, 1406, 1587, 1763, 1830, 1379, 668, 183, 1024, 636, 902, 1194, 464, 257, 1644]. Relative-Risk [445, 902, 257]. Relatives [798, 346]. Release [1890, 1637]. Release- [1890]. Reliability [263, 995]. Remarks [319]. Reminiscence [321]. REML [1403]. Removal [752, 849, 1716, 1103]. Renal [719]. Renewable [1516]. Repair [1267]. Reparameterization [140, 38, 141]. Répartition [483]. Repeated [273, 310, 1057, 1411, 590, 1071, 1682, 1678, 1868, 218, 1861, 1559, 580, 906, 1008, 1312, 458, 1815, 1035, 1015, 1367, 1759, 1647, 1812, 1724, 1679, 1449, 1275, 4, 462]. Repeated-Measures [590, 218, 580, 1312, 1724]. Repeatedly [1732]. Replace [52]. Replacement [1156]. Replacing [631]. Replication [660]. Report [390, 554, 768, 1300, 1530, 1899, 365, 982, 366, 983, 63]. Reports [1616, 1617]. Reproducibility [1729]. Reproductive [1376]. Reprot [999]. Resampling [700]. Research [551, 764, 1222, 828, 1888, 699, 384, 923, 321, 716, 334, 710, 300, 756, 1429, 1427, 1059, 1792, 1611, 97, 757, 879, 1388, 530, 170, 1043, 646, 1045, 1383]. Research-Doctorate [764]. Resemblance [899]. Residual [967, 775, 1539, 569, 1403, 777]. Residuals [1773, 1727, 702]. Resighting [1506, 628]. Resistance [342]. Resistant [1809, 412]. Resources [1516]. Respond [136, 1546, 5]. Responders [87]. Response [1111, 343, 229, 1824, 974, 1440, 688, 261, 1191, 1192, 1238, 519, 1873, 1241, 192, 1007, 682, 1679, 1690, 957, 584, 1036, 1365, 1505, 144, 1311, 1139]. Response-Curve [688]. Response-Surface [682]. Responses [1107, 1281, 1321, 86, 607, 1010, 1207, 1492, 1820, 157, 1007, 159, 1174]. Restricted [779, 263, 674, 1635, 780]. Restrictions [1104, 1182]. Results [1105, 853, 146, 1068, 724, 356]. Résumé [1174, 1175, 1176, 1180, 1181, 1182, 1183]. Rethink [231]. Retrospective [1327, 196]. Reversibility [125]. Review [1430, 602, 239, 994, 174, 173, 697, 945, 1342, 100, 305, 176, 249, 306, 432, 551, 387, 552, 388, 492, 493, 553, 764, 765, 766, 840, 841, 1091, 1092, 1127, 1093, 1090, 1094, 1128, 1171, 1054, 1346, 1298, 1347, 1255, 1254, 1256, 1344, 1216, 1295, 1299, 1348, 1217, 1219, 1220, 1345, 1296, 1258, 1297, 1349, 1391, 1394, 1517, 1522, 1392, 1432, 1433, 1434, 1519, 1520, 1395, 1396, 1397, 1467, 1470, 1393, 1398, 1399, 1523, 1521, 1471, 1472, 1473, 1435, 1469, 1474, 1475, 1436, 1570, 1663, 1702]. Review [1703, 1569, 1664, 1667, 1618, 1571, 1574, 1665, 1572, 1622, 1705, 1666, 1704, 1575, 1576, 1619, 1620, 1623, 1573, 1577, 1621, 1837, 1749, 1838, 1840, 1795, 1796, 1750, 1751, 1797, 1839, 1893, 1894, 1798, 1842, 1752, 1753, 484, 698, 1380, 1512, 290, 536, 702, 931, 1611, 1658, 377, 755, 1833, 749, 420, 937, 939, 374, 531, 95, 934, 421, 649, 705, 468, 472, 550, 936, 1246, 1740, 1741, 1464, 97, 821, 1794, 246, 375, 1052, 595, 1738, 98, 1338, 647, 1084, 243, 236, 535, 611, 1656, 612, 244, 1169, 237, 700, 288]. Review [389, 757, 1124, 1292, 1461, 289, 1129, 938, 1087, 1343, 64, 600, 1468, 597, 879, 1088, 1745, 883, 992, 1340, 828, 1293, 1892, 539, 942, 1388, 62, 248, 606, 701, 1460, 655, 609, 886, 823, 1251, 1462, 1565, 303, 423, 30, 285, 1086, 530, 767, 297, 422, 170, 544, 943, 1253, 1888, 61, 695, 1165, 479, 833, 1040, 889, 941, 1252, 1653, 1428, 1424, 247, 651, 932, 175, 652, 476, 607, 835, 1887, 543, 1386, 299, 172, 541, 604, 829, 429, 382, 245, 301, 1041, 1218, 1215, 1743, 1513, 425, 707, 598, 117, 703, 63, 204, 593, 989]. Review [820, 292, 881, 1420, 1746, 1044, 1514, 1515, 1788, 706, 656, 946, 1123, 101, 295, 470, 610, 699, 1043, 1166, 1250, 1249, 1510, 1583, 199, 1885, 653, 129, 1121, 207, 201, 474, 940, 1049, 1744, 761, 1290, 475, 758, 836, 1422, 426, 750, 762, 1390, 608, 893, 894, 654, 1789, 1742, 378, 548, 481, 759, 825, 1047, 1122, 613, 996, 128, 376, 1463, 282, 1085, 709, 538, 1891, 287, 1288, 1889, 1508, 753, 1213, 489, 752, 477, 990, 882, 1615, 995, 304, 286, 381, 547, 876, 834, 1167, 1168, 1285, 1210, 1341, 1381, 1423, 1431, 1382, 1655, 1563, 1701, 1613, 1832, 1737]. Review [1793, 469, 171, 384, 529, 997, 1747, 546, 891, 439, 1883, 1882, 532, 1212, 165, 1083, 1509, 1661, 642, 877, 1884, 751, 528, 822, 1286, 371, 998, 545, 473, 644, 826, 1089, 1050, 1211, 486, 887, 540, 1659, 206, 385, 648, 241, 467, 125, 238, 645, 1421, 1051, 200, 754, 991, 1662, 1614, 603, 1214, 838, 133, 1385, 646, 824, 1289, 1700, 1291, 1466, 605, 478, 168, 485, 839, 1336, 1389, 1748, 424, 819, 130, 293, 537, 1082, 1836, 1739, 302, 1383, 1287, 131, 373, 933, 1120, 482, 710, 59, 930, 1890, 205, 419, 380, 827]. Review [711, 763, 1886, 696, 166, 1881, 885, 250, 379, 534, 760, 643, 1126, 1425, 1337, 1660, 427, 1834, 725, 96, 704, 781, 708, 599, 126, 944, 888, 208, 430, 1516, 242, 300, 1339, 831, 296, 1229, 1294, 99, 1564, 132, 818, 549, 837, 431, 533, 488, 596, 756, 1042, 1429, 1387, 1048, 203, 383, 1617, 294, 1046, 1654, 1567, 29, 650, 1427, 1791, 832, 480, 1426, 487, 890, 169, 601, 892, 880, 830, 291, 298, 1045, 1566, 594, 542, 127, 94, 283, 284, 202, 490, 929, 993, 1568, 307, 1657, 209, 428, 491, 878, 935, 1835]. Review [235, 1125, 1790, 1792, 884, 1384, 1081, 167, 418]. Reviews [1170, 1095, 1172, 1221, 1257, 1350, 1518, 1524, 1525, 1706, 1895, 1612, 372, 1465, 60, 483, 1248, 1247, 1511, 1787, 240, 1211, 386, 1616, 471]. Rey [603]. Reyment [1122]. Rhythm [845]. Rhythmometry [1628]. Richardson [1082]. Richness [813, 559, 794]. Riedl [1092]. Riedwyl [936, 829]. Right [773, 1896, 1901, 340, 110, 1030, 677, 1545, 1819, 1902, 1905]. Right-Censored [110, 677]. Rijckevorsel [1744]. Rind [542]. Rindskopf [1702]. Ring [1853, 1807]. Ring-Recovery [1853]. Ringing [153]. Ripley [1614, 824]. Rise [1511]. Risikoextrapolation [1837]. Risk [1223, 1704, 1373, 162, 445, 889, 575, 1831, 150, 278, 270, 193, 1585, 188, 1830, 486, 902, 860, 1194, 1856, 402, 1875, 464, 1555, 985, 257]. Risks [1452, 1888, 1205, 1600, 861, 1012, 725, 1699, 1038]. Ritual [869]. Rizvi [1572]. Robbins [1344]. Robert [839]. Robertson [285]. Robinson [250]. Robson [173]. Robust [1181, 1893, 448, 350, 1308, 759, 959, 1872, 734, 1102, 280, 730, 397, 1612, 1612, 1613]. Robustness [401, 458, 781, 1090, 1296]. ROC [1609]. Roche [205]. Rochel [934]. Rock [659]. Rodent [1725].

Roger [888]. Rogers [1577, 946]. Rohatgi [1171]. Rohlf [467]. Role [328, 520, 1584, 168, 1772]. Rome [374]. Ronchetti [1612]. Rose [1790]. Rosen [30]. Rosenzweig [243]. Roskam [1398]. Rossum [654]. Roth [1519]. Rothamsted [716]. Rounded [90]. Rousseeuw [1612]. Route [781]. Routine [585]. Roux [1248]. Roy [1880]. Rubin [1653, 1747]. Rubinstein [605]. Rule [526]. Rules [917, 896, 1863]. Run [1131, 1028, 1132]. Runs [1131, 71, 1028, 1160, 1877, 1132]. Russwurm [756]. Rustagi [1343]. Rybova [1087]. Ryman [1891].

S [1430, 173, 176, 840, 841, 1221, 1345, 1297, 1394, 1519, 1666, 1575, 1573, 1749, 1796, 1797, 1894, 1798, 702, 705, 1246, 535, 612, 1343, 879, 828, 423, 61, 479, 833, 1040, 889, 607, 835, 1887, 543, 429, 881, 1510, 1885, 1121, 761, 836, 426, 762, 548, 825, 996, 287, 1889, 882, 1285, 1248, 1210, 1511, 1737, 1787, 469, 891, 642, 822, 645, 646, 302, 1126, 1425, 96, 781, 599, 208, 296, 818, 1042, 383, 471, 1427, 1791, 480, 169, 291, 1566, 490, 1568, 1081, 351]. Sachs [1467, 832]. Sacks [1393]. Sacrifice [311, 1233, 41, 342, 1366, 786]. Safety [1183, 1116, 1458]. Salem [880]. Salmon [71]. Sample [1180, 1353, 1224, 1573, 1897, 227, 1821, 965, 1643, 1374, 73, 232, 1269, 1780, 53, 352, 624, 1453, 1068, 1492, 637, 633, 268, 1825, 858, 447, 1450, 905, 1721, 672, 1206, 1236, 1266, 1671, 1550, 1594, 809, 1830, 685, 47, 416, 1688, 960, 1455, 1359, 1322, 1723, 634, 1537, 1101, 621, 731, 54, 524, 1275, 457, 43, 1151, 793, 258, 1323, 1845, 1225, 1184, 1301, 1670]. Sample-Size [232, 685, 634]. Sample-Sizes [633]. Sampled [312, 112, 313]. Samples [1903, 1591, 1498, 1638, 513, 949, 1029, 968, 1647, 678, 1268, 11, 679, 628, 1904]. Sampling [1330, 140, 1518, 1595, 1501, 363, 1023, 42, 1357, 853, 1327, 199, 1851, 38, 737, 1416, 1408, 752, 728, 732, 1627, 1025, 862, 190, 514, 1764, 141, 1189, 173, 171, 302, 430, 431]. Sampson [1432]. San [421]. Saporta [297]. Sarndal [1619]. SAS [1342, 893]. Satisfying [1454]. Save [385]. Scale [1601, 1283]. Scaling [383]. Schachtel [1170]. Scheiber [1741]. Scheines [1739]. Schemes [853, 190]. Schlesselman [708]. Schmitt [548]. Schools [767]. Schubö [894]. Schwager [1510]. Schwartz [299]. Science [945, 101, 129, 1836, 869, 837, 601, 715]. Sciences [1518, 1380, 1246, 833, 1040, 520, 282, 1210, 1737, 1787, 642, 1042, 1384, 388, 1221, 1216, 1397, 1569, 597, 606, 997, 826, 206, 130, 1383, 307]. Scientific [531, 606, 1248, 486, 839, 1616, 1617, 1090, 385]. Scientists [1622]. Score [1239, 686, 950, 964, 1718]. Scores [1419]. Scott [754]. Scotto [943]. Screen [619, 1872]. Screened [1368]. Screening [1451, 904, 783, 903, 1815, 82, 1814, 226]. Screens [266]. Search [791, 384, 1842]. Searching [1779]. Searle [1510, 762, 169]. Season [1584]. Seasonal [1810]. Seasonality [1692]. Seber [72, 628, 886, 1422, 912, 153]. Second [1666, 969, 295, 10]. Second-Order [969, 10]. Secondary [1002, 799]. Sectional [339]. Sections [626]. Sedransk [1081]. Seed [661, 155]. Seed-Purity [661]. Seeds [1709]. Segment [1449]. Selected [100, 1344, 1668, 989, 513, 512, 1257, 1524, 1257, 1524]. Selecting [1752, 904, 7, 835, 1806]. Selection [27, 140, 1572, 577, 229, 852, 951, 588, 1263, 1156, 989, 1632, 75, 509, 233, 38, 219, 874, 660, 788, 8, 79, 89, 1238, 80, 1230, 631, 508, 1731, 1426, 396, 395, 141, 1738, 285]. Selective [227]. Self [798, 1109, 662, 1552, 346]. Self-Contained [1552]. Self-Fertilization [1109, 346]. Self-Incompatible [662]. selon [483]. Semiparametric [1145, 1852, 1366]. Sen [1470, 1380, 834, 1287]. Sensitive [1180, 1101, 1184]. Sensitivity [619, 1146, 270]. Seppal [1384]. Sequence [1783, 1456, 891]. Sequences [1402, 364, 1016, 1179, 1175]. Sequencing [1792]. Sequential [779, 771, 1469, 15, 1549, 1491, 1728, 674, 1155, 1817, 903, 1490, 1777, 152, 1826, 1025, 353, 763, 1193, 190, 918, 1372, 155, 579, 1262, 1709, 780, 778, 1189]. Sequentially [1476, 1414, 1779, 1477]. Sequentielle [1741]. Serfling [305]. Serial [1223, 150, 580, 952, 957, 214]. Serial-Dilution [952]. Series [137, 306, 1095, 1255, 1254, 1256, 1753, 811, 410, 1689, 117, 230, 21, 1766, 1024, 860, 534, 680, 1683, 174, 876, 59, 250]. Serra [651]. Serum [413]. Set [1533, 11, 1646]. Sets [1539, 627, 858, 790, 884]. Several [627, 263, 573, 1030, 1770, 1201]. Severity [74]. Sex [1002, 665, 799, 851, 1791]. Sex-Linked [665]. Sexe [483]. Shape [1182, 1104, 626]. Shapes [397, 1382, 1381]. Shaping [1082]. Shapiro [479, 825]. Shared [1827, 1813]. Shellfish [1158]. Shift [1643, 36]. Shirley [1201]. Shorter [366, 983]. Should [52, 815]. Shukla [1472]. Sib [588, 851, 661, 1446]. Sib-Sib [851]. Sibship [664, 1633]. Sibships [190]. Side [594]. Sided [1549]. Siegel [695]. Siegmund [1344]. Sightability [1761]. Sigler [1087]. Signals [1806]. Signed [1727]. Significance [1057, 266, 1734, 1717, 921, 1405, 1367, 1360, 1598, 1273, 1814, 1823]. Silverman [1654]. Similarity [1898, 1593, 1813]. Simon [1291]. Simple [1769, 53, 618, 847, 809, 1863, 863, 1455, 256, 1241, 1607, 1003]. Simplified [783, 367, 1103]. Simpson [246, 1199]. Simulated [161, 1698]. Simulation [1622, 1496, 1068, 1889, 871, 1234, 1299, 1751, 1168, 1614, 605]. Simulation-Based [1496]. Simuliids [847]. Simultaneity [1064]. Simultaneous [1011, 1418, 1778, 1151]. Simultaneous-Equation [1778]. Singer [100]. Singh [1573, 1619]. Single [1897, 260, 1865, 1594, 1493, 395]. Single-Consent [1493]. Single-Mark [1865]. Single-Recapture [1865]. Singly [1180, 1101, 1184]. Singular [1551]. Sinusoid [845]. Siotani [1423]. SIR [1667]. SIR/DBMS [1667]. Sire [521, 18]. Site [1714, 1710]. Sites [1852, 1141, 790]. situ [1775]. Situations [963]. Six [805]. Size [1353, 1224, 1301, 1897, 1821, 1501, 1483, 1762, 1374, 232, 1357, 853, 1780, 624, 201, 1492, 1763, 1721, 1206, 1266, 1594, 809, 1830, 685, 1864, 960, 256, 1322, 734, 634, 1537, 629, 621, 154, 1633, 793, 1323, 1186, 1845, 1225]. Sizes [1670, 523, 73, 53, 633, 1550, 809, 1455, 1723, 54, 1275, 1671]. Skellam [93]. Skewed [795]. Skewness [1583]. Skin [943, 846]. Skin-Snip [846]. Skipworth [1088]. Slope [1318, 1008]. Slopes [443]. Slowly [1806]. Small [1301, 1591, 1498, 1453, 1068, 1319, 513, 446, 858, 447, 1450, 905, 672, 1206, 1236, 1830, 1864, 47, 1688, 126, 1283, 914, 1014, 1151, 30, 1619]. Small- [905]. Small-Sample [1453, 1068, 858, 447, 672, 1236, 1688, 1151]. Small-Scale [1283]. Smirnov [110]. Smith [1289, 131, 1337, 1342, 1298, 1392, 1622, 232, 489, 838, 533, 294]. Smoking [1778]. Smolt [71]. Smooth [718]. Smoothers [1812]. Smoothing [1850, 1264, 1806]. Snedecor [389]. Snell [1661, 428]. Snip [846]. Social [388, 1216, 1397, 531, 606, 129, 1045]. Society [117, 485, 487, 134, 177, 210, 390, 554, 433, 494, 614, 768, 842, 999, 1130, 1300, 1530, 1672, 1899, 102, 66]. Software [432, 1841, 1347]. Soil [742]. Sokal [467]. Sole [1827]. Solomon [1664]. Solution [1119]. Some [246, 1112, 1308, 676, 872, 146, 439, 336, 870, 668, 1803, 1652, 412, 678, 451, 272, 396, 731, 438, 326]. Sound [1636]. Source [275]. Sources [293]. Sozialwissenschaften [894, 831]. Space [1387]. Sparse [1762, 1119, 1013, 1232]. Späth [1214]. Spatial [42, 1308, 1717, 224, 1556, 949, 1307, 186, 163, 623, 1063, 1691, 914, 1409, 307, 1658, 241, 824, 1336, 379]. Spatially [622, 1106]. Spatz [539]. Spearman [639, 183, 1413]. Special [583, 402]. Specialities [1461]. Species [1621, 813, 910, 662, 559, 268, 848, 794, 1813, 1187, 731]. Species-Diversity [731]. Specific [153, 1684, 108, 914, 786, 628]. Specificity [619, 1146]. Spectra [735, 909]. Spectral [1811]. Speech [955]. Speech-Organization [955]. Spells [721]. Sperm [402]. Spheres [882]. Spherical [1615]. Spirtes [1739]. Spline [1371]. Splines [1265, 1588]. Split [230, 1559]. Split-Plot [230, 1559]. Spriet [1291]. Spring [117]. Square [369, 932, 1776, 513, 1115, 512, 862]. Squared [1359, 1444]. Squares [1314, 663, 1119, 19, 805, 1782, 1533, 1442, 46]. SRI [1616, 1617]. Srivastava [841, 1472]. St [1787]. Stability [1823, 1228, 1360, 1603]. Stabilized [1906, 1486]. Stable [1236, 519]. Stablein [827]. Stage [1451, 917, 1068, 1279, 660, 1872, 1489, 81, 1770]. Stahel [1612]. Stallard [1428]. Stand [351]. Standard [1034, 381, 258]. Standardization [441]. Standardized [1452, 618, 1317]. Standardizing [586]. Stands [736]. Stanley [757]. Staquet [1123]. State [546, 1575]. Statements [390, 554, 768, 999, 1300, 1530, 1899]. States [1316, 764, 943]. Stationary [974, 250]. Statistic [1670, 1780, 874, 641, 1550, 1117, 81, 151, 1548, 1323, 1845, 1671]. Statistica [386, 386]. Statistical [239, 994, 432, 1127, 1347, 1299, 1219, 1345, 1296, 1518, 1396, 1467, 1469, 1663, 1571, 1572, 1751, 1246, 1227, 726, 1461, 955, 1343, 1457, 1460, 50, 1165, 676, 833, 1040, 1776, 1887, 872, 1109, 1788, 656, 129, 1500, 1535, 265, 1544, 268, 538, 1307, 1248, 1210, 1511, 1423, 1737, 1787, 171, 413, 439, 1083, 1509, 642, 321, 691, 644, 334, 1700, 1748, 1484, 727, 348, 126, 626, 451, 1828, 272, 818, 596, 1042, 625, 518, 601, 1036, 1045, 594, 1074, 491, 935, 1097, 1128, 1171, 1349, 1576, 1611, 595, 389, 64, 606, 530, 170, 479, 1653]. Statistical [301, 820, 706, 1043, 1166, 1249, 1510, 1615, 304, 286, 1341, 891, 1421, 133, 1289, 96, 300, 480, 307, 1835]. Statistician [120]. Statisticians [117, 120, 118, 765, 1088]. Statistics [945, 1342, 249, 388, 492, 1091, 1172, 1346, 1254, 1256, 1257, 1350, 1392, 1470, 1472, 1524, 1525, 1570, 1664, 1705, 1706, 1668, 1895, 1136, 1380, 220, 937, 939, 1305, 1304, 288, 289, 1468, 597, 539, 767, 328, 247, 520, 101, 1049, 761, 513, 477, 876, 834, 1167, 1511, 1431, 1563, 1661, 620, 206, 512, 646, 537, 1767, 631, 888, 1189, 1654, 832, 396, 585, 428, 320, 1257, 1524, 1703, 1247, 305, 176, 841, 1296, 1257, 1520, 1521, 1524, 1702, 1618, 1574, 1665, 1666, 1619, 1796, 1612, 236, 535, 1252, 593, 292, 295, 1290, 762]. Statistics [1789, 1508, 1285, 529, 877, 545, 824, 1389, 99, 837, 533, 1081, 1094, 1297, 1435, 1620, 531, 757, 879, 699, 882, 877, 1385, 646, 1466, 643, 1660, 880, 209, 235, 418, 1306]. Statistik [1841, 1740, 1467, 936, 830, 894, 830]. Statistik-Programm-System [894]. Statistik-Software [1841]. Statistikprogramme [831]. Statistique [938, 883, 101]. Statistiques [382, 528, 603]. Statistische [1623, 95, 98, 1741, 1474, 1565]. StatLab [1700]. Status [923]. Stayer [1608, 1829]. Steel [288]. Steinber [126]. Stemming [798]. Step [1110]. Stephens [1567]. Stepwise [233, 80]. Stereological [1307, 372, 372]. Stereology [109, 1307]. Steroid [486]. Stigler [1511]. Stimulating [1097]. Stimulation [1026]. Stirling [1700]. Stiteler [422]. Stochastic [1221, 1518, 16, 733, 736, 1805, 1866, 1803, 1629, 125, 427, 1856, 871, 1827, 1106, 1798, 1614]. Stock [930]. Stone [940]. Stoodley [1091]. Stopping [779, 1549, 674, 917, 896, 1193, 780]. Storage [1709]. Stoyan [1221, 1798]. Strata [1269, 1543, 1587, 1787, 1232, 1018]. Strategies [385, 822]. Stratification [1353, 1562, 1266, 816, 578, 1637]. Stratified [1648, 149, 1408]. Stratum [1158]. Strauss [1337]. Strobeck [1129]. Stroup [840]. Structural [1795]. Structure [563, 1861, 1558, 240, 815, 1188, 1724, 541, 1739]. Structured [703, 1119, 1312]. Structures [606, 1307, 1750]. Stuart [876, 1563]. Studi [386, 386, 386]. Studies [1110, 1223, 445, 1648, 1439, 622, 1309, 1060, 1098, 688, 575, 1586, 1694, 1327, 1245, 150, 1543, 1851, 1874, 270, 1639, 446, 618, 1369, 666, 1316, 1236, 187, 591, 1693, 1161, 788, 259, 962, 1360, 168, 1070, 1322, 901, 1231, 1489, 271, 1230, 1854, 1772, 708, 1283, 1649, 272, 1791, 625, 480, 621, 1270, 1059, 1481, 1323, 1384, 1843, 1847, 931, 828, 300]. Study [311, 343, 1457, 218, 670, 1606, 161, 1109, 76, 973, 401, 1811, 447, 41, 1540, 1495, 485, 1234, 487, 1018, 325, 326]. Studying [678, 1444]. Subject [577, 574, 1867, 1077]. Subject-by-Formulation [1867]. Subject-Years [574]. Subjects [136, 5, 1319, 958]. Subpopulation [1034]. Subsequences [1711]. Subset [1546]. Subsets [1900, 1058, 1552, 1870]. Substitute [691]. Subtle [1376]. Successive [975]. Suchindran [1659]. Sukhatme [890]. Sum [1533]. Sum-of-Squares [1533]. Summaries [815]. Summary [343, 34]. Sums [1400, 2]. Supplement [1431]. Supplementary [1262]. Surface [974, 682]. Surfaces [1824, 144]. Survey [1703, 1573, 732, 1714]. Surveys [1158, 587, 1024, 1356, 1771, 1696, 1761, 1747, 430, 1429]. Survie [1154]. Survival [1144, 403, 1110, 311, 1353, 1084, 341, 1019, 726, 787, 515, 1708, 1233, 1371, 516, 1034, 221, 504, 1406, 339, 874, 1719, 191, 447, 1720, 1640, 1316, 1600, 41, 342, 1540, 1266, 1159, 620, 898, 259, 57, 1367, 1030, 153, 1209, 1070, 1767, 223, 1589, 1715, 1890, 1366, 827, 978, 354, 897, 84, 621, 257, 571, 786, 1265, 258, 628, 478]. Survival/Sacrifice [311, 1233, 41, 342, 1366, 786]. Survivor [1124].

Survivors [516]. Survivorship [45]. Susanne [205]. Susceptibility [265]. Sváb [551]. Swartzman [1889]. Swath [1558]. Swimming [423]. Sylvester [1123]. Symmetric [1817]. Symposium [945, 374]. Synergism [1162]. Synthesis [789]. Synthetic [1231, 1854, 1258, 126]. Syski [62]. System [567, 1827, 894]. Systems [1520, 1840, 950, 1228, 549, 29, 1444, 303, 128, 287, 1384].

T [1430, 1127, 1090, 1344, 1666, 536, 595, 535, 244, 992, 530, 943, 1165, 652, 1515, 295, 470, 1166, 825, 489, 1615, 1168, 1248, 1511, 1423, 546, 165, 887, 1385, 419, 1126, 1425]. Table [403, 387, 1659, 1051]. Tables [1257, 1224, 1301, 1524, 1400, 1669, 1668, 563, 2, 232, 1037, 1419, 1722, 347, 633, 791, 149, 1235, 1650, 1381, 1382, 1686, 406, 1206, 1599, 685, 1204, 1099, 960, 1607, 1532, 1225, 61, 491, 1752, 1088, 838]. Tadikamalla [1299]. Tag [1540, 628]. tag-resighting [628]. Tagged [1715]. Taguchi [1893]. Tailed [924, 1652]. Taillie [606, 286, 131]. Tails [1042]. Tails-Limit [1042]. Tales [539]. Tamhane [1881]. Tampere [1666]. Tan [1613]. Tango [1335, 1208]. Tarone [1234]. Taste [582]. Tau [804]. Taux [1154]. Taxonomic [361, 1577]. Taylor [1794, 1454]. Teaching [1521, 1789, 1700, 119, 767]. Technical [1616, 1617, 1465]. Techniques [189, 1461, 1659, 480, 1277, 1835, 479, 654, 1567, 832, 384]. Technology [877]. Tell [765]. Temperature [924, 36, 402]. Temperature-Based [924]. Temporal [241, 1691]. Temporary [1710]. Teratological [399, 1007, 1505]. Teratology [1192, 272, 1561]. Term [516, 1075, 785, 1709]. Terms [1105, 1359, 735, 1139]. Terrestrial [754]. Terry [581]. Test [138, 1003, 1180, 1131, 1476, 227, 965, 582, 1728, 797, 267, 1643, 515, 1146, 232, 1037, 1817, 1780, 110, 950, 1816, 265, 633, 580, 1279, 1414, 1202, 905, 1456, 115, 197, 620, 660, 1141, 1485, 1325, 79, 863, 916, 1860, 1028, 1370, 407, 873, 1160, 1879, 1193, 367, 1677, 897, 81, 1101, 1877, 1698, 918, 913, 155, 1548, 1201, 258, 1554, 1845, 961, 1477, 1132, 1184, 1845, 63]. Test-Based [897]. Testing [635, 344, 1000, 1055, 1177, 1469, 536, 1203, 266, 1011, 1631, 1374, 854, 1868, 1113, 851, 260, 352, 1058, 1419, 347, 1243, 1406, 1712, 1279, 191, 231, 1022, 510, 1857, 1560, 25, 409, 1015, 921, 1405, 1367, 1066, 1142, 1329, 1726, 255, 1241, 902, 807, 1598, 724, 1078, 1649, 1503, 359, 408, 1036, 1810, 395, 679, 1004, 1040]. TESTRAND [536]. Tests [779, 1669, 1057, 1591, 577, 162, 1539, 777, 569, 1822, 1546, 674, 676, 1155, 1364, 689, 111, 810, 966, 55, 278, 686, 1377, 1645, 1823, 26, 1450, 1720, 1721, 672, 1443, 1490, 1777, 406, 1599, 964, 1159, 1815, 962, 87, 163, 1863, 1360, 1718, 1652, 1767, 1061, 1812, 1073, 1267, 1137, 1138, 1273, 1234, 1814, 269, 584, 1147, 258, 780, 775, 469]. Tetrachoric [681, 875]. Textbook [452]. Their [1796, 1892, 248, 1711, 1659, 168, 565, 890, 731]. Theorem [1580, 1196, 1581]. Theorems [305, 1042]. Theoretic [1549]. Theoretical [372, 30, 1431, 1714, 1427, 1791]. Theorie [604, 528]. Theories [429, 1475]. Theory [1128, 1399, 1664, 649, 530, 61, 835, 1510, 1068, 128, 876, 1563, 1214, 131, 1834, 351, 296, 818, 283, 929, 1790, 721, 320, 1095, 306, 1348, 1518, 1469, 1842, 1563, 760, 1568, 1095, 1573, 1213]. Therapeutic [648, 1162]. Therapies [1808]. Therapiestudien [709]. Therapy [974]. There [87, 1030]. Thermal [400, 377]. Thibodeau [1656]. Thinking [1511]. Third [870, 869]. Thoday [94]. Thom [990]. Thompson [94, 1338, 1124, 541]. Thought [1082]. Thoughts [1652]. Three [227, 1457, 806, 117, 460, 630, 679, 1215]. Three-Mode [1215]. Three-Period [1457, 806]. Three-Sample [227]. Threshold [1440, 1456, 1876]. Throughout [767]. Tiku [1613, 1296]. Timber [885]. Time [239, 137, 1095, 1255, 1254, 1256, 341, 1019, 1784, 1105, 404, 1606, 1145, 1315, 1809, 21, 906, 394, 876, 924, 1869, 986, 46, 1873, 59, 897, 1387, 84, 455, 784, 621, 257, 786, 1208, 680, 1765, 1683, 628, 652, 250, 306, 998, 534]. Time- [876, 1765]. Time-Dependent [1784, 1606, 84]. Time-Points [906]. Time-Series [680, 250]. time-specific [628]. Times [1753, 1155, 277, 40, 1866, 1012]. TIMESLAB [1753]. Tissue [1267, 625]. Titterington [1289]. Tolerance [1859]. Tomassone [883, 473]. Tome [298]. Tong [476]. Tool [1605, 167]. Topics [880]. Torrie [288]. Total [1495]. Toxicants [314, 148]. Toxicity [1108]. Toxicological [519]. Toxicology [1440]. Tracking [215, 216, 690, 214]. Traffic [1597]. Training [120, 118, 117]. Trait [509, 18, 1277]. Traitement [240]. Traits [1055, 1276, 854, 1142, 508]. Trampisch [1424]. Transect [1033, 1182, 410, 1023, 1357, 24, 199, 561, 1104, 977, 1764]. Transfer [277, 144]. Transform [345]. Transformation [1001, 449, 1781]. Transformation/Weighting [1781]. Transformations [1417, 1108, 1420]. Transformer [1267]. Transient [16]. Transition [792]. Transitions [954]. Transitivity [1027]. Transmission [349]. Transplants [719]. Transport [1087]. Trapping [1103]. Trawl [1024]. Treated [1450]. Treatment [136, 1591, 15, 857, 222, 1203, 1457, 1546, 1783, 1058, 5, 1640, 1205, 1857, 921, 1077, 927]. Treatment-Induced [1591]. Treatments [806, 805, 1770, 457]. Tree [1807]. Trees [1534, 940, 993]. Trend [1113, 1559, 1455, 359, 408]. Trends [1428, 1198, 1381, 1382, 1560]. Trial [1862, 1148, 917, 966, 578, 648, 1778, 692, 1193, 354, 457, 1262, 1274, 1548]. Trials [1361, 583, 675, 771, 774, 1259, 1670, 1057, 1111, 15, 857, 222, 1549, 14, 52, 335, 1491, 1734, 1494, 1553, 322, 117, 352, 896, 1123, 3, 1632, 120, 1375, 122, 415, 1149, 1492, 118, 825, 996, 1490, 279, 152, 1550, 157, 1141, 334, 1495, 119, 368, 353, 1438, 724, 1077, 262, 1818, 1641, 1814, 1770, 739, 585, 4, 1372, 1651, 579, 1260, 1671, 778, 310, 299, 1291, 482, 763]. Trials-Issues [825]. Triangle [582, 1701]. Triangular [771, 1695, 579, 778]. Tribute [318, 107]. Trieste [1126]. Trimmed [1487]. Trio [1695]. True [358]. Truncated [1033, 83, 1604, 1197]. Truncation [75, 508]. Tsay [418]. Tufte [1041]. Tukey [1749, 881, 759, 1381, 1382, 781]. Tumor [1591, 343, 1113, 1590, 1725, 40, 1317, 1869, 197, 1366]. Tumor-Growth [197]. Tumor-Induced [1591]. Tumors [1113]. Turnover [866]. Two [583, 1110, 310, 314, 493, 1173, 1180, 1259, 1498, 1821, 1502, 1539, 360, 222, 74, 965, 1451, 1148, 1439, 1457, 1717, 1643, 1031, 1642, 806, 917, 1783, 670, 1780, 146, 1645, 1068, 415, 1149, 673, 116, 1279, 959, 866, 1408, 1450, 1443, 868, 809, 867, 1610, 1863, 1872, 79, 984, 416, 1029, 1284, 1241, 692, 1489, 909, 690, 1879, 1193, 367, 1771, 508, 81, 459, 1101, 518, 1770, 1698, 54, 457, 113, 148, 4, 1274, 1548, 1845, 1260, 1184]. Two-Compartment [866]. Two-Dimensional [909]. Two-Group [79]. Two-Period [310, 1259, 1148, 1783, 415, 1149, 518, 4, 1274, 1548, 1260]. Two-Phase [1408, 1771]. Two-Sample [1180, 965, 1643, 1780, 1450, 416, 1101, 457, 1845, 1184]. Two-Stage [1451, 917, 1068, 1279, 1872, 1489, 81, 1770]. Two-Step [1110]. Two-Treatment [1457, 1783]. Two-Treatment-Group [222]. Two-Way [1439, 1443, 867, 692]. Tygstrup [648]. Type [404, 1768, 1268]. Types [1181, 73, 560, 1102]. Typologique [429].

U [1221, 1667, 1575, 1289]. Uehlinger [1841, 893]. Ullah [701]. Ultrasound [1133, 956]. Umphrey [1518, 1518]. UMPU [686]. Unbalanced [1020, 1312, 1443, 892, 1448]. Unbiased [109, 521, 1307, 1865, 728, 18]. Uncatchable [1330]. Uncensored [1180, 1101, 1184]. Uncertain [840, 978]. Uncertainty [1511]. Undergoing [798]. Understanding [1389, 759]. Unequal [1181, 1483, 20, 1102]. Unfinished [1436]. Unified [1630, 351, 1046]. Uniform [1662]. Uniformity [623]. Unimodal [1139, 1621]. Unique [1775]. United [764, 943]. Units [910]. Univariate [580, 1248, 914, 721]. Universe [1436]. Università [386]. Universities [1097]. Unkelbach [1515, 1514]. Unknown [1764]. Unknowns [567]. Unrelated [852]. Unreplicated [230]. Untersuchungen [934]. Upton [1336]. Usage [1763]. Use [456, 1471, 969, 1498, 343, 1452, 1155, 516, 1697, 1376, 1816, 458, 1009, 1693, 685, 1860, 849, 860, 631, 1877, 359, 408, 518, 54, 155, 1687, 1362]. Used [1235, 412]. Useful [762]. User [1882]. Using [139, 311, 775, 1001, 1400, 1227, 1501, 1062, 569, 449, 7, 247, 1689, 670, 1774, 161, 1327, 671, 1552, 1034, 572, 737, 559, 949, 45, 847, 1692, 1773, 1636, 41, 406, 1533, 152, 581, 1560, 898, 952, 46, 1629, 1367, 845, 722, 1812, 1188, 44, 871, 351, 1775, 1429, 1813, 1273, 81, 84, 11, 625, 1265, 1548, 1561, 1323, 1588, 1609, 2, 777, 1389]. Utilization [1124, 1238]. Utter [1891]. Uwe [378].

V [492, 764, 1171, 1295, 1392, 1570, 1749, 1842, 1741, 757, 938, 609, 767, 61, 1165, 1249, 1885, 474, 1085, 287, 1248, 1832, 165, 1884, 838, 1389, 885, 831, 431, 533, 596, 1387, 890]. Vajda [836]. Validation [1632, 1445]. Validatory [88]. Validity [187, 678]. Vallin [1292]. Value [1252, 1605, 1863, 82]. Values [1650, 546, 121, 1532, 1876, 214]. Vandaele [480]. Vandermeer [373]. Varga [468]. Variability [1630, 1305, 1304, 925, 1306]. Variable [1057, 1410, 1263, 1562, 1121, 233, 874, 804, 79, 1264, 80, 1356, 465, 725, 1106, 1877, 1038, 1413, 1633]. Variable-Area [1356]. Variable-Grid [1264]. Variables [635, 1140, 556, 1112, 1098, 1263, 369, 357, 78, 510, 620, 89, 1635, 849, 271, 351, 1649, 1731, 1551, 493]. Variance [1005, 1176, 1570, 919, 505, 563, 1631, 502, 717, 664, 1100, 20, 1020, 1443, 1236, 1488, 83, 867, 1459, 1035, 186, 1021, 687, 1143, 636, 1415, 1232, 1875, 11, 524, 680, 1687, 1178, 537, 169]. Variance-Component [664]. Variances [1173, 1031, 717, 218]. Variate [627, 1859, 1662]. Variates [152]. Variation [244, 1376, 1310, 560, 1853, 94]. Variations [1002, 799]. Various [1721]. Vark [1047]. Vegetation [1125]. Velleman [696]. Ventilation [1026, 17]. Ventilation/Perfusion [17]. Venus [420]. Verfahren [1741]. Verification [577]. Verma [1472]. Version [1453, 1202]. Versuchsplane [95]. Versus [1559, 897, 1072, 457]. Verteilungsfreie [472, 550]. Verwandtes [829]. Vessel [1237]. VI [1247]. via [1604, 1759]. Victor [709]. View [118]. Viewpoint [120]. Vincze [1094]. Vinod [701]. Visual [1041, 427]. Vital [1305, 1304, 617, 1306]. Vital-Rates [1305]. Vito [374]. Vleck [169]. Vol [1254, 1256, 1257, 1350, 1518, 1470, 1524, 1525, 1668, 1895, 1246, 289, 372, 606, 1040, 1250, 1049, 876, 1248, 1210, 1247, 1737, 1787, 384, 642, 1211, 1336, 386, 537, 300, 1042, 1384]. Voleske [1170]. Vollmar [1514, 1837, 1044]. Vols [1393, 1247]. Volterra [374, 1463]. Volume [106, 254, 499, 714, 948, 1135, 1303, 1479, 1664, 1674, 1749, 1849, 937, 939, 1563, 659, 1197, 1380]. volvulus [847]. vs [1115, 1458].

W [174, 432, 387, 493, 1170, 1127, 1094, 1298, 1347, 1221, 1220, 1663, 1569, 1749, 1840, 1798, 1612, 931, 1833, 649, 705, 1738, 98, 236, 1169, 389, 1087, 1460, 422, 593, 881, 946, 470, 1789, 759, 1047, 1463, 489, 1248, 1247, 1381, 1431, 1382, 1701, 1613, 384, 1882, 1212, 644, 826, 1089, 486, 206, 133, 1700, 1383, 131, 827, 781, 208, 242, 1294, 1564, 203, 1654, 480, 1426, 1566, 127]. W. [1081]. Wall [387]. Walter [981]. Walters [1516]. Waltman [991]. Wampler [827]. Warden [548]. Wardlaw [1468]. Ware [1656]. Waring [721]. Warner [371]. Warwick [1083]. Wasserman [1460, 1089]. Waters [242]. Watson [753, 882]. Wattelar [483]. Way [1439, 1443, 867, 692, 680]. Wayne [707]. Wearden [879]. Weather [1558, 304]. Weber [1740, 1426]. Wechselwirkung [427]. Wegman [304]. Weibel [372, 372]. Weibull [1108]. Weight [1778, 158]. Weighted [1822, 46, 1767]. Weighting [1781]. Weights [88, 1160]. Weinberg [176, 111, 1712, 25, 1485]. Weir [891]. Weisberg [1394, 702, 471, 480]. Weldon [1435]. Welsch [293]. Wertz [1094]. Western [117, 659]. Wetherill [1742, 538]. Whale [1186]. Whether [1857]. White [64, 1088, 752, 924, 1890, 608]. White-Tailed [924]. Whitehead [763]. Whitfield [1093]. Whitney [1037, 1768]. Whittemore [889]. Whittle [932]. Wichern [1663, 644]. Wilcoxon [1643, 1037, 1768, 1370, 151]. Wildlife [24, 1356, 1761, 1464]. Wilks [1797]. Will [136, 5]. William [107, 716]. Williams [1202, 1431]. Wilson [828]. Winfree [652]. Wirkungs [1515]. Withdrawal [504]. Withdrawals [122, 368]. Within [1156, 631, 159, 226]. Within-Family [1156, 631]. within-Individual [226]. without [660]. Wm [204]. Wohlzogen [1741]. Wolf [1515]. Women [770, 264, 36, 776]. Wonnacott [295, 295]. Wood [208]. Woodward [1347]. Wool [96]. Word [630]. Worf [172]. Work [606, 1248, 551, 422, 286]. Worker [274, 121]. Workers [1267]. Working [120]. Works [1749].

Workshop [126]. World [1172, 1350, 1525, 1706, 1895, 1463, 870, 869, 1172, 1350, 1525, 1706, 1895, 840, 767, 1172, 1350, 1525, 1706, 1895]. Wörz [245]. Wörz-Busekros [245]. Wricke [1426]. Wrigley [307]. Write [839]. Wrong [1737]. Wu [1090, 96]. Wynn [1393].

X [1741, 625]. X-Chromosome [625].

Yates [1743, 430, 54]. Year [405, 1853]. Years [574]. Yeats [1088]. Yield [1632]. Young [1436]. Yourself [247]. Yu [1221].

Z [1469, 468, 1835]. Zahn [594]. Zanakis [1297]. Zeitereihen [1474]. Zelen [675, 1493]. Zero [1201, 792]. Zero-Dose [1201]. Zeroes [121]. Zimmerman [1431]. Zinder [1471]. Zucchini [1738]. Zuchtwertschätzung [542]. Zupan [884]. zur [1467, 1837]. zwischen [604]. Zyskind [1787].

References

Hinkelmann:1971:EHE

[1] Klaus Hinkelmann. Estimation of heritability from experiments with inbred and related individuals. Biometrics, 27(1):183–190, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528936. See corrections [1578].

Brunden:1972:NAN

[2] Marshall N. Brunden. 333. Note: The analysis of non-independent 2 × 2 tables and 2×C tables using rank sums. Biometrics, 28(2):603–607, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556175. See correction [1400].

Gail:1974:PCD

[3] Mitchell Gail. Power computations for designing comparative Poisson trials. Biometrics, 30(2):231–237, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2529645. See corrections [774].

Wallenstein:1977:ATP

[4] Sylvan Wallenstein and Alan C. Fisher. The analysis of the two-period repeated measurements crossover design with application to clinical trials. Biometrics, 33(1):261–269, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2529321. See correction [310].

Good:1979:DTE

[5] Phillip I. Good. Detection of a treatment effect when not all experimental subjects will respond to treatment. Biometrics, 35(2):483–489, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530351. See correction

[136].

Anonymous:1980:FMa

[6] Anonymous. Front matter. Biometrics, 36(1):i–iv, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530488.

Curnow:1980:SCU

[7] R. N. Curnow. Selecting crosses using information from a diallel cross. Biometrics, 36(1):1–8, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530489.

Manly:1980:NMS

[8] B. F. J. Manly. A note on a model for selection experiments. Biometrics, 36(1):9–18, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530490.

Donner:1980:EIC

[9] Allan Donner and John J. Koval. The estimation of intraclass correlation in the analysis of family data. Biometrics, 36(1):19–25, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530491.

Hodge:1980:SOA

[10] Susan E. Hodge, David A. Greenberg, Jerome I. Rotter, and Kenneth L. Lange. Second-order approximations of ascertainment probabilities. Biometrics, 36(1):27–33, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530492.

Stokes:1980:EVU

[11] S. L. Stokes. Estimation of variance using judgment ordered ranked set samples. Biometrics, 36(1):35–42, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530493.

Lefkovitch:1980:CC

[12] L. P. Lefkovitch. Conditional clustering. Biometrics, 36(1):43–58, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530494.

Agresti:1980:GOR

[13] Alan Agresti. Generalized odds ratios for ordinal data. Biometrics, 36(1): 59–67, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530495.

Brown:1980:CEC

[14] Byron Wm. Brown, Jr. The crossover experiment for clinical trials. Biometrics, 36(1):69–79, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530496.

Begg:1980:TAP

[15] Colin B. Begg and Boris Iglewicz. A treatment allocation procedure for sequential clinical trials. Biometrics, 36(1):81–90, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530497.

Dunn:1980:SMO

[16] Judith K. Dunn and Robert J. Hardy. A stochastic model for the occurrence of transient ischemic attacks. Biometrics, 36(1):91–103, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530498.

Ryan:1980:EVP

[17] D. M. Ryan and A. J. Scott. Estimating ventilation/perfusion distributions from inert gas data: a Bayesian approach. Biometrics, 36(1):105– 115, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530499.

Quaas:1980:CTS

[18] R. L. Quaas and L. D. Van Vleck. Categorical trait sire evaluation by best linear unbiased prediction of future progeny category frequencies. Biometrics, 36(1):117–122, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530500.

Goodnight:1980:CEM

[19] J. H. Goodnight and F. M. Speed. Computing expected mean squares. Biometrics, 36(1):123–125, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530501.

GarciaBen:1980:MAV

[20] M. S. Garcia Ben and V. J. Yohai. Multivariate analysis of variance in a randomized blocks design when covariance matrices are unequal. Biometrics, 36(1):127–133, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530502.

Glasbey:1980:NRA

[21] C. A. Glasbey. Nonlinear regression with autoregressive time series errors. Biometrics, 36(1):135–139, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2530503. See correction [137].

Otis:1980:ECR

[22] David L. Otis. An extension of the change-in-ratio method. Biometrics, 36(1):141–147, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530504.

Darroch:1980:NCR

[23] J. N. Darroch and D. Ratcliff. A note on capture–recapture estimation. Biometrics, 36(1):149–153, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530505.

Gates:1980:IBA

[24] C. E. Gates and P. W. Smith. An implementation of the Burnham– Anderson distribution-free method of estimating wildlife densities from line transect data. Biometrics, 36(1):155–160, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530506.

Ledwina:1980:THW

[25] Teresa Ledwina and Stanisław Gnot. Testing for Hardy–Weinberg equilibrium. Biometrics, 36(1):161–165, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530507.

Hui:1980:EER

[26] S. L. Hui and S. D. Walter. Estimating the error rates of diagnostic tests. Biometrics, 36(1):167–171, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530508.

Aitkin:1980:NSL

[27] Murray Aitkin. A note on the selection of log-linear models. Biometrics, 36(1):173–178, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530509.

Anonymous:1980:Aa

[28] Anonymous. Abstracts. Biometrics, 36(1):179–190, March 1980. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530510.

Steinhorst:1980:BRI

[29] R. Kirk Steinhorst. Book review: An Introduction to Systems Analysis: With Ecological Applications, by J. N. R. Jeffers. Biometrics, 36(1):191, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530511.

Curnow:1980:BRP

[30] R. N. Curnow. Book review: Progress in Theoretical Biology, by R. Rosen, F. M. Small. Biometrics, 36(1):191–192, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530512.

Anonymous:1980:NAa

[31] Anonymous. News and announcements. Biometrics, 36(1):193, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530513.

Anonymous:1980:PPBa

[32] Anonymous. Papers to be published in Biometrics. Biometrics, 36(1): 194, March 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530514.

Anonymous:1980:FMb

[33] Anonymous. Front matter. Biometrics, 36(2):i–iv, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529970.

Freeman:1980:SR

[34] Daniel H. Freeman, Jr. and Theodore R. Holford. Summary rates. Biometrics, 36(2):195–205, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529971.

Kraemer:1980:EKC

[35] Helena Chmura Kraemer. Extension of the kappa coefficient. Biometrics, 36(2):207–216, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529972.

Royston:1980:OMD

[36] J. P. Royston and R. M. Abrams. An objective method for detecting the shift in basal body temperature in women. Biometrics, 36(2):217– 224, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529973.

Volund:1980:MB

[37] Aage Vølund. Multivariate bioassay. Biometrics, 36(2):225–236, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529974. See correction [772].

Hayes:1980:RGS

[38] J. F. Hayes and W. G. Hill. A reparameterization of a genetic selection index to locate its sampling properties. Biometrics, 36(2):237–248, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529975. See correction [140].

Becker:1980:ECM

[39] Niels Becker. An epidemic chain model. Biometrics, 36(2):249–254, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529976.

Gail:1980:ACC

[40] M. H. Gail, T. J. Santner, and C. C. Brown. An analysis of comparative carcinogenesis experiments based on multiple times to tumor. Biometrics, 36(2):255–266, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529977.

Kodell:1980:IDM

[41] R. L. Kodell and C. J. Nelson. An illness-death model for the study of the carcinogenic process using survival/sacrifice data. Biometrics, 36(2): 267–277, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529978. See

corrections [311].

Byth:1980:SSP

[42] Karen Byth and B. D. Ripley. On sampling spatial patterns by distance methods. Biometrics, 36(2):279–284, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529979.

Walter:1980:LSF

[43] S. D. Walter. Large sample formulae for the expected number of matches in a category matched design. Biometrics, 36(2):285–291, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529980.

Rubin:1980:BRU

[44] Donald B. Rubin. Bias reduction using Mahalanobis-metric matching. Biometrics, 36(2):293–298, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2529981. See correction [139].

Holford:1980:ARS

[45] Theodore R. Holford. The analysis of rates and of survivorship using loglinear models. Biometrics, 36(2):299–305, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529982.

Lubin:1980:AUC

[46] Jay H. Lubin. Analysis under Cox’s failure time model using weighted least squares. Biometrics, 36(2):307–312, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529983.

Minder:1980:NIM

[47] Ch. E. Minder. A note on an indirect method for estimating the number of small particles in a blood sample. Biometrics, 36(2):313–316, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529984.

Morgan:1980:MMI

[48] Byron J. T. Morgan and Simon A. Watts. On modelling microbial infections. Biometrics, 36(2):317–321, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529985.

Armitage:1980:EN

[49] P. Armitage. Editor’s note. Biometrics, 36(2):322, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529986.

Dambrosia:1980:SCM

[50] James M. Dambrosia and Jonas H. Ellenberg. Statistical considerations for a medical data base. Biometrics, 36(2):323–332, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529987.

Starmer:1980:CIC

[51] C. Frank Starmer, Kerry L. Lee, Frank E. Harrell, and Robert A. Rosati. On the complexity of investigating chronic illness. Biometrics, 36(2):333– 335, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529988.

Byar:1980:WDB

[52] David P. Byar. Why data bases should not replace randomized clinical trials. Biometrics, 36(2):337–342, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529989.

Fleiss:1980:SAC

[53] Joseph L. Fleiss, Alex Tytun, and Hans K. Ury. A simple approximation for calculating sample sizes for comparing independent proportions. Biometrics, 36(2):343–346, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529990.

Ury:1980:ASS

[54] Hans K. Ury and Joseph L. Fleiss. On approximate sample sizes for comparing two independent proportions with the use of Yates’ correction. Biometrics, 36(2):347–351, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529991.

Fisher:1980:QDB

[55] Richard Fisher. [A qualitative discrepancy between censored data rank tests]. Biometrics, 36(2):353, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2529992. See correction [138].

Simon:1980:E

[56] Richard M. Simon, Robert W. Makuch, Ross L. Prentice, and Patrick Marek. [Editorial]. Biometrics, 36(2):354–355, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529993.

Mantel:1980:CKS

[57] Nathan Mantel, Mitchell H. Gail, and James H. Ware. Comparisons with a known survival curve. Biometrics, 36(2):355–356, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529994.

Anonymous:1980:OJC

[58] Anonymous. Obituary: Jerome Cornfield, 1912–1979. Biometrics, 36(2): 357–358, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529995.

Poole:1980:BRP

[59] Robert W. Poole. Book review: Practical Experiences with Modelling and Forecasting Time Series, by G. M. Jenkins. Biometrics, 36(2):359, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529996.

Dagnelie:1980:BRC

[60] P. Dagnelie. Book reviews: Classification Automatique pour l’Analyse des Données. 1. Méthodes et Algorithmes, by M. Jambu; Classification Automatique pour l’Analyse des Données. 2. Logiciels, by M. Jambu, M. O. Lebeaux. Biometrics, 36(2):359–360, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529997.

Darroch:1980:BRI

[61] J. N. Darroch. Book review: Information Theory and Contingency Tables, by D. V. Gokhale, S. Kullback. Biometrics, 36(2):360, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529998.

Cooke:1980:BRR

[62] D. Cooke. Book review: Random Processes: a First Look, by R. Syski. Biometrics, 36(2):360–361, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529999.

Finney:1980:BRR

[63] D. J. Finney. Book review: Report on the LD50 Test, by Advisory Committee on the Administration of the Cruelty to Animals Act 1876. Biometrics, 36(2):361–362, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530000.

Chapman:1980:BRS

[64] Douglas G. Chapman. Book review: Statistical Inference from Capture Data on Closed Animal Populations, by D. L. Otis, K. P. Burnham, G. C. White, D. R. Anderson. Biometrics, 36(2):362, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530001.

Anonymous:1980:Ab

[65] Anonymous. Abstracts. Biometrics, 36(2):363–371, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530002.

Sandland:1980:BS

[66] R. L. Sandland. The Biometric Society. Biometrics, 36(2):373, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530003.

Anonymous:1980:NAb

[67] Anonymous. News and announcements. Biometrics, 36(2):375–378, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530004.

Anonymous:1980:PPBb

[68] Anonymous. Papers to be published in Biometrics. Biometrics, 36(2): 379, June 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530005.

Anonymous:1980:FMc

[69] Anonymous. Front matter. Biometrics, 36(3):i–iv, September 1980. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530208.

Mantel:1980:BIP

[70] Nathan Mantel. A biometrics invited paper. Assessing laboratory evidence for neoplastic activity. Biometrics, 36(3):381–399, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530209.

Macdonald:1980:MRE

[71] P. D. M. Macdonald and Howard D. Smith. Mark–recapture estimation of salmon smolt runs. Biometrics, 36(3):401–417, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530210.

Buckland:1980:MAJ

[72] Stephen T. Buckland. A modified analysis of the jolly-seber capture– recapture model. Biometrics, 36(3):419–435, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530211.

Dahiya:1980:EPS

[73] Ram C. Dahiya. Estimating the population sizes of different types of organisms in a plankton sample. Biometrics, 36(3):437–446, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530212.

Bonaiti:1980:DHT

[74] Catherine Bona¨ ıti and Florence Demenais. Diseases having two classes of severity: Estimation of the proportions of each manifestation by maximum likelihood. Biometrics, 36(3):447–460, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530213.

Ghai:1980:PEF

[75] G. L. Ghai. Probability of extinction in a finite population with truncation selection. Biometrics, 36(3):461–472, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530214.

Hartigan:1980:CNS

[76] P. M. Hartigan. Contact networks and the study of contagion. Biometrics, 36(3):473–485, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530215.

Win:1980:IBD

[77] Kyi Win and A. Dey. Incomplete block designs for parallel-line assays. Biometrics, 36(3):487–492, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530216.

Krzanowski:1980:MCC

[78] W. J. Krzanowski. Mixtures of continuous and categorical variables in discriminant analysis. Biometrics, 36(3):493–499, September 1980. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530217.

McLachlan:1980:RBF

[79] G. J. McLachlan. On the relationship between the F test and the overall error rate for variable selection in two-group discriminant analysis. Biometrics, 36(3):501–510, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530218.

Peduzzi:1980:SVS

[80] P. N. Peduzzi, R. J. Hardy, and T. R. Holford. A stepwise variable selection procedure for nonlinear regression models. Biometrics, 36(3):511– 516, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530219.

Spurrier:1980:TST

[81] John D. Spurrier and John E. Hewett. Two-stage test of independence using Kendall’s statistic. Biometrics, 36(3):517–522, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530220.

McHugh:1980:PPV

[82] Richard McHugh, Timothy Church, Jack Mandel, and Christine White. The parameterization of predictive value for multisite screening. Biometrics, 36(3):523–529, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530221.

Lee:1980:MVP

[83] Larry Lee and R. G. Krutchkoff. Mean and variance of partiallytruncated distributions. Biometrics, 36(3):531–536, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530222.

Stablein:1980:SAD

[84] Donald M. Stablein, Walter H. Carter, Jr., and Galen L. Wampler. Survival analysis of drug combinations using a hazards model with timedependent covariates. Biometrics, 36(3):537–546, September 1980. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530223.

Sprent:1980:QGM

[85] P. Sprent and G. R. Dolby. Query: The geometric mean functional relationship. Biometrics, 36(3):547–550, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530224.

Davies:1980:NRC

[86] O. L. Davies. Note on regression with correlated responses. Biometrics, 36(3):551–552, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530225.

Mantel:1980:RTW

[87] Nathan Mantel. Randomization tests when there are non-responders. Biometrics, 36(3):552–553, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530226.

Kshirsagar:1980:CVC

[88] A. M. Kshirsagar, M. Stone, and L. Jensen. Cross-validatory choice of weights for inter- and intra-block estimation in balanced incomplete block designs. Biometrics, 36(3):553, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530227.

McLachlan:1980:SVD

[89] Geoff McLachlan. Selection of variables in discriminant analysis. Biometrics, 36(3):554, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530228.

Kruskal:1980:RD

[90] W. Kruskal. Rounded data. Biometrics, 36(3):554–555, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530229.

Anonymous:1980:OGL

[91] Anonymous. Obituary: G. Lefort, 1921–1979. Biometrics, 36(3):557–558, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1980:O

[92] Anonymous. Obituaries. Biometrics, 36(3):557–559, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530230.

Anonymous:1980:OJG

[93] Anonymous. Obituary: J. G. Skellam, 1914–1979. Biometrics, 36(3): 558–559, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Weir:1980:BRQ

[94] B. S. Weir. Book review: Quantitative Genetic Variation, by James N. Thompson, Jr., J. M. Thoday. Biometrics, 36(3):561–562, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530231.

Bauer:1980:BRL

[95] P. Bauer. Book review: Lineare Statistische Modelle und Optimale Versuchsplane, by Olaf Krafft. Biometrics, 36(3):562, September 1980. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530233.

Robertson:1980:BRS

[96] Alan Robertson. Book review: Statistical Genetics (Breeding for Milk, Meat, Eggs and Wool), by J. S. Wu. Biometrics, 36(3):562, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530232.

Berry:1980:BRM

[97] G. Berry. Book review: Mainland’s Notes on Biometry in Medical Research, by Donald Mainland, M. Clinton Miller, III. Biometrics, 36 (3):562–563, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530234.

Bock:1980:BRE

[98] H. H. Bock. Book review: Elementare Statistische Methoden, by A. Linder, W. Berchtold. Biometrics, 36(3):563, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530236.

Shotter:1980:BRM

[99] M. V. Shotter. Book review: Mainland’s Notes from a Laboratory of Medical Statistics, by Donald Mainland, M. Clinton Miller, III. Biometrics, 36(3):563, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530235.

Anonymous:1980:BRD

[100] Anonymous. Book review: Distribution-Free Methods for Non-Parametric Methods: a Classified and Selected Bibliography, by B. Singer. Biometrics, 36(3):564, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530238.

Freeman:1980:BRS

[101] G. H. Freeman. Book review: Statistique et Informatique en Agronomie/ Statistics and Computer Science in Agriculture, by P. Cantillon, P. Dagnelie. Biometrics, 36(3):564, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530237.

Bliss:1980:BS

[102] Chester I. Bliss. The Biometric Society. Biometrics, 36(3):565–566, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530239.

Anonymous:1980:NAc

[103] Anonymous. News and announcements. Biometrics, 36(3):567, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530240.

Anonymous:1980:PPBc

[104] Anonymous. Papers to be published in Biometrics. Biometrics, 36(3): 569, September 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530241.

Anonymous:1980:FMd

[105] Anonymous. Front matter. Biometrics, 36(4):i–iv, December 1980. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556110.

Anonymous:1980:VI

[106] Anonymous. Volume information. Biometrics, 36(4):vii–xxii, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556109.

Anderson:1980:WGC

[107] R. L. Anderson. William Gemmell Cochran, 1909–1980: a personal tribute. Biometrics, 36(4):574–578, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556111.

Reed:1980:OAS

[108] William J. Reed. Optimum age-specific harvesting in a nonlinear population model. Biometrics, 36(4):579–593, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556112.

Cruz-Orive:1980:BLU

[109] Luis-M. Cruz-Orive. Best linear unbiased estimators for stereology. Biometrics, 36(4):595–605, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2556113.

Fleming:1980:MKS

[110] Thomas R. Fleming, Judith R. O’Fallon, Peter C. O’Brien, and David P. Harrington. Modified Kolmogorov–Smirnov test procedures with application to arbitrarily right-censored data. Biometrics, 36(4):607–625, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556114.

Emigh:1980:CTH

[111] Ted H. Emigh. A comparison of tests for Hardy–Weinberg equilibrium. Biometrics, 36(4):627–642, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2556115.

Gregorius:1980:PLA

[112] Hans-Rolf Gregorius. The probability of losing an allele when diploid genotypes are sampled. Biometrics, 36(4):643–652, December 1980. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556116. See corrections [312, 313].

Wahrendorf:1980:BBI

[113] Jürgen Wahrendorf and Charles C. Brown. Bootstrapping a basic inequality in the analysis of joint action of two drugs. Biometrics, 36 (4):653–657, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556117.

Hasselblad:1980:MPA

[114] Victor Hasselblad, Andrew G. Stead, and John P. Creason. Multiple probit analysis with a nonzero background. Biometrics, 36(4):659–663, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556118.

Klotz:1980:MCF

[115] Jerome Klotz. A modified Cochran–Friedman test with missing observations and ordered categorical data. Biometrics, 36(4):665–670, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556119.

Hewett:1980:CTP

[116] John E. Hewett and Zuhdi Lababidi. Comparison of two populations with multivariate data. Biometrics, 36(4):671–675, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556120.

Feigl:1980:BRT

[117] Polly Feigl. Book review: Training of Statisticians for Clinical Trials: a Series of Three Papers Presented at the Spring Meeting of the Western North American Region of the Biometric Society, 26–28 June 1979, Los Angeles, California. Biometrics, 36(4):677–678, December 1980. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556121.

Hammond:1980:TCT

[118] Denman Hammond. The training of clinical trials statisticians: a clinician’s view. Biometrics, 36(4):679–685, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556122.

Peterson:1980:TPC

[119] Arthur V. Peterson, Jr. and Lloyd D. Fisher. Teaching the principles of clinical trials design and management. Biometrics, 36(4):687–697, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556123.

Gehan:1980:TSC

[120] Edmund A. Gehan. The training of statisticians for cooperative clinical trials: a working statistician’s viewpoint. Biometrics, 36(4):699–706, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556124.

Owen:1980:EML

[121] W. J. Owen and T. A. DeRouen. Estimation of the mean for lognormal data containing zeroes and left- censored values, with applications to the measurement of worker exposure to air contaminants. Biometrics, 36(4):707–719, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556125.

Gould:1980:NAA

[122] A. Lawrence Gould. A new approach to the analysis of clinical drug trials with withdrawals. Biometrics, 36(4):721–727, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556126.

Jansen:1980:CCC

[123] A. A. M. Jansen and C. M. Theobald. Comparative calibration and congeneric measurements. Biometrics, 36(4):729–734, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556127.

Anonymous:1980:Ac

[124] Anonymous. Abstracts. Biometrics, 36(4):735–737, December 1980. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556128.

Moran:1980:BRR

[125] P. A. P. Moran. Book review: Reversibility and Stochastic Networks, by F. P. Kelly. Biometrics, 36(4):739, December 1980. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556130.

Rowntree:1980:BRS

[126] J. A. Rowntree. Book review: Synthetic Estimates for Small Areas: Statistical Workshop Papers and Discussion, by J. Steinber. Biometrics, 36(4):739, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556129.

Weir:1980:BRM

[127] B. S. Weir. Book review: Mathematical Population Genetics, by W. J. Ewens. Biometrics, 36(4):739–740, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556131.

Hinde:1980:BRS

[128] J. P. Hinde. Book review: Systems Theory in Immunology, by C. Bruni. Biometrics, 36(4):740–741, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2556132.

Goldstein:1980:BRD

[129] H. Goldstein. Book review: Discrete Statistical Models with Social Science Applications, by E. B. Andersen. Biometrics, 36(4):741, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556133.

Patterson:1980:BRD

[130] H. D. Patterson. Book review: Design and Analysis of Experiments in the Animal and Medical Sciences, by J. L. Gill. Biometrics, 36(4):741–742, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556134.

Pielou:1980:BRE

[131] E. C. Pielou. Book review: Ecological Diversity in Theory and Practice, by J. F. Grassle, G. P. Patil, W. Smith, C. Taillie. Biometrics, 36(4):742– 743, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2556135.

Sibson:1980:BRP

[132] Robin Sibson. Book review: Practical Geostatistics, by Isobel Clark. Biometrics, 36(4):743, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2556136.

Nelder:1980:BRS

[133] J. A. Nelder. Book review: Statistical Computing, by W. J. Kennedy, Jr., J. E. Gentle. Biometrics, 36(4):743–744, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556137.

Anonymous:1980:BS

[134] Anonymous. The Biometric Society. Biometrics, 36(4):745–747, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556138.

Anonymous:1980:NAd

[135] Anonymous. News and announcements. Biometrics, 36(4):749–750, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556139.

Anonymous:1980:CCD

[136] Anonymous. Corrections: Correction to ‘Detection of a treatment effect when not all experimental subjects will respond to treatment’. Biometrics, 36(4):751, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2556140. See [5].

Anonymous:1980:CCN

[137] Anonymous. Corrections: Correction to ‘Nonlinear regression with autoregressive time series errors’. Biometrics, 36(4):751, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556142. See [21].

Anonymous:1980:CCQ

[138] Anonymous. Corrections: Correction to ‘A qualitative discrepancy between censored data rank tests’. Biometrics, 36(4):751, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556141. See [55].

Anonymous:1980:CCB

[139] Anonymous. Corrections: Correction to ‘Bias reduction using Mahalanobis-metric matching’. Biometrics, 36(4):752, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556144. See [44].

Anonymous:1980:CCR

[140] Anonymous. Corrections: Correction to ‘A reparameterization of a genetic selection index to locate its sampling properties’. Biometrics, 36(4): 752, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2556143. See

[38].

Hayes:1980:CRG

[141] J. F. Hayes and W. G. Hill. Correction to: “A reparameterization of a genetic selection index to locate its sampling properties”. Biometrics, 36(4):752, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [38].

Anonymous:1980:PPBd

[142] Anonymous. Papers to be published in Biometrics. Biometrics, 36(4): 753, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2556145.

Anonymous:1981:FMa

[143] Anonymous. Front matter. Biometrics, 37(1):i–iv, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530516.

Wood:1981:ITE

[144] Constance L. Wood and Foster B. Cady. Intersite transfer of estimated response surfaces. Biometrics, 37(1):1–10, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530517.

Selwyn:1981:BAB

[145] Murray R. Selwyn, Arthur P. Dempster, and Nancy R. Hall. A Bayesian approach to bioequivalence for the 2 × 2 changeover design. Biometrics, 37(1):11–21, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530518.

Ganesalingam:1981:SER

[146] S. Ganesalingam and G. J. McLachlan. Some efficiency results for the estimation of the mixing proportion in a mixture of two normal distributions. Biometrics, 37(1):23–33, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530519.

Symons:1981:CCM

[147] M. J. Symons. Clustering criteria and multivariate normal mixtures. Biometrics, 37(1):35–43, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530520.

Wahrendorf:1981:ODA

[148] Jürgen Wahrendorf, Reinhard Zentgraf, and Charles C. Brown. Optimal designs for the analysis of interactive effects of two carcinogens or other toxicants. Biometrics, 37(1):45–54, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530521. See correction [314].

Heilbron:1981:ARO

[149] David C. Heilbron. The analysis of ratios of odds ratios in stratified contingency tables. Biometrics, 37(1):55–66, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530522.

Gail:1981:ESC

[150] Mitchell H. Gail. Evaluating serial cancer marker studies in patients at risk of recurrent disease. Biometrics, 37(1):67–78, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530523. See corrections [1223].

Tarone:1981:DML

[151] Robert E. Tarone. On the distribution of the maximum of the logrank statistic and the modified Wilcoxon statistic. Biometrics, 37(1):79–85, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530524.

Lachin:1981:SCT

[152] John M. Lachin. Sequential clinical trials for normal variates using interval composite hypotheses. Biometrics, 37(1):87–101, March 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530525.

North:1981:SME

[153] Philip M. North and Richard M. Cormack. On Seber’s method for estimating age-specific bird survival rates from ringing recoveries. Biometrics, 37(1):103–112, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530526.

Upton:1981:MIC

[154] Graham J. G. Upton and Gordon A. Lampitt. A model for interyear change in the size of bird populations. Biometrics, 37(1):113–127, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530527.

Whitehead:1981:USP

[155] John Whitehead. The use of the sequential probability ratio test for monitoring the percentage germination of accessions in seed banks. Biometrics, 37(1):129–136, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530528.

Katz:1981:BAA

[156] D. Katz, S. P. Azen, and A. Schumitzky. Bayesian approach to the analysis of nonlinear models: Implementation and evaluation. Biometrics, 37(1):137–142, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530529.

Langenberg:1981:CMC

[157] P. Langenberg and R. Srinivasan. On the colton model for clinical trials with delayed observations — normally-distributed responses. Biometrics, 37(1):143–148, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530530.

Uche:1981:HWN

[158] P. I. Uche. The height and weight of Nigerian children: The fit of Ehrenberg’s equation. Biometrics, 37(1):149–151, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530531.

Segreti:1981:EML

[159] Anthony C. Segreti and Albert E. Munson. Estimation of the median lethal dose when responses within a litter are correlated. Biometrics, 37(1):153–156, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530532.

Wolfe:1981:MDR

[160] Robert A. Wolfe, Larry D. Roi, and Elizabeth H. Margosches. Monotonic dichotomous regression estimates: a burn care example. Biometrics, 37(1):157–167, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530533.

Farewell:1981:SSH

[161] Vernon T. Farewell and Giulio J. D’Angio. A simulated study of historical controls using real data. Biometrics, 37(1):169–176, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530534.

Bennett:1981:LTH

[162] B. M. Bennett. On LR tests of hypotheses concerning relative risk. Biometrics, 37(1):177–178, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530535.

Marriott:1981:TRS

[163] F. H. C. Marriott. Tests for randomness in spatial patterns. Biometrics, 37(1):178, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530536.

Anonymous:1981:Aa

[164] Anonymous. Abstracts. Biometrics, 37(1):179–198, March 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530537.

Lachenbruch:1981:BRM

[165] P. A. Lachenbruch. Book review: Multivariate Analysis, by K. V. Mardia, J. T. Kent, J. M. Bibby. Biometrics, 37(1):199, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530538.

Rasch:1981:BRK

[166] Dieter Rasch. Book review: Klassifikation und Erkenntnis III: Numerische Klassifikation, by H. H. Bock. Biometrics, 37(1):199–200, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530539.

Young:1981:BRR

[167] J. C. Young. Book review: Regression Methods: a Tool for Data Analysis, by R. J. Freund, P. D. Minton. Biometrics, 37(1):200, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530540.

Oldham:1981:BRD

[168] P. D. Oldham. Book review: The Design and Analysis of Longitudinal Studies: Their Role in the Measurement of Change, by H. Goldstein. Biometrics, 37(1):200–201, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530541.

Thompson:1981:BRV

[169] R. Thompson. Book review: Variance Components and Animal Breeding, by L. D. Van Vleck, S. R. Searle. Biometrics, 37(1):201–202, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530542.

Daldry:1981:BRS

[170] S. J. Daldry. Book review: Statistical Exercises in Medical Research, by J. F. Osborn. Biometrics, 37(1):202, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530543.

Kempton:1981:BRS

[171] R. A. Kempton. Book review: Sampling Design and Statistical Methods for Environmental Biologists, by R. H. Green. Biometrics, 37(1):202– 203, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530544.

Elliott:1981:BRB

[172] J. M. Elliott. Book review: Biological Monitoring for Environmental Effects, by D. L. Worf. Biometrics, 37(1):203, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530545.

Anderson:1981:BRS

[173] D. R. Anderson. Book review: Sampling Biological Populations, by R. M. Cormack, G. P. Patil, D. S. Robson. Biometrics, 37(1):203–204, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530546.

Anderson:1981:BRB

[174] A. J. B. Anderson. Book review: BMDP-79: Biomedical Computer Programs P -Series, by W. J. Dixon, M. B. Brown. Biometrics, 37(1):204, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530547.

Downham:1981:BRM

[175] D. Y. Downham. Book review: Mathematical Models in Cell Biology and Cancer Chemotheraphy, by M. Eisen. Biometrics, 37(1):204–205, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530548.

Anonymous:1981:BRI

[176] Anonymous. Book review: Introduction to Dental Statistics, by R. Weinberg, S. L. Cheuk. Biometrics, 37(1):205, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530549.

Anonymous:1981:BSa

[177] Anonymous. The Biometric Society. Biometrics, 37(1):207–208, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530550.

Anonymous:1981:NAa

[178] Anonymous. News and announcements. Biometrics, 37(1):209–210, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530551.

Anonymous:1981:EN

[179] Anonymous. Editor’s note. Biometrics, 37(1):210, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530552.

Anonymous:1981:PPBa

[180] Anonymous. Papers to be published in Biometrics. Biometrics, 37(1): 211, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530553.

Anonymous:1981:FMb

[181] Anonymous. Front matter. Biometrics, 37(2):i–iv, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530411.

Mandallaz:1981:CDM

[182] Daniel Mandallaz and Jochen Mau. Comparison of different methods for decision-making in bioequivalence assessment. Biometrics, 37(2):213– 222, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530412.

Morton:1981:GSE

[183] R. Morton. Generalized Spearman estimators of relative dose. Biometrics, 37(2):223–233, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530413.

Cressie:1981:ADH

[184] N. A. C. Cressie and D. D. Keightley. Analysing data from hormonereceptor assays. Biometrics, 37(2):235–249, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530414.

Becker:1981:GCB

[185] Niels Becker. A general chain binomial model for infectious diseases. Biometrics, 37(2):251–258, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530415.

Manton:1981:VCA

[186] Kenneth G. Manton, Max A. Woodbury, and Eric Stallard. A variance components approach to categorical data models with heterogeneous cell populations: Analysis of spatial gradients in lung cancer mortality rates in North Carolina counties. Biometrics, 37(2):259–269, June 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530416.

Kupper:1981:MES

[187] Lawrence L. Kupper, John M. Karon, David G. Kleinbaum, Hal Morgenstern, and Donald K. Lewis. Matching in epidemiologic studies: Validity and efficiency considerations. Biometrics, 37(2):271–291, June 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530417.

Leung:1981:CCI

[188] Hoi M. Leung and Lawrence L. Kupper. Comparisons of confidence intervals for attributable risk. Biometrics, 37(2):293–302, June 1981.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530418.

Bellhouse:1981:AEP

[189] D. R. Bellhouse. Area estimation by point-counting techniques. Biometrics, 37(2):303–312, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530419.

Thompson:1981:OSP

[190] E. A. Thompson. Optimal sampling for pedigree analysis: Sequential schemes for sibships. Biometrics, 37(2):313–325, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530420.

Joe:1981:CPT

[191] H. Joe, James A. Koziol, and A. John Petkau. Comparison of procedures for testing the equality of survival distributions. Biometrics, 37(2):327– 340, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530421.

Rai:1981:GMD

[192] Kamta Rai and John Van Ryzin. A generalized multihit dose-response model for low-dose extrapolation. Biometrics, 37(2):341–352, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530422.

Krewski:1981:CRA

[193] Daniel Krewski and Charles Brown. Carcinogenic risk assessment: a guide to the literature. Biometrics, 37(2):353–366, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530423.

Cox:1981:RNN

[194] Trevor F. Cox. Reflexive nearest neighbours. Biometrics, 37(2):367– 369, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530424.

Sprott:1981:MLA

[195] D. A. Sprott. Maximum likelihood applied to a capture–recapture model. Biometrics, 37(2):371–375, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530425.

Wiggins:1981:EPR

[196] R. D. Wiggins and M. Slater. Estimating probabilities from retrospective data with an application to cot death in Lambeth. Biometrics, 37(2): 377–382, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530426.

Koziol:1981:DFT

[197] James A. Koziol, Donna A. Maxwell, Matsuro Fukushima, M. E. M. Colmerauer, and Yosef H. Pilch. A distribution-free test for tumorgrowth curve analyses with application to an animal tumor immunotherapy experiment. Biometrics, 37(2):383–390, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530427.

Henderson:1981:BMR

[198] Harold V. Henderson and Paul F. Velleman. Building multiple regression models interactively. Biometrics, 37(2):391–411, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530428.

Gates:1981:BRE

[199] Charles E. Gates. Book review: Estimation of Density from Line Transect Sampling of Biological Populations, by K. P. Burnham, D. R. Anderson, J. L. Laake. Biometrics, 37(2):413, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530429.

Morgan:1981:BRD

[200] Byron J. T. Morgan. Book review: Diffusion and Ecological Problems: Mathematical Models, by A. Okubo. Biometrics, 37(2):413–414, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530430.

Gordon:1981:BRD

[201] A. D. Gordon. Book review: The Distribution of the Size of the Maximum Cluster of Points on a Line, by N. D. Neff, J. I. Naus. Biometrics, 37(2): 414, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530431.

Weir:1981:BRP

[202] B. S. Weir. Book review: Principles of Population Genetics, by D. L. Hartl. Biometrics, 37(2):414–415, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530432.

Solomon:1981:BRM

[203] Daniel L. Solomon. Book review: Mathematical Modelling in Biology and Ecology. Proceedings, 1979, by W. M. Getz. Biometrics, 37(2):415– 416, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530433.

Finney:1981:BRB

[204] D. J. Finney. Book review: Biostastistics Casebook, by R. G. Miller, B. Efron, B. Wm Brown, Jr., L. E. Moses. Biometrics, 37(2):416, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530434.

Preece:1981:BRH

[205] M. A. Preece. Book review: Human Physical Growth and Maturation: Methodologies and Factors, by F. E. Johnston, A. F. Roche, C. Susanne. Biometrics, 37(2):416–417, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530435.

McLachlan:1981:BRM

[206] G. J. McLachlan. Book review: Mathematics and Statistics for the Bio-Sciences, by G. Eason, C. W. Coles, G. Gettinby. Biometrics, 37(2): 417, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530436.

Gordon:1981:BRC

[207] A. D. Gordon. Book review: Cluster Analysis, by B. Everitt. Biometrics, 37(2):417–418, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530437.

Rutherford:1981:BRF

[208] Andrew A. Rutherford. Book review: Fitting Equations to Data, by C. Daniel, F. S. Wood, J. W. Gorman. Biometrics, 37(2):418–419, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530438.

Williams:1981:BRS

[209] E. R. Williams. Book review: Statistics and Experimental Design, by G. M. Clarke. Biometrics, 37(2):419, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530439.

Anonymous:1981:BSb

[210] Anonymous. The Biometric Society. Biometrics, 37(2):421, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530440.

Anonymous:1981:NAb

[211] Anonymous. News and announcements. Biometrics, 37(2):423–424, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530441.

Anonymous:1981:PPBb

[212] Anonymous. Papers to be published in Biometrics. Biometrics, 37(2): 425, June 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530442.

Anonymous:1981:FMc

[213] Anonymous. Front matter. Biometrics, 37(3):i–iv, September 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530555.

Ware:1981:TPF

[214] James H. Ware and Margaret C. Wu. Tracking: Prediction of future values from serial measurements. Biometrics, 37(3):427–437, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530556.

Foulkes:1981:ITL

[215] M. A. Foulkes and C. E. Davis. An index of tracking for longitudinal data. Biometrics, 37(3):439–446, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530557.

McMahan:1981:IT

[216] C. Alex McMahan. An index of tracking. Biometrics, 37(3):447–455, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530558.

Ashford:1981:GMJ

[217] J. R. Ashford. General models for the joint action of mixtures of drugs. Biometrics, 37(3):457–474, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530559.

Elashoff:1981:RMB

[218] Janet Dixon Elashoff. Repeated-measures bioassay with correlated errors and heterogeneous variances: a Monte Carlo study. Biometrics, 37(3):475–482, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530560.

Hayes:1981:MEP

[219] J. F. Hayes and W. G. Hill. Modification of estimates of parameters in the construction of genetic selection indices (’Bending’). Biometrics, 37(3):483–493, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530561.

Arnold:1981:SNP

[220] Jonathan Arnold. Statistics of natural populations. I. Estimating an allele probability in cryptic fathers with a fixed number of offspring. Biometrics, 37(3):495–504, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530562.

Gross:1981:POS

[221] Alan J. Gross and Chan F. Lam. Paired observations from a survival distribution. Biometrics, 37(3):505–511, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530563.

Bernstein:1981:EPH

[222] L. Bernstein, J. Anderson, and M. C. Pike. Estimation of the proportional hazard in two-treatment-group clinical trials. Biometrics, 37 (3):513–519, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530564.

Pollock:1981:CRM

[223] K. H. Pollock. Capture–recapture models allowing for age-dependent survival and capture rates. Biometrics, 37(3):521–529, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530565.

Diggle:1981:BMS

[224] Peter J. Diggle. Binary mosaics and the spatial pattern of heather. Biometrics, 37(3):531–539, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530566.

Kolakowski:1981:MGP

[225] Donald Kolakowski and R. Darrell Bock. A multivariate generalization of probit analysis. Biometrics, 37(3):541–551, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530567.

Wilson:1981:SDW

[226] P. David Wilson, J. Richard Hebel, and Roger Sherwin. Screening and diagnosis when within-individual observations are Markov-dependent. Biometrics, 37(3):553–565, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530568.

Baskerville:1981:TST

[227] J. C. Baskerville and J. J. Koval. A three-sample test for selective pairing. Biometrics, 37(3):567–573, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530569.

Dunsmore:1981:APC

[228] Ian R. Dunsmore. Analysis of preferences in crossover designs. Biometrics, 37(3):575–578, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530570.

Bennett:1981:ACS

[229] G. L. Bennett. Analysis of curvilinear selection response. Biometrics, 37(3):579–583, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530571.

Freeman:1981:QSU

[230] G. H. Freeman. Query: Series of unreplicated split-plot lattice experiments. Biometrics, 37(3):585–587, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530572.

Kirkwood:1981:BTN

[231] Thomas B. L. Kirkwood and W. J. Westlake. Bioequivalence testing — a need to rethink. Biometrics, 37(3):589–594, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530573.

Diegert:1981:NIC

[232] Carl Diegert and Kathleen Vickers Diegert. Note on inversion of Casagrande–Pike–Smith approximate sample-size formula for Fisher– Irwin test on 2 × 2 tables. Biometrics, 37(3):595, September 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530574.

Harrell:1981:SVS

[233] Frank E. Harrell, Jr., Kerry L. Lee, P. N. Peduzzi, R. J. Hardy, and T. R. Holford. A stepwise variable selection procedure for nonlinear regression models. Biometrics, 37(3):595–596, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530575.

Anonymous:1981:Ab

[234] Anonymous. Abstracts. Biometrics, 37(3):597–620, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530576.

Williamson:1981:BRS

[235] Mark Williamson. Book review: Statistics for Biologists, by D. J. Finney. Biometrics, 37(3):621, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530577.

Brown:1981:BRP

[236] B. M. Brown. Book review: Practical Non-Parametric Statistics, by W. J. Conover. Biometrics, 37(3):621–622, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530578.

Bulusu:1981:BRM

[237] L. Bulusu. Book review: Manual of Mortality Analysis, by H. Campbell. Biometrics, 37(3):622, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530580.

Moran:1981:BRM

[238] M. A. Moran. Book review: Multivariate Analysis, by M. G. Kendall. Biometrics, 37(3):622, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530579.

Aitkin:1981:BRS

[239] Murray Aitkin. Book review: The Statistical Analysis of Failure Time Data, by J. D. Kalbfleisch, R. L. Prentice. Biometrics, 37(3):622–624, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530581.

Lebreton:1981:BRE

[240] J. D. Lebreton. Book reviews: Ecologie Numérique, 1. Le Traitement Multiple des Données Écologiques, by O. Legendre, L. Legendre; Ecologie Numérique, 2. La Structure des Données Écologiques, by O. Legendre, L. Legendre. Biometrics, 37(3):624–625, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530582.

Mead:1981:BRS

[241] R. Mead. Book review: Spatial and Temporal Analysis in Ecology, by R. M. Cormack, J. K. Ord. Biometrics, 37(3):625, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530583.

Sandland:1981:BRE

[242] R. L. Sandland. Book review: Environmental Biomonitoring, Assessment, Prediction, and Management, by J. Cairns, G. P. Patil, W. E. Waters. Biometrics, 37(3):625–626, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530584.

Brown:1981:BRC

[243] David Brown. Book review: Contemporary Quantitative Ecology and Related Ecometrics, by G. P. Patil, M. L. Rosenzweig. Biometrics, 37(3):626–627, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530585.

Bulmer:1981:BRE

[244] Michael Bulmer. Book review: Evolution and Variation of Multigene Families, by T. Ohta. Biometrics, 37(3):627, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530587.

Etherington:1981:BRA

[245] I. M. H. Etherington. Book review: Algebras in Genetics, by Angelika Wörz-Busekros. Biometrics, 37(3):627, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530586.

Blackith:1981:BRW

[246] R. E. Blackith. Book review: Why and How: Some Problems and Methods in Historical Biology, by George Gaylord Simpson. Biometrics, 37(3):627–628, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530588.

Diggle:1981:BRA

[247] Peter J. Diggle. Book review: APLSTAT: a Do-it-Yourself Guide to Computational Statistics Using APL., by J. B. Ramsey, G. L. Musgrave. Biometrics, 37(3):628, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530589.

Cooke:1981:BRF

[248] Dennis Cooke. Book review: Finite Markov Processes and Their Applications, by M. Iosifescu. Biometrics, 37(3):628–629, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530590.

Anonymous:1981:BRR

[249] Anonymous. Book review: The Readable Maths and Statistics Book, by Barry Edwards. Biometrics, 37(3):629, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530592.

Renshaw:1981:BRP

[250] Eric Renshaw. Book review: Physical Applications of Stationary Time-Series, by E. A. Robinson. Biometrics, 37(3):629, September 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530591.

Anonymous:1981:NAc

[251] Anonymous. News and announcements. Biometrics, 37(3):631, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530593.

Anonymous:1981:PPBc

[252] Anonymous. Papers to be published in Biometrics. Biometrics, 37(3): 633, September 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530594.

Anonymous:1981:FMd

[253] Anonymous. Front matter. Biometrics, 37(4):i–iv, December 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530146.

Anonymous:1981:VI

[254] Anonymous. Volume information. Biometrics, 37(4):ix–xxviii, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530145.

Piazza:1981:MTG

[255] A. Piazza, P. Menozzi, and L. Cavalli-Sforza. The making and testing of geographic gene-frequency maps. Biometrics, 37(4):635–659, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530147.

North:1981:MSC

[256] Philip M. North. Models for a simple capture–recapture procedure to estimate population size. Biometrics, 37(4):661–672, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530148.

Thomas:1981:GRR

[257] Duncan C. Thomas. General relative-risk models for survival time and matched case-control analysis. Biometrics, 37(4):673–686, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530149.

Woolson:1981:RTO

[258] Robert F. Woolson. Rank tests and a one-sample logrank test for comparing observed survival data to a standard population. Biometrics, 37(4):687–696, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530150.

Lustbader:1981:DCS

[259] Edward D. Lustbader and Samuel Litwin. Design of censored survival studies. Biometrics, 37(4):697–704, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530151.

Doornbos:1981:TSO

[260] R. Doornbos. Testing for a single outlier in a linear model. Biometrics, 37(4):705–711, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530152.

Disch:1981:BNI

[261] Damon Disch. Bayesian nonparametric inference for effective doses in a quantal-response experiment. Biometrics, 37(4):713–722, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530153.

Simon:1981:CRD

[262] Richard Simon. Composite randomization designs for clinical trials. Biometrics, 37(4):723–731, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530154.

Carter:1981:RML

[263] Randy L. Carter. Restricted maximum likelihood estimation of bias and reliability in the comparison of several measuring methods. Biometrics, 37(4):733–741, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530155.

Carter:1981:BCP

[264] Randy L. Carter and Barry J. N. Blight. A Bayesian change-point problem with an application to the prediction and detection of ovulation in women. Biometrics, 37(4):743–751, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530156. See correction [770, 776].

Grimson:1981:STC

[265] Roger C. Grimson and Katherine H. Chavigny. A statistical test for classification, with applications to the characterization of pathogens according to antibiotic susceptibility patterns. Biometrics, 37(4):753–761, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530157.

Brown:1981:ESL

[266] Charles C. Brown and Thomas R. Fears. Exact significance levels for multiple binomial testing with application to carcinogenicity screens. Biometrics, 37(4):763–774, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530158.

Collings:1981:ABD

[267] Bruce J. Collings, Barry H. Margolin, and Gary W. Oehlert. Analyses for binomial data, with application to the fluctuation test for mutagenicity. Biometrics, 37(4):775–794, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530159.

Holgate:1981:SES

[268] P. Holgate. The statistical entropy of a sample from a community of species. Biometrics, 37(4):795–799, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530160.

Thibodeau:1981:EDT

[269] L. A. Thibodeau. Evaluating diagnostic tests. Biometrics, 37(4):801–804, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530161.

Halperin:1981:DSE

[270] Max Halperin and Norman J. Johnson. Design and sensitivity evaluation of follow-up studies for risk factor assessment. Biometrics, 37(4):805–810, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530162.

Raynor:1981:CMC

[271] William J. Raynor, Jr. and Lawrence L. Kupper. Category-matching of continuous variables in case-control studies. Biometrics, 37(4):811–817, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530163.

Shirley:1981:ESS

[272] Eryl A. C. Shirley and R. Hickling. An evaluation of some statistical methods for analysing numbers of abnormalities found amongst litters in teratology studies. Biometrics, 37(4):819–829, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530164.

Aitkin:1981:QRM

[273] Murray Aitkin. Query: Regression models for repeated measurements. Biometrics, 37(4):831–832, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530165.

Mantel:1981:LAD

[274] Nathan Mantel, W. J. Owen, and T. A. DeRouen. Lognormal analyses of data on worker exposure to air contaminants. Biometrics, 37(4):833–834, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530166.

Healy:1981:SAH

[275] M. J. R. Healy. A source of assay heterogeneity. Biometrics, 37(4):834– 835, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530167.

Snee:1981:GDA

[276] Ronald D. Snee. Graphical display and assessment of means. Biometrics, 37(4):835–836, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530168.

Faddy:1981:CMT

[277] M. J. Faddy, Ralph L. Kodell, and Cornelius J. Nelson. Compartmental model transfer times. Biometrics, 37(4):837, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530169.

Gart:1981:TRR

[278] John J. Gart and B. M. Bennett. Tests for relative risk. Biometrics, 37(4):837–838, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530170.

Kjelsberg:1981:COC

[279] Marcus O. Kjelsberg and Robert S. Gordon. Course offerings in clinical trials methodology. Biometrics, 37(4):839, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530171.

Raab:1981:RCR

[280] G. M. Raab. Robust calibration and radioimmunoassay. Biometrics, 37(4):839–841, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530172.

Anonymous:1981:Ac

[281] Anonymous. Abstract. Biometrics, 37(4):843–853, December 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530173.

Holland:1981:BRM

[282] Derek A. Holland. Book review: Mathematics for the Biological Sciences — from Graphs Through Calculus to Differential Equations, by J. C. Newby. Biometrics, 37(4):855, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530174.

Weir:1981:BRMa

[283] B. S. Weir. Book review: The Mathematical Theory of Quantitative Genetics, by M. G. Bulmer. Biometrics, 37(4):855–856, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530175.

Weir:1981:BRMb

[284] B. S. Weir. Book review: Mathematics of Genetic Diversity, by J. F. C. Kingman. Biometrics, 37(4):857, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530176.

Curnow:1981:BRS

[285] R. N. Curnow. Book review: Selection Experiments in Laboratory and Domestic Animals, by A. Robertson. Biometrics, 37(4):857–858, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530177.

Kemp:1981:BRS

[286] A. W. Kemp. Book review: Statistical Distributions in Ecological Work, by J. K. Ord, G. P. Patil, C. Taillie. Biometrics, 37(4):858, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530178.

Jeffers:1981:BRS

[287] J. N. R. Jeffers. Book review: Systems Analysis of Ecosystems, by G. S. Innis, R. V. O’Neill. Biometrics, 37(4):858–859, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530179.

Calinski:1981:BRP

[288] T. Caliński. Book review: Principles and Procedures of Statistics: a Biometrical Approach, by R. G. D. Steel, J. H. Torrie. Biometrics, 37(4):859–860, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530180.

Cannings:1981:BRD

[289] C. Cannings. Book review: Developments in Statistics, Vol. 3, by Paruchuri R. Krishnaiah. Biometrics, 37(4):860, December 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530181.

Atkinson:1981:BRI

[290] A. C. Atkinson. Book review: Identification of Outliers, by D. M. Hawkins. Biometrics, 37(4):860–861, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530182.

Titterington:1981:BRF

[291] D. M. Titterington. Book review: Finite Mixture Distributions, by B. S. Everitt, D. J. Hand. Biometrics, 37(4):861, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530183.

Fisher:1981:BRO

[292] N. I. Fisher. Book review: Order Statistics, by Herbert A. David. Biometrics, 37(4):861–862, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530184.

Patterson:1981:BRR

[293] H. D. Patterson. Book review: Regression Diagnostics: Identifying Influential Data and Sources of Collinearity, by D. A. Belsley, E. Kuh, R. E. Welsch. Biometrics, 37(4):862–863, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530185.

Sprent:1981:BRA

[294] P. Sprent. Book review: Applied Regression Analysis, by N. R. Draper, H. Smith. Biometrics, 37(4):863, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530186.

Freeman:1981:BRR

[295] G. H. Freeman. Book review: Regression: a Second Course in Statistics, by T. H. Wonnacott, R. J. Wonnacott. Biometrics, 37(4):863–864, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530187.

Searle:1981:BRT

[296] S. R. Searle. Book review: The Theory of Linear Models and Multivariate Analysis, by S. F. Arnold. Biometrics, 37(4):864, December 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530188.

Dagnelie:1981:BRL

[297] P. Dagnélie. Book review: L’Analyse des Données, by J. M. Bouroche, G. Saporta. Biometrics, 37(4):864–865, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530189.

Tomassone:1981:BRO

[298] R. Tomassone. Book review: Optimisation en Classification Automatique, Tome 1 et 2, by E. Diday. Biometrics, 37(4):865, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530190.

Ellenberg:1981:BRC

[299] Jonas H. Ellenberg. Book review: Clinical Trials, by D. Schwartz, R. Flamant, J. Lellouch, M. J. R. Healy. Biometrics, 37(4):865–866, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530191.

Schmitz:1981:BRS

[300] P. I. M. Schmitz. Book review: Statistical Methods in Cancer Research. Vol. 1: The Analysis of Case-Control Studies, by N. E. Breslow, N. E. Day. Biometrics, 37(4):866–867, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530192.

Everitt:1981:BRS

[301] B. S. Everitt. Book review: Statistical Methods for Rates and Proportions, by J. L. Fleiss. Biometrics, 37(4):867, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530193.

Pearson:1981:BRS

[302] J. C. G. Pearson. Book review: Sampling for Health Professionals, by P. A. Levy, S. Lemeshow. Biometrics, 37(4):867–868, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530194.

Covall:1981:BRM

[303] David Covall. Book review: Modelling and Optimization of Complex Systems, by G. I. Marchuk. Biometrics, 37(4):868, December 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530195.

Katz:1981:BRS

[304] R. W. Katz. Book review: Statistical Analysis of Weather Modification Experiments, by E. J. Wegman, D. J. DePriest. Biometrics, 37(4):868, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530196.

Anonymous:1981:BRA

[305] Anonymous. Book review: Approximation Theorems of Mathematical Statistics, by R. J. Serfling. Biometrics, 37(4):869, December 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530199.

Anonymous:1981:BRT

[306] Anonymous. Book review: Time Series: Data Analysis and Theory, by D. R. Brillinger. Biometrics, 37(4):869, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530198.

White:1981:BRS

[307] Gary C. White. Book review: Statistical Applications in the Spatial Sciences, by N. Wrigley. Biometrics, 37(4):869, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530197.

Anonymous:1981:NAd

[308] Anonymous. News and announcements. Biometrics, 37(4):873–874, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530200.

Anonymous:1981:AAL

[309] Anonymous. Addition to assessing laboratory evidence for neoplastic activity. Biometrics, 37(4):875, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530203.

Anonymous:1981:CAT

[310] Anonymous. Correction: Analysis of the Two-Period Repeated Measurements Crossover Design with Application to Clinical Trials. Biometrics, 37(4):875, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530201. See [4].

Anonymous:1981:CID

[311] Anonymous. Corrections: An illness-death model for the study of the carcinogenic process using survival/sacrifice data. Biometrics, 37(4):875, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530202. See [41].

Anonymous:1981:CPL

[312] Anonymous. Correction: The probability of losing an allele when diploid genotypes are sampled. Biometrics, 37(4):875, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530204. See [112].

Gregorius:1981:CPL

[313] H.-R. Gregorius. Correction to: “The probability of losing an allele when diploid genotypes are sampled”. Biometrics, 37(4):875, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See

Anonymous:1981:COD

[314] Anonymous. Correction: Optimal designs for the analysis of interactive effects of two carcinogens or other toxicants. Biometrics, 37(4):876, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530205. See [148].

Anonymous:1981:PPBd

[315] Anonymous. Papers to be published in Biometrics. Biometrics, 37(4): 877, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530206.

Anonymous:1982:MJC

[316] Anonymous. In memoriam: Jerome Cornfield, 1912–1979. Biometrics, 38(p1):v–vi, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529845.

Seigel:1982:F

[317] Daniel G. Seigel. Foreword. Biometrics, 38(p1):1–2, March 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529846.

Greenhouse:1982:T

[318] Samuel W. Greenhouse. A tribute. Biometrics, 38(p1):3–6, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529847.

Fredrickson:1982:R

[319] Donald S. Fredrickson. Remarks. Biometrics, 38(p1):7, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529848.

Zelen:1982:CJC

[320] Marvin Zelen. The contributions of Jerome Cornfield to the theory of statistics. Biometrics, 38(p1):11–15, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529849.

Mantel:1982:JCS

[321] Nathan Mantel. Jerome Cornfield and statistical applications to laboratory research: a personal reminiscence. Biometrics, 38(p1):17–23, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529850.

Ederer:1982:JCC

[322] Fred Ederer. Jerome Cornfield’s contributions to the conduct of clinical trials. Biometrics, 38(p1):25–32, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529851.

Greenhouse:1982:JCC

[323] Samuel W. Greenhouse. Jerome Cornfield’s contributions to epidemiology. Biometrics, 38(p1):33–45, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529852.

Anonymous:1982:JCP

[324] Anonymous. Jerome Cornfield’s publications. Contributions to the literature as they appeared from 1941 through 1980. Biometrics, 38(p1): 47–52, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529853.

Feinleib:1982:BNM

[325] Manning Feinleib, Peter M. Lambert, Truls Zeiner-Henriksen, Eugene Rogot, Barbara M. Hunt, and Lillian Ingster-Moore. The British– Norwegian Migrant Study — analysis of parameters of mortality differentials associated with angina. Biometrics, 38(p1):55–71, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529854. See comments [326].

Haenszel:1982:DSC

[326] William Haenszel. Discussion: Some comments on the British–Norwegian Migrant Study. Biometrics, 38(p1):71–74, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529855. See [325].

Geisser:1982:APE

[327] Seymour Geisser. Aspects of the predictive and estimative approaches in the determination of probabilities. Biometrics, 38(p1):75–85, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529856. See comments [328]

and rejoinder [329].

DeGroot:1982:DCR

[328] Morris H. DeGroot. Discussion: Comments on the role of parameters in the predictive approach to statistics. Biometrics, 38(p1):86–91, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529857. See [327] and rejoinder

[329].

Geisser:1982:RD

[329] Seymour Geisser. Rejoinder to discussant. Biometrics, 38(p1):92–93, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529858. See [327,

328].

Stamler:1982:DCH

[330] Jeremiah Stamler. Diet and coronary heart disease. Biometrics, 38(p1): 95–114, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529859.

Wynder:1982:DCB

[331] Ernst L. Wynder. Discussion: Cornfield and biological plausibility. Biometrics, 38(p1):114–118, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529860.

Armitage:1982:ALD

[332] P. Armitage. The assessment of low-dose carcinogenicity. Biometrics, 38(p1):119–129, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529861.

vanRyzin:1982:DAL

[333] John van Ryzin. Discussion: [the assessment of low-dose carcinogenicity]. Biometrics, 38(p1):130–139, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529862.

Meier:1982:CRS

[334] Paul Meier. Current research in statistical methodology for clinical trials. Biometrics, 38(p1):141–150, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529863.

Chalmers:1982:DCD

[335] Thomas C. Chalmers. Discussion: Combination of data from randomized control trials. Biometrics, 38(p1):150–153, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529864.

Lilienfeld:1982:SAC

[336] Abraham M. Lilienfeld. Some aspects of cancer epidemiology. Biometrics, 38(p1):155–160, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529865.

Anonymous:1982:FMa

[337] Anonymous. Front matter. Biometrics, 38(p1):i–53, March 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529844.

Anonymous:1982:FMb

[338] Anonymous. Front matter. Biometrics, 38(1):i–iv, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530283.

Hazelrig:1982:PSA

[339] J. B. Hazelrig, M. E. Turner, and E. H. Blackstone. Parametric survival analysis combining longitudinal and cross-sectional-censored and interval-censored data with concomitant information. Biometrics, 38(1): 1–15, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530284.

Emerson:1982:NCI

[340] John D. Emerson. Nonparametric confidence intervals for the median in the presence of right censoring. Biometrics, 38(1):17–27, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530285. See correction [773].

Brookmeyer:1982:CIM

[341] Ron Brookmeyer and John Crowley. A confidence interval for the median survival time. Biometrics, 38(1):29–41, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530286.

Kodell:1982:NJE

[342] Ralph L. Kodell, Gary W. Shaw, and Alan M. Johnson. Nonparametric joint estimators for disease resistance and survival functions in survival/ sacrifice experiments. Biometrics, 38(1):43–58, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530287.

Begg:1982:SUP

[343] Colin B. Begg and Martin Larson. A study of the use of the probabilityof-being-in-response function as a summary of tumor response data. Biometrics, 38(1):59–66, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530288.

Andersen:1982:TGF

[344] Per Kragh Andersen. Testing goodness of fit of Cox’s regression and life model. Biometrics, 38(1):67–77, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2530289. See corrections [1004].

Lange:1982:CED

[345] Kenneth Lange. Calculation of the equilibrium distribution for a deleterious gene by the finite Fourier transform. Biometrics, 38(1):79–86, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530290.

Ghai:1982:CAR

[346] G. L. Ghai. Covariances among relatives in populations under mixed selffertilization and random mating. Biometrics, 38(1):87–92, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530291.

Haber:1982:TII

[347] Michael Haber. Testing for independence in intraclass contingency tables. Biometrics, 38(1):93–103, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530292.

Rosner:1982:SMO

[348] Bernard Rosner. Statistical methods in ophthalmology: an adjustment for the intraclass correlation between eyes. Biometrics, 38(1):105–114, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530293.

Longini:1982:HCT

[349] Ira M. Longini, Jr. and James S. Koopman. Household and community transmission parameters from final distributions of infections in households. Biometrics, 38(1):115–126, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530294.

Byth:1982:RDB

[350] Karen Byth. On robust distance-based intensity estimators. Biometrics, 38(1):127–135, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530295.

Schreuder:1982:TUD

[351] Hans T. Schreuder, Helen T. Bhattacharyya, and Joe P. McClure. Towards a unified distribution theory for stand variables using the S BBB distribution. Biometrics, 38(1):137–142, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530296.

Fleming:1982:OSM

[352] Thomas R. Fleming. One-sample multiple testing procedure for Phase II clinical trials. Biometrics, 38(1):143–151, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530297.

Pocock:1982:IAR

[353] Stuart J. Pocock. Interim analyses for randomized clinical trials: The group sequential approach. Biometrics, 38(1):153–162, March 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530298.

Schoenfeld:1982:NCN

[354] David A. Schoenfeld and Jane R. Richter. Nomograms for calculating the number of patients needed for a clinical trial with survival as an endpoint. Biometrics, 38(1):163–170, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530299.

Abdelbasit:1982:EDJ

[355] K. M. Abdelbasit and R. L. Plackett. Experimental design for joint action. Biometrics, 38(1):171–179, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530300.

Volund:1982:CMB

[356] Aage Vølund. Combination of multivariate bioassay results. Biometrics, 38(1):181–190, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530301.

Knoke:1982:DAD

[357] James D. Knoke. Discriminant analysis with discrete and continuous variables. Biometrics, 38(1):191–200, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530302. See correction [556].

Bithell:1982:ETL

[358] John F. Bithell and Jürgen Wahrendorf. Estimation of the true length of broken molecules. Biometrics, 38(1):201–213, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530303.

Tarone:1982:UHCa

[359] Robert E. Tarone. The use of historical control information in testing for a trend in proportions. Biometrics, 38(1):215–220, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530304.

Berkey:1982:CTL

[360] Catherine S. Berkey. Comparison of two longitudinal growth models for preschool children. Biometrics, 38(1):221–234, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530305.

Gart:1982:RTD

[361] John J. Gart, Andrew F. Siegel, and Rebecca Z. German. Rarefaction and taxonomic diversity. Biometrics, 38(1):235–241, March 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530306.

Patil:1982:ONE

[362] S. A. Patil, J. L. Kovner, and K. P. Burnham. Optimum nonparametric estimation of population density based on ordered distances. Biometrics, 38(1):243–248, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530307.

Bucher:1982:ENS

[363] K. D. Bucher and H. G. Schrott. The effect of nonrandom sampling on familial correlations. Biometrics, 38(1):249–253, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530308.

Finney:1982:ICS

[364] D. J. Finney. Intervention and correlated sequences of observations. Biometrics, 38(1):255–267, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530309.

Armitage:1982:ER

[365] Peter Armitage. Editor’s report, 1980–1981. Biometrics, 38(1):269–271, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530310.

Gart:1982:RES

[366] John J. Gart. Report of the Editor of shorter communications, 1980– 1981. Biometrics, 38(1):273–274, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530311.

Schenzle:1982:STI

[367] Dieter Schenzle, J. Wahrendorf, and C. C. Brown. A simplified test for the independence of the joint action of two drugs. Biometrics, 38(1):275– 276, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530312.

Pledger:1982:WDT

[368] Gordon Pledger, David Hall, and A. Lawrence Gould. Withdrawals from drug trials. Biometrics, 38(1):276–278, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530313.

Davis:1982:DLC

[369] Charles S. Davis. The distribution of a linear combination of chi square variables. Biometrics, 38(1):279, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530314.

Anonymous:1982:Aa

[370] Anonymous. Abstracts. Biometrics, 38(1):281–282, March 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530315.

Lutz:1982:BRC

[371] W. Lutz. Book review: Computer-Assisted Medical Decision-Making, by Homer R. Warner. Biometrics, 38(1):283, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530316.

Clarke:1982:BRS

[372] K. R. Clarke. Book reviews: Stereological Methods. Vol. 1: Practical Methods for Biological Morphometry, by E. R. Weibel; Stereological Methods. Vol. 2: Theoretical Foundations by E. R. Weibel. Biometrics, 38(1):283–284, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530317.

Pielou:1982:BRE

[373] E. C. Pielou. Book review: Elementary Mathematical Ecology, by John Vandermeer. Biometrics, 38(1):284–285, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530318.

Bartlett:1982:BRP

[374] M. S. Bartlett. Book review: Proceedings of the Vito Volterra Symposium on Mathematical Models in Biology, Rome 1979, by C. Barigozzi. Biometrics, 38(1):285–286, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530319.

Blackith:1982:BRP

[375] R. E. Blackith. Book review: Perspectives in Primate Biology, by E. H. Ashton, R. L. Holmes. Biometrics, 38(1):286, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530320.

Holgate:1982:BRG

[376] Philip Holgate. Book review: The Geometry of Population Genetics, by E. Akin. Biometrics, 38(1):286–287, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530321.

Baines:1982:BRT

[377] C. R. Baines. Book review: Thermal Properties of Foods and Agricultural Materials, by N. N. Mohsenin. Biometrics, 38(1):287, March 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530323.

Hawkes:1982:BRA

[378] A. G. Hawkes. Book review: Analysis of Neural Networks, by Uwe Heiden. Biometrics, 38(1):287, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530322.

Renshaw:1982:BRS

[379] Eric Renshaw. Book review: Spatial Processes, Models and Applications, by A. D. Cliff, J. K. Ord. Biometrics, 38(1):287–288, March 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530324.

Preece:1982:BRE

[380] D. A. Preece. Book review: Experiments with Mixtures: Designs, Models, and the Analysis of Mixture Data, by J. A. Cornell. Biometrics, 38(1): 288, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530325.

Kemp:1982:BRAa

[381] A. W. Kemp. Book review: Analysis with Standard Contagious Distributions, by J. B. Douglas. Biometrics, 38(1):289, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530326.

Escoufier:1982:BRI

[382] Y. Escoufier. Book review: Introduction aux Analyses Statistiques Multidimensionnelles, by J. Lefebvre. Biometrics, 38(1):289–290, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530327.

Southward:1982:BRA

[383] G. Morris Southward. Book review: Analysis of Categorical Data: Dual Scaling and Its Applications, by S. Nishisato. Biometrics, 38(1):290, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530328.

Kempton:1982:BRT

[384] R. A. Kempton. Book review: Techniques in Operational Research, Vol. 2: Models, Search and Randomization, by B. W. Conolly. Biometrics, 38(1):290–291, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530329.

McPherson:1982:BRS

[385] Klim McPherson. Book review: Scientific Strategies to Save Your Life, by D. J. Irwin Bross. Biometrics, 38(1):291, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530330.

Parmeggiani:1982:BRS

[386] A. Parmeggiani. Book reviews: Studi in Onore di Paolo Fortunati. Vol. I: Statistica, Statistica Economica, Demografia by Università Degli Studi di Bologna-Istituto di Statistica; Studi in Onore di Paolo Fortunati. Vol. II: Economia, Diritto, Altri Argomenti by Università Degli Studi di Bologna — Istituto di Statistica. Biometrics, 38(1):291–292, March 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530331.

Anonymous:1982:BRF

[387] Anonymous. Book review: The Fourfold Table up to N = 80, by H.-P. Kruger, W. Lehmacher, K.-D. Wall. Biometrics, 38(1):292, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530333.

Anonymous:1982:BRI

[388] Anonymous. Book review: Introduction to Statistics: a Nonparametric Approach for the Social Sciences, by C. Leach. Biometrics, 38(1):292, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530334.

Campbell:1982:BRS

[389] R. C. Campbell. Book review: Statistical Methods, by G. W. Snedecor, W. G. Cochran. Biometrics, 38(1):292, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530332.

Anonymous:1982:BSFa

[390] Anonymous. The Biometric Society financial statements and Auditor’s report January 31, 1981. Biometrics, 38(1):293–295, March 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530335.

Anonymous:1982:NAa

[391] Anonymous. News and announcements. Biometrics, 38(1):297–298, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530336.

Anonymous:1982:PPBa

[392] Anonymous. Papers to be published in Biometrics. Biometrics, 38(1): 299, March 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530337.

Anonymous:1982:FMc

[393] Anonymous. Front matter. Biometrics, 38(2):i–iv, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530444.

Jolly:1982:MRM

[394] G. M. Jolly. Mark–recapture models with parameters constant in time. Biometrics, 38(2):301–321, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530445.

Watterson:1982:TSS

[395] G. A. Watterson. Testing selection at a single locus. Biometrics, 38(2): 323–331, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530446.

Tong:1982:SAI

[396] Y. L. Tong. Some applications of inequalities for extreme order statistics to a genetic selection problem. Biometrics, 38(2):333–339, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530447.

Siegel:1982:RCB

[397] Andrew F. Siegel and Richard H. Benson. A robust comparison of biological shapes. Biometrics, 38(2):341–350, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530448.

Buser:1982:DNC

[398] Mauro W. Buser and Cesare Baroni-Urbani. A direct nondimensional clustering method for binary data. Biometrics, 38(2):351–360, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530449.

Paul:1982:APA

[399] S. R. Paul. Analysis of proportions of affected foetuses in teratological experiments. Biometrics, 38(2):361–370, June 1982. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530450.

Sundararaj:1982:EET

[400] N. Sundararaj, C. Le, and R. McHugh. Estimation of the efficacy of thermal microbial disinfection. Biometrics, 38(2):371–376, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530451.

Hill:1982:RRL

[401] MaryAnn Hill and W. J. Dixon. Robustness in real life: a study of clinical laboratory data. Biometrics, 38(2):377–396, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530452.

Royston:1982:BBT

[402] J. P. Royston. Basal body temperature, ovulation and the risk of conception, with special reference to the lifetimes of sperm and egg. Biometrics, 38(2):397–406, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530453.

Anderson:1982:ACI

[403] J. R. Anderson, L. Bernstein, and M. C. Pike. Approximate confidence intervals for probabilities of survival and quantiles in life-table analysis. Biometrics, 38(2):407–416, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530454.

Dinse:1982:NEP

[404] Gregg E. Dinse. Nonparametric estimation for partially-complete time and type of failure data. Biometrics, 38(2):417–431, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530455.

James:1982:MMA

[405] I. R. James and M. R. Segal. On a method of mortality analysis incorporating age-year interaction, with application to prostate cancer mortality. Biometrics, 38(2):433–443, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530456.

Korn:1982:AET

[406] Edward L. Korn. The asymptotic efficiency of tests using misclassified data in contingency tables. Biometrics, 38(2):445–450, June 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530457.

Oakes:1982:CTI

[407] David Oakes. A concordance test for independence in the presence of censoring. Biometrics, 38(2):451–455, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530458.

Tarone:1982:UHCb

[408] Robert E. Tarone. The use of historical control information in testing for a trend in Poisson means. Biometrics, 38(2):457–462, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530459.

Matthews:1982:TCH

[409] David E. Matthews and Vernon T. Farewell. On testing for a constant hazard against a change-point alternative. Biometrics, 38(2):463–468, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530460. See correc-

tions [1177].

Buckland:1982:NFS

[410] Stephen T. Buckland. A note on the Fourier series model for analysing line transect data. Biometrics, 38(2):469–477, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530461.

Lloyd:1982:RAC

[411] F. Thomas Lloyd, H. David Muse, and William L. Hafley. A regression application for comparing growth potential of environments at different points in the growth cycle. Biometrics, 38(2):479–484, June 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530462.

Pregibon:1982:RFS

[412] Daryl Pregibon. Resistant fits for some commonly used logistic models with medical applications. Biometrics, 38(2):485–498, June 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530463.

Kimball:1982:SMF

[413] A. W. Kimball, H. L. Lau, M. G. Maguire, W. C. Oppel, and S. E. Linkins. Statistical methods for fetal monitoring with alpha-fetoprotein in maternal serum. Biometrics, 38(2):499–509, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530464.

Aitkin:1982:IRM

[414] Murray Aitkin, Brian Francis, Paul F. Velleman, and Harold V. Henderson. Interactive regression modelling. Biometrics, 38(2):511–516, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530465.

Grieve:1982:TPC

[415] A. P. Grieve. The two-period changeover design in clinical trials. Biometrics, 38(2):517, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530466.

Moreau:1982:EHR

[416] T. Moreau. Estimation of hazards-ratio in the two-sample problem. Biometrics, 38(2):518–519, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530467.

Anonymous:1982:Ab

[417] Anonymous. Abstracts. Biometrics, 38(2):521–522, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530468.

Young:1982:BRS

[418] Peter A. Young. Book review: Statistics in the Pharmaceutical Industry, by C. R. Buncher, J.-Y. Tsay. Biometrics, 38(2):523–524, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530469.

Preece:1982:BRA

[419] D. A. Preece. Book review: Agricultural Experimentation: Design and Analysis, by T. M. Little, F. J. Hills. Biometrics, 38(2):524–525, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530470.

Barnes:1982:BRB

[420] A. Barnes. Book review: The Biometry of Plant Growth, by D. R. Causton, J. C. Venus. Biometrics, 38(2):525–526, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530471.

Beddington:1982:BRM

[421] J. R. Beddington. Book review: Management and Analysis of Biological Populations, by Bean-San Goh. Biometrics, 38(2):526, June 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530472.

Dagnelie:1982:BRM

[422] P. Dagnélie. Book review: Multivariate Methods in Ecological Work, by L. Orloci, C. R. Rao, W. M. Stiteler. Biometrics, 38(2):526–527, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530473.

Craik:1982:BRM

[423] A. D. D. Craik. Book review: Mechanics of Swimming and Flying, by S. Childress. Biometrics, 38(2):527, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530474.

Paterson:1982:BRC

[424] J. Y. F. Paterson. Book review: Carbohydrate Metabolism, by C. Cobelli, R. N. Bergman. Biometrics, 38(2):527–528, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530475.

Falconer:1982:BRE

[425] D. S. Falconer. Book review: Eléménts de Génétique Quantitative, by L. Ollivier. Biometrics, 38(2):528, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530476.

Green:1982:BRN

[426] P. J. Green. Book review: Nonparametric Geostatistics, by S. Henley. Biometrics, 38(2):528, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530477.

Ripley:1982:BRF

[427] B. D. Ripley. Book review: Figurale Wechselwirkung im Gesichtfeld (Stochastic Interactions in the Visual Field), by Günter Lehmann. Biometrics, 38(2):529, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530478.

Williams:1982:BRA

[428] David Williams. Book review: Applied Statistics: Principles and Examples, by D. R. Cox, E. J. Snell. Biometrics, 38(2):529, June 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530479.

Escoufier:1982:BRA

[429] Y. Escoufier. Book review: Analyse Typologique. Theories et Applications, by J. L. Chandon, S. Pinson. Biometrics, 38(2):529–530, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530480.

Rutherford:1982:BRS

[430] A. A. Rutherford. Book review: Sampling Methods for Censuses and Surveys, by F. Yates. Biometrics, 38(2):530, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530481.

Skinner:1982:BRS

[431] C. J. Skinner. Book review: Sampling from a Finite Population, by J. Hajek, V. Dupac. Biometrics, 38(2):530–531, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530482.

Anonymous:1982:BRBa

[432] Anonymous. Book review: BMDP Statistical Software 1981, by W. J. Dixon. Biometrics, 38(2):531, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530483.

Anonymous:1982:BSa

[433] Anonymous. The Biometric Society. Biometrics, 38(2):533, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530484.

Anonymous:1982:NAb

[434] Anonymous. News and announcements. Biometrics, 38(2):535, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530485.

Anonymous:1982:PPN

[435] Anonymous. Papers to be published in the next issue. Biometrics, 38(2): 537, June 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530486.

Armitage:1982:I

[436] Peter Armitage. Introduction. Biometrics, 38(3):540, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530038.

Bryce:1982:PGE

[437] G. Rex Bryce. Preface by the guest editor. Biometrics, 38(3):540, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530039.

Cox:1982:BIP

[438] D. R. Cox and P. McCullagh. A Biometrics invited paper with discussion. Some aspects of analysis of covariance. Biometrics, 38(3):541–561, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530040.

Koch:1982:RSS

[439] Gary G. Koch, Ingrid A. Amara, Gordon W. Davis, and Dennis B. Gillings. A review of some statistical methods for covariance analysis of categorical data. Biometrics, 38(3):563–595, September 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530041.

Quade:1982:NAC

[440] Dana Quade. Nonparametric analysis of covariance by matching. Biometrics, 38(3):597–611, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530042.

Lane:1982:ACS

[441] Peter W. Lane and John A. Nelder. Analysis of covariance and standardization as instances of prediction. Biometrics, 38(3):613–621, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530043.

Henderson:1982:ACM

[442] Charles R. Henderson, Jr. Analysis of covariance in the mixed model: Higher-level, nonhomogeneous, and random regressions. Biometrics, 38(3):623–640, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530044.

Hendrix:1982:CAH

[443] Leland J. Hendrix, Melvin W. Carter, and Del T. Scott. Covariance analyses with heterogeneity of slopes in fixed models. Biometrics, 38(3):641– 650, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530045.

Urquhart:1982:ACW

[444] N. Scott Urquhart. Adjustment in covariance when one factor affects the covariate. Biometrics, 38(3):651–660, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530046.

Breslow:1982:CAR

[445] N. Breslow. Covariance adjustment of relative-risk estimates in matched studies. Biometrics, 38(3):661–672, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530047.

Holford:1982:CAC

[446] Theodore R. Holford. Covariance analysis for case-control studies with small blocks. Biometrics, 38(3):673–683, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530048.

Johnson:1982:CAS

[447] Mark E. Johnson, H. Dennis Tolley, Maurice C. Bryson, and Aaron S. Goldman. Covariate analysis of survival data: a small-sample study of Cox’s model. Biometrics, 38(3):685–698, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530049.

Birch:1982:RAC

[448] Jeffrey B. Birch and Raymond H. Myers. Robust analysis of covariance. Biometrics, 38(3):699–713, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530050.

Conover:1982:ACU

[449] W. J. Conover and Ronald L. Iman. Analysis of covariance using the rank transformation. Biometrics, 38(3):715–724, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530051. See correction [1001].

Hemmerle:1982:ACA

[450] William J. Hemmerle. Analysis of covariance algorithms. Biometrics, 38(3):725–736, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530052.

Searle:1982:SDF

[451] S. R. Searle and G. F. S. Hudson. Some distinctive features of output from statistical computing packages for analysis of covariance. Biometrics, 38(3):737–745, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530053.

Bingham:1982:TAC

[452] Christopher Bingham and Stephen E. Fienberg. Textbook analysis of covariance — is it correct? Biometrics, 38(3):747–753, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530054.

Smouse:1982:MAH

[453] Peter E. Smouse and Robert C. Williams. Multivariate analysis of HLAdisease associations. Biometrics, 38(3):757–768, September 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530055.

Farewell:1982:EGF

[454] V. T. Farewell. The estimation of gene frequency, based on a particular mixed leukocyte culture experiment. Biometrics, 38(3):769–775, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530056.

Staudte:1982:PBM

[455] Robert G. Staudte. Planning blocked mitosis experiments for efficient estimation of population-doubling time and cell-cycle time. Biometrics, 38(3):777–785, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530057.

Allen:1982:UPI

[456] David M. Allen and David C. Jordan. The use of prior information for prediction. Biometrics, 38(3):787–799, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530058.

Wacholder:1982:PVT

[457] Sholom Wacholder and Clarice Ring Weinberg. Paired versus two-sample design for a clinical trial of treatments with dichotomous outcome: Power considerations. Biometrics, 38(3):801–812, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530059.

Jensen:1982:ERU

[458] D. R. Jensen. Efficiency and robustness in the use of repeated measurements. Biometrics, 38(3):813–825, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530060.

Spurrier:1982:CTR

[459] John D. Spurrier, John E. Hewett, and Zuhdi Lababidi. Comparison of two regression lines over a finite interval. Biometrics, 38(3):827–836, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530061.

Hewett:1982:CTR

[460] John E. Hewett and Zuhdi Lababidi. Comparison of three regression lines over a finite interval. Biometrics, 38(3):837–841, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530062.

Schenzle:1982:PDE

[461] Dieter Schenzle. Problems in drawing epidemiological inferences by fitting epidemic chain models to lumped data. Biometrics, 38(3):843–847, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530063.

Wallenstein:1982:RMR

[462] Sylvan Wallenstein, F. Yates, and Murray Aitkin. Regression models for repeated measurements. Biometrics, 38(3):849–853, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530064.

Anonymous:1982:OHO

[463] Anonymous. Obituary: Herman Otto Hartley, 1912–1980. Biometrics, 38(3):855–856, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530065.

Seigel:1982:ERR

[464] Daniel Seigel, Samuel W. Greenhouse, Lawrence L. Kupper, and John M. Karon. Estimation of relative risk. Biometrics, 38(3):857–859, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530066.

Ricker:1982:LRN

[465] W. E. Ricker, P. Sprent, and G. R. Dolby. Linear regression for naturally variable data. Biometrics, 38(3):859–860, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530067.

Anonymous:1982:Ac

[466] Anonymous. Abstracts. Biometrics, 38(3):861–862, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530068.

Mead:1982:BRB

[467] R. Mead. Book review: Biometry, by R. R. Sokal, F. J. Rohlf. Biometrics, 38(3):863–864, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530069.

Belagyi:1982:BRI

[468] J. Belagyi. Book review: Introduction to Biomathematics, by J. Izsak, P. Juhasz-Nagy, Z. Varga. Biometrics, 38(3):864, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530070.

Kempthorne:1982:BRR

[469] Oscar Kempthorne. Book review: Randomization Tests, by Eugene S. Edgington. Biometrics, 38(3):864–867, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530071.

Freeman:1982:BRF

[470] G. H. Freeman. Book review: Factorial Designs, by B. L. Raktoe, A. Hedayat, W. T. Federer. Biometrics, 38(3):867, September 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530072.

Sprent:1982:BRR

[471] P. Sprent. Book reviews: Regression Analysis and its Applications, by R. F. Gunst, R. L. Mason; Applied Linear Regression by S. Weisberg. Biometrics, 38(3):867–868, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530073.

Berchtold:1982:BRVa

[472] W. Berchtold. Book review: Verteilungsfreie Methoden in der Biostatistik, Band II, by G. A. Lienert. Biometrics, 38(3):868, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530074.

Marriott:1982:BRA

[473] F. H. C. Marriott. Book review: Analyse de Données et Informatique, by R. Tomassone. Biometrics, 38(3):868–869, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530075.

Gordon:1982:BRI

[474] A. D. Gordon. Book review: Interpreting Multivariate Data, by V. Barnett. Biometrics, 38(3):869, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530076.

Gower:1982:BRM

[475] J. C. Gower. Book review: Morphometrics: The Multivariate Analysis of Biological Data, by R. A. Pimentel. Biometrics, 38(3):869–870, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530077.

Dunn:1982:BRP

[476] O. J. Dunn. Book review: Probability Inequalities in Multivariate Distributions, by Y. L. Tong. Biometrics, 38(3):870, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530078.

Jupp:1982:BRC

[477] P. E. Jupp. Book review: Circular Statistics in Biology, by E. Batschelet. Biometrics, 38(3):870, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530079.

Oakes:1982:BRS

[478] D. Oakes. Book review: Survival Analysis, by R. G. Miller. Biometrics, 38(3):871, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530080.

Davies:1982:BRS

[479] Paul Davies. Book review: Statistical Modelling Techniques, by S. S. Shapiro, A. J. Gross. Biometrics, 38(3):871–872, September 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530081.

Sutherland:1982:BRS

[480] I. Sutherland. Book review: Statistical Methods for Comparative Studies: Techniques for Bias Reduction, by S. Anderson, A. Auquier, W. W. Hauck, D. Oakes, W. Vandaele, H. I. Weisberg. Biometrics, 38(3):872, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530082.

Heady:1982:BRF

[481] J. A. Heady. Book review: Foundations of Epidemiology, by A. M. Lilienfeld, D. E. Lilienfeld. Biometrics, 38(3):872–873, September 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530083.

Pocock:1982:BRF

[482] S. J. Pocock. Book review: Fundamentals of Clinical Trials, by L. M. Friedman, C. D. Furberg, D. L. Demets. Biometrics, 38(3):873, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530084.

Diamond:1982:BRE

[483] Ian Diamond. Book reviews: Un Essai de Modélisation de la Répartition des Décès selon l’Âge et la Cause dans les Pays Industrialisés, by Josianne Duchene; Perspectives Démographiques par Sexe et par Âge by Christine Wattelar. Biometrics, 38(3):873–874, September 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530085.

Appleton:1982:BRM

[484] David R. Appleton. Book review: Mathematical Modelling and Computers in Endocrinology, by J. E. A. McIntosh, R. P. McIntosh. Biometrics, 38(3):874–875, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530086.

Oldham:1982:BRB

[485] P. D. Oldham. Book review: Blood Pressure Study, 1979, by Society of Actuaries and Association of Life Insurance Medical Directors of America. Biometrics, 38(3):875–876, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530087.

Marshall:1982:BRS

[486] T. F. De C. Marshall. Book review: Steroid Contraception and the Risk of Neoplasia, by W. H. O. Scientific Group. Biometrics, 38(3):876, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530089.

Tanner:1982:BRB

[487] J. M. Tanner. Book review: Build Study, 1979, by Society of Actuaries and Association of Life Insurance Medical Directors of America. Biometrics, 38(3):876, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530088.

Smith:1982:BRD

[488] R. H. Smith. Book review: Deterministic Mathematical Models in Population Ecology, by H. I. Freedman. Biometrics, 38(3):876–877, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530090.

Jolly:1982:BRD

[489] G. M. Jolly. Book review: Dynamics of Large Mammal Populations, by C. W. Fowler, T. D. Smith. Biometrics, 38(3):877, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530091.

Weir:1982:BRG

[490] B. S. Weir. Book review: The Genetics of Altruism, by S. A. Boorman, P. R. Levitt. Biometrics, 38(3):877–878, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530092.

Williams:1982:BRE

[491] D. A. Williams. Book review: Elementary Statistical Tables, by H. R. Neave. Biometrics, 38(3):878–879, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530093.

Anonymous:1982:BRM

[492] Anonymous. Book review: Modern Statistics: Methods and Applications, by R. V. Hogg. Biometrics, 38(3):879, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530094.

Anonymous:1982:BRPa

[493] Anonymous. Book review: The Product of Two Normally Distributed Random Variables, by W. Q. Meeker, Jr., L. W. Cornwell, L. Aroian. Biometrics, 38(3):879, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530095.

Anonymous:1982:BSb

[494] Anonymous. The Biometric Society. Biometrics, 38(3):881, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530096.

Anonymous:1982:NAc

[495] Anonymous. News and announcements. Biometrics, 38(3):882, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530097.

Anonymous:1982:PPBb

[496] Anonymous. Papers to be published in Biometrics. Biometrics, 38(3): 883, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530098.

Anonymous:1982:FMd

[497] Anonymous. Front matter. Biometrics, 38(3):i–754, September 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530037.

Anonymous:1982:FMe

[498] Anonymous. Front matter. Biometrics, 38(4):i–iv, December 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529868.

Anonymous:1982:VI

[499] Anonymous. Volume information. Biometrics, 38(4):x–xxxiii, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529867.

Morgan:1982:MP

[500] Byron J. T. Morgan. Modelling polyspermy. Biometrics, 38(4):885–898, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529869.

Whittemore:1982:RAC

[501] Alice S. Whittemore, Jerry W. McLarty, Noble Fortson, and Keaven Anderson. Regression analysis of cytopathological data. Biometrics, 38(4):899–905, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529870.

Currie:1982:EMM

[502] David J. Currie. Estimating Michaelis–Menten parameters: Bias, variance and experimental design. Biometrics, 38(4):907–919, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529871.

Dinse:1982:NEL

[503] Gregg E. Dinse and S. W. Lagakos. Nonparametric estimation of lifetime and disease onset distributions from incomplete observations. Biometrics, 38(4):921–932, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529872.

Hakulinen:1982:CSC

[504] Timo Hakulinen. Cancer survival corrected for heterogeneity in patient withdrawal. Biometrics, 38(4):933–942, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529873.

Breslow:1982:VMH

[505] N. E. Breslow and K. Y. Liang. The variance of the Mantel–Haenszel estimator. Biometrics, 38(4):943–952, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529874. See corrections [1005].

Berkey:1982:BAN

[506] Catherine S. Berkey. Bayesian approach for a nonlinear growth model. Biometrics, 38(4):953–961, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529875.

Laird:1982:REM

[507] Nan M. Laird and James H. Ware. Random-effects models for longitudinal data. Biometrics, 38(4):963–974, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529876.

Smith:1982:OTP

[508] S. P. Smith and R. L. Quaas. Optimal truncation points for independent culling-level selection involving two traits. Biometrics, 38(4):975–980, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529877.

Gunsett:1982:EPE

[509] F. C. Gunsett, K. N. Andriano, and J. J. Rutledge. Estimating the precision of estimates of genetic parameters realized from multiple-trait selection experiments. Biometrics, 38(4):981–989, December 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529878.

Krzanowski:1982:MCC

[510] W. J. Krzanowski. Mixtures of continuous and categorical variables in discriminant analysis: a hypothesis-testing approach. Biometrics, 38(4):991–1002, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529879.

Brittain:1982:OAC

[511] Erica Brittain and James J. Schlesselman. Optimal allocation for the comparison of proportions. Biometrics, 38(4):1003–1009, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529880.

Miller:1982:MSC

[512] Rupert Miller and David Siegmund. Maximally selected chi square statistics. Biometrics, 38(4):1011–1016, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529881.

Halpern:1982:MSC

[513] Jerry Halpern. Maximally selected chi square statistics for small samples. Biometrics, 38(4):1017–1023, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529882.

Walker:1982:AAS

[514] Alexander M. Walker. Anamorphic analysis: Sampling and estimation for covariate effects when both exposure and disease are known. Biometrics, 38(4):1025–1032, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529883.

Cuzick:1982:EPT

[515] Jack Cuzick. The efficiency of the proportions test and the logrank test for censored survival data. Biometrics, 38(4):1033–1039, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529884.

Farewell:1982:UMM

[516] V. T. Farewell. The use of mixture models for the analysis of survival data with long-term survivors. Biometrics, 38(4):1041–1046, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529885.

Davies:1982:MAM

[517] Mark Davies and Joseph L. Fleiss. Measuring agreement for multinomial data. Biometrics, 38(4):1047–1051, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529886.

Taulbee:1982:NUN

[518] John D. Taulbee. A note on the use of nonparametric methods in the statistical analysis of the two-period changeover design. Biometrics, 38 (4):1053–1055, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529887.

Pasternack:1982:ADR

[519] Bernard S. Pasternack and Roy E. Shore. Analysis of dichotomous response data from toxicological experiments involving stable laboratory mouse populations. Biometrics, 38(4):1057–1067, December 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529888.

Dolby:1982:RSM

[520] G. R. Dolby. The role of statistics in the methodology of the life sciences. Biometrics, 38(4):1069–1083, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529889.

Gianola:1982:QSE

[521] Daniel Gianola, Bruno Goffinet, and Michael G. Bulmer. Query: Sire evaluation with best linear unbiased predictors. Biometrics, 38(4):1085– 1088, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529890.

Beal:1982:BAN

[522] Stuart L. Beal, Darryl Katz, Stanley P. Azen, and Alan Schumitzky. Bayesian analysis of nonlinear models. Biometrics, 38(4):1089–1092, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529891.

Campbell:1982:CRG

[523] M. J. Campbell, Joseph L. Fleiss, Alex Tytun, and Hans K. Ury. The choice of relative group sizes for the comparison of independent proportions. Biometrics, 38(4):1093–1094, December 1982. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529892.

Ury:1982:HAL

[524] Hans K. Ury. Hauck’s approximate large-sample variance of the Mantel– Haenszel estimator. Biometrics, 38(4):1094–1095, December 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529893.

Finney:1982:PPC

[525] David J. Finney. Presentation of papers at conferences. Biometrics, 38(4):1096, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529894.

Bross:1982:GRE

[526] Irwin D. J. Bross. The Galilean rule and editorial decisions. Biometrics, 38(4):1097, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529895.

Anonymous:1982:Ad

[527] Anonymous. Abstracts. Biometrics, 38(4):1099–1120, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529896.

Lebreton:1982:BRT

[528] Jean-Dominique Lebreton. Book review: Théorie et Méthodes Statistiques: Exercices, by P. Dagnelie. Biometrics, 38(4):1121, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529897.

Kerridge:1982:BRI

[529] D. F. Kerridge. Book review: Ideas in Statistics, by J. L. Folks. Biometrics, 38(4):1121–1122, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529898.

Currie:1982:BRS

[530] I. D. Currie. Book review: Statistical Theory and Inference in Research, by T. A. Bancroft, Chien-Pai Han. Biometrics, 38(4):1122, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529899.

Bartlett:1982:BRS

[531] M. S. Bartlett. Book review: Statistics in Britain 1865–1930. The Social Construction of Scientific Knowledge, by G. A. MacKenzie. Biometrics, 38(4):1122–1123, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529900.

Krzanowski:1982:BRA

[532] W. J. Krzanowski. Book review: Applied Probability, by F. A. Haight. Biometrics, 38(4):1123, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529901.

Smith:1982:BRB

[533] Cedric A. B. Smith. Book review: Bayesian Statistics, by M. H. de Groot, J. M. Bernardo, D. V. Lindley, A. F. M. Smith. Biometrics, 38(4):1123–1124, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529902.

Renshaw:1982:BRT

[534] E. Renshaw. Book review: Time Series Models, by A. C. Harvey. Biometrics, 38(4):1124, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529903.

Brown:1982:BRM

[535] R. A. Brown. Book review: Mathematical Programming in Statistics, by T. S. Arthanari, Y. Dodge. Biometrics, 38(4):1124–1125, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529904.

Atkinson:1982:BRH

[536] A. C. Atkinson. Book review: The Handbook of Random Number Generation and Testing with TESTRAND Computer Code, by E. J. Dudewicz, T. G. Ralley. Biometrics, 38(4):1125–1126, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529905.

Pearce:1982:BRH

[537] S. C. Pearce. Book review: Handbook of Statistics, Vol. 1: Analysis of Variance, by P. R. Krishnaiah. Biometrics, 38(4):1126, December 1982.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529906.

Inkson:1982:BRI

[538] R. H. E. Inkson. Book review: Intermediate Statistical Methods, by G. Barrie Wetherill. Biometrics, 38(4):1126–1127, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529907.

Constable:1982:BRB

[539] R. L. Constable. Book review: Basic Statistics: Tales of Distributions, by C. Spatz, J. O. Johnston. Biometrics, 38(4):1127, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529908.

McCullagh:1982:BRA

[540] P. McCullagh. Book review: The Analysis of Categorical Data, by R. L. Plackett. Biometrics, 38(4):1127, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529909.

Elston:1982:BRG

[541] R. C. Elston. Book review: Genealogical and Genetic Structure, by C. Cannings, E. A. Thompson. Biometrics, 38(4):1128, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529911.

Vidovic:1982:BRZ

[542] Vitomir Vidovic. Book review: Zuchtwertschätzung beim Rind mit eimer ausführlichen Darstellung der BLUP-Metode, by L. Dempfle. Biometrics, 38(4):1128, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529910.

Edwards:1982:BRI

[543] A. W. F. Edwards. Book review: Introduction to Quantitative Genetics, by D. S. Falconer. Biometrics, 38(4):1128–1129, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529912.

Darby:1982:BRP

[544] S. C. Darby. Book review: Public Program Analysis. A New Categorical Data Approach, by R. N. Forthofer, R. G. Lehnen. Biometrics, 38 (4):1129–1130, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529913.

Machin:1982:BRP

[545] David Machin. Book review: Perspectives in Medical Statistics, by J. F. Bithell, R. Coppi. Biometrics, 38(4):1130–1131, December 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529914.

Kirkwood:1982:BRR

[546] T. B. L. Kirkwood. Book review: Reference Values in Laboratory Medicine: The Current State of the Art, by R. Gräsbeck, T. Alström. Biometrics, 38(4):1131, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529915.

Kemp:1982:BRAb

[547] A. W. Kemp. Book review: Applicable Mathematics of Non-Physical Phenomena, by F. Oliviera-Pinto, B. Conolly. Biometrics, 38(4):1131– 1132, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529916.

Hawkes:1982:BRO

[548] A. G. Hawkes. Book review: The Organization of the Cerebral Cortex, by G. Adelman, S. G. Dennis, F. O. Schmitt, F. G. Warden. Biometrics, 38(4):1132–1133, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529917.

Sinclair:1982:BRE

[549] C. D. Sinclair. Book review: Environmental Systems Optimization, by D. A. Haith. Biometrics, 38(4):1133, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529918.

Berchtold:1982:BRVb

[550] W. Berchtold. Book review: Verteilungsfreie Methoden in der Biostatistik, by G. A. Lienert. Biometrics, 38(4):1133–1134, December 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529919.

Anonymous:1982:BRBb

[551] Anonymous. Book review: Biometric Methods in Research Work, by J. Sváb. Biometrics, 38(4):1134, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529922.

Anonymous:1982:BRH

[552] Anonymous. Book review: Handbook of the Normal Distribution, by J. K. Patel, C. B. Read. Biometrics, 38(4):1134, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529920.

Anonymous:1982:BRPb

[553] Anonymous. Book review: Principes d’Experimentation, by P. Dagnelie. Biometrics, 38(4):1134, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529921.

Anonymous:1982:BSFb

[554] Anonymous. The Biometric Society financial statements and auditor’s report January 31, 1982. Biometrics, 38(4):1135–1137, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529923.

Anonymous:1982:NAd

[555] Anonymous. News and announcements. Biometrics, 38(4):1139–1141, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529924.

Anonymous:1982:CDA

[556] Anonymous. Corrections: Discriminant analysis with discrete and continuous variables. Biometrics, 38(4):1143, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529925. See [357].

Anonymous:1982:PPBc

[557] Anonymous. Papers to be published in Biometrics. Biometrics, 38(4): 1144, December 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529926.

Anonymous:1983:FMa

[558] Anonymous. Front matter. Biometrics, 39(1):i–iv, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530801.

Heltshe:1983:ESR

[559] James F. Heltshe and Nancy E. Forrester. Estimating species richness using the jackknife procedure. Biometrics, 39(1):1–11, March 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530802.

Manly:1983:APV

[560] B. F. J. Manly. Analysis of polymorphic variation in different types of habitat. Biometrics, 39(1):13–27, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530803.

Hayes:1983:RDM

[561] R. J. Hayes and S. T. Buckland. Radial-distance models for the line-transect method. Biometrics, 39(1):29–42, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530804.

Arya:1983:CPP

[562] A. S. Arya. Circulant plans for partial diallel crosses. Biometrics, 39(1): 43–52, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530805.

Brien:1983:AVT

[563] C. J. Brien. Analysis of variance tables based on experimental structure. Biometrics, 39(1):53–59, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530806.

Robinson:1983:BAC

[564] P. J. Robinson, A. N. Pettitt, J. Zornig, and L. Bass. A Bayesian analysis of capillary heterogeneity in the intact pig liver. Biometrics, 39(1):61– 69, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530807.

Strenio:1983:EBE

[565] Judith F. Strenio, Herbert I. Weisberg, and Anthony S. Bryk. Empirical Bayes estimation of individual growth-curve parameters and their relationship to covariates. Biometrics, 39(1):71–86, March 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530808.

Nordheim:1983:IEI

[566] E. V. Nordheim, A. Tsiatis, and T. J. Shapas. Incorporating extra information in experimental design for bioassay. Biometrics, 39(1):87–96, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530809.

Hochberg:1983:ICU

[567] Yosef Hochberg, Irit Marom, Ruth Keret, and Shulamith Peleg. On improved calibrations of unknowns in a system of quality-controlled assays. Biometrics, 39(1):97–107, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530810.

Aranda-Ordaz:1983:EPH

[568] Francisco J. Aranda-Ordaz. An extension of the proportional-hazards model for grouped data. Biometrics, 39(1):109–117, March 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530811.

Chen:1983:TMM

[569] Yuan-Yan Chen, Myles Hollander, and Naftali A. Langberg. Tests for monotone mean residual life, using randomly censored data. Biometrics, 39(1):119–127, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530812. See corrections [775, 777].

Hanley:1983:NEM

[570] James A. Hanley and Milton N. Parnes. Nonparametric estimation of a multivariate distribution in the presence of censoring. Biometrics, 39(1): 129–139, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530813.

Tibshirani:1983:FPA

[571] Robert J. Tibshirani and Antonio Ciampi. A family of proportional- and additive-hazards models for survival data. Biometrics, 39(1):141–147, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530814.

Goodman:1983:ADC

[572] Leo A. Goodman. The analysis of dependence in cross-classifications having ordered categories, using log-linear models for frequencies and loglinear models for odds. Biometrics, 39(1):149–160, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530815.

Gilbert:1983:ESM

[573] Ethel S. Gilbert. An evaluation of several methods for assessing the effects of occupational exposure to radiation. Biometrics, 39(1):161–171, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530816.

Berry:1983:AMS

[574] G. Berry. The analysis of mortality by the subject-years method. Biometrics, 39(1):173–184, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530817.

Denman:1983:IEA

[575] Daniel W. Denman, Iii, and James J. Schlesselman. Interval estimation of the attributable risk for multiple exposure levels in case-control studies. Biometrics, 39(1):185–192, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530818.

Symons:1983:CRE

[576] M. J. Symons, R. C. Grimson, and Y. C. Yuan. Clustering of rare events. Biometrics, 39(1):193–205, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530819.

Begg:1983:ADT

[577] Colin B. Begg and Robert A. Greenes. Assessment of diagnostic tests when disease verification is subject to selection bias. Biometrics, 39(1): 207–215, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530820.

McHugh:1983:PSR

[578] Richard McHugh and John Matts. Post-stratification in the randomized clinical trial. Biometrics, 39(1):217–225, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530821.

Whitehead:1983:GSC

[579] John Whitehead and Irene Stratton. Group sequential clinical trials with triangular continuation regions. Biometrics, 39(1):227–236, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530822. See correction [771, 778].

Hearne:1983:TSC

[580] Erwin M. Hearne III, Gary M. Clark, and John P. Hatch. A test for serial correlation in univariate repeated-measures analysis. Biometrics, 39(1): 237–243, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530823.

Lancaster:1983:REP

[581] J. F. Lancaster and Dana Quade. Random effects in paired-comparison experiments using the Bradley–Terry model. Biometrics, 39(1):245–249, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530824.

Cash:1983:MTT

[582] William S. Cash. Modification of the triangle–taste–test model to allow for nonforced guessing. Biometrics, 39(1):251–255, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530825.

Anbar:1983:EDB

[583] Dan Anbar. On estimating the difference between two probabilities, with special reference to clinical trials. Biometrics, 39(1):257–262, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530826.

Thomas:1983:NET

[584] Duncan C. Thomas. Nonparametric estimation and tests of fit for doseresponse relations. Biometrics, 39(1):263–268, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530827.

Wahrendorf:1983:RSC

[585] Jürgen Wahrendorf. Routine statistics and clinical trials. Biometrics, 39(1):269–273, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530828.

Finch:1983:DRN

[586] Peter D. Finch. Do rates need standardizing for comparative purposes? Biometrics, 39(1):275–279, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530829.

Pennington:1983:EEA

[587] Michael Pennington. Efficient estimators of abundance, for fish and plankton surveys. Biometrics, 39(1):281–286, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530830.

Curnow:1983:QIM

[588] R. N. Curnow. Query: Inbreeding in modified full-sib selection. Biometrics, 39(1):287–289, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530831.

Anonymous:1983:OIJ

[589] Anonymous. Obituary: Iréneusz Juvancz 1910–1982. Biometrics, 39(1): 291, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530832.

Buonaccorsi:1983:RMB

[590] John P. Buonaccorsi and Janet D. Elashoff. Repeated-measures bioassay. Biometrics, 39(1):293, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530833.

Le:1983:APM

[591] Chap T. Le, N. E. Breslow, and N. E. Day. The analysis of pair-matched case-control studies. Biometrics, 39(1):294–295, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530834.

Mead:1983:ACN

[592] R. Mead. Analysis of covariance and neighbouring-plot adjustment. Biometrics, 39(1):295, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530835.

Finney:1983:BRC

[593] D. J. Finney. Book review: Contributions to Statistics, by W. G. Cochran. Biometrics, 39(1):297, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530836.

VanBelle:1983:BRH

[594] G. Van Belle. Book review: The Human Side of Statistical Consulting, by J. R. Boen, D. A. Zahn. Biometrics, 39(1):297–298, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530837.

Blight:1983:BRS

[595] B. J. N. Blight. Book review: Statistical Methods in Biology, by N. T. J. Bailey. Biometrics, 39(1):298, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530838.

Smith:1983:BRC

[596] C. A. B. Smith. Book review: Comparative Statistical Inference, by V. Barnett. Biometrics, 39(1):298–299, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530839.

Clarke:1983:BRP

[597] G. M. Clarke. Book review: Probability and Statistics for Engineering and the Sciences, by J. L. Devore. Biometrics, 39(1):299, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530841.

Federer:1983:BRD

[598] W. T. Federer. Book review: Design and Analysis of Experiments, by M. N. Das, N. C. Giri. Biometrics, 39(1):299, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530840.

Rostron:1983:BRP

[599] John Rostron. Book review: A Primer in Data Reduction, by A. S. C. Ehrenberg. Biometrics, 39(1):299–300, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530842.

Chessel:1983:BRQ

[600] D. Chessel. Book review: Qu’est-ce que l’Analyse des Données?, by Jean-Pierre Fenelon. Biometrics, 39(1):300, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530844.

Thompson:1983:BRE

[601] E. A. Thompson. Book review: Essays in Statistical Science, Papers in Honour of P. A. P. Moran, by J. Gani, E. J. Hannan. Biometrics, 39(1): 300, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530843.

Aitken:1983:BRK

[602] C. G. G. Aitken. Book review: Kernel Discriminant Analysis, by D. J. Hand. Biometrics, 39(1):300–301, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530845.

Morineau:1983:BRL

[603] Alain Morineau. Book review: Logiciels d’Analyses Statistiques Multidimensionnelles sur Micro-Ordinateurs, by J. Lefebvre, M. Boitard, J.-F. Rey. Biometrics, 39(1):301, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530846.

Enderlein:1983:BRN

[604] G. Enderlein. Book review: Numerische und Nicht-Numerische Klassifikation zwischen Theorie und Praxis, by P. Ihm, I. Dahlberg. Biometrics, 39(1):301–302, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530847.

ODonovan:1983:BRS

[605] Thomas M. F. O’Donovan. Book review: Simulation and the Monte Carlo Method, by R. Y. Rubinstein. Biometrics, 39(1):302, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530848.

Cooke:1983:BRS

[606] D. Cooke. Book review: Statistical Distributions in Scientific Work. Vol. 4: Models, Structures and Characterisations. Vol. 5: Inferential Problems and Properties. Vol. 6 Applications in Physical, Social and Life Sciences, by C. Taillie, G. P. Patil, B. A. Baldessari. Biometrics, 39(1): 302–303, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530849.

Dunnett:1983:BRI

[607] C. W. Dunnett. Book review: The Interpretation of Quantal Responses in Biology, by P. S. Hewlett, R. L. Plackett. Biometrics, 39(1):303, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530850.

Griffiths:1983:BRC

[608] D. Griffiths. Book review: Compartmental Analysis of Ecosystem Models, by J. H. Matis, B. C. Patten, G. C. White. Biometrics, 39(1):303–304, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530851.

Cormack:1983:BRQ

[609] R. M. Cormack. Book review: Quantitative Population Dynamics, by D. G. Chapman, V. F. Gallucci. Biometrics, 39(1):304, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530852.

Freeman:1983:BRM

[610] G. H. Freeman. Book review: The Mathematics of Photosynthesis and Productivity, by D. A. Charles-Edwards. Biometrics, 39(1):304, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530853.

Brown:1983:BRE

[611] R. A. Brown. Book review: Electric Fields of the Brain, by P. L. Nunez. Biometrics, 39(1):305, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530855.

Buckland:1983:BRP

[612] Stephen Buckland. Book review: Predatory Birds, Pesticides and Pollution, by A. S. Cooke, A. A. Bell, M. B. Haas. Biometrics, 39(1):305, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530854.

Hill:1983:BRB

[613] W. G. Hill. Book review: Biometrical Genetics, by K. Mather, J. L. Jinks. Biometrics, 39(1):305–306, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530856.

Anonymous:1983:BS

[614] Anonymous. The Biometric Society. Biometrics, 39(1):307, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530857.

Anonymous:1983:NAa

[615] Anonymous. News and announcements. Biometrics, 39(1):309, March 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530858.

Anonymous:1983:FMb

[616] Anonymous. Front matter. Biometrics, 39(2):i–iv, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531003.

Holford:1983:EAP

[617] Theodore R. Holford. The estimation of age, period and cohort effects for vital rates. Biometrics, 39(2):311–324, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531004.

Howe:1983:CIE

[618] G. R. Howe. Confidence interval estimation for the ratio of simple and standardized rates in cohort studies. Biometrics, 39(2):325–331, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531005.

Chinchilli:1983:ESS

[619] Vernon M. Chinchilli. Estimates of sensitivity and specificity in a multistage screen for medical diagnosis. Biometrics, 39(2):333–340, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531006.

Lee:1983:CTS

[620] Kerry L. Lee, Frank E. Harrell, Jr., H. Dennis Tolley, and Robert A. Rosati. A comparison of test statistics for assessing the effects of concomitant variables in survival analysis. Biometrics, 39(2):341–350, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531007.

Taulbee:1983:SSD

[621] John D. Taulbee and Michael J. Symons. Sample size and duration for cohort studies of survival time with covariables. Biometrics, 39(2):351– 360, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531008.

Cook:1983:MRG

[622] D. G. Cook and S. J. Pocock. Multiple regression in geographical mortality studies, with allowance for spatially correlated errors. Biometrics, 39(2):361–371, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531009.

Modjeska:1983:SCA

[623] Janet Shy Modjeska and J. O. Rawlings. Spatial correlation analysis of uniformity data. Biometrics, 39(2):373–384, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531010.

Garsd:1983:SSE

[624] Armando Garsd, Gary E. Ford, George O. Waring Iii, and Leon S. Rosenblatt. Sample size for estimating the quantiles of endothelial cell-area distribution. Biometrics, 39(2):385–394, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531011.

Stone:1983:GSM

[625] Mervyn Stone. A general statistical model for clone-tissue studies, using X-chromosome inactivation data. Biometrics, 39(2):395–409, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531012.

Sampson:1983:SAA

[626] Paul D. Sampson. Statistical analysis of arch shape with conic sections. Biometrics, 39(2):411–423, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531013.

Campbell:1983:CVA

[627] N. A. Campbell and J. A. Tomenson. Canonical variate analysis for several sets of data. Biometrics, 39(2):425–435, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531014.

Brownie:1983:ETS

[628] C. Brownie and D. S. Robson. Estimation of time-specific survival rates from tag-resighting samples: a generalization of the Jolly–Seber model. Biometrics, 39(2):437–453, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531015.

Swindel:1983:CSN

[629] Benee F. Swindel. Choice of size and number of quadrats to estimate density from frequency in Poisson and binomially dispersed populations. Biometrics, 39(2):455–464, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531016.

Jacquard:1983:HOW

[630] A. Jacquard. Heritability: One word, three concepts. Biometrics, 39(2): 465–477, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531017.

Roux:1983:RPM

[631] C. Z. Roux. Replacing parents by meritorious offspring: Maximizing within-family selection gains by the use of order statistics. Biometrics, 39(2):479–483, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531018.

Emigh:1983:NOC

[632] Ted H. Emigh. On the number of observed classes from a multinomial distribution. Biometrics, 39(2):485–491, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531019.

Haber:1983:SSE

[633] Michael Haber. Sample-sizes for the exact test of ‘no interaction’ in 2×2×2 tables. Biometrics, 39(2):493–498, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531020.

Schoenfeld:1983:SSF

[634] David A. Schoenfeld. Sample-size formula for the proportional-hazards regression model. Biometrics, 39(2):499–503, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531021.

Agresti:1983:TMH

[635] Alan Agresti. Testing marginal homogeneity for ordinal categorical variables. Biometrics, 39(2):505–510, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531022.

Pepper:1983:CPM

[636] W. D. Pepper. Choosing plant-mating design allocations to estimate genetic variance components in the absence of prior knowledge of the relative magnitudes. Biometrics, 39(2):511–521, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531023.

Guilbaud:1983:LSD

[637] O. Guilbaud and Walter Hauck. On the large-sample distribution of the Mantel–Haenszel odds-ratio estimator. Biometrics, 39(2):523–525, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531024.

Greenberg:1983:JOI

[638] Bernard G. Greenberg. Joseph Oscar Irwin 1898–1982. Biometrics, 39(2): 527–528, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531025.

Mantel:1983:MGS

[639] Nathan Mantel and R. Morton. Morton’s generalization of Spearman estimates. Biometrics, 39(2):529–530, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531026.

Allison:1983:MCM

[640] Paul D. Allison. A model for cancer mortality rates. Biometrics, 39(2): 530–532, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531027.

Klotz:1983:PGF

[641] Jerome Klotz. Priority for a generalized Friedman statistic. Biometrics, 39(2):532, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531028.

Lawless:1983:BRE

[642] J. F. Lawless. Book review: Encyclopedia of Statistical Sciences, Vol. 1. A-Circular Probable Error, by S. Kotz, N. L. Johnson, C. B. Read. Biometrics, 39(2):533, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531029.

Renshaw:1983:BRS

[643] E. Renshaw. Book review: Statistics and Probability (Essays in Honour of C. R. Rao), by G. Kallianpur, P. R. Krishnaiah, J. K. Ghosh. Biometrics, 39(2):533–534, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531030.

Marriott:1983:BRA

[644] F. H. C. Marriott. Book review: Applied Multivariate Statistical Analysis, by R. A. Johnson, D. W. Wichern. Biometrics, 39(2):534, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531032.

Moran:1983:BRA

[645] M. A. Moran. Book review: Applied Multivariate Analysis, by S. James Press. Biometrics, 39(2):534, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531031.

Newell:1983:BRS

[646] D. J. Newell. Book review: Statistics in Practice: Statistics in Question and Statistics and Ethics in Medical Research, by D. G. Altman, S. M. Gore. Biometrics, 39(2):534–535, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531033.

Brascamp:1983:BRP

[647] E. W. Brascamp. Book review: Population Genetics in Animal Breeding, by F. Prichner. Biometrics, 39(2):535, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531035.

McPherson:1983:BRR

[648] Klim McPherson. Book review: The Randomized Clinical Trial and Therapeutic Decisions, by N. Tygstrup, J. M. Lachin, E. Juhl. Biometrics, 39(2):535, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531034.

Beddington:1983:BRMa

[649] J. R. Beddington. Book review: The Mathematical Theory of the Dynamics of Biological Populations, II, by R. W. Hiorns, D. Cooke. Biometrics, 39(2):536, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531037.

Stene:1983:BRO

[650] Jon Stene. Book review: Outline of Genetic Epidemiology, by N. E. Morton. Biometrics, 39(2):536, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531036.

Diggle:1983:BRI

[651] P. J. Diggle. Book review: Image Analysis and Mathematical Morphology, by J. Serra. Biometrics, 39(2):536–537, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531038.

Downs:1983:BRG

[652] Thomas Downs. Book review: The Geometry of Biological Time, by A. T. Winfree. Biometrics, 39(2):537, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531039.

Glasbey:1983:BRP

[653] C. A. Glasbey. Book review: Plant Growth Curves-the Functional Approach to Plant Growth Analysis, by R. Hunt. Biometrics, 39(2):537, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531040.

Harpley:1983:BRP

[654] F. W. Harpley. Book review: Pharmacokinetics during Drug Development: Data Analysis and Evaluation Techniques, by G. Bozler, J. M. van Rossum. Biometrics, 39(2):537–538, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531041.

Cormack:1983:BRK

[655] R. M. Cormack. Book review: Kinetic Data Analysis: Design and Analysis of Enzyme and Pharmacokinetic Experiments, by L. Endrenyi. Biometrics, 39(2):538, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531042.

Franklin:1983:BRB

[656] M. F. Franklin. Book review: Basic Statistical Computing, by D. Cooke, A. H. Craven, G. M. Clarke. Biometrics, 39(2):538–539, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531043.

Anonymous:1983:NAb

[657] Anonymous. News and announcements. Biometrics, 39(2):541, June 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531044.

Anonymous:1983:FMc

[658] Anonymous. Front matter. Biometrics, 39(3):i–iv, September 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531081.

Maller:1983:DMF

[659] R. A. Maller, E. S. de Boer, L. M. Joll, D. A. Anderson, and J. P. Hinde. Determination of the maximum foregut volume of western rock lobsters (Panulirus cygnus) from field data. Biometrics, 39(3):543–551, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531082.

Lin:1983:MAD

[660] C. S. Lin and G. Poushinsky. A modified augmented design for an early stage of plant selection involving a large number of test lines without replication. Biometrics, 39(3):553–561, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531083.

Samaniego:1983:ESP

[661] Francisco J. Samaniego and Pere Arús. On estimating the sib proportion in seed-purity determinations. Biometrics, 39(3):563–572, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531084.

Finney:1983:GES

[662] D. J. Finney. Genetic equilibria of species with self-incompatible polymorphisms. Biometrics, 39(3):573–585, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531085.

Dahm:1983:GLS

[663] P. Frederick Dahm, Bryan E. Melton, and Wayne A. Fuller. Generalized least squares estimation of a genotypic covariance matrix. Biometrics, 39(3):587–597, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531086.

Donner:1983:VCE

[664] Allan Donner and John J. Koval. Variance-component estimation from human sibship data. Biometrics, 39(3):599–605, September 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531087.

Carbonell:1983:SLM

[665] E. A. Carbonell, W. E. Nyquist, and A. E. Bell. Sex-linked and maternal effects in the Eberhart–Gardner general genetics model. Biometrics, 39(3):607–619, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531088.

Katz:1983:EDE

[666] Darryl Katz and David Z. D’Argenio. Experimental design for estimating integrals by numerical quadrature, with applications to pharmacokinetic studies. Biometrics, 39(3):621–628, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531089.

Allen:1983:PEN

[667] David M. Allen. Parameter estimation for nonlinear models with emphasis on compartmental models. Biometrics, 39(3):629–637, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531090.

Martin:1983:RES

[668] George L. Martin. The relative efficiency of some forest growth estimators. Biometrics, 39(3):639–649, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531091.

James:1983:ANA

[669] Ian R. James. Analysis of nonagreements among multiple raters. Biometrics, 39(3):651–657, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531092.

Elton:1983:CEM

[670] R. A. Elton and S. W. Duffy. Correcting for the effect of misclassification bias in a case-control study using data from two different questionnaires. Biometrics, 39(3):659–663, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531093.

Frome:1983:ARU

[671] E. L. Frome. The analysis of rates using Poisson regression models. Biometrics, 39(3):665–674, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531094.

Kellerer:1983:SSP

[672] A. M. Kellerer and D. Chmelevsky. Small-sample properties of censoreddata rank tests. Biometrics, 39(3):675–682, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531095.

Hand:1983:CTM

[673] D. J. Hand. A comparison of two methods of discriminant analysis applied to binary data. Biometrics, 39(3):683–694, September 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531096.

Dambrosia:1983:ESS

[674] James M. Dambrosia and Samuel W. Greenhouse. Early stopping for sequential restricted tests of binomial distributions. Biometrics, 39(3):695– 710, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531097. See

correction [779, 780].

Anbar:1983:REZ

[675] Dan Anbar. The relative efficiency of Zelen’s prerandomization design for clinical trials. Biometrics, 39(3):711–718, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531098.

Davies:1983:SST

[676] Paul Davies. Some statistical tests for constrained occupancy problems. Biometrics, 39(3):719–725, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531099.

Woolson:1983:RAC

[677] Robert F. Woolson and Peter A. Lachenbruch. Rank analysis of covariance with right-censored data. Biometrics, 39(3):727–733, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531100.

Quesenberry:1983:SMS

[678] C. P. Quesenberry, F. G. Giesbrecht, and J. C. Burns. Some methods for studying the validity of normal model assumptions for multiple samples. Biometrics, 39(3):735–739, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531101.

Whaley:1983:ETI

[679] Fredrick S. Whaley. The equivalence of three independently derived permutation procedures for testing the homogeneity of multidimensional samples. Biometrics, 39(3):741–745, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531102.

Yang:1983:OWA

[680] Mark C. K. Yang and Randy L. Carter. One-way analysis of variance with time-series data. Biometrics, 39(3):747–751, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531103.

Digby:1983:ATC

[681] P. G. N. Digby. Approximating the tetrachoric correlation coefficient. Biometrics, 39(3):753–757, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531104.

Stablein:1983:CRC

[682] Donald M. Stablein, Walter H. Carter, Jr., and Galen L. Wampler. Confidence regions for constrained optima in response-surface experiments. Biometrics, 39(3):759–763, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531105.

Metz:1983:DEB

[683] J. A. J. Metz, M. Wedel, and A. F. Angulo. Discovering an epidemic before it has reached a certain level of prevalence. Biometrics, 39(3):765– 770, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531106.

Donner:1983:ICC

[684] Allan Donner and Shelley Bull. Inferences concerning a common intraclass correlation coefficient. Biometrics, 39(3):771–775, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531107.

Mantel:1983:EUB

[685] Nathan Mantel. Extended use of binomial sample-size tables. Biometrics, 39(3):777–779, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531108.

Gart:1983:RBS

[686] John J. Gart, Robert E. Tarone, N. E. Day, and D. P. Byar. The relation between score tests and approximate UMPU tests in exponential models common in biometry. Biometrics, 39(3):781–786, September 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531109.

OBrien:1983:AAV

[687] P. C. O’Brien. The appropriateness of analysis of variance and multiplecomparison procedures. Biometrics, 39(3):787–788, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531110.

Dawkins:1983:MCM

[688] H. C. Dawkins. Multiple comparisons misused: Why so frequently in response-curve studies? Biometrics, 39(3):789–790, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531111.

Duncan:1983:ATM

[689] D. B. Duncan and L. J. Brant. Adaptive t tests for multiple comparisons. Biometrics, 39(3):790–794, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531112.

Rogosa:1983:CTI

[690] David R. Rogosa, John B. Willett, Mary A. Foulkes, C. E. Davis, and C. Alex McMahan. Comparing two indices of tracking. Biometrics, 39(3):795–796, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531113.

Mantel:1983:SDC

[691] Nathan Mantel, David M. Allen, and David C. Jordan. Statistical devices cannot substitute for model propriety. Biometrics, 39(3):797–798, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531114.

Poloniecki:1983:ITW

[692] J. D. Poloniecki and A. C. Pearce. Interaction in the two-way crossover trial. Biometrics, 39(3):798, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531115.

Nelder:1983:PPC

[693] John A. Nelder. Presentation of papers at conferences. Biometrics, 39(3): 799, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531116.

Anonymous:1983:Aa

[694] Anonymous. Abstracts. Biometrics, 39(3):801–813, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531117.

Darzentas:1983:BRM

[695] J. Darzentas. Book review: Modern Data Analysis, by R. L. Launer, A. F. Siegel. Biometrics, 39(3):815, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531119.

Radford:1983:BRA

[696] P. J. Radford. Book review: Applications, Basics, and Computing of Exploratory Data Analysis, by P. F. Velleman, D. C. Hoaglin. Biometrics, 39(3):815, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531118.

Anderson:1983:BRI

[697] A. J. B. Anderson. Book review: An Introduction to Genstat, by N. Alvey, N. Galwey, P. Lane. Biometrics, 39(3):816, September 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531121.

Appleton:1983:BRM

[698] D. R. Appleton. Book review: Multivariate Data Analysis in Industrial Practice, by P. J. Lewi. Biometrics, 39(3):816, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531120.

Freeman:1983:BRS

[699] P. R. Freeman. Book review: Statistics in Research and Development, by R. Caulcutt. Biometrics, 39(3):816, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531122.

Burnham:1983:BRJ

[700] K. P. Burnham. Book review: The Jackknife, the Bootstrap and Other Resampling Plans, by B. Efron. Biometrics, 39(3):816–817, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531123.

Copas:1983:BRR

[701] J. B. Copas. Book review: Recent Advances in Regression Methods, by H. D. Vinod, A. Ullah. Biometrics, 39(3):817–818, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531124.

Atkinson:1983:BRR

[702] A. C. Atkinson. Book review: Residuals and Influence in Regression, by R. D. Cook, S. Weisberg. Biometrics, 39(3):818, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531125.

Felsenstein:1983:BRE

[703] J. Felsenstein. Book review: Evolution in Age-Structured Populations, by B. Charlesworth. Biometrics, 39(3):818–819, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531126.

Rodolphe:1983:BRM

[704] F. Rodolphe. Book review: Modeles Dynamiques Deterministes en Biologie, by J. D. Lebreton, C. Miller. Biometrics, 39(3):819, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531127.

Beddington:1983:BRMb

[705] J. R. Beddington. Book review: Modelling Fluctuating Populations, by R. M. Nisbet, W. S. C. Gurney. Biometrics, 39(3):819–820, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531128.

Francis:1983:BRS

[706] B. Francis. Book review: Statistical Models and Methods for Lifetime Data, by J. F. Lawless. Biometrics, 39(3):820, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531129.

Farley:1983:BRA

[707] T. M. M. Farley. Book review: Applied Life Data Analysis, by Wayne Nelson. Biometrics, 39(3):820–821, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531130.

Rossiter:1983:BRC

[708] C. E. Rossiter. Book review: Case-Control Studies. Design, Conduct, Analysis, by J. J. Schlesselman. Biometrics, 39(3):821–822, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531131.

Immich:1983:BRT

[709] H. Immich. Book review: Therapiestudien, by N. Victor. Biometrics, 39(3):822, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531132.

Pocock:1983:BRE

[710] S. J. Pocock. Book review: Epidemiologic Research: Principles and Quantitative Methods, by D. G. Kleinbaum, L. L. Kupper, H. Morgenstern. Biometrics, 39(3):822–823, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531133.

Prescott:1983:BRB

[711] R. J. Prescott. Book review: Biostatistics in Medicine, by Edmond A. Murphy. Biometrics, 39(3):823–824, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531134.

Anonymous:1983:NAc

[712] Anonymous. News and announcements. Biometrics, 39(3):825, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531135.

Anonymous:1983:FMd

[713] Anonymous. Front matter. Biometrics, 39(4):i–iv, December 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531317.

Anonymous:1983:VI

[714] Anonymous. Volume information. Biometrics, 39(4):xi–xxx, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531316.

Zelen:1983:BIP

[715] Marvin Zelen. A biometrics invited paper with discussion. Biostatistical science as a discipline: a look into the future. Biometrics, 39(4):827–837, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531318.

Margolin:1983:WGC

[716] Barry H. Margolin. William Gemmell Cochran: The influence of Rothamsted and designed experiments on his research. Biometrics, 39(4):839–848, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531319.

Denis:1983:EDM

[717] Jean-Baptiste Denis. Extension du modele additif d’Analyse de variance par modelisation multiplicative des variances. Biometrics, 39(4):849–856, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531320.

Prunty:1983:CFS

[718] Lyle Prunty. Curve fitting with smooth functions that are piecewiselinear in the limit. Biometrics, 39(4):857–866, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531321.

Smith:1983:MRT

[719] A. F. M. Smith and M. West. Monitoring renal transplants: an application of the multiprocess Kalman filter. Biometrics, 39(4):867–878, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531322.

Morgan:1983:OQA

[720] Byron J. T. Morgan. Observations on quantit analysis. Biometrics, 39(4):879–886, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531323.

Xekalaki:1983:UGW

[721] Evdokia Xekalaki. The univariate generalized Waring distribution in relation to accident theory: Proneness, spells or contagion? Biometrics, 39(4):887–895, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531324.

Portier:1983:LDR

[722] C. Portier and D. Hoel. Low-dose-rate extrapolation using the multistage model. Biometrics, 39(4):897–906, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531325.

Kalbfleisch:1983:EMM

[723] J. D. Kalbfleisch, J. F. Lawless, and W. M. Vollmer. Estimation in Markov models from aggregate data. Biometrics, 39(4):907–919, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531326.

Seigel:1983:FRM

[724] Daniel Seigel and Roy C. Milton. Further results on a multiple-testing procedure for clinical trials. Biometrics, 39(4):921–928, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531327.

Risch:1983:EMR

[725] N. Risch. Estimating morbidity risks with variable age of onset: Review of methods and a maximum likelihood approach. Biometrics, 39(4):929– 939, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531328.

Brown:1983:SCR

[726] Charles C. Brown. The statistical comparison of relative survival rates. Biometrics, 39(4):941–948, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531329.

Risch:1983:SAM

[727] N. Risch and K. Lange. Statistical analysis of multilocus recombination. Biometrics, 39(4):949–963, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531330.

Kaiser:1983:UEL

[728] Lee Kaiser. Unbiased estimation in line-intercept sampling. Biometrics, 39(4):965–976, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531331.

Lo:1983:MLE

[729] Hing-P. P. Lo and Jagannath K. Wani. Maximum likelihood estimation of the parameters of the invariant abundance distributions. Biometrics, 39(4):977–986, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531332.

Reed:1983:CEE

[730] William J. Reed. Confidence estimation of ecological aggregation indices based on counts-a robust procedure. Biometrics, 39(4):987–998, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531333.

Tong:1983:SDP

[731] Y. L. Tong. Some distribution properties of the sample species-diversity indices and their applications. Biometrics, 39(4):999–1008, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531334.

Nedelman:1983:NBM

[732] Jerry Nedelman. A negative binomial model for sampling mosquitoes in a malaria survey. Biometrics, 39(4):1009–1020, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531335.

Dupont:1983:SCE

[733] William D. Dupont. A stochastic catch-effort method for estimating animal abundance. Biometrics, 39(4):1021–1033, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531336.

Pollock:1983:REP

[734] Kenneth H. Pollock and Mark C. Otto. Robust estimation of population size in closed animal populations from capture–recapture experiments. Biometrics, 39(4):1035–1049, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531337.

Prentice:1983:MLL

[735] I. C. Prentice and R. W. Parsons. Maximum likelihood linear calibration of pollen spectra in terms of forest composition. Biometrics, 39(4):1051– 1057, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531338.

Garcia:1983:SDE

[736] Oscar García. A stochastic differential equation model for the height growth of forest stands. Biometrics, 39(4):1059–1072, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531339.

Heltshe:1983:EDU

[737] J. F. Heltshe and N. E. Forrester. Estimating diversity using quadrat sampling. Biometrics, 39(4):1073–1076, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531340.

Miller:1983:WPK

[738] Rupert G. Miller, Jr. What price Kaplan–Meier? Biometrics, 39 (4):1077–1081, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531341.

Titterington:1983:CBR

[739] D. M. Titterington. On constrained balance randomization for clinical trials. Biometrics, 39(4):1083–1086, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531342.

Laska:1983:OCD

[740] Eugene Laska, Morris Meisner, and H. B. Kushner. Optimal crossover designs in the presence of carryover effects. Biometrics, 39(4):1087–1091, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531343.

Koopman:1983:DNP

[741] P. A. R. Koopman. On the determination of the number of particles in a liquid by counting and differentiation. Biometrics, 39(4):1093–1096, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531344.

Roberts:1983:AMC

[742] E. A. Roberts and J. M. Raison. An analysis of the moisture content of soil cores in a designed experiment. Biometrics, 39(4):1097–1105, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531345.

OFallon:1983:O

[743] Judith Rich O’Fallon, R. M. Cormack, and J. F. Bithell. Obituaries. Biometrics, 39(4):1107–1111, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531346.

Goldstein:1983:APC

[744] Harvey Goldstein and Ian James. Age, period and cohort effects. Biometrics, 39(4):1113, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531347.

Hauck:1983:RAC

[745] Walter W. Hauck, A. S. Whittemore, J. W. McLarty, N. Fortson, and K. Anderson. Regression analysis of cytopathological data. Biometrics, 39(4):1114, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531348.

Fazal:1983:OAC

[746] Sadrudin S. Fazal. Optimal allocation for the comparison of proportions. Biometrics, 39(4):1115, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531349.

Preece:1983:CCD

[747] D. A. Preece. Criss-cross designs. Biometrics, 39(4):1115–1116, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531350.

Anonymous:1983:Ab

[748] Anonymous. Abstracts. Biometrics, 39(4):1117–1120, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531351.

Barnard:1983:BRN

[749] G. A. Barnard. Book review: Neyman from Life, by C. Reid. Biometrics, 39(4):1121, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531352.

Green:1983:BRMa

[750] R. H. Green. Book review: Modelling, by J. N. R. Jeffers. Biometrics, 39(4):1121–1122, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531353.

LeBoulenge:1983:BRI

[751] E. Le Boulenge and J. J. Claustriaux. Book review: Introduction aux Modèles Mathématiques en Biologie, by E. Jolivet. Biometrics, 39(4): 1122, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531354.

Jolly:1983:BRC

[752] G. M. Jolly. Book review: Capture–Recapture and Removal Methods for Sampling Closed Populations, by G. C. White, D. R. Anderson, K. P. Burnham, D. L. Otis. Biometrics, 39(4):1122–1123, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531355.

Johnson:1983:BRA

[753] Douglas H. Johnson and James W. Grier. Book review: Animal Population Dynamics, by R. Moss, A. Watson, J. Ollason. Biometrics, 39(4): 1123, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531356.

Morgan:1983:BRE

[754] Byron J. T. Morgan and Philip M. North. Book review: Estimating Numbers of Terrestrial Birds, by C. J. Ralph, J. M. Scott. Biometrics, 39(4):1123–1124, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531357.

Balas:1983:BRM

[755] E. A. Balas. Book review: Medical Biometrics, by I. Juvancz, A. Paksy. Biometrics, 39(4):1125, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531359.

Smith:1983:BRF

[756] G. L. Smith. Book review: Food Research and Data Analysis, by H. Martens, H. Russwurm, Jr. Biometrics, 39(4):1125, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531358.

Campbell:1983:BRS

[757] M. J. Campbell. Book review: Statistics in Medical Research, by V. Miké, K. E. Stanley. Biometrics, 39(4):1125–1126, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531360.

Gower:1983:BRC

[758] J. C. Gower. Book review: Classification: Methods for the Exploratory Analysis of Multivariate Data, by A. D. Gordon. Biometrics, 39(4):1126, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531361.

Healy:1983:BRU

[759] M. J. R. Healy. Book review: Understanding Robust and Exploratory Data Analysis, by D. C. Hoaglin, F. Mosteller, J. W. Tukey. Biometrics, 39(4):1126–1127, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531362.

Renshaw:1983:BRA

[760] Eric Renshaw. Book review: Applications of Queueing Theory, by G. F. Newell. Biometrics, 39(4):1127, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531363.

Gore:1983:BRB

[761] Sheila M. Gore. Book review: Biomedical Statistics in Computing, by M. H. Regier, R. N. Mohapatra, S. N. Mohapatra. Biometrics, 39(4):1127– 1128, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531364.

Green:1983:BRMb

[762] P. J. Green. Book review: Matrix Algebra Useful for Statistics, by S. R. Searle. Biometrics, 39(4):1128, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531366.

Prescott:1983:BRD

[763] R. J. Prescott. Book review: The Design and Analysis of Sequential Clinical Trials, by J. Whitehead. Biometrics, 39(4):1128, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531365.

Anonymous:1983:BRA

[764] Anonymous. Book review: An Assessment of Research-Doctorate Programs in the United States, by L. V. Jones, G. Lindzey, P. E. Coggeshall. Biometrics, 39(4):1129, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531370.

Anonymous:1983:BRH

[765] Anonymous. Book review: How to Tell the Liars from the Statisticians, by R. Hooke. Biometrics, 39(4):1129, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531369.

Anonymous:1983:BRQ

[766] Anonymous. Book review: Quotes, Damned Quotes, and …, by John Bibby. Biometrics, 39(4):1129, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531368.

Currie:1983:BRT

[767] I. Currie. Book review: Teaching Statistics in Schools Throughout the World, by V. Barnett. Biometrics, 39(4):1129, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531367.

Anonymous:1983:BSF

[768] Anonymous. The Biometric Society financial statements and auditor’s report. January 31, 1983. Biometrics, 39(4):1131–1133, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531371.

Anonymous:1983:NAd

[769] Anonymous. News and announcements. Biometrics, 39(4):1135–1136, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531372.

Anonymous:1983:CBC

[770] Anonymous. Correction: A Bayesian change-point problem with an application to the prediction and detection of ovulation in women. Biometrics, 39(4):1137, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531375. See [264].

Anonymous:1983:CGS

[771] Anonymous. Correction: Group sequential clinical trials with triangular continuation regions. Biometrics, 39(4):1137, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531378. See [579].

Anonymous:1983:CMB

[772] Anonymous. Corrections: Multivariate bioassay. Biometrics, 39(4):1137, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531374. See [37].

Anonymous:1983:CNC

[773] Anonymous. Correction: Nonparametric confidence intervals for the median in the presence of right censoring. Biometrics, 39(4):1137, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531376. See [340].

Anonymous:1983:CPC

[774] Anonymous. Corrections: Power computations for designing comparative Poisson trials. Biometrics, 39(4):1137, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531373. See [3].

Anonymous:1983:CTM

[775] Anonymous. Correction: Tests for monotone mean residual life, using randomly censored data. Biometrics, 39(4):1137, December 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531377. See [569].

Carter:1983:CBC

[776] R. L. Carter and B. J. N. Blight. Correction to: “A Bayesian change-point problem with an application to the prediction and detection of ovulation in women” [Biometrics 37 (1981), no. 4, 743–751; MR0675315 (83m:62171)]. Biometrics, 39(4):1137, December 1983. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [264].

Chen:1983:CTM

[777] Yuan-Yan Chen, Myles Hollander, and Naftali A. Langberg. Corrections of: “Tests for monotone mean residual life, using randomly censored data”. Biometrics, 39(4):1137, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [569].

Whitehead:1983:CGS

[778] John Whitehead and Irene Stratton. Correction to: “Group sequential clinical trials with triangular continuation regions”. Biometrics, 39(4): 1137, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [579].

Anonymous:1983:CES

[779] Anonymous. Correction: Early stopping for sequential restricted tests of binomial distributions. Biometrics, 39(4):1138, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531379. See [674].

Dambrosia:1983:CES

[780] J. M. Dambrosia and S. W. Greenhouse. Correction to: “Early stopping for sequential restricted tests of binomial distributions”. Biometrics, 39 (4):1138, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [674].

Rosenberger:1983:RCP

[781] J. L. Rosenberger. Review: Configural Polysampling: A Route to Practical Robustness, by S. Morgenthaler and J. W. Tukey. Biometrics, 39(4): 1284, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532280.

Anonymous:1984:FMa

[782] Anonymous. Front matter. Biometrics, 40(1):i–iv, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530738.

Day:1984:SMS

[783] Nicholas E. Day and Stephen D. Walter. Simplified models of screening for chronic disease: Estimation procedures from mass screening programmes. Biometrics, 40(1):1–13, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530739.

Tango:1984:DDC

[784] Toshiro Tango. The detection of disease clustering in time. Biometrics, 40(1):15–26, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530740.

Sawyer:1984:CCP

[785] Charles Sawyer, Richard Peto, Leslie Bernstein, and Malcolm C. Pike. Calculation of carcinogenic potency from long-term animal carcinogenesis experiments. Biometrics, 40(1):27–40, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530741.

Turnbull:1984:NED

[786] Bruce W. Turnbull and Toby J. Mitchell. Nonparametric estimation of the distribution of time to onset for specific diseases in survival/sacrifice experiments. Biometrics, 40(1):41–50, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530742.

Buckley:1984:AMM

[787] J. D. Buckley. Additive and multiplicative models for relative survival rates. Biometrics, 40(1):51–62, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530743.

Lubin:1984:BSC

[788] Jay H. Lubin and Mitchell H. Gail. Biased selection of controls for casecontrol analyses of cohort studies. Biometrics, 40(1):63–75, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530744.

Deaton:1984:MSC

[789] Michael L. Deaton, James J. Higgins, and Joseph E. Smith. A model for the synthesis of chemical compounds in red blood cells, with an application to diabetes. Biometrics, 40(1):77–84, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530745.

Wetter:1984:FDC

[790] Th. Wetter and R. Repges. Flow distribution in closed networks: Minimal sets of measuring sites. Biometrics, 40(1):85–94, March 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530746.

Havranek:1984:PMS

[791] Tomáš Havránek. A procedure for model search in multidimensional contingency tables. Biometrics, 40(1):95–100, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530747.

deJong:1984:FOM

[792] Piet de Jong and Malcolm Greig. First-order Markov chains with a zero diagonal transition matrix. Biometrics, 40(1):101–107, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530748.

Willson:1984:MEC

[793] Linda J. Willson, J. Leroy Folks, and J. H. Young. Multistage estimation compared with fixed-sample-size estimation of the negative binomial parameter k. Biometrics, 40(1):109–117, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530749.

Smith:1984:NES

[794] Eric P. Smith and Gerald van Belle. Nonparametric estimation of species richness. Biometrics, 40(1):119–129, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530750.

Higgins:1984:ACA

[795] Stewart S. Higgins, Robert B. Bendel, and Richard N. Mack. Assessing competition among skewed distributions of plant biomass: an application of the jackknife. Biometrics, 40(1):131–137, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530751.

Griffiths:1984:FGA

[796] David Griffiths and Ron Sandland. Fitting generalized allometric models to multivariate growth data. Biometrics, 40(1):139–150, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530752.

Chinchilli:1984:LRT

[797] Vernon M. Chinchilli and Walter H. Carter, Jr. A likelihood ratio test for a patterned covariance matrix in a multivariate growth-curve model. Biometrics, 40(1):151–156, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530753.

Cockerham:1984:CRS

[798] C. Clark Cockerham and B. S. Weir. Covariances of relatives stemming from a population undergoing mixed self and random mating. Biometrics, 40(1):157–164, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530754.

Crouchley:1984:MIL

[799] R. Crouchley and A. R. Pickles. Methods for the identification of Lexian, Poisson and Markovian variations in the secondary sex ratio. Biometrics, 40(1):165–175, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530755. See corrections [1002].

Khamis:1984:LLM

[800] H. J. Khamis and K. Hinkelmann. Log-linear-model analysis of the association between disease and genotype. Biometrics, 40(1):177–188, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530756.

Haber:1984:LLM

[801] Michael Haber. Log-linear models for linked loci. Biometrics, 40(1):189– 198, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530757.

Clayton:1984:RPM

[802] Gareth Clayton. On the random-pairs method of density estimation. Biometrics, 40(1):199–202, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530758.

Jarrett:1984:CRP

[803] R. G. Jarrett and B. J. T. Morgan. A critical reappraisal of the patchgap model. Biometrics, 40(1):203–207, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530759.

Korn:1984:KTB

[804] Edward L. Korn. Kendall’s tau with a blocking variable. Biometrics, 40(1):209–214, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530760.

Hall:1984:DAS

[805] John W. Hall. Designs for the addition of six further treatments to latin squares of order six. Biometrics, 40(1):215–218, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530761.

Ebbutt:1984:TPC

[806] A. F. Ebbutt. Three-period crossover designs for two treatments. Biometrics, 40(1):219–224, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530762.

Rocke:1984:TB

[807] David M. Rocke. On testing for bioequivalence. Biometrics, 40(1):225– 230, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530763.

Pearce:1984:EI

[808] S. C. Pearce and R. N. Edmondson. Experimenting with intercrops. Biometrics, 40(1):231–237, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530764.

Lee:1984:QSA

[809] Young Jack Lee. Quick and simple approximation of sample sizes for comparing two independent binomial distributions: Different-sample-size case. Biometrics, 40(1):239–241, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530765.

Falkenhagen:1984:QIP

[810] E. R. Falkenhagen, W. G. Warren, and G. H. Freeman. Query: The interpretation of provenance tests. Biometrics, 40(1):243–247, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530766.

Bowman:1984:EMS

[811] Kimiko O. Bowman. Extended moment series and the parameters of the negative binomial distribution. Biometrics, 40(1):249–252, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530767.

daSilva:1984:OEC

[812] J. Gilberto Corrêa da Silva. Obituary: Edilberto Da Costa Amaral 1911– 1983. Biometrics, 40(1):253, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530768.

Burnham:1984:JES

[813] Kenneth P. Burnham. Jackknife estimate of species richness. Biometrics, 40(1):255, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530769.

Aranda-Ordaz:1984:DFM

[814] F. J. Aranda-Ordaz, Robert Tibshirani, and Antonio Ciampi. Degrees of freedom in model-fitting. Biometrics, 40(1):255–256, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530770.

Peace:1984:DLS

[815] Karl E. Peace. Data lists and summaries should reflect experimental structure. Biometrics, 40(1):256, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530771.

Mantel:1984:PSP

[816] Nathan Mantel, R. McHugh, and J. Matts. Pre-stratification or post-stratification. Biometrics, 40(1):256–258, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530772.

Anonymous:1984:Aa

[817] Anonymous. Abstracts. Biometrics, 40(1):259–276, March 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530773.

Silvey:1984:BRA

[818] Valerie Silvey. Book review: The Agricultural Field Experiment. A Statistical Examination of Theory and Practice, by S. C. Pearce. Biometrics, 40 (1):277, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530774.

Paterson:1984:BRC

[819] L. Paterson. Book review: Connectivity in Multi-Factor Designs, by L. Butz. Biometrics, 40(1):277–278, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530775.

Finney:1984:BRS

[820] D. J. Finney. Book review: Statistical Methods in Agriculture and Experimental Biology, by R. Mead, R. N. Curnow. Biometrics, 40(1):278, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530776.

Berthet:1984:BRA

[821] P. Berthet. Book review: Analyse de l’Écologie des Espèces dans les Communautés, by Ph. Daget, M. Godron. Biometrics, 40(1):278–279, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530777.

Lebreton:1984:BRS

[822] Jean-Dominique Lebreton. Book review: Stratégies d’Échantillonnage en Écologie, by S. Frontier. Biometrics, 40(1):279–280, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530778.

Cormack:1984:BRN

[823] R. M. Cormack. Book review: Numerical Ecology, by L. Legendre, P. Legendre. Biometrics, 40(1):280, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530779.

Newton:1984:BRS

[824] J. Newton. Book review: Spatial Statistics, by B. D. Ripley. Biometrics, 40(1):280–281, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530780.

Healy:1984:BRC

[825] M. J. R. Healy. Book review: Clinical Trials-Issues and Approaches, by S. H. Shapiro, T. A. Louis. Biometrics, 40(1):281, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530781.

Marriott:1984:BRB

[826] F. H. C. Marriott. Book review: Biostatistics: a Foundation for Analysis in the Health Sciences, by W. W. Daniel. Biometrics, 40(1):281–282, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530782.

Prentice:1984:BRR

[827] Ross Prentice. Book review: Regression Analysis of Survival Data in Cancer Chemotherapy, by W. H. Carter, Jr., G. L. Wampler, D. L. Stablein. Biometrics, 40(1):282–283, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530783.

Clayton:1984:BRG

[828] David G. Clayton. Book review: Generalised Linear Models in Epidemiological Research: Case-Control Studies, by M. A. Adena, S. R. Wilson. Biometrics, 40(1):283, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530784.

Enke:1984:BRR

[829] H. Enke. Book review: Regressionsgerade und Verwandtes, by H. Riedwyl. Biometrics, 40(1):283, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530785.

Thoni:1984:BRS

[830] H. Thöni. Book review: Statistik: Lehr- und Handbuch der angewandten Statistik, by J. Hartung, B. Elpelt, K.-H. und Klösener. Biometrics, 40(1):283–285, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530786.

Schupbach:1984:BRS

[831] M. Schüpbach. Book review: Statistikprogramme für die Bio-, Humanund Sozialwissenschaften, by G. Bollinger, A. Herrmann, V. Montmann. Biometrics, 40(1):285, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530787.

Stern:1984:BRA

[832] R. D. Stern. Book review: Applied Statistics: a Handbook of Techniques, by L. Sachs. Biometrics, 40(1):285–286, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530788.

Dawid:1984:BRE

[833] A. P. Dawid. Book review: Encyclopedia of Statistical Sciences 2, by S. Kotz, N. L. Johnson, C. B. Read. Biometrics, 40(1):286, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530789.

Kemp:1984:BRC

[834] A. W. Kemp. Book review: Contributions to Statistics: Essays in Honour of Norman Lloyd Johnson, by P. K. Sen. Biometrics, 40(1):286–287, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530790.

Dunnett:1984:BRM

[835] C. W. Dunnett. Book review: Multiple Decision Procedures: Theory and Methodology of Selecting and Ranking Populations, by S. S. Gupta, S. Panchapakesan. Biometrics, 40(1):287–288, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530791.

Gower:1984:BRL

[836] J. C. Gower. Book review: Linear Programing (Algorithms and Applications), by S. Vajda. Biometrics, 40(1):288, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530793.

Sinclair:1984:BRC

[837] C. D. Sinclair. Book review: Computer Science and Statistics, by J. F. Gentleman. Biometrics, 40(1):288, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530792.

Naylor:1984:BRT

[838] J. C. Naylor. Book review: Tables of Numerical Integration, by H. V. Smith. Biometrics, 40(1):289, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530794.

Oldham:1984:BRH

[839] P. D. Oldham. Book review: How to Write and Publish a Scientific Paper, by Robert A. Day. Biometrics, 40(1):289, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530795.

Anonymous:1984:BRE

[840] Anonymous. Book review: Educated Guessing: How to Cope in an Uncertain World, by S. Kotz, D. E. Stroup. Biometrics, 40(1):289–290, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530796.

Anonymous:1984:BRI

[841] Anonymous. Book review: An Introduction to Applied Multivariate Statistics, by M. S. Srivastava, E. M. Carter. Biometrics, 40(1):290, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530797.

Anonymous:1984:BS

[842] Anonymous. The Biometric Society. Biometrics, 40(1):291, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530798.

Anonymous:1984:NA

[843] Anonymous. News and announcements. Biometrics, 40(1):293–294, March 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530799.

Anonymous:1984:FMb

[844] Anonymous. Front matter. Biometrics, 40(2):i–iv, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531381.

Pettitt:1984:FSB

[845] A. N. Pettitt. Fitting a sinusoid to biological rhythm data using ranks. Biometrics, 40(2):295–300, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531382.

Yanagawa:1984:MEI

[846] Takashi Yanagawa, Fumiyoshi Kasagi, and Takesumi Yoshimura. A method for estimating incidence rates of onchocerciasis from skin-snip biopsies with consideration of false negatives. Biometrics, 40(2):301– 311, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531383.

Jerwood:1984:MMO

[847] David Jerwood, Trevor Lewis, and Funmilayo W. O. Saporu. Modelling the migration of Onchocerca volvulus in simuliids, using a simple compartmental process. Biometrics, 40(2):313–322, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531384.

Lewins:1984:BEN

[848] W. A. Lewins and D. N. Joanes. Bayesian estimation of the number of species. Biometrics, 40(2):323–328, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531385.

Pollock:1984:UAV

[849] Kenneth H. Pollock, James E. Hines, and James D. Nichols. The use of auxiliary variables in capture–recapture and removal experiments. Biometrics, 40(2):329–340, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531386.

Green:1984:DFM

[850] Jack Green and Heng Chin Low. The distributions of N and f: Measures of HLA haplotype concordance. Biometrics, 40(2):341–348, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531387.

Donner:1984:TES

[851] Allan Donner, John J. Koval, and Shelley Bull. Testing the effect of sex differences on sib-sib correlations. Biometrics, 40(2):349–356, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531388.

Burrows:1984:IUSa

[852] Peter M. Burrows. Inbreeding under selection from unrelated families. Biometrics, 40(2):357–366, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531389.

Ewens:1984:EPF

[853] W. J. Ewens and B. Asaba. Estimating parameters of the family-size distribution in ascertainment sampling schemes: Numerical results. Biometrics, 40(2):367–374, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531390.

Curnow:1984:PTA

[854] R. N. Curnow. Progeny testing for all-or-none traits when a multifactorial model applies. Biometrics, 40(2):375–382, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531391. See corrections [1055].

Im:1984:PME

[855] S. Im. Plan mixte equilibré à deux facteurs hiérarchiques pour les données binaires. Biometrics, 40(2):383–392, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531392.

Harville:1984:MMP

[856] David A. Harville and Robert W. Mee. A mixed-model procedure for analyzing ordered categorical data. Biometrics, 40(2):393–408, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531393.

Begg:1984:TAN

[857] Colin B. Begg and Leslie A. Kalish. Treatment allocation for nonlinear models in clinical trials: The logistic model. Biometrics, 40(2):409– 420, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531394.

Jewell:1984:SSB

[858] Nicholas P. Jewell. Small-sample bias of point estimators of the odds ratio from matched sets. Biometrics, 40(2):421–435, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531395.

Rosenbaum:1984:DRA

[859] Paul R. Rosenbaum and Donald B. Rubin. Difficulties with regression analyses of age-adjusted rates. Biometrics, 40(2):437–443, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531396.

Prentice:1984:UCS

[860] Ross L. Prentice, William M. Vollmer, and John D. Kalbfleisch. On the use of case series to identify disease risk factors. Biometrics, 40(2):445– 458, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531397.

Larson:1984:CAC

[861] Martin G. Larson. Covariate analysis of competing-risks data with loglinear models. Biometrics, 40(2):459–469, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531398.

Smith:1984:MCC

[862] Kimberly C. Smith, N. E. Savin, and Jacqueline L. Robertson. A Monte Carlo comparison of maximum likelihood and minimum chi square sampling distributions in logit analysis. Biometrics, 40(2):471–482, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531399.

Nagelkerke:1984:STG

[863] N. J. D. Nagelkerke, J. Oosting, and A. A. M. Hart. A simple test for goodness of fit of Cox’s proportional hazards model. Biometrics, 40(2): 483–486, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531400. See

corrections [1003].

Michalek:1984:LRP

[864] Joel E. Michalek and Daniel Mihalko. Linear rank procedures on littermatched data. Biometrics, 40(2):487–491, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531401.

Cain:1984:ACI

[865] Kevin C. Cain and Nicholas T. Lange. Approximate case influence for the proportional hazards regression model with censored data. Biometrics, 40(2):493–499, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531402.

Hughes:1984:ITC

[866] T. H. Hughes and J. H. Matis. An irreversible two-compartment model with age-dependent turnover rates. Biometrics, 40(2):501–505, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531403.

Lee:1984:VCE

[867] Kwan R. Lee and C. H. Kapadia. Variance components estimators for the balanced two-way mixed model. Biometrics, 40(2):507–512, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531404.

Koopman:1984:CIR

[868] P. A. R. Koopman. Confidence intervals for the ratio of two binomial proportions. Biometrics, 40(2):513–517, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531405.

Preece:1984:BTW

[869] D. A. Preece. Biometry in the third world: Science not ritual. Biometrics, 40(2):519–523, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531406.

Macfarlane:1984:SOB

[870] Sarah B. J. Macfarlane. Some opportunities for biometry in promoting child health in the third world. Biometrics, 40(2):525–533, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531407.

Ruppert:1984:OUS

[871] David Ruppert, Rodney L. Reish, Richard B. Deriso, and Raymond J. Carroll. Optimization using stochastic approximation and Monte Carlo simulation (with application to harvesting of Atlantic menhaden). Biometrics, 40(2):535–545, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531408.

Farewell:1984:SSM

[872] V. T. Farewell and S. Dahlberg. Some statistical methodology for the analysis of HLA data. Biometrics, 40(2):547–560, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531409.

Peace:1984:DFN

[873] Karl E. Peace and Eryl Howard. Degrees of freedom in a nonparametric test. Biometrics, 40(2):561, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531410.

Hopkins:1984:RSV

[874] Alan Hopkins, K. L. Lee, F. E. Harrell, H. D. Tolley, and R. A. Rosati. Rao’s statistic for variable selection in Cox’s survival model. Biometrics, 40(2):561–563, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531411.

Edwards:1984:ATC

[875] J. H. Edwards and A. W. F. Edwards. Approximating the tetrachoric correlation coefficient. Biometrics, 40(2):563, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531412.

Kemp:1984:BRA

[876] C. D. Kemp. Book review: The Advanced Theory of Statistics, Vol. 3. Design and Analysis and Time- Series, by M. G. Kendall, A. Stuart, J.

K. Ord. Biometrics, 40(2):565, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531413.

Lawrence:1984:BRS

[877] C. J. Lawrence. Book review: Statistics for Technology — a Course in Applied Statistics, by C. Chatfield. Biometrics, 40(2):565–566, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531414.

Williams:1984:BRG

[878] D. A. Williams. Book review: Generalized Linear Models, by P. McCullagh, J. A. Nelder. Biometrics, 40(2):566, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531415.

Clarke:1984:BRS

[879] G. M. Clarke. Book review: Statistics for Research, by S. Dowdy, S. Wearden. Biometrics, 40(2):566–567, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531416.

Thompson:1984:BRS

[880] R. Thompson. Book review: Statistics and Related Topics, by M. Csörgo, D. A. Dawson, J. N. K. Rao, A. K. Salem. Biometrics, 40(2):567, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531417.

Fisher:1984:BRG

[881] N. I. Fisher. Book review: Graphical Methods for Data Analysis, by J. M. Chambers, W. S. Cleveland, B. Kleiner, P. A. Tukey. Biometrics, 40(2): 567–568, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531418.

Jupp:1984:BRS

[882] P. E. Jupp. Book review: Statistics on Spheres, by G. S. Watson. Biometrics, 40(2):568, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531419.

Claustriaux:1984:BRRa

[883] J. J. Claustriaux. Book review: La Regression: Nouveaux Regards sur une Ancienne Methode Statistique, by R. Tomassone, E. Lesquoy, C.

Millier. Biometrics, 40(2):569, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531420.

Wishart:1984:BRC

[884] D. Wishart. Book review: Clustering of Large Data Sets, by J. Zupan. Biometrics, 40(2):569, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531421.

Rennolls:1984:BRT

[885] K. Rennolls. Book review: Timber Management — a Quantitative Approach, by J. L. Clutter, J. C. Fortson, L. V. Pienaar, G. H. Brister, R. L. Bailey. Biometrics, 40(2):569–570, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531422.

Cormack:1984:BRE

[886] R. M. Cormack. Book review: The Estimation of Animal Abundance and Related Parameters, by G. A. F. Seber. Biometrics, 40(2):570, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531423.

Marshall:1984:BRB

[887] T. F. De C. Marshall. Book review: The Biomathematics of Malaria, by N. T. J. Bailey. Biometrics, 40(2):570, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531424.

Russell:1984:BRH

[888] I. T. Russell. Book review: Hospital Statistics in Europe, by P. M. Lambert, F. H. Roger. Biometrics, 40(2):570–571, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531425.

Day:1984:BRE

[889] N. E. Day. Book review: Environmental Epidemiology: Risk Assessment, by R. L. Prentice, A. S. Whittemore. Biometrics, 40(2):571, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531426.

Taylor:1984:BRN

[890] J. Taylor. Book review: Newer Concepts in Nutrition and Their Implications for Policy, by P. V. Sukhatme. Biometrics, 40(2):571, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531427.

Kirkwood:1984:BRS

[891] T. B. L. Kirkwood. Book review: Statistical Analysis of DNA Sequence Data, by B. S. Weir. Biometrics, 40(2):572, June 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531428.

Thompson:1984:BRA

[892] R. Thompson. Book review: Analysis of Unbalanced Data: a Pre-Program Introduction, by C. C. Li. Biometrics, 40(2):572, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531429.

Hackl:1984:BRD

[893] P. Hackl. Book review: Datenverarbeitung und Datenanalyse mit SAS: Eine problemorientierte Einführung, by H.-M. Uehlinger. Biometrics, 40(2):572–573, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531430.

Hansert:1984:BRS

[894] E. Hansert. Book review: Statistik-Programm-System für die Sozialwissenschaften, by P. Beutel, M. Schubö. Biometrics, 40(2):573, June 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531431.

Anonymous:1984:FMc

[895] Anonymous. Front matter. Biometrics, 40(3):i–iv, September 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530901.

Freedman:1984:SRC

[896] L. S. Freedman, D. Lowe, and P. Macaskill. Stopping rules for clinical trials incorporating clinical opinion. Biometrics, 40(3):575–586, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530902.

Slud:1984:CRV

[897] Eric V. Slud, David P. Byar, and Sylvan B. Green. A comparison of reflected versus test-based confidence intervals for the median survival time, based on censored data. Biometrics, 40(3):587–600, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530903.

Link:1984:CIS

[898] Carol L. Link. Confidence intervals for the survival function using Cox’s proportional- hazard model with covariates. Biometrics, 40(3):601–609, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530904.

Lathrope:1984:PAF

[899] G. Mark Lathrope, J. M. Lalouel, and A. Jacquard. Path analysis of family resemblance and gene-environment interaction. Biometrics, 40(3):611–625, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530905.

Borgan:1984:EED

[900] Ørnulf Borgan, Knut Liestøl, and Peter Ebbesen. Efficiencies of experimental designs for an illness-death model. Biometrics, 40(3):627–638, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530906.

Pregibon:1984:DAM

[901] Daryl Pregibon. Data analytic methods for matched case-control studies. Biometrics, 40(3):639–651, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530907.

Prentice:1984:HDR

[902] Ross L. Prentice, Rainer Storb, K. Stephen Brown, and Mark W. Mason. HLA and disease: Relative-risk regression methods and multiple testing considerations. Biometrics, 40(3):653–661, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530908.

Geertsema:1984:ASP

[903] J. C. Geertsema and C. J. Reinecke. The application of a sequential procedure with elimination as a method for the screening for metabolic diseases. Biometrics, 40(3):663–673, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530909.

Brownie:1984:SSC

[904] C. Brownie and J.-P. Habicht. Selecting a screening cut-off point or diagnostic criterion for comparing prevalences of disease. Biometrics, 40(3):675–684, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530910.

Jolicoeur:1984:PCF

[905] Pierre Jolicoeur. Principal components, factor analysis, and multivariate allometry: a small- sample direction test. Biometrics, 40(3):685–690, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530911.

Hui:1984:CFR

[906] Siu L. Hui. Curve fitting for repeated measurements made at irregular time-points. Biometrics, 40(3):691–697, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530912.

Carter:1984:GCM

[907] E. M. Carter and J. J. Hubert. A growth-curve model approach to multivariate quantal bioassay. Biometrics, 40(3):699–706, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530913.

Ramsey:1984:DD

[908] Fred L. Ramsey and Christopher P. Marsh. Diet dissimilarity. Biometrics, 40(3):707–715, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530914.

Renshaw:1984:CEL

[909] Eric Renshaw. Competition experiments for light in a plant monoculture: an analysis based on two-dimensional spectra. Biometrics, 40(3):717–728, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530915.

Engen:1984:ABP

[910] Steinar Engen. The association between pairs of species on discrete habitable units. Biometrics, 40(3):729–738, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530916.

Conroy:1984:GMM

[911] M. J. Conroy and B. K. Williams. A general methodology for maximum likelihood inference from band-recovery data. Biometrics, 40(3):739–748, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530917.

Manly:1984:OCL

[912] Bryan F. J. Manly. Obtaining confidence limits on parameters of the Jolly–Seber model for capture–recapture data. Biometrics, 40(3):749– 758, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530918.

Upton:1984:MTP

[913] Graham J. G. Upton. On Mead’s test for pattern. Biometrics, 40(3):759– 766, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530919.

Schweder:1984:IGS

[914] Tore Schweder. Investigating group-specific clustering in univariate spatial data with many small groups. Biometrics, 40(3):767–776, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530920.

Bartels:1984:EBM

[915] Robert Bartels. Estimation in a bidirectional mixture of von Mises distributions. Biometrics, 40(3):777–784, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530921.

Nordheim:1984:PLR

[916] Erik V. Nordheim, David M. O’Malley, and Shein-Chung Chow. On the performance of a likelihood ratio test for genetic linkage. Biometrics, 40(3):785–790, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530922.

Elashoff:1984:TSC

[917] Janet D. Elashoff and Terry J. Reedy. Two-stage clinical trial stopping rules. Biometrics, 40(3):791–795, September 1984. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530923.

Tsiatis:1984:ECI

[918] Anastasios A. Tsiatis, Gary L. Rosner, and Cyrus R. Mehta. Exact confidence intervals following a group sequential test. Biometrics, 40(3):797– 803, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530924.

Bedrick:1984:EVE

[919] Edward J. Bedrick. Estimating the variance of empirical logits and contrasts in empirical log probabilities. Biometrics, 40(3):805–809, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530925.

Buckland:1984:MCC

[920] S. T. Buckland. Monte Carlo confidence intervals. Biometrics, 40(3):811– 817, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530926.

Mehta:1984:EST

[921] Cyrus R. Mehta, Nitin R. Patel, and Anastasios A. Tsiatis. Exact significance testing to establish treatment equivalence with ordered categorical data. Biometrics, 40(3):819–825, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530927.

Federer:1984:CEB

[922] Walter T. Federer. Cutting edges in biometry. Biometrics, 40(3):827–839, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530928.

Lauckner:1984:SBA

[923] F. B. Lauckner. The status of biometry in agricultural research in the CARICOM region. Biometrics, 40(3):841–848, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530929.

Kienzler:1984:TBE

[924] J. M. Kienzler, P. F. Dahm, W. A. Fuller, and A. F. Ritter. Temperaturebased estimation for the time of death in white-tailed deer. Biometrics, 40(3):849–854, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530930.

Weinberg:1984:IMM

[925] C. R. Weinberg and M. A. Pfeifer. An improved method for measuring heart-rate variability: Assessment of cardiac autonomic function. Biometrics, 40(3):855–861, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530931.

Perry:1984:NBM

[926] Joe Perry and Jerry Nedelman. Negative binomial model for mosquitoes. Biometrics, 40(3):863–864, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530932.

vanLier:1984:DTB

[927] Henk J. J. van Lier. Disadvantages of treatment blinding. Biometrics, 40(3):864, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530933.

Anonymous:1984:Ab

[928] Anonymous. Abstracts. Biometrics, 40(3):865–867, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530934.

Weir:1984:BRN

[929] B. S. Weir. Book review: The Natural Theory of Molecular Evolution, by M. Kimura. Biometrics, 40(3):869, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530935.

Pope:1984:BRF

[930] J. A. Pope. Book review: Fish Stock Assessment — A Manual of Basic Methods, by J. A. Gulland. Biometrics, 40(3):869–870, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530936.

Bailey:1984:BRP

[931] N. T. J. Bailey. Book review: Planning and Analysis of Observational Studies, by W. G. Cochran. Biometrics, 40(3):870–871, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530937.

Diggle:1984:BRP

[932] P. J. Diggle. Book review: Prediction and Regulation, by Linear Least-Square Methods, by P. Whittle. Biometrics, 40(3):871, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530938.

Pike:1984:BRC

[933] D. J. Pike. Book review: A Course in Linear Models, by A. M. Kshirsagar. Biometrics, 40(3):871–872, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530939.

Bauer:1984:BRP

[934] P. Bauer. Book review: Planung und Auswertung von Untersuchungen im Rahmen des Allgemeinen linearen Modells, by H. Rochel. Biometrics, 40(3):872, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530941.

Williams:1984:BRI

[935] D. A. Williams. Book review: An Introduction to Statistical Modelling, by Annette J. Dobson. Biometrics, 40(3):872, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530940.

Berchtold:1984:BRA

[936] W. Berchtold. Book review: Angewandte multivariate Statistik, by B. Flury, H. Riedwyl. Biometrics, 40(3):872–873, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530942.

Barnett:1984:BRCa

[937] V. Barnett. Book review: The Chronological Annotated Bibliography of Order Statistics, Volume 1: Pre-1950, by H. L. Harter. Biometrics, 40(3): 873, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530944.

Caussinus:1984:BRE

[938] H. Caussinus. Book review: Exercices Commentés de Mathématiques pour l’Analyse Statistique des Données, by J.-P. Nakache, A. Chevalier, V. Morice. Biometrics, 40(3):873, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530943.

Barnett:1984:BRCb

[939] V. Barnett. Book review: The Chronological Annotated Bibliography of Order Statistics, Volume 2: 1950-59, by H. L. Harter. Biometrics, 40(3):873–874, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530945.

Gordon:1984:BRC

[940] A. D. Gordon. Book review: Classification and Regression Trees, by L. Breiman, J. H. Friedman, R. A. Olshen, C. J. Stone. Biometrics, 40(3): 874, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530946.

Day:1984:BRP

[941] N. E. Day. Book review: Probability Models and Cancer, by L. Le Cam, J. Neyman. Biometrics, 40(3):874–875, September 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530947.

Cook:1984:BRD

[942] D. G. Cook. Book review: A Dictionary of Epidemiology, by J. M. Last. Biometrics, 40(3):875, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530949.

Darby:1984:BRI

[943] S. C. Darby. Book review: Incidence of Nonmelanoma Skin Cancer in the United States, by J. Scotto, T. R. Fears, J. F. Fraumeni. Biometrics, 40(3):875, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530948.

Rubenstein:1984:BRM

[944] L. Rubenstein. Book review: Mathematical Methods in Medicine, by R. Bellman. Biometrics, 40(3):875–876, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530950. See corrections [1054].

Anderson:1984:BRC

[945] A. J. B. Anderson. Book review: Computer Science and Statistics: Proceedings of the Fifteenth Symposium on the Interface, by J. E. Gentle.

Biometrics, 40(3):876, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530952.

Franklin:1984:BRB

[946] M. F. Franklin. Book review: Basic Microcomputing and Biostatistics, by D. W. Rogers. Biometrics, 40(3):876, September 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530951.

Anonymous:1984:FMd

[947] Anonymous. Front matter. Biometrics, 40(4):i–iv, December 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531138.

Anonymous:1984:VI

[948] Anonymous. Volume information. Biometrics, 40(4):ix–xxx, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531137.

Heltshe:1984:SPD

[949] J. F. Heltshe and T. A. Ritchey. Spatial pattern detection using quadrat samples. Biometrics, 40(4):877–885, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531139.

Gart:1984:STP

[950] John J. Gart and Jun mo Nam. A score test for the possible presence of recessive alleles in generalized ABO-like genetic systems. Biometrics, 40(4):887–894, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531140.

Burrows:1984:IUSb

[951] Peter M. Burrows. Inbreeding under selection from related families. Biometrics, 40(4):895–906, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531141.

Loyer:1984:IED

[952] Milton W. Loyer and Martin A. Hamilton. Interval estimation of the density of organisms using a serial-dilution experiment. Biometrics, 40(4):907–916, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531142.

Henstridge:1984:MGP

[953] J. D. Henstridge and R. L. Tweedie. A model for the growth pattern of muttonbirds. Biometrics, 40(4):917–925, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531143.

Klein:1984:BMM

[954] John P. Klein, Jerome H. Klotz, and Michael R. Grever. A biological marker model for predicting disease transitions. Biometrics, 40(4):927– 936, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531144.

Carter:1984:SES

[955] Randy L. Carter, Paul Satz, and Miles Hohenegger. On the statistical estimation of speech-organization distributions from aphasia data. Biometrics, 40(4):937–946, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531145.

Oman:1984:EFA

[956] Samuel D. Oman and Yohanan Wax. Estimating fetal age by ultrasound measurements: an example of multivariate calibration. Biometrics, 40(4):947–960, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531146. See correction [1133].

Stiratelli:1984:REM

[957] Robert Stiratelli, Nan Laird, and James H. Ware. Random-effects models for serial observations with binary response. Biometrics, 40(4):961–971, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531147.

OConnell:1984:GOA

[958] Dianne L. O’Connell and Annette J. Dobson. General observeragreement measures on individual subjects and groups of subjects. Biometrics, 40(4):973–983, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531148.

Hill:1984:RCT

[959] N. J. Hill and A. R. Padmanabhan. Robust comparison of two regression lines and biomedical applications. Biometrics, 40(4):985–994, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531149.

Munoz:1984:PSS

[960] Alvaro Muñoz and Bernard Rosner. Power and sample size for a collection of 2×2 tables. Biometrics, 40(4):995–1004, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531150. See corrections [1224, 1225].

Duffy:1984:AEP

[961] Stephen W. Duffy. Asymptotic and exact power for the McNemar test and its analogue with R controls per case. Biometrics, 40(4):1005–1015, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531151.

Lustbader:1984:TNH

[962] Edward D. Lustbader, Suresh H. Moolgavkar, and David J. Venzon. Tests of the null hypothesis in case-control studies. Biometrics, 40(4):1017– 1024, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531152.

Rosner:1984:MMO

[963] Bernard Rosner. Multivariate methods in ophthalmology with application to other paired-data situations. Biometrics, 40(4):1025–1035, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531153.

Lagakos:1984:PPH

[964] S. W. Lagakos and D. A. Schoenfeld. Properties of proportional-hazards score tests under misspecified regression models. Biometrics, 40(4):1037– 1048, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531154.

Breslow:1984:TSC

[965] N. E. Breslow, L. Edler, and J. Berger. A two-sample censored-data rank test for acceleration. Biometrics, 40(4):1049–1062, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531155.

Fidler:1984:CCT

[966] Vaclav Fidler. Change-over clinical trial with binary data: Mixed-model-based comparison of tests. Biometrics, 40(4):1063–1070, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531156.

Abeyasekera:1984:DAR

[967] Savitri Abeyasekera and R. N. Curnow. The desirability of adjusting for residual effects in a crossover design. Biometrics, 40(4):1071–1078, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531157.

OBrien:1984:PCS

[968] Peter C. O’Brien. Procedures for comparing samples with multiple endpoints. Biometrics, 40(4):1079–1087, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531158.

Baddeley:1984:CEU

[969] A. J. Baddeley and B. W. Silverman. A cautionary example on the use of second-order methods for analyzing point patterns. Biometrics, 40(4):1089–1093, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531159.

Thaler:1984:ACC

[970] Howard T. Thaler and Henry I. Braun. On the apparent clustering of clonal albumin production and enzyme activity levels. Biometrics, 40 (4):1095–1102, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531160.

Selwyn:1984:BMB

[971] Murray R. Selwyn and Nancy R. Hall. On Bayesian methods for bioequivalence. Biometrics, 40(4):1103–1108, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531161.

Berry:1984:ADD

[972] D. A. Berry, M. L. Eaton, B. P. Ekholm, and T. L. Fox. Assessing differential drug effect. Biometrics, 40(4):1109–1115, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531162.

Hauck:1984:CSC

[973] Walter W. Hauck. A comparative study of conditional maximum likelihood estimation of a common odds ratio. Biometrics, 40(4):1117–1123, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531163.

Carter:1984:CIA

[974] Walter H. Carter, Jr., Vernon M. Chinchilli, Eleanor D. Campbell, and Galen L. Wapler. Confidence interval about the response at the stationary point of a response surface, with an application to preclinical cancer therapy. Biometrics, 40(4):1125–1130, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531164.

Ulm:1984:CBS

[975] K. Ulm. Classification on the basis of successive observations. Biometrics, 40(4):1131–1136, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531165.

Oakenfull:1984:NEM

[976] Elizabeth Oakenfull, Ilan Salbe, and Ben Zehnwirth. A note on the estimation of the mean intensity of insect infestation. Biometrics, 40 (4):1137–1143, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531166.

Otten:1984:LTE

[977] A. Otten and P. G. de Vries. On line-transect estimators for population density, based on elliptic flushing curves. Biometrics, 40(4):1145–1150, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531167.

Racine-Poon:1984:NES

[978] Amy H. Racine-Poon and David G. Hoel. Nonparametric estimation of the survival function when cause of death is uncertain. Biometrics, 40(4):1151–1158, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531168.

Nelder:1984:MRP

[979] J. A. Nelder and E. L. Frome. Models for rates with Poisson errors. Biometrics, 40(4):1159–1162, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531169.

Puri:1984:AJN

[980] Prem S. Puri. Appreciation. Jerzy Neyman 1894–1981. Biometrics, 40 (4):1163–1167, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531170.

Jesdinsky:1984:OEW

[981] Hans Joachim Jesdinsky. Obituary: Edward Walter 1925–1984. Biometrics, 40(4):1169–1170, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531171.

Armitage:1984:ER

[982] Peter Armitage. Editor’s report, 1979–1984. Biometrics, 40(4):1171– 1172, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531172.

Gart:1984:RES

[983] John J. Gart. Report of the Editor of shorter communications, 1979– 1984. Biometrics, 40(4):1173, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531173.

Mee:1984:CBD

[984] Robert W. Mee and Dan Anbar. Confidence bounds for the difference between two probabilities. Biometrics, 40(4):1175–1176, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531174.

Sturt:1984:MRL

[985] Elizabeth Sturt and Neil Risch. Morbidity risk for late-onset disease. Biometrics, 40(4):1176–1178, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531175.

Levin:1984:DDC

[986] Bruce Levin, Jenny Kline, and Toshiro Tango. Detection of disease clustering in time. Biometrics, 40(4):1179–1180, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531176.

Peace:1984:EE

[987] Karl E. Peace. ’Evaluable’ or ’evaluatable’? Biometrics, 40(4):1180– 1181, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531177.

Anonymous:1984:Ac

[988] Anonymous. Abstracts. Biometrics, 40(4):1183–1207, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531178.

Finney:1984:BRN

[989] D. J. Finney. Book review: Natural Selection, Heredity, and Eugenics, Including Selected Correspondence of R. A. Fisher with Leonard Darwin and Others, by J. H. Bennett, R. A. Fisher, Leonard Darwin. Biometrics, 40(4):1209, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531179.

Jupp:1984:BRM

[990] P. E. Jupp. Book review: Mathematical Models of Morphogenesis, by R. Thom. Biometrics, 40(4):1209, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531180.

Morgan:1984:BRC

[991] B. J. T. Morgan. Book review: Competition Models in Population Biology, by P. Waltman. Biometrics, 40(4):1209–1210, December 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531181.

Claustriaux:1984:BRRb

[992] J. J. Claustriaux. Book review: Régression Linéaire sur Micro-Ordinateurs, by T. Foucart, J.-Y. Lafaye. Biometrics, 40(4):1210, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531183.

Weir:1984:BRT

[993] B. S. Weir. Book review: Trees and Networks in Biological Models, by N. MacDonald. Biometrics, 40(4):1210, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531182.

Aitkin:1984:BRL

[994] M. Aitkin. Book review: Linear Statistical Analysis of Discrete Data, by M. Aickin. Biometrics, 40(4):1210–1211, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531184.

Kalbfleisch:1984:BRR

[995] J. D. Kalbfleisch. Book review: Reliability in the Acquisitions Process, by D. J. De Priest, R. L. Launer. Biometrics, 40(4):1211, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531185.

Hill:1984:BRC

[996] I. D. Hill. Book review: Clinical Trials: a Practical Approach, by S. J. Pocock. Biometrics, 40(4):1211–1212, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531186.

Kerridge:1984:BRI

[997] D. F. Kerridge. Book review: Introductory Biostatistics for the Health Sciences, by R. C. Duncan, R. G. Knapp, C. M. Miller. Biometrics, 40(4): 1212, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531188.

Lutz:1984:BRR

[998] W. Lutz. Book review: Real-Time Clinical Computing, by I. R. Perry. Biometrics, 40(4):1212, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531187.

Anonymous:1984:BSF

[999] Anonymous. The Biometric Society financial statements and auditors reprot January 31, 1984. Biometrics, 40(4):1213–1215, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531189.

Andersen:1984:CTG

[1000] Per Kragh Andersen. Correction to: “Testing goodness of fit of Cox’s regression and life model” [Biometrics 38 (1982), no. 1, 67–77; MR0677963 (83m:62169)]. Biometrics, 40(4):1217, December 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See

Anonymous:1984:CAC

[1001] Anonymous. Corrections: Analysis of covariance using the rank transformation. Biometrics, 40(4):1217, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531191. See [449].

Anonymous:1984:CMI

[1002] Anonymous. Corrections: Methods for the identification of Lexian, Poisson and Markovian variations in the secondary sex ratio. Biometrics, 40(4):1217, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531193. See [799].

Anonymous:1984:CST

[1003] Anonymous. Corrections: Simple test for goodness of fit of Cox’s proportional hazards model. Biometrics, 40(4):1217, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531194. See [863].

Anonymous:1984:CTG

[1004] Anonymous. Corrections: Testing goodness of fit of Cox’s regression and life model. Biometrics, 40(4):1217, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531190. See [344].

Anonymous:1984:CVM

[1005] Anonymous. Corrections: The variance of the Mantel–Haenszel estimator. Biometrics, 40(4):1217, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531192. See [505].

Anonymous:1985:FMa

[1006] Anonymous. Front matter. Biometrics, 41(1):i–iv, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530636.

Rai:1985:DRM

[1007] Kamta Rai and John Van Ryzin. A dose-response model for teratological experiments involving quantal responses. Biometrics, 41(1):1–9, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530637.

Hui:1985:MSR

[1008] Siu L. Hui and Saul H. Rosenberg. Multivariate slope ratio assay with repeated measurements. Biometrics, 41(1):11–18, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530638.

Kenward:1985:UFH

[1009] M. G. Kenward. The use of fitted higher-order polynomial coefficients as covariates in the analysis of growth curves. Biometrics, 41(1):19–28, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530639.

Foutz:1985:MCR

[1010] Robert V. Foutz, D. R. Jensen, and Gary W. Anderson. Multiple comparisons in the randomization analysis of designed experiments with growth curve responses. Biometrics, 41(1):29–37, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530640. See corrections [1174].

Calinski:1985:CMA

[1011] T. Caliński and L. C. A. Corsten. Clustering means in ANOVA by simultaneous testing. Biometrics, 41(1):39–48, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530641.

Lubin:1985:CCM

[1012] Jay H. Lubin. Case-control methods in the presence of multiple failure times and competing risks. Biometrics, 41(1):49–54, March 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530642.

Greenland:1985:ECE

[1013] Sander Greenland and James M. Robins. Estimation of a common effect parameter from sparse follow-up data. Biometrics, 41(1):55–68, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530643.

Tsutakawa:1985:ECM

[1014] Robert K. Tsutakawa. Estimation of cancer mortality rates: a Bayesian analysis of small frequencies. Biometrics, 41(1):69–79, March 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530644.

McClish:1985:IEP

[1015] Donna McClish and Dana Quade. Improving estimates of prevalence by repeated testing. Biometrics, 41(1):81–89, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530645.

Muenz:1985:MMC

[1016] Larry R. Muenz and Lawrence V. Rubinstein. Markov models for covariate dependence of binary sequences. Biometrics, 41(1):91–101, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530646. See correction

[1175, 1179].

Rosenbaum:1985:BDI

[1017] Paul R. Rosenbaum and Donald B. Rubin. The bias due to incomplete matching. Biometrics, 41(1):103–116, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530647.

Weinberg:1985:PAS

[1018] Clarice R. Weinberg. On pooling across strata when frequency matching has been followed in a cohort study. Biometrics, 41(1):117–127, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530648.

Brookmeyer:1985:EAF

[1019] Ron Brookmeyer and Steven G. Self. The effect of additional followup and accrual on survival time data. Biometrics, 41(1):129–136, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530649.

Harville:1985:CIV

[1020] David A. Harville and Alan P. Fenech. Confidence intervals for a variance ratio, or for heritability, in an unbalanced mixed linear model. Biometrics, 41(1):137–152, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530650.

Meyer:1985:MLE

[1021] K. Meyer. Maximum likelihood estimation of variance components for a multivariate mixed model with equal design matrices. Biometrics, 41(1): 153–165, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530651.

Konishi:1985:THA

[1022] Sadanori Konishi. Testing hypotheses about interclass correlations from familial data. Biometrics, 41(1):167–176, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530652.

Buckland:1985:PDM

[1023] S. T. Buckland. Perpendicular distance models for line transect sampling. Biometrics, 41(1):177–195, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530653.

Pennington:1985:ERA

[1024] Michael Pennington. Estimating the relative abundance of fish from a series of trawl surveys. Biometrics, 41(1):197–202, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530654.

Plant:1985:BMS

[1025] Richard E. Plant and L. T. Wilson. A Bayesian method for sequential sampling and forecasting in agricultural pest management. Biometrics, 41(1):203–214, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530655.

Morgan:1985:MAL

[1026] Byron J. T. Morgan and Philip M. North. A model for avian lung ventilation, and the effect of accelerating stimulation in Japanese quail embryos. Biometrics, 41(1):215–226, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530656.

Colgan:1985:EAF

[1027] Patrick W. Colgan and J. Terry Smith. Experimental analysis of food preference transitivity in fish. Biometrics, 41(1):227–236, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530657.

OBrien:1985:RTB

[1028] Peter C. O’Brien and Peter J. Dyck. A runs test based on run lengths. Biometrics, 41(1):237–244, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2530658. See correction [1131, 1132].

Moreau:1985:EHR

[1029] T. Moreau, M. Le Minor, P. Myquel, and J. Lellouch. Estimation of the hazards ratio in two grouped samples. Biometrics, 41(1):245–252, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530659.

Moeschberger:1985:CSM

[1030] M. L. Moeschberger and John P. Klein. A comparison of several methods of estimating the survival function when there is extreme right censoring. Biometrics, 41(1):253–259, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530660.

Cox:1985:IER

[1031] C. Philip Cox. Interval estimates for the ratio of the means of two normal populations with variances related to the means. Biometrics, 41(1):261– 265, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530661. See

corrections [1173].

Spezzaferri:1985:NMC

[1032] F. Spezzaferri. A note on multivariate calibration experiments. Biometrics, 41(1):267–272, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530662.

Alldredge:1985:LTE

[1033] J. Richard Alldredge and Charles E. Gates. Line transect estimators for left-truncated distributions. Biometrics, 41(1):273–280, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530663.

Geramita:1985:UCS

[1034] Joan M. Geramita and J. Terry Smith. Using the correct standard errors in comparing subpopulation survival rates in marked populations. Biometrics, 41(1):281–285, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530664.

Mansour:1985:MLE

[1035] Hussein Mansour, Erik V. Nordheim, and J. J. Rutledge. Maximum likelihood estimation of variance components in repeated measures designs assuming autoregressive errors. Biometrics, 41(1):287–294, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530665.

Tukey:1985:TSC

[1036] J. W. Tukey, J. L. Ciminera, and J. F. Heyse. Testing the statistical certainty of a response to increasing doses of a drug. Biometrics, 41(1): 295–301, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530666.

Emerson:1985:NWM

[1037] John D. Emerson and Lincoln E. Moses. A note on the Wilcoxon–Mann– Whitney test for 2 × k ordered tables. Biometrics, 41(1):303–309, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530667.

Sturt:1985:EMR

[1038] Elizabeth Sturt and Neil Risch. Estimating morbidity risks with variable age of onset. Biometrics, 41(1):311–313, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530668.

Anonymous:1985:Aa

[1039] Anonymous. Abstracts. Biometrics, 41(1):315–339, March 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530669.

Dawid:1985:BRE

[1040] A. P. Dawid. Book review: Encyclopedia of Statistical Sciences, Vol. 3. Faà di Bruno’s Formula — Hypothesis Testing, by S. Kotz, N. L. Johnson, C. B. Read. Biometrics, 41(1):341–342, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530670.

Everitt:1985:BRV

[1041] B. S. Everitt. Book review: The Visual Display of Quantitative Information, by E. R. Tufte. Biometrics, 41(1):342, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530672.

Smith:1985:BRE

[1042] G. L. Smith. Book review: Encyclopedia of Statistical Sciences, Vol. 4. Icing the Tails-Limit Theorems, by S. Kotz, N. L. Johnson, C. B. Read.

Biometrics, 41(1):342, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530671.

Freeman:1985:BRSa

[1043] G. H. Freeman. Book review: Statistical Procedures for Agricultural Research, by K. A. Gomez, A. A. Gomez. Biometrics, 41(1):342–343, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530673.

Fluhler:1985:BRB

[1044] H. Flühler. Book review: Biometrie in der Chemisch-Pharmazeutischen Industrie 1, by J. Vollmar. Biometrics, 41(1):343–344, March 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530674.

Upton:1985:BRL

[1045] G. J. G. Upton. Book review: Linear Statistical Models and Related Methods: With Applications to Social Research, by J. Fox. Biometrics, 41 (1):344, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530675.

Sprent:1985:BRN

[1046] P. Sprent. Book review: Nonlinear Regression Modeling — A Unified Practical Approach, by D. A. Ratkowsky. Biometrics, 41(1):344–345, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530676.

Healy:1985:BRMa

[1047] M. J. R. Healy. Book review: Multivariate Methods in Physical Anthropology, by G. N. van Vark, W. W. Howells. Biometrics, 41(1):345, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530677.

Sokal:1985:BRC

[1048] R. R. Sokal. Book review: Cluster Analysis and Data Analysis, by M. Jambu, M.-O. Lebeaux. Biometrics, 41(1):345, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530678.

Gordon:1985:BRH

[1049] A. D. Gordon. Book review: Handbook of Statistics, Vol. 2. Classification, Pattern Recognition and Reduction of Dimensionality, by P. R.

Krishnaiah, L. N. Kanal. Biometrics, 41(1):345–346, March 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530679.

Marriott:1985:BRI

[1050] F. H. C. Marriott. Book review: Interpretation of Ecological Data. A Primer on Classification and Ordination, by E. C. Pielou. Biometrics, 41 (1):346, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530680.

Moreno-Navarro:1985:BRL

[1051] L. Moreno-Navarro. Book review: The Life Table and Its Applications, by C. L. Chiang. Biometrics, 41(1):346–347, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530681.

Blackith:1985:BRO

[1052] R. E. Blackith. Book review: The Order of Man: a Biomathematical Anatomy of the Primates, by C. Oxnard. Biometrics, 41(1):347, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530682.

Dagnelie:1985:BEC

[1053] Pierre Dagnelie and Herbert A. David. Biometrics editorial changes. Biometrics, 41(1):349, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530683.

Anonymous:1985:CBR

[1054] Anonymous. Corrections: Book review of Mathematical Methods in Medicine. Biometrics, 41(1):351, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530685. See [944].

Anonymous:1985:CPT

[1055] Anonymous. Corrections: Progeny testing for all-or-none traits when a multifactorial model applies. Biometrics, 41(1):351, March 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530684. See [854].

Anonymous:1985:FMb

[1056] Anonymous. Front matter. Biometrics, 41(2):i–iv, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530860.

Armitage:1985:RST

[1057] Peter Armitage, Irene M. Stratton, and Helen V. Worthington. Repeated significance tests for clinical trials with a fixed number of patients and variable follow-up. Biometrics, 41(2):353–359, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530861.

Gail:1985:TQI

[1058] M. Gail and R. Simon. Testing for qualitative interactions between treatment effects and patient subsets. Biometrics, 41(2):361–372, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530862.

Whitehead:1985:DPI

[1059] John Whitehead. Designing Phase II studies in the context of a programme of clinical research. Biometrics, 41(2):373–383, June 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530863.

Crosbie:1985:PMC

[1060] S. F. Crosbie and B. F. J. Manly. Parsimonious modelling of capture– mark–recapture studies. Biometrics, 41(2):385–398, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530864.

Pollock:1985:GFT

[1061] Kenneth H. Pollock, James E. Hines, and James D. Nichols. Goodnessof-fit tests for open capture–recapture models. Biometrics, 41(2):399– 410, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530865.

Brownie:1985:AMC

[1062] Cavell Brownie and Kenneth H. Pollock. Analysis of multiple capture– recapture data using band-recovery methods. Biometrics, 41(2):411– 420, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530866.

Ogata:1985:EIP

[1063] Yosihiko Ogata and Masaharu Tanemura. Estimation of interaction potentials of marked spatial point patterns through the maximum likelihood method. Biometrics, 41(2):421–433, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530867.

Wilson:1985:SIE

[1064] Berry K. Wilson. Simultaneity and its impact on ecological regression applications. Biometrics, 41(2):435–445, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530868.

Nedelman:1985:EMM

[1065] Jerry Nedelman. Estimation for a model of multiple malaria infections. Biometrics, 41(2):447–453, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530869.

Nam:1985:MET

[1066] Jun mo Nam and John J. Gart. The ML estimation and testing of generalized ABO-like data with no observed double recessives. Biometrics, 41(2):455–466, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530870.

James:1985:EDD

[1067] Alan T. James and Robert A. J. Conyers. Estimation of a derivative by a difference quotient: Its application to hepatocyte lactate metabolism. Biometrics, 41(2):467–476, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530871.

Giesbrecht:1985:TSA

[1068] F. G. Giesbrecht and J. C. Burns. Two-stage analysis based on a mixed model: Large-sample asymptotic theory and small-sample simulation results. Biometrics, 41(2):477–486, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530872.

Davis:1985:GMH

[1069] Linda June Davis. Generalization of the Mantel–Haenszel estimator to nonconstant odds ratios. Biometrics, 41(2):487–495, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530873.

Palta:1985:IMP

[1070] Mari Palta. Investigating maximum power losses in survival studies with nonstratified randomization. Biometrics, 41(2):497–504, June 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530874.

Crepeau:1985:AIM

[1071] H. Crepeau, J. Koziol, N. Reid, and Y. S. Yuh. Analysis of incomplete multivariate data from repeated measurement experiments. Biometrics, 41(2):505–514, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530875.

Laake:1985:PMD

[1072] Petter Laake, Knut Laake, and Rolf Aaberge. On the problem of measuring the distance between distribution functions: Analysis of hospitalization versus mortality. Biometrics, 41(2):515–523, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530876.

Ryan:1985:EAA

[1073] Louise M. Ryan. Efficiency of age-adjusted tests in animal carcinogenicity experiments. Biometrics, 41(2):525–531, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530877.

Velu:1985:NSM

[1074] Raja Velu and Maurice McInerney. A note on statistical methods adjusting for intraclass correlation. Biometrics, 41(2):533–538, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530878.

Machado:1985:AIB

[1075] Stella G. Machado and Kent R. Bailey. Assessment of interaction between carcinogens in long-term factorially designed animal experiments. Biometrics, 41(2):539–545, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530879.

Laska:1985:MB

[1076] Eugene M. Laska, Howard B. Kushner, Morris Meisner, and Aage Volund. Multivariate bioassay. Biometrics, 41(2):547–554, June 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530880.

Senn:1985:ETE

[1077] Stephen J. Senn, Richard A. Brown, and K. E. James. Estimating treatment effects in clinical trials subject to regression to the mean. Biometrics, 41(2):555–560, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530881.

Selwyn:1985:TB

[1078] Murray R. Selwyn, Nancy R. Hall, and Arthur P. Dempster. On testing for bioequivalence. Biometrics, 41(2):561, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530882.

Anderson:1985:E

[1079] Sharon Anderson, Walter W. Hauck, and David M. Rocke. [Editorial]. Biometrics, 41(2):561–563, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530883.

Anonymous:1985:Ab

[1080] Anonymous. Abstracts. Biometrics, 41(2):565–590, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530884.

Yates:1985:BRW

[1081] F. Yates. Book review: W. G. Cochran’s Impact on Statistics, by P. S. R. S. Rao, J. Sedransk. Biometrics, 41(2):591–592, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530885.

Pearce:1985:BRM

[1082] S. C. Pearce. Book review: Models of Reality: Shaping Thought and Action, by J. Richardson. Biometrics, 41(2):592, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530886.

Lachenbruch:1985:BRM

[1083] P. A. Lachenbruch. Book review: Multivariate Descriptive Statistical Analysis, by L. Lebart, A. Morineau, K. M. Warwick. Biometrics, 41(2): 592–593, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530887.

Breslow:1985:BRA

[1084] N. Breslow. Book review: Analysis of Survival Data, by D. R. Cox, D. Oakes. Biometrics, 41(2):593–594, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530888.

Hunt:1985:BRM

[1085] R. D. Hunt. Book review: Mathematical Approach to Pattern and Form in Plant Growth, by R. V. Jean. Biometrics, 41(2):594–595, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530889.

Curnow:1985:BRM

[1086] R. N. Curnow. Book review: Methods in Genetic Epidemiology (Contributions to Epidemiology and Biostatistics), by N. E. Morton, D. C. Rao, J.-M. Lalouel. Biometrics, 41(2):595, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530890.

Causton:1985:BRM

[1087] D. R. Causton. Book review: Membrane Transport in Plants, by W. J. Cram, K. Janacek, R. Rybova, K. Sigler. Biometrics, 41(2):595–596, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530891.

Clarke:1985:BRT

[1088] G. M. Clarke. Book review: Tables for Statisticians, by J. White, A. Yeats, G. Skipworth. Biometrics, 41(2):596, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530892.

Marriott:1985:BRA

[1089] F. H. C. Marriott. Book review: Applied Linear Regression Models, by J. Neter, W. Wasserman, M. H. Kutner. Biometrics, 41(2):596, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530893.

Anonymous:1985:BRSa

[1090] Anonymous. Book review: Scientific Inference, Data Analysis and Robustness, by G. E. P. Box, T. Leonard, C.-F. Wu. Biometrics, 41(2): 596–597, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530894.

Anonymous:1985:BRA

[1091] Anonymous. Book review: Applied and Computational Statistics: a First Course, by K. Stoodley. Biometrics, 41(2):597, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530898.

Anonymous:1985:BRB

[1092] Anonymous. Book review: Biology of Knowledge: The Evolutionary Basis of Reason, by R. Riedl. Biometrics, 41(2):597, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530895.

Anonymous:1985:BRN

[1093] Anonymous. Book review: Neurocommunications: an Introduction, by I. C. Whitfield. Biometrics, 41(2):597, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530897.

Anonymous:1985:BRSb

[1094] Anonymous. Book review: Statistics and Probability, by J. Mogyoródi, I. Vincze, W. Wertz. Biometrics, 41(2):597, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530896.

Anonymous:1985:BRT

[1095] Anonymous. Book reviews: Theory of Probability, by H. Jeffreys; An Introduction to Probability Theory, by P. A. P. Moran; The Analysis of Time Series: an Introduction, by C. Chatfield. Biometrics, 41(2): 597, June 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530899.

Anonymous:1985:FMc

[1096] Anonymous. Front matter. Biometrics, 41(3):i–iv, September 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531279.

Wilson:1985:SSP

[1097] I. M. Wilson, Savitri Abeyasekera, R. D. Stern, and V. K. Samaranayake. Stimulating statistical progress in a developing country by a link between universities. Biometrics, 41(3):599–607, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531280.

Cuzick:1985:MAC

[1098] Jack Cuzick. A method for analysing case-control studies with ordinal exposure variables. Biometrics, 41(3):609–621, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531281.

Mickey:1985:GMH

[1099] Ruth M. Mickey and Robert M. Elashoff. A generalization of the Mantel– Haenszel estimator of partial association for 2×J ×K tables. Biometrics, 41(3):623–635, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531282.

Flanders:1985:NVE

[1100] W. Dana Flanders. A new variance estimator for the Mantel–Haenszel odds ratio. Biometrics, 41(3):637–642, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531283. See correction [1176, 1178].

Stablein:1985:TST

[1101] Donald M. Stablein and I. A. Koutrouvelis. A two-sample test sensitive to crossing hazards in uncensored and singly censored data. Biometrics, 41(3):643–652, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531284. See corrections [1180].

Pollock:1985:NCR

[1102] K. H. Pollock, R. A. Lancia, M. C. Conner, and B. L. Wood. A new change-in-ratio procedure robust to unequal catchability of types of animal. Biometrics, 41(3):653–662, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531285. See corrections [1181].

Soms:1985:SPI

[1103] Andrew P. Soms. Simplified point and interval estimation for removal trapping. Biometrics, 41(3):663–668, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531286.

Johnson:1985:LTM

[1104] E. G. Johnson and R. D. Routledge. The line transect method: a nonparametric estimator based on shape restrictions. Biometrics, 41(3):669– 679, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531287. See

corrections [1182].

Cowan:1985:BPR

[1105] Richard Cowan. Branching process results in terms of moments of the generation-time distribution. Biometrics, 41(3):681–689, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531288.

Stedinger:1985:SMI

[1106] Jery R. Stedinger, Christine A. Shoemaker, and Robert F. Tenga. A stochastic model of insect phenology for a population with spatially variable development rates. Biometrics, 41(3):691–701, September 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531289.

Carter:1985:APL

[1107] E. M. Carter and J. J. Hubert. Analysis of parallel-line assays with multivariate responses. Biometrics, 41(3):703–710, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531290.

Christensen:1985:GNM

[1108] Erik R. Christensen and Chung-Yuan Chen. A general noninteractive multiple toxicity model including probit, logit, and Weibull transformations. Biometrics, 41(3):711–725, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531291.

Farewell:1985:SMS

[1109] V. T. Farewell. A statistical model for the study of self-fertilization in plants. Biometrics, 41(3):727–732, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531292.

Anderson:1985:AET

[1110] James R. Anderson and Leslie Bernstein. Asymptotically efficient twostep estimators of the hazards ratio for follow-up studies and survival data. Biometrics, 41(3):733–739, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531293.

Atkinson:1985:CLP

[1111] E. Neely Atkinson and Barry W. Brown. Confidence limits for probability of response in multistage Phase II clinical trials. Biometrics, 41(3):741– 744, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531294.

Bonett:1985:SEL

[1112] Douglas G. Bonett, J. Arthur Woodward, and P. M. Bentler. Some extensions of a linear model for categorical variables. Biometrics, 41(3):745– 750, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531295.

Dinse:1985:TTT

[1113] Gregg E. Dinse. Testing for a trend in tumor prevalence rates: I. Nonlethal tumors. Biometrics, 41(3):751–770, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531296.

Chiang:1985:WB

[1114] Chin Long Chiang and Marvin Zelen. What is biostatistics? Biometrics, 41(3):771–775, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531297.

Mantel:1985:MLV

[1115] Nathan Mantel, Kimberly C. Smith, Jacqueline L. Robertson, and N. E. Savin. Maximum likelihood vs. minimum chi-square. Biometrics, 41 (3):777–783, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531298.

Bross:1985:WPS

[1116] Irwin D. Bross. Why proof of safety is much more difficult than proof of hazard. Biometrics, 41(3):785–793, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531299. See corrections [1183].

Smeeton:1985:EHK

[1117] Nigel C. Smeeton. Early history of the kappa statistic. Biometrics, 41(3): 795, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531300.

Anonymous:1985:Ac

[1118] Anonymous. Abstracts. Biometrics, 41(3):797–809, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531301.

Golub:1985:SLS

[1119] G. H. Golub. Solution of large sparse structured least-squares problem. Biometrics, 41(3):799, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Plackett:1985:BRA

[1120] R. L. Plackett. Book review: Analysis of Ordinal Categorical Data, by A. Agresti. Biometrics, 41(3):811, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531302.

Goldstein:1985:BRI

[1121] H. Goldstein. Book review: An Introduction to Latent Variable Models, by B. S. Everitt. Biometrics, 41(3):811–812, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531303.

Healy:1985:BRMb

[1122] M. J. R. Healy. Book review: Multivariate Morphometrics, 2nd Edition, by R. A. Reyment, R. E. Blackith, N. A. Campbell. Biometrics, 41(3): 812, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531304.

Freedman:1985:BRC

[1123] L. S. Freedman. Book review: Cancer Clinical Trials: Methods and Practice, by M. E. Buyse, M. J. Staquet, R. J. Sylvester. Biometrics, 41(3):812–813, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531305.

Campbell:1985:BRA

[1124] M. J. Campbell. Book review: Atomic Bomb Survivor Data: Utilization and Analysis, by R. L. Prentice, D. J. Thompson. Biometrics, 41(3):813– 814, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531306.

Williamson:1985:BRI

[1125] M. H. Williamson. Book review: Information Analysis of Vegetation Data, by E. Feoli, M. Lagonegro, L. Orloci. Biometrics, 41(3):814, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531307.

Renshaw:1985:BRM

[1126] E. Renshaw. Book review: Mathematical Ecology: Proceedings, Trieste, 1982, by S. A. Levin, T. G. Hallam. Biometrics, 41(3):814–815, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531308.

Anonymous:1985:BRI

[1127] Anonymous. Book review: An Introduction to Multivariate Statistical Analysis, 2nd Edition, by T. W. Anderson. Biometrics, 41(3):815, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531310.

Anonymous:1985:BRSc

[1128] Anonymous. Book review: Statistical Theory and Data Analysis, by K. Matusita. Biometrics, 41(3):815, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531311.

Cannings:1985:BRP

[1129] C. Cannings. Book review: Population Biology: Proceedings, Edmonton, 1982, by H. I. Freedman, C. Strobeck. Biometrics, 41(3):815, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531309.

Anonymous:1985:BS

[1130] Anonymous. The Biometric Society. Biometrics, 41(3):817–819, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531312.

Anonymous:1985:CRT

[1131] Anonymous. Correction: A runs test based on run lengths. Biometrics, 41(3):821, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531313. See [1028].

OBrien:1985:CRT

[1132] P. C. O’Brien and P. J. Dyck. Correction to: “A runs test based on run lengths” [Biometrics 41 (1985), no. 1, 237–244]. Biometrics, 41(3):821, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1028].

Anonymous:1985:CEF

[1133] Anonymous. Correction: Estimating fetal age by ultrasound measurements: an example of multivariate calibration. Biometrics, 41(3):821– 822, September 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531314. See

[1133].

Anonymous:1985:FMd

[1134] Anonymous. Front matter. Biometrics, 41(4):i–iv, December 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530955.

Anonymous:1985:VI

[1135] Anonymous. Volume information. Biometrics, 41(4):ix–xxxiv, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530954.

Armitage:1985:BMS

[1136] Peter Armitage. Biometry and medical statistics. Biometrics, 41(4):823– 833, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530956.

Sinclair:1985:TCD

[1137] Dennis F. Sinclair, James E. Mosimann, and Duane A. Meeter. Tests for character displacement. Biometrics, 41(4):835–846, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530957.

Sinclair:1985:TD

[1138] Dennis F. Sinclair, James E. Mosimann, and Duane A. Meeter. Tests for deletion. Biometrics, 41(4):847–857, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530958.

terBraak:1985:CAI

[1139] Cajo J. F. ter Braak. Correspondence analysis of incidence and abundance data: Properties in terms of a unimodal response model. Biometrics, 41(4):859–873, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530959.

Anderson:1985:GCM

[1140] J. A. Anderson and J. D. Pemberton. The grouped continuous model for multivariate ordered categorical variables and covariate adjustment. Biometrics, 41(4):875–885, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530960.

Lindgren:1985:CEN

[1141] Dag Lindgren. Cost-efficient number of test sites for ranking entries in field trials. Biometrics, 41(4):887–893, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530961.

Narain:1985:PTA

[1142] Prem Narain. Progeny testing with auxiliary traits. Biometrics, 41 (4):895–907, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530962.

Pantula:1985:NAV

[1143] Sastry G. Pantula and Kenneth H. Pollock. Nested analysis of variance with autocorrelated errors. Biometrics, 41(4):909–920, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530963.

Andersen:1985:CRM

[1144] Per Kragh Andersen, Knut Borch-Johnsen, Torsten Deckert, Anders Green, Philip Hougaard, Niels Keiding, and Svend Kreiner. A Cox regression model for the relative mortality and its application to Diabetes mellitus survival data. Biometrics, 41(4):921–932, December 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530964.

Finkelstein:1985:SMR

[1145] Dianne M. Finkelstein and Robert A. Wolfe. A semiparametric model for regression analysis of interval-censored failure time data. Biometrics, 41(4):933–945, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530965.

DeLong:1985:SSM

[1146] Elizabeth R. DeLong, W. Benson Vernon, and R. Randal Bollinger. Sensitivity and specificity of a monitoring test. Biometrics, 41(4):947–958, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530966.

Vacek:1985:ECD

[1147] Pamela M. Vacek. The effect of conditional dependence on the evaluation of diagnostic tests. Biometrics, 41(4):959–968, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530967.

Castellana:1985:ATP

[1148] J. V. Castellana and H. I. Patel. Analysis of two-period crossover design in a multicenter clinical trial. Biometrics, 41(4):969–977, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530968.

Grieve:1985:BAT

[1149] A. P. Grieve. A Bayesian analysis of the two-period crossover design for clinical trials. Biometrics, 41(4):979–990, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530969. See correction [1259, 1260].

Beitler:1985:MEM

[1150] Paula J. Beitler and J. Richard Landis. A mixed-effects model for categorical data. Biometrics, 41(4):991–1000, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530970. See corrections [1352].

Westfall:1985:SSS

[1151] Peter Westfall. Simultaneous small-sample multivariate Bernoulli confidence intervals. Biometrics, 41(4):1001–1013, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530971.

Racine-Poon:1985:BAN

[1152] A. Racine-Poon. A Bayesian approach to nonlinear random effects models. Biometrics, 41(4):1015–1023, December 1985. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530972.

Shanmugam:1985:IPD

[1153] Ramalingam Shanmugam. An intervened Poisson distribution and its medical application. Biometrics, 41(4):1025–1029, December 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530973.

Clobert:1985:DFM

[1154] J. Clobert and J. D. Lebreton. Dependance de facteurs de milieu dans les estimations de taux de survie par capture–recapture. Biometrics, 41(4):1031–1037, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530974.

DeMets:1985:ULT

[1155] David L. DeMets and Mitchell H. Gail. Use of logrank tests and group sequential methods at fixed calendar times. Biometrics, 41(4):1039–1044, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530975.

David:1985:RPM

[1156] H. A. David and In-Hye Ha. Replacement of parents by meritorious offspring in within-family selection. Biometrics, 41(4):1045–1048, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530976.

Hall:1985:CMI

[1157] Peter Hall. Counting methods for inference in binary mosaics. Biometrics, 41(4):1049–1052, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530977.

Iachan:1985:OSB

[1158] Ronaldo Iachan. Optimum stratum boundaries for shellfish surveys. Biometrics, 41(4):1053–1062, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530978.

Lan:1985:RTS

[1159] K. K. Gordon Lan and Janet T. Wittes. Rank tests for survival analysis: a comparison by analogy with games. Biometrics, 41(4):1063–1069, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530979.

Potthoff:1985:CWO

[1160] Richard F. Potthoff, Peter C. O’Brien, and Peter J. Dyck. Choice of weights for the O’Brien–Dyck runs test. Biometrics, 41(4):1071–1073, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530980.

Longford:1985:ICE

[1161] Nicholas Longford and Siu L. Hui. Intrasubject covariance estimation in longitudinal studies. Biometrics, 41(4):1075–1076, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530981.

Stablein:1985:TS

[1162] Donald M. Stablein, Walter H. Carter, Vernon M. Chinchilli, Eleanor D. Campbell, and Galen L. Wampler. Therapeutic synergism. Biometrics, 41(4):1076–1077, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530982.

Richardson:1985:PP

[1163] S. C. Richardson. “Polychotomous” or “Polytomous”? Biometrics, 41(4):1077, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530983.

Anonymous:1985:Ad

[1164] Anonymous. Abstracts. Biometrics, 41(4):1079–1098, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530984.

David:1985:BRO

[1165] H. A. David. Book review: Outliers in Statistical Data, by V. Barnett, T. Lewis. Biometrics, 41(4):1099, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530985.

Freeman:1985:BRSb

[1166] G. H. Freeman. Book review: Statistical Inference Based on Ranks, by T. P. Hettmansperger. Biometrics, 41(4):1099–1100, December 1985.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530986.

Kemp:1985:BRA

[1167] A. W. Kemp. Book review: Applied Statistics Algorithms, by P. Griffiths, I. D. Hill. Biometrics, 41(4):1100, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530987.

Kemp:1985:BRE

[1168] C. D. Kemp. Book review: Elements of Simulation, by B. J. T. Morgan. Biometrics, 41(4):1100–1101, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530988.

Bulmer:1985:BRM

[1169] M. G. Bulmer. Book review: Manual of Quantitative Genetics, by W. A. Becker. Biometrics, 41(4):1101, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530989.

Anonymous:1985:BRG

[1170] Anonymous. Book reviews: Grundbegriffe der Biometrie, by R. J. Lorenz; Biometrie, by W. Kohler, G. Schachtel, P. Voleske. Biometrics, 41 (4):1101–1102, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530990.

Anonymous:1985:BRSd

[1171] Anonymous. Book review: Statistical Inference, by V. K. Rohatgi. Biometrics, 41(4):1102, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530991.

Anonymous:1985:BRW

[1172] Anonymous. Book reviews: World Health Statistics Annual 1984, by World Health Organization; World Health Statistics Quarterly 1984/ 1985, by World Health Organization. Biometrics, 41(4):1102, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530992.

Anonymous:1985:CIE

[1173] Anonymous. Corrections: Interval estimates for the ratio of the means of two normal populations with variances related to the means. Biometrics, 41(4):1103, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530994. See [1031].

Anonymous:1985:CRAa

[1174] Anonymous. Corrections: Résumé to accompany Multiple Comparisons in the Randomization Analysis of Designed Experiments with Growth Curve Responses. Biometrics, 41(4):1103, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530995. See [1010].

Anonymous:1985:CRAb

[1175] Anonymous. Corrections: Résumé to accompany Markov Models for Covariate Dependence of Binary Sequences. Biometrics, 41(4):1103, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530996. See [1016].

Anonymous:1985:CRAc

[1176] Anonymous. Corrections: Résumé to accompany A New Variance Estimator for the Mantel–Haenszel Odds Ratio. Biometrics, 41(4):1103, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530997. See

[1100].

Anonymous:1985:CTC

[1177] Anonymous. Corrections: On testing for a constant hazard against a change-point alternative. Biometrics, 41(4):1103, December 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530993. See [409].

Flanders:1985:CNV

[1178] W. Dana Flanders. Correction: “A new variance estimator for the Mantel–Haenszel odds ratio” [Biometrics 41 (1985), no. 3, 637–642]. Biometrics, 41(4):1103, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1100].

Muenz:1985:CMM

[1179] Larry R. Muenz and Lawrence V. Rubinstein. Correction: “Markov models for covariate dependence of binary sequences” [Biometrics 41 (1985), no. 1, 91–101]. Biometrics, 41(4):1103, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1016].

Anonymous:1985:CRAd

[1180] Anonymous. Corrections: Résumé to accompany A Two-Sample Test Sensitive to Crossing Hazards in Uncensored and Singly Censored Data. Biometrics, 41(4):1104, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530998. See [1101].

Anonymous:1985:CRAe

[1181] Anonymous. Corrections: Résumé to accompany A New Change-in-Ratio Procedure Robust to Unequal Catchability of Types of Animal. Biometrics, 41(4):1104, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530999. See [1102].

Anonymous:1985:CRAf

[1182] Anonymous. Corrections: Résumé to accompany The Line Transect Method: a Nonparametric Estimator Based on Shape Restrictions. Biometrics, 41(4):1104, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531000. See [1104].

Anonymous:1985:CRAg

[1183] Anonymous. Corrections: Résumé to accompany Why Proof of Safety is Much More Difficult Than Proof of Hazard. Biometrics, 41(4):1104, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531001. See

[1116].

Stablein:1985:ATS

[1184] Donald M. Stablein and I. A. Koutrouvelis. Addendum: “A two-sample test sensitive to crossing hazards in uncensored and singly censored data”. Biometrics, 41(4):1104, December 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1986:FMa

[1185] Anonymous. Front matter. Biometrics, 42(1):i–iv, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531238.

Zeh:1986:MME

[1186] Judith E. Zeh, Daijin Ko, Bruce D. Krogman, and Ronald Sonntag. A multinomial model for estimating the size of a whale population from incomplete census data. Biometrics, 42(1):1–14, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531239.

Sutherland:1986:MIM

[1187] R. A. Sutherland. A method for inferring the minimum number of genes controlling the reactions of cultivars of a host plant species to isolates of a pathogen under a gene-for-gene model. Biometrics, 42(1):15–24, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531240.

Ritland:1986:JML

[1188] Kermit Ritland. Joint maximum likelihood estimation of genetic and mating structure using open-pollinated progenies. Biometrics, 42(1):25– 43, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531241.

Sobel:1986:SNP

[1189] Marc J. Sobel, Jonathan Arnold, and Milton Sobel. Statistics of natural populations. III. Sequential sampling plans for the estimation of gene frequencies. Biometrics, 42(1):45–65, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531242.

Gimelfarb:1986:OPG

[1190] A. Gimelfarb. Offspring-parent genotypic regression: How linear is it? Biometrics, 42(1):67–71, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531243.

Frome:1986:MLE

[1191] E. L. Frome and R. J. DuFrain. Maximum likelihood estimation for cytogenetic dose-response curves. Biometrics, 42(1):73–84, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531244.

Kupper:1986:ILE

[1192] Lawrence L. Kupper, Christopher Portier, Michael D. Hogan, and Eiji Yamamoto. The impact of litter effects on dose-response modeling in teratology. Biometrics, 42(1):85–98, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531245.

Samuel-Cahn:1986:STC

[1193] Ester Samuel-Cahn and Yohanan Wax. A sequential test for comparing two infection rates in a randomized clinical trial, and incorporation of data accumulated after stopping. Biometrics, 42(1):99–108, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531246.

Prentice:1986:ALR

[1194] Ross L. Prentice and Mark W. Mason. On the application of linear relative risk regression models. Biometrics, 42(1):109–120, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531247.

Zeger:1986:LDA

[1195] Scott L. Zeger and Kung-Yee Liang. Longitudinal data analysis for discrete and continuous outcomes. Biometrics, 42(1):121–130, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531248.

Srivastava:1986:MBC

[1196] M. S. Srivastava. Multivariate bioassay, combination of bioassays, and Fieller’s theorem. Biometrics, 42(1):131–141, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531249. See corrections [1580, 1581].

McLaren:1986:AVR

[1197] Christine E. McLaren, Gary M. Brittenham, and Victor Hasselblad. Analysis of the volume of red blood cells: Application of the expectationmaximization algorithm to grouped data from the doubly-truncated lognormal distribution. Biometrics, 42(1):143–158, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531250.

Hudson:1986:ETM

[1198] H. M. Hudson. Evaluation of trends in middle ear disease among Australian Aborigines. Biometrics, 42(1):159–169, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531251.

Lyons:1986:ESD

[1199] N. I. Lyons and K. Hutcheson. Estimation of Simpson’s diversity when counts follow a Poisson distribution. Biometrics, 42(1):171–176, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531252.

Oakes:1986:ALP

[1200] David Oakes. An approximate likelihood procedure for censored data. Biometrics, 42(1):177–182, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531253.

Williams:1986:NSN

[1201] D. A. Williams. A note on Shirley’s nonparametric test for comparing several dose levels with a zero-dose control. Biometrics, 42(1):183–186, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531254.

House:1986:NVW

[1202] Dennis E. House. A nonparametric version of Williams’ test for a randomized block design. Biometrics, 42(1):187–190, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531255.

Boos:1986:TTE

[1203] Dennis D. Boos and Cavell Brownie. Testing for a treatment effect in the presence of nonresponders. Biometrics, 42(1):191–197, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531256.

Mantel:1986:AEC

[1204] Nathan Mantel and Walter W. Hauck. Alternative estimators of the common odds ratio for 2 × 2 tables. Biometrics, 42(1):199–202, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531257.

Kay:1986:TEC

[1205] Richard Kay. Treatment effects in competing-risks analysis of prostate cancer data. Biometrics, 42(1):203–211, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531258.

Korn:1986:SST

[1206] Edward L. Korn. Sample size tables for bounding small proportions. Biometrics, 42(1):213–216, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2531259. See correction [1301].

Gianola:1986:REM

[1207] Daniel Gianola and Rohan L. Fernando. Random-effects models for binary responses. Biometrics, 42(1):217–218, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531260.

Whittemore:1986:TID

[1208] Alice Whittemore and Joseph B. Keller. On Tango’s index for disease clustering in time. Biometrics, 42(1):218, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531261.

OQuigley:1986:CIS

[1209] John O’Quigley and Carol L. Link. Confidence intervals for the survival function in the presence of covariates. Biometrics, 42(1):219–220, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531262.

Kemp:1986:BRE

[1210] A. W. Kemp. Book review: Encyclopedia of Statistical Sciences, Vol. 6. Multivariate Analysis- Plackett and Burman Designs, by S. Kotz, N. L. Johnson, C. B. Read. Biometrics, 42(1):221–222, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531263.

Marriott:1986:BRC

[1211] F. H. C. Marriott. Book reviews: Canonical Analysis: a Review with Applications in Ecology; Biomathematics, Vol. 12, by R. Gittins. Biometrics, 42(1):222, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531264.

Krzanowski:1986:BRM

[1212] W. J. Krzanowski. Book review: Multivariate Analysis-Methods and Applications, by W. R. Dillon, M. Goldstein. Biometrics, 42(1):222–223, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531265.

Jolliffe:1986:BRT

[1213] I. T. Jolliffe. Book review: Theory and Applications of Correspondence Analysis, by M. J. Greenacre. Biometrics, 42(1):223, March 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531266.

Murtagh:1986:BRC

[1214] F. Murtagh. Book review: Cluster Dissection and Analysis: Theory, Fortran Programs, Examples, by H. Späth. Biometrics, 42(1):224, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531267.

Everitt:1986:BRT

[1215] B. S. Everitt. Book review: Three-Mode Principal Component Analysis, by P. M. Kroonenberg. Biometrics, 42(1):224–225, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531268.

Anonymous:1986:BRMa

[1216] Anonymous. Book review: Mathematics in the Social and Life Sciences, by M. A. Ball. Biometrics, 42(1):225, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531270.

Anonymous:1986:BRNa

[1217] Anonymous. Book review: Nonparametric Methods for Quantitative Analysis, 2nd Edition, by J. D. Gibbons. Biometrics, 42(1):225, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531271.

Everitt:1986:BRH

[1218] B. S. Everitt. Book review: Homogeneity Analysis of Incomplete Data, by J. Meulman. Biometrics, 42(1):225, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531269.

Anonymous:1986:BRNb

[1219] Anonymous. Book review: Nonparametric Statistical Inference, 2nd Edition, by J. D. Gibbons. Biometrics, 42(1):225–226, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531272.

Anonymous:1986:BRPa

[1220] Anonymous. Book review: Population, Disease and Land in Early Japan, 645–900, by W. W. Farris. Biometrics, 42(1):226, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531273.

Anonymous:1986:BRHa

[1221] Anonymous. Book reviews: Handbook of Stochastic Methods for Physics, Chemistry and the Natural Sciences, by C. W. Gardiner; Elements of Applied Stochastic Processes, 2nd Edition, by U. N. Bhat; Comparison Methods for Queues and Other Stochastic Models, by D. Stoyan, D. J. Daley; Parameter Estimation for Stochastic Processes, by Yu A. Kutoyants, B. L. S. Prakasa Rao. Biometrics, 42(1):226–227, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531274.

Anonymous:1986:MBR

[1222] Anonymous. Models for biomedical research, by committee on models for biomedical research. Biometrics, 42(1):227, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531275.

Anonymous:1986:CESa

[1223] Anonymous. Corrections: Evaluating serial cancer marker studies in patients at risk of recurrent disease. Biometrics, 42(1):229, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531276. See [150].

Anonymous:1986:CPS

[1224] Anonymous. Corrections: Power and sample size for a collection of 2 × 2 tables. Biometrics, 42(1):229, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531277. See [960].

Munoz:1986:CPS

[1225] Alvaro Muñoz and Bernard Rosner. Correction to: “Power and sample size for a collection of 2×2 tables” [Biometrics 40 (1984), no. 4, 995–1004; MR0786176 (86d:62090)]. Biometrics, 42(1):229, March 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [960].

Anonymous:1986:FMb

[1226] Anonymous. Front matter. Biometrics, 42(2):i–iv, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531046.

Besag:1986:SAF

[1227] Julian Besag and Rob Kempton. Statistical analysis of field experiments using neighbouring plots. Biometrics, 42(2):231–251, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531047.

Mead:1986:SCI

[1228] Roger Mead, Janet Riley, Keith Dear, and S. P. Singh. Stability comparison of intercropping and monocropping systems. Biometrics, 42(2): 253–266, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531048.

Seber:1986:REA

[1229] G. A. F. Seber. A review of estimating animal abundance. Biometrics, 42(2):267–292, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531049.

Robins:1986:MBS

[1230] James M. Robins, Mitchell H. Gail, and Jay H. Lubin. More on “Biased selection of controls for case-control analyses of cohort studies”. Biometrics, 42(2):293–299, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531050.

Prentice:1986:DSC

[1231] Ross L. Prentice. On the design of synthetic case-control studies. Biometrics, 42(2):301–310, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531051.

Robins:1986:EMH

[1232] James Robins, Norman Breslow, and Sander Greenland. Estimators of the Mantel–Haenszel variance consistent in both sparse data and largestrata limiting models. Biometrics, 42(2):311–323, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531052.

Dewanji:1986:NMS

[1233] Anup Dewanji and J. D. Kalbfleisch. Nonparametric methods for survival/sacrifice experiments. Biometrics, 42(2):325–341, June 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531053.

Tamura:1986:IHC

[1234] Roy N. Tamura and S. Stanley Young. The incorporation of historical control information in tests of proportions: Simulation study of Tarone’s procedure. Biometrics, 42(2):343–349, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531054.

Jewell:1986:BCU

[1235] Nicholas P. Jewell. On the bias of commonly used measures of association for 2×2 tables. Biometrics, 42(2):351–358, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531055.

Kraemer:1986:MAS

[1236] Helena Chmura Kraemer. A measure of 2 × 2 association with stable variance and approximately normal small-sample distribution: Planning cost-effective studies. Biometrics, 42(2):359–370, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531056.

McMahan:1986:EDB

[1237] C. Alex McMahan, L. C. Maxwell, and A. P. Shepherd. Estimation of the distribution of blood vessel diameters from the arteriovenous passage of microspheres. Biometrics, 42(2):371–380, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531057.

Muir:1986:ERS

[1238] William M. Muir. Estimation of response to selection and utilization of control populations for additional information and accuracy. Biometrics, 42(2):381–391, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531058.

Clerget-Darpoux:1986:EMG

[1239] Françoise Clerget-Darpoux, Catherine Bonaïti-Pellié, and Joëlle Hochez. Effects of misspecifying genetic parameters in lod score analysis. Biometrics, 42(2):393–399, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531059.

Donner:1986:CCI

[1240] Allan Donner and George Wells. A comparison of confidence interval methods for the intraclass correlation coefficient. Biometrics, 42(2):401– 412, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531060. See cor-

rections [1351, 1526, 1354, 1528].

Piegorsch:1986:TSI

[1241] Walter W. Piegorsch, Clarice R. Weinberg, and Joseph K. Haseman. Testing for simple independent action between two factors for dichotomous response data. Biometrics, 42(2):413–419, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531061.

Meisner:1986:CMB

[1242] Morris Meisner, H. B. Kushner, and Eugene M. Laska. Combining multivariate bioassays. Biometrics, 42(2):421–427, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531062.

Haber:1986:TPI

[1243] Michael Haber. Testing for pairwise independence. Biometrics, 42(2): 429–435, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531063.

Giesbrecht:1986:ADI

[1244] F. G. Giesbrecht. Analysis of data from incomplete block designs. Biometrics, 42(2):437–448, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531064. See correction [1624, 1625].

Fleiss:1986:MCS

[1245] Joseph L. Fleiss. On multiperiod crossover studies. Biometrics, 42(2): 449–450, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531065.

Berg:1986:BRE

[1246] S. Berg. Book review: Encyclopedia of Statistical Sciences. Vol. 5: Lindeberg Condition to Multitrait-Multimethod Matrices, by S. Kotz, N. L.

Johnson, C. B. Read. Biometrics, 42(2):451–452, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531066.

Kemp:1986:BRH

[1247] A. W. Kemp. Book reviews: Handbook of Applicable Mathematics. Vol. II: Probability, by E. Lloyd, W. Ledermann; Handbook of Applicable Mathematics. Vols. VI(A), VI(B): Statistics, by E. Lloyd, W. Ledermann. Biometrics, 42(2):452–453, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531067.

Kemp:1986:BRD

[1248] C. D. Kemp. Book reviews: Dictionary and Classified Bibliography of Statistical Distributions in Scientific Work. Vol. 1: Discrete Models, by G. P. Patil, M. T. Boswell, S. W. Joshi, M. V. Ratnaparkhi; Dictionary and Classified Bibliography of Statistical Distributions in Scientific Work. Vol. 2: Continuous Univariate Models, by G. P. Patil, M. T. Boswell, M. V. Ratnaparkhi; Dictionary and Classified Bibliography of Statistical Distributions in Scientific Work. vol. 3: Multivariate Models, by G. P. Patil, M. T. Boswell, M. V. Ratnaparkhi, J. J. J. Roux. Biometrics, 42(2):453–454, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531068.

Freeman:1986:BRS

[1249] P. R. Freeman. Book review: Statistical Methods for Meta-Analysis, by L. V. Hedges, I. Olkin. Biometrics, 42(2):454, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531069.

Freeman:1986:BRA

[1250] G. H. Freeman. Book review: Analysis of Messy Data. Vol. 1: Designed Experiments, by G. A. Milliken, D. E. Johnson. Biometrics, 42(2):454– 455, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531070.

Cormack:1986:BRI

[1251] R. M. Cormack. Book review: Introduction to Data Analysis, with Examples from Population and Community Ecology, by L. Orloci, N. C. Kenkel. Biometrics, 42(2):455, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531071.

Derr:1986:BRA

[1252] J. A. Derr. Book review: Applied Extreme Value Statistics, by R. R. Kinnison. Biometrics, 42(2):455–456, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531072.

Darby:1986:BRH

[1253] S. C. Darby. Book review: A Handbook of Radioactivity Measurements Procedures, 2nd Edition, by National Council on Radiation Protection and Measurements. Biometrics, 42(2):456, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531073.

Anonymous:1986:BRHb

[1254] Anonymous. Book review: Handbook of Statistics. Vol. 3: Time Series in the Frequency Domain, by D. R. Brillinger, P. R. Krishnaiah. Biometrics, 42(2):456–457, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531074.

Anonymous:1986:BRE

[1255] Anonymous. Book review: Essays in Time Series and Allied Processes, by J. Gani, M. Priestley. Biometrics, 42(2):457, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531076.

Anonymous:1986:BRHc

[1256] Anonymous. Book review: Handbook of Statistics. Vol. 5: Time Series in the Time Domain, by E. J. Hannan, P. R. Krishnaiah, M. M. Rao. Biometrics, 42(2):457, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531075.

Anonymous:1986:BRSa

[1257] Anonymous. Book reviews: Selected Tables in Mathematical Statistics, Vol. 8, by Institute of Mathematical Statistics; Selected Tables in Mathematical Statistics, Vol. 9, by Institute of Mathematical Statistics. Biometrics, 42(2):457–458, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531077.

Anonymous:1986:BRSb

[1258] Anonymous. Book review: Synthetic Membrane Processes, by G. Belfort. Biometrics, 42(2):458, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531078.

Anonymous:1986:CBA

[1259] Anonymous. Correction: A Bayesian analysis of the two-period crossover design for clinical trials. Biometrics, 42(2):459, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531079. See [1149].

Grieve:1986:CBA

[1260] A. P. Grieve. Correction to: “A Bayesian analysis of the two-period crossover design for clinical trials” [Biometrics 41 (1985), no. 4, 979–990]. Biometrics, 42(2):459, June 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1149].

Anonymous:1986:FMc

[1261] Anonymous. Front matter. Biometrics, 42(3):i–iv, September 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531196.

Whitehead:1986:SAC

[1262] John Whitehead. Supplementary analysis at the conclusion of a sequential clinical trial. Biometrics, 42(3):461–471, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531197.

Daudin:1986:SVM

[1263] J.-J. J. Daudin. Selection of variables in mixed-variable discriminant analysis. Biometrics, 42(3):473–481, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531198.

Musmeci:1986:VGA

[1264] F. Musmeci and D. Vere-Jones. A variable-grid algorithm for smoothing clustered data. Biometrics, 42(3):483–494, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531199.

Whittemore:1986:SEU

[1265] Alice S. Whittemore and Joseph B. Keller. Survival estimation using splines. Biometrics, 42(3):495–506, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531200.

Lachin:1986:ESS

[1266] John M. Lachin and Mary A. Foulkes. Evaluation of sample size and power for analyses of survival with allowance for nonuniform patient entry, losses to follow-up, noncompliance, and stratification. Biometrics, 42(3):507–519, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531201. See corrections [1353].

Self:1986:LRT

[1267] Steven G. Self and Elizabeth A. Grossman. Linear rank tests for intervalcensored data with application to PCB levels in adipose tissue of transformer repair workers. Biometrics, 42(3):521–530, September 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531202.

Schneider:1986:IBT

[1268] H. Schneider and L. Weissfeld. Inference based on Type II censored samples. Biometrics, 42(3):531–536, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531203.

Donner:1986:LSR

[1269] Allan Donner and Walter W. Hauck. The large-sample relative efficiency of the Mantel–Haenszel estimator in the fixed-strata case. Biometrics, 42(3):537–545, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531204.

Weinberg:1986:BGD

[1270] Clarice Ring Weinberg and Beth C. Gladen. The beta-geometric distribution applied to comparative fecundability studies. Biometrics, 42(3):547– 560, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531205.

Brownie:1986:CPC

[1271] Cavell Brownie, James E. Hines, and James D. Nichols. Constantparameter capture–recapture models. Biometrics, 42(3):561–574, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531206.

Ratkowsky:1986:CNL

[1272] David A. Ratkowsky and Terry J. Reedy. Choosing near-linear parameters in the four-parameter logistic model for radioligand and related assays. Biometrics, 42(3):575–582, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531207.

Snapinn:1986:TSU

[1273] Steven M. Snapinn and Robert D. Small. Tests of significance using regression models for ordered categorical data. Biometrics, 42(3):583– 592, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531208.

Willan:1986:CTP

[1274] Andrew R. Willan and Joseph L. Pater. Carryover and the two-period crossover clinical trial. Biometrics, 42(3):593–599, September 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531209.

Vonesh:1986:SSM

[1275] Edward F. Vonesh and M. Anthony Schork. Sample sizes in the multivariate analysis of repeated measurements. Biometrics, 42(3):601–610, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531210.

Bonney:1986:RLM

[1276] George Ebow Bonney. Regressive logistic models for familial disease and other binary traits. Biometrics, 42(3):611–625, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531211.

Weller:1986:MLT

[1277] J. I. Weller. Maximum likelihood techniques for the mapping and analysis of quantitative trait loci with the aid of genetic markers. Biometrics, 42(3):627–640, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531212.

Williams:1986:IEM

[1278] D. A. Williams. Interval estimation of the median lethal dose. Biometrics, 42(3):641–645, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531213. See correction [1527, 1529].

Hewett:1986:TST

[1279] John E. Hewett and Eugenia Bair. A two-stage test for ordered means in the Poisson case with an example from mutagenicity testing. Biometrics, 42(3):647–651, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531214.

Munoz:1986:RAP

[1280] Alvaro Muñoz, Bernard Rosner, and Vincent Carey. Regression analysis in the presence of heterogeneous intraclass correlations. Biometrics, 42(3):653–658, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531215.

Cobby:1986:DEE

[1281] J. M. Cobby, P. F. Chapman, and D. J. Pike. Design of experiments for estimating inverse quadratic polynomial responses. Biometrics, 42 (3):659–664, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531216.

Weir:1986:LLM

[1282] B. S. Weir, S. R. Wilson, and Michael Haber. Log-linear models for linked loci. Biometrics, 42(3):665–670, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531217.

Salsburg:1986:AHE

[1283] David Salsburg. Alternative hypotheses for the effects of drugs in smallscale clinical studies. Biometrics, 42(3):671–674, September 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531218.

Nurminen:1986:CIR

[1284] Markku Nurminen. Confidence intervals for the ratio and difference of two binomial proportions. Biometrics, 42(3):675–676, September 1986.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531219.

Kemp:1986:BRC

[1285] A. W. Kemp. Book review: A Celebration of Statistics, by A. C. Atkinson, S. E. Fienberg. Biometrics, 42(3):677–678, September 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531220.

Loynes:1986:BRA

[1286] R. M. Loynes. Book review: The Analysis of Linear Models, by R. R. Hocking. Biometrics, 42(3):678, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531221.

Pettitt:1986:BRN

[1287] A. N. Pettitt. Book review: Nonparametric Methods in General Linear Models, by M. L. Puri, P. K. Sen. Biometrics, 42(3):678–679, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531222.

Jeffers:1986:BRL

[1288] J. N. R. Jeffers. Book review: Linear Models in Biology, by M. R. Cullen. Biometrics, 42(3):679, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531223.

Newton:1986:BRS

[1289] J. Newton. Book review: Statistical Analysis of Finite Mixture Distributions, by D. M. Titterington, A. F. M. Smith, U. E. Makov. Biometrics, 42(3):679–680, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531224.

Gore:1986:BRC

[1290] S. M. Gore. Book review: A Companion to Medical Statistics, by E. A. Murphy. Biometrics, 42(3):680–681, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531225.

North:1986:BRM

[1291] P. M. North. Book review: Methodology of Clinical Drug Trials, by A. Spriet, P. Simon. Biometrics, 42(3):681–682, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531226.

Campbell:1986:BRM

[1292] M. J. Campbell. Book review: Methodologies for the Collection and Analysis of Mortality Data, by J. Vallin, J. H. Pollard, L. Heligman. Biometrics, 42(3):682, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531227.

Cohen:1986:BRP

[1293] G. R. Cohen. Book review: Population and Biology, by N. Keyfitz. Biometrics, 42(3):682–683, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531228.

Shepherd:1986:BRB

[1294] J. G. Shepherd. Book review: Bioeconomic Modelling and Fisheries Management, by C. W. Clark. Biometrics, 42(3):683–684, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531229.

Anonymous:1986:BRMb

[1295] Anonymous. Book review: Making Decisions, by D. V. Lindley. Biometrics, 42(3):684, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531230.

Anonymous:1986:BRR

[1296] Anonymous. Book review: Robustness of Statistical Methods and Nonparametric Statistics, by D. Rasch, M. L. Tiku. Biometrics, 42(3):684, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531231.

Anonymous:1986:BRSc

[1297] Anonymous. Book review: Statistics and Optimization: The Interface, by B. L. Golden, A. A. Assad, S. H. Zanakis. Biometrics, 42(3):684, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531232.

Anonymous:1986:BRC

[1298] Anonymous. Book review: Chaos, Fractiles and Dynamics, by P. Fischer, W. R. Smith. Biometrics, 42(3):685, September 1986. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531234.

Anonymous:1986:BRMc

[1299] Anonymous. Book review: Modern Statistical Methods in Digital Simulation, by P. R. Tadikamalla. Biometrics, 42(3):685, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531233.

Anonymous:1986:BSF

[1300] Anonymous. The Biometric Society financial statements and auditors’ report January 31, 1986. Biometrics, 42(3):687–689, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531235.

Anonymous:1986:CSS

[1301] Anonymous. Correction: Sample size tables for bounding small proportions. Biometrics, 42(3):691, September 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531236. See [1206].

Anonymous:1986:FMd

[1302] Anonymous. Front matter. Biometrics, 42(4):i–iv, December 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530688.

Anonymous:1986:VI

[1303] Anonymous. Volume information. Biometrics, 42(4):xiii–xxx, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530687.

Brillinger:1986:BIP

[1304] David R. Brillinger. A biometrics invited paper with discussion: The natural variability of vital rates and associated statistics. Biometrics, 42(4):693–734, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530689. With discussion and a reply by the author.

Breslow:1986:NVV

[1305] N. Breslow, C. L. Chiang, J. M. Hoem, P. Jagers, N. Keiding, L. Kish, K. G. Manton, and J. W. Tukey. The natural variability of vital-rates and associated statistics — discussion. Biometrics, 42(4):712–732, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1986:CNV

[1306] John W. Tukey. Comments on The natural variability of vital rates and associated statistics. Biometrics, 42(4):729–732, December 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530689.

Jensen:1986:SMS

[1307] E. B. Jensen and R. Sundberg. Statistical models for stereological inference about spatial structures: On the applicability of best linear unbiased estimators in stereology. Biometrics, 42(4):735–751, December 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530690.

Clayton:1986:SRD

[1308] Gareth Clayton and Trevor F. Cox. Some robust density estimators for spatial point processes. Biometrics, 42(4):753–767, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530691.

Cowan:1986:BAM

[1309] Richard Cowan and Robert Staudte. The bifurcating autoregression model in cell lineage studies. Biometrics, 42(4):769–783, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530692.

Jolicoeur:1986:LNV

[1310] Pierre Jolicoeur and Alfred A. Heusner. Log-normal variation belts for growth curves. Biometrics, 42(4):785–794, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530693.

Zerbe:1986:MCR

[1311] Gary O. Zerbe and James R. Murphy. On multiple comparisons in the randomization analysis of growth and response curves. Biometrics, 42(4):795–804, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530694.

Jennrich:1986:URM

[1312] Robert I. Jennrich and Mark D. Schluchter. Unbalanced repeatedmeasures models with structured covariance matrices. Biometrics, 42 (4):805–820, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530695.

Fresen:1986:NFM

[1313] John L. Fresen and June M. Juritz. A note on Foss’s method of obtaining initial estimates for exponential curve fitting by numerical integration. Biometrics, 42(4):821–827, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530696.

Aastveit:1986:AII

[1314] Are Halvor Aastveit and Harald Martens. ANOVA interactions interpreted by partial least squares regression. Biometrics, 42(4):829–844, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530697.

Finkelstein:1986:PHM

[1315] Dianne M. Finkelstein. A proportional hazards model for intervalcensored failure time data. Biometrics, 42(4):845–854, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530698.

Kay:1986:MMA

[1316] Richard Kay. A Markov model for analysing cancer markers and disease states in survival studies. Biometrics, 42(4):855–865, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530699.

Kodell:1986:STR

[1317] R. L. Kodell, D. W. Gaylor, and J. J. Chen. Standardized tumor rates for chronic bioassays. Biometrics, 42(4):867–873, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530700.

Buonaccorsi:1986:DSR

[1318] John P. Buonaccorsi. Designs for slope ratio assays. Biometrics, 42 (4):875–882, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530701.

Gross:1986:KCA

[1319] Shulamith T. Gross. The kappa coefficient of agreement for multiple observers when the number of subjects is small. Biometrics, 42(4):883– 893, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530702.

Simonoff:1986:AEP

[1320] Jeffrey S. Simonoff, Yosef Hochberg, and Benjamin Reiser. Alternative estimation procedures for Pr(X < Y) in categorized data. Biometrics, 42(4):895–907, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530703. See correction [1799, 1801].

Dale:1986:GCR

[1321] Jocelyn R. Dale. Global cross-ratio models for bivariate, discrete, ordered responses. Biometrics, 42(4):909–917, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530704.

Parker:1986:SSI

[1322] R. A. Parker and D. J. Bregman. Sample size for individually matched case-control studies. Biometrics, 42(4):919–926, December 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530705.

Woolson:1986:SSC

[1323] Robert F. Woolson, Judy A. Bean, and Patricio B. Rojas. Sample size for case-control studies using Cochran’s statistic. Biometrics, 42(4):927–932, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530706.

Plant:1986:MCE

[1324] Richard E. Plant. A method for computing the elements of the Leslie matrix. Biometrics, 42(4):933–939, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530707.

Loukas:1986:IDT

[1325] S. Loukas and C. D. Kemp. The index of dispersion test for the bivariate Poisson distribution. Biometrics, 42(4):941–948, December 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530708.

Breslow:1986:MEL

[1326] N. E. Breslow and J. Cologne. Methods of estimation in log odds ratio regression models. Biometrics, 42(4):949–954, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530709.

Fears:1986:LRM

[1327] Thomas R. Fears and Charles C. Brown. Logistic regression methods for retrospective case-control studies using complex sampling procedures. Biometrics, 42(4):955–960, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530710.

Yee:1986:CPR

[1328] K. F. Yee. The calculation of probabilities in rejecting bioequivalence. Biometrics, 42(4):961–965, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530711.

Pack:1986:HTP

[1329] Simon E. Pack. Hypothesis testing for proportions with overdispersion. Biometrics, 42(4):967–972, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530712.

Aebischer:1986:EPU

[1330] N. J. Aebischer. Estimating the proportion of uncatchable animals in a population by double- sampling. Biometrics, 42(4):973–979, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530713.

Gill:1986:PAA

[1331] J. L. Gill, G. F. Ericsson, and I. S. Helland. Precision of assessing anthelmintic efficacy. Biometrics, 42(4):981–987, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530714.

Binet:1986:FNB

[1332] Francis E. Binet. Fitting the negative binomial distribution. Biometrics, 42(4):989–992, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530715. See corrections [1579].

Morgan:1986:OCP

[1333] Timothy M. Morgan, S. W. Lagakos, and D. A. Schoenfeld. Omitting covariates from the proportional hazards model. Biometrics, 42(4):993– 995, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530716.

Im:1986:MIM

[1334] Sotan Im, Paula J. Beitler, and J. Richard Landis. Mutual independence in a mixed-effects model. Biometrics, 42(4):997–998, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530717.

Tango:1986:ADT

[1335] Toshiro Tango. On the asymptotic distribution of Tango’s clustering index. Biometrics, 42(4):998–999, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530718.

Ord:1986:BRS

[1336] J. K. Ord. Book review: Spatial Data Analysis, by Example, Vol. 1: Point Pattern and Quantitative Data, by G. Upton, B. Fingleton. Biometrics, 42(4):1001–1002, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530719.

Reyment:1986:BRM

[1337] R. A. Reyment. Book review: Morphometrics in Evolutionary Biology, by F. L. Bookstein, B. Chernoff, R. L. Elder, J. M. Humphries, Jr., G. R. Smith, R. E. Strauss. Biometrics, 42(4):1002–1003, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530720.

Bonney:1986:BRP

[1338] G. E. Bonney. Book review: Pedigree Analysis in Human Genetics, by E. A. Thompson. Biometrics, 42(4):1003, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530722.

Schneider:1986:BRM

[1339] B. Schneider. Book review: Mathematical Models in Immunology, by G. I. Marchuk. Biometrics, 42(4):1003, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530721.

Claustriaux:1986:BRP

[1340] J. J. Claustriaux. Book review: La Preparation des Fichiers pour les Analyses Multivariables sur Micro- Ordinateurs, by P. Laval. Biometrics, 42(4):1004, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530724.

Kemp:1986:BRS

[1341] A. W. Kemp. Book review: Statistical Computation, by J. H. Maindonald. Biometrics, 42(4):1004, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530723.

Anderson:1986:BRA

[1342] A. J. B. Anderson. Book review: Applied Statistics and the SAS Programming Language, by R. P. Cody, J. K. Smith. Biometrics, 42(4):1004–1005, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530725.

Causton:1986:BRI

[1343] D. R. Causton. Book review: Introduction to Statistical Methods, by J. S. Rustagi. Biometrics, 42(4):1005–1006, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530726.

Anonymous:1986:BRHd

[1344] Anonymous. Book review: Herbert Robbins, Selected Papers, by T. L. Lai, D. Siegmund. Biometrics, 42(4):1006, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530727.

Anonymous:1986:BRPb

[1345] Anonymous. Book review: Pacific Statistical Congress 1985 Proceedings, by I. S. Francis, B. F. J. Manly, F. C. Lam. Biometrics, 42(4):1006–1007, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530728.

Anonymous:1986:BRA

[1346] Anonymous. Book review: Asymptotic Statistics 2, by P. Mandl, M. Huskova. Biometrics, 42(4):1007, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530729.

Anonymous:1986:BRD

[1347] Anonymous. Book review: Directory of Statistical Microcomputer Software, by W. A. Woodward, A. C. Elliott, H. L. Gray. Biometrics, 42(4): 1007, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530731.

Anonymous:1986:BRMd

[1348] Anonymous. Book review: Modern Probability Theory, by B. R. Bhat. Biometrics, 42(4):1007, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530732.

Anonymous:1986:BRSd

[1349] Anonymous. Book review: Statistical Packages for the IBM PC Family, by L. O’Keefe, J. Klagge. Biometrics, 42(4):1007, December 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530730.

Anonymous:1986:BRW

[1350] Anonymous. Book reviews: World Health Statistics Annual 1985, by World Health Organization; World Health Statistics Quarterly 1985, Vol. 38 (2), (3), (4); 1986, Vol. 39 (1), (2), by World Health Organization. Biometrics, 42(4):1007–1008, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530733.

Anonymous:1986:CCC

[1351] Anonymous. Corrections: A comparison of confidence intervals for the intraclass correlation coefficient. Biometrics, 42(4):1009, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530735. See [1240].

Anonymous:1986:CCM

[1352] Anonymous. Corrections: Correction to A Mixed-Effects Model for Categorical Data. Biometrics, 42(4):1009, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530734. See [1150].

Anonymous:1986:CESb

[1353] Anonymous. Corrections: Evaluation of sample size and power for analyses of survival with allowance for nonuniform patient entry, losses to follow-up, noncompliance, and stratification. Biometrics, 42(4):1009, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530736. See

[1266].

Donner:1986:CCC

[1354] A. Donner and G. Wells. Correction to: “A comparison of confidence intervals for the intraclass correlation coefficient” [Biometrics 42 (1986), no. 2, 401–412]. Biometrics, 42(4):1009, December 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1240].

Anonymous:1987:FMa

[1355] Anonymous. Front matter. Biometrics, 43(1):i–iv, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531942.

Ramsey:1987:CAE

[1356] Fred L. Ramsey, Valerie Wildman, and John Engbring. Covariate adjustments to effective area in variable-area wildlife surveys. Biometrics, 43 (1):1–11, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531943.

Drummer:1987:SBL

[1357] Thomas D. Drummer and Lyman L. McDonald. Size bias in line transect sampling. Biometrics, 43(1):13–21, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531944.

Kimura:1987:MED

[1358] Daniel K. Kimura and Shiro Chikuni. Mixtures of empirical distributions: an iterative application of the age- length key. Biometrics, 43(1):23–35, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531945.

ONeill:1987:CPO

[1359] M. E. O’Neill, L. H. Prasetyo, A. C. Kirby, and F. W. Nicholas. A comparison of parent-offspring correlation estimators in terms of largesample mean squared error and bias. Biometrics, 43(1):37–44, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531946.

Nassar:1987:SEP

[1360] R. Nassar and M. Hühn. Studies on estimation of phenotypic stability: Tests of significance for nonparametric measures of phenotypic stability.

Biometrics, 43(1):45–53, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531947.

Ainsley:1987:MPE

[1361] A. E. Ainsley, L. J. Paterson, and H. D. Patterson. A method for predicting the efficiency of incomplete-block trials. Biometrics, 43(1):55–59, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531948.

Helland:1987:IUR

[1362] Inge S. Helland. On the interpretation and use of R^{2} in regression analysis. Biometrics, 43(1):61–69, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531949.

Halperin:1987:DFC

[1363] Max Halperin, Peter R. Gilbert, and John M. Lachin. Distribution-free confidence intervals for Pr(X_{1} < X_{2}). Biometrics, 43(1):71–80, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531950.

Duffy:1987:CIB

[1364] Diane E. Duffy and Thomas J. Santner. Confidence intervals for a binomial parameter based on multistage tests. Biometrics, 43(1):81–93, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531951.

VanRyzin:1987:DRM

[1365] John Van Ryzin and Kamta Rai. A dose-response model incorporating nonlinear kinetics. Biometrics, 43(1):95–105, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531952.

Portier:1987:SAT

[1366] Christopher J. Portier and Gregg E. Dinse. Semiparametric analysis of tumor incidence rates in survival/sacrifice experiments. Biometrics, 43(1):107–114, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531953.

Mihalko:1987:RST

[1367] Daniel Mihalko. Repeated significance testing in survival analysis using a litter-matched design. Biometrics, 43(1):115–125, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531954.

Flehinger:1987:NHL

[1368] Betty J. Flehinger and Marek Kimmel. The natural history of lung cancer in a periodically screened population. Biometrics, 43(1):127–144, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531955.

Kalish:1987:EED

[1369] Leslie A. Kalish and Colin B. Begg. Evaluation of efficient designs for observational epidemiologic studies. Biometrics, 43(1):145–167, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531956.

OBrien:1987:PPW

[1370] Peter C. O’Brien and Thomas R. Fleming. A paired Prentice–Wilcoxon test for censored paired data. Biometrics, 43(1):169–180, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531957.

Etezadi-Amoli:1987:EHR

[1371] Jamshid Etezadi-Amoli and Antonio Ciampi. Extended hazard regression for censored survival data with covariates: a spline approximation for the baseline hazard function. Biometrics, 43(1):181–192, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531958.

Wang:1987:AOO

[1372] Samuel K. Wang and Anastasios A. Tsiatis. Approximately optimal oneparameter boundaries for group sequential trials. Biometrics, 43(1):193– 199, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531959.

Bailey:1987:CLR

[1373] B. J. R. Bailey. Confidence limits to the risk ratio. Biometrics, 43(1): 201–205, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531960.

Connor:1987:SST

[1374] Robert J. Connor. Sample size for testing differences in proportions for the paired-sample design. Biometrics, 43(1):207–211, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531961.

Geller:1987:IAR

[1375] Nancy L. Geller and Stuart J. Pocock. Interim analyses in randomized clinical trials: Ramifications and guidelines for practitioners. Biometrics, 43(1):213–223, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531962.

Girard:1987:UMC

[1376] D. M. Girard and D. B. Sager. The use of Markov chains to detect subtle variation in reproductive cycling. Biometrics, 43(1):225–234, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531963.

Gart:1987:EAA

[1377] John J. Gart, Robert E. Tarone, and Louise M. Ryan. On the efficiency of age-adjusted tests in animal carcinogenicity experiments. Biometrics, 43(1):235–244, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531964.

Glasbey:1987:EFD

[1378] C. A. Glasbey, D. A. Elston, Alan T. James, and Robert A. J. Conyers. An example of fitting a differential equation to data. Biometrics, 43(1): 245–248, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531965.

Mantel:1987:REM

[1379] Nathan Mantel, Allan Donner, and Walter W. Hauck. On the relative efficiency of the Mantel–Haenszel estimator. Biometrics, 43(1):249–250, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531966.

Armitage:1987:BRB

[1380] P. Armitage. Book review: Biostatistics: Statistics in Biomedical, Public Health and Environmental Sciences, the Bernard G. Greenberg Volume, by P. K. Sen. Biometrics, 43(1):251–252, March 1987. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531967.

Kemp:1987:BRE

[1381] A. W. Kemp. Book review: Exploring Data Tables, Trends, and Shapes, by D. C. Hoaglin, F. Mosteller, J. W. Tukey. Biometrics, 43(1):252, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531968.

Kemp:1987:RED

[1382] A. W. Kemp. Review: Exploring Data Tables, Trends, and Shapes, by D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Biometrics, 43(1):252, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531968.

Perry:1987:BRR

[1383] J. N. Perry. Book review: Research Data Management in the Ecological Sciences, by W. K. Michener. Biometrics, 43(1):252–253, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531969.

Wright:1987:BRS

[1384] H. L. Wright. Book review: Systems Analysis in Forestry and Forest Industries. Studies in the Management Sciences, Vol. 21, by M. Kallio, A. E. Andersson, R. Seppal, A. Morgan. Biometrics, 43(1):253, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531970.

Nelder:1987:BRS

[1385] J. A. Nelder. Book review: Statistics in Ornithology, by B. J. T. Morgan, P. M. North. Biometrics, 43(1):253–254, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531971.

Edwards:1987:BRA

[1386] A. W. F. Edwards. Book review: Analysis of Human Genetic Linkage, by J. Ott. Biometrics, 43(1):254–255, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531972.

Smith:1987:BRS

[1387] C. A. B. Smith. Book review: Space, Time and Life: The Probabilistic Pathways of Evolution, by V. V. Nalimov. Biometrics, 43(1):255, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531973.

Cook:1987:BRC

[1388] D. G. Cook. Book review: Clinical Epidemiology: The Architecture of Clinical Research, by A. R. Feinstein. Biometrics, 43(1):255–256, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531974.

Osborn:1987:BRU

[1389] J. F. Osborn. Book review: Using and Understanding Medical Statistics, by D. E. Matthews, V. Farewell. Biometrics, 43(1):256, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531975.

Gregson:1987:BRB

[1390] K. Gregson. Book review: Biostatistical Microcomputing in Pascal, by J. H. Klotz, R. D. Meyer. Biometrics, 43(1):256–257, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531976.

Anonymous:1987:BRAa

[1391] Anonymous. Book review: The Analysis of Cross-Classified Data Having Ordered Categories, by L. A. Goodman. Biometrics, 43(1):257, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531978.

Anonymous:1987:BRBa

[1392] Anonymous. Book review: Bayesian Statistics 2, by J. M. Bernardo, M. H. DeGroot, D. V. Lindley, A. F. M. Smith. Biometrics, 43(1):257, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531979.

Anonymous:1987:BRJ

[1393] Anonymous. Book review: Jack Carl Kiefer: Collected Papers, Vols I, II, III, by L. D. Brown, Jack Carl Kiefer, I. Olkin, J. Sacks, H. P. Wynn. Biometrics, 43(1):257, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531977.

Anonymous:1987:BRAb

[1394] Anonymous. Book review: Applied Linear Regression, 2nd Edition, by S. Weisberg. Biometrics, 43(1):258, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531984.

Anonymous:1987:BRFa

[1395] Anonymous. Book review: Flow Cytometry: Instrumentation and Data Analysis, by M. A. Van Dilla, P. N. Dean, O. E. Laerum, M. R. Melamed. Biometrics, 43(1):258, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531985.

Anonymous:1987:BRFb

[1396] Anonymous. Book review: The Frontiers of Modern Statistical Inference Procedures, by E. J. Dudewicz. Biometrics, 43(1):258, March 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531980.

Anonymous:1987:BRGa

[1397] Anonymous. Book review: The General Linear Model: Data Analysis in the Social and Behavioral Sciences, by R. L. Horton. Biometrics, 43(1): 258, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531983.

Anonymous:1987:BRMa

[1398] Anonymous. Book review: Measurement and Personality Assessment, by E. E. Roskam. Biometrics, 43(1):258, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531981.

Anonymous:1987:BRMb

[1399] Anonymous. Book review: The Mathematical Theory of Quantitative Genetics, by M. G. Bulmer. Biometrics, 43(1):258, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531982.

Anonymous:1987:CNA

[1400] Anonymous. Correction: 333. Note: The analysis of non-independent 2×2 tables using rank sums. Biometrics, 43(1):259, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531986. See [2].

Anonymous:1987:FMb

[1401] Anonymous. Front matter. Biometrics, 43(2):i–iv, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531810.

Barry:1987:ADB

[1402] Daniel Barry and J. A. Hartigan. Asynchronous distance between homologous DNA sequences. Biometrics, 43(2):261–276, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531811.

Gleeson:1987:RML

[1403] Alan C. Gleeson and Brian R. Cullis. Residual maximum likelihood (REML) estimation of a neighbour model for field experiments. Biometrics, 43(2):277–287, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531812.

Liang:1987:EMH

[1404] Kung-Yee Liang. Extended Mantel–Haenszel estimating procedure for multivariate logistic regression models. Biometrics, 43(2):289–299, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531813.

Meng:1987:BAM

[1405] Cliff Y. K. Meng and Arthur P. Dempster. A Bayesian approach to the multiplicity problem for significance testing with binomial data. Biometrics, 43(2):301–311, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531814.

Hakulinen:1987:TER

[1406] Timo Hakulinen, Leena Tenkanen, Kamal Abeywickrama, and Lassi Päivärinta. Testing equality of relative survival patterns based on aggregated data. Biometrics, 43(2):313–325, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531815.

Lindsey:1987:MCA

[1407] Jane C. Lindsey, Agnes M. Herzberg, and Donald G. Watts. A method for cluster analysis based on projections and quantile-quantile plots. Biometrics, 43(2):327–341, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531816.

Jinn:1987:OTP

[1408] J. H. Jinn, J. Sedransk, and Philip Smith. Optimal two-phase stratified sampling for estimation of the age composition of a fish population. Biometrics, 43(2):343–353, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531817.

Thompson:1987:DFO

[1409] Steven K. Thompson and Fred L. Ramsey. Detectability functions in observing spatial point processes. Biometrics, 43(2):355–362, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531818.

Buckland:1987:VCP

[1410] S. T. Buckland. On the variable circular plot method of estimating animal density. Biometrics, 43(2):363–384, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531819.

Berk:1987:CIR

[1411] Kenneth Berk. Computing for incomplete repeated measures. Biometrics, 43(2):385–398, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531820.

Badger:1987:MMC

[1412] Gary J. Badger, Pamela M. Vacek, and Richard C. Reichman. A Markov model for a clinical episode of recurrent genital herpes. Biometrics, 43(2): 399–408, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531821.

Taylor:1987:KSC

[1413] Jeremy M. G. Taylor. Kendall’s and Spearman’s correlation coefficients in the presence of a blocking variable. Biometrics, 43(2):409–416, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531822.

Holland:1987:ISR

[1414] Burt S. Holland and Margaret DiPonzio Copenhaver. An improved sequentially rejective Bonferroni test procedure. Biometrics, 43(2):417– 423, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531823. See cor-

rection [1476, 1477].

Phillips:1987:EVM

[1415] Abraham Phillips and Paul W. Holland. Estimators of the variance of the Mantel–Haenszel log-odds-ratio estimate. Biometrics, 43(2):425– 431, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531824.

Jannarone:1987:EIA

[1416] Robert J. Jannarone, Caroline A. Macera, and Carol Z. Garrison. Evaluating interrater agreement through “Case-Control” sampling. Biometrics, 43(2):433–437, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531825.

Berry:1987:LTA

[1417] Donald A. Berry. Logarithmic transformations in ANOVA. Biometrics, 43(2):439–456, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531826.

Fuchs:1987:SCI

[1418] Camil Fuchs and Allan R. Sampson. Simultaneous confidence intervals for the general linear model. Biometrics, 43(2):457–469, June 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531827.

Graubard:1987:CCS

[1419] Barry I. Graubard and Edward L. Korn. Choice of column scores for testing independence in ordered 2×K contingency tables. Biometrics, 43(2): 471–476, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531828.

Fisher:1987:BRP

[1420] N. I. Fisher. Book review: Plots, Transformations, and Regression: an Introduction to Graphical Methods of Diagnostic Regression Analysis, by A. C. Atkinson. Biometrics, 43(2):477–478, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531829.

Moran:1987:BRS

[1421] M. A. Moran. Book review: Statistical Inference in Linear Models, by H. Bunke, O. Bunke. Biometrics, 43(2):478, June 1987. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531830.

Gower:1987:BRM

[1422] J. C. Gower. Book review: Multivariate Observations, by G. A. F. Seber. Biometrics, 43(2):478–479, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531831.

Kemp:1987:BRMa

[1423] A. W. Kemp. Book review: Modern Multivariate Statistical Analysis: a Graduate Course and Handbook, by M. Siotani, T. Hayakawa, Y. Fujikoshi. Biometrics, 43(2):479–480, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531832.

Dietz:1987:BRC

[1424] E. Dietz. Book review: Cluster-analyse und Diskriminanzanalyse, by G. Deischsel, H. J. Trampisch. Biometrics, 43(2):480, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531833.

Renshaw:1987:BRM

[1425] E. Renshaw. Book review: Mathematical Ecology: an Introduction. Lecture Notes in Biomathematics, by T. G. Hallam, S. A. Levin. Biometrics, 43(2):480–481, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531834.

Sutherland:1987:BRQ

[1426] R. A. Sutherland. Book review: Quantitative Genetics and Selection in Plant Breeding, by G. Wricke, W. E. Weber. Biometrics, 43(2):481– 482, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531835.

Stene:1987:BRT

[1427] J. Stene. Book review: Theoretical Epidemiology: Principles of Occurrence Research in Medicine, by O. S. Miettinen. Biometrics, 43(2):482, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531836.

Diamond:1987:BRR

[1428] I. D. Diamond. Book review: Recent Trends in Mortality Analysis, by K. G. Manton, E. Stallard. Biometrics, 43(2):483, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531837.

Smith:1987:BRN

[1429] G. L. Smith. Book review: Nutrient Adequacy: Assessment Using Food Consumption Surveys, by National Research Council. Biometrics, 43(2): 483–484, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531838.

Abakuks:1987:BRM

[1430] A. Abakuks. Book review: Markov Processes: Characterization and Convergence, by S. N. Ethier, T. G. Kurtz. Biometrics, 43(2):484, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531839.

Kemp:1987:BRMb

[1431] A. W. Kemp. Book review: Modern Elementary Statistics, with Theoretical Supplement and BASIC Programming, by D. W. Zimmerman, R. H. Williams. Biometrics, 43(2):485, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531840.

Anonymous:1987:BRBb

[1432] Anonymous. Book review: Biological Information Processing, by J. R. Sampson. Biometrics, 43(2):485–486, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531841.

Anonymous:1987:BRBc

[1433] Anonymous. Book review: On Becoming a Biologist, by J. Janovy, Jr. Biometrics, 43(2):486, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531845.

Anonymous:1987:BRBd

[1434] Anonymous. Book review: The Bicentennial Census: New Directions for Methodology in 1990, by C. F. Citro, M. L. Cohen. Biometrics, 43(2): 486, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531842.

Anonymous:1987:BRSa

[1435] Anonymous. Book review: Statistics: a Conceptual Approach, by K. L. Weldon. Biometrics, 43(2):486, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531843.

Anonymous:1987:BRU

[1436] Anonymous. Book review: The Unfinished Universe, by L. B. Young. Biometrics, 43(2):486, June 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531844.

Anonymous:1987:FMc

[1437] Anonymous. Front matter. Biometrics, 43(3):i–iv, September 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531988.

Pocock:1987:AME

[1438] Stuart J. Pocock, Nancy L. Geller, and Anastasios A. Tsiatis. The analysis of multiple endpoints in clinical trials. Biometrics, 43(3):487–498, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531989.

Chen:1987:ATW

[1439] J. J. Chen and R. L. Kodell. Analyses of two-way studies. Biometrics, 43(3):499–509, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531990.

Cox:1987:TDR

[1440] Christopher Cox. Threshold dose-response models in toxicology. Biometrics, 43(3):511–523, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531991.

Kokoska:1987:ACC

[1441] Stephen M. Kokoska. The analysis of cancer chemoprevention experiments. Biometrics, 43(3):525–534, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531992.

Lu:1987:CIR

[1442] Tai-Fang C. Lu, Franklin A. Graybill, and Richard K. Burdick. Confidence intervals on the ratio of expected mean squares (θ_{1} + dθ_{2})/θ_{3}. Biometrics, 43(3):535–543, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531993.

Khuri:1987:ETM

[1443] A. I. Khuri and R. C. Littell. Exact tests for the main effects variance components in an unbalanced random two-way model. Biometrics, 43(3):545–560, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531994.

Wallach:1987:MSE

[1444] D. Wallach and B. Goffinet. Mean squared error of prediction in models for studying ecological and agronomic systems. Biometrics, 43(3):561– 573, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531995.

Krzanowski:1987:CVP

[1445] W. J. Krzanowski. Cross-validation in principal component analysis. Biometrics, 43(3):575–584, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531996.

Thompson:1987:PSL

[1446] E. A. Thompson and T. R. Meagher. Parental and sib likelihoods in genealogy reconstruction. Biometrics, 43(3):585–600, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531997.

Smith:1987:MLE

[1447] Philip J. Smith. Maximum likelihood estimation of the distribution of length at age. Biometrics, 43(3):601–615, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531998.

Vonesh:1987:EIR

[1448] Edward F. Vonesh and Randy L. Carter. Efficient inference for randomcoefficient growth curve models with unbalanced data. Biometrics, 43 (3):617–628, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531999.

Stewart:1987:LSC

[1449] Paul W. Stewart. Line-segment confidence bands for repeated measures. Biometrics, 43(3):629–640, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532000.

Johnson:1987:TSR

[1450] Richard A. Johnson, Steve Verrill, and Dan H. Moore, II. Two-sample rank tests for detecting changes that occur in a small proportion of the treated population. Biometrics, 43(3):641–655, September 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532001.

Brookmeyer:1987:TSM

[1451] Ron Brookmeyer and N. E. Day. Two-stage models for the analysis of cancer screening data. Biometrics, 43(3):657–669, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532002.

Clayton:1987:EBE

[1452] David Clayton and John Kaldor. Empirical Bayes estimates of agestandardized relative risks for use in disease mapping. Biometrics, 43 (3):671–681, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532003.

Geurts:1987:SSP

[1453] Johannes H. J. Geurts. On the small-sample performance of Efron’s and of Gill’s version of the product limit estimator under nonproportional hazards. Biometrics, 43(3):683–692, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532004.

Kemp:1987:FDD

[1454] A. W. Kemp. Families of discrete distributions satisfying Taylor’s power law. Biometrics, 43(3):693–699, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532005.

Nam:1987:SAC

[1455] Jun mo Nam. A simple approximation for calculating sample sizes for detecting linear trend in proportions. Biometrics, 43(3):701–705, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532006.

Kimball:1987:TTO

[1456] A. W. Kimball. A test for a threshold in an ordered sequence of correlated proportions. Biometrics, 43(3):707–712, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532007.

Ciminera:1987:SET

[1457] J. L. Ciminera, J. A. Bolognese, and M. H. Gregg. The statistical evaluation of a three-period two-treatment crossover pharmacokinetic drug interaction study. Biometrics, 43(3):713–718, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532008.

Millard:1987:PSV

[1458] Steven P. Millard and Irwin D. Bross. Proof of safety vs proof of hazard. Biometrics, 43(3):719–725, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532009.

Liao:1987:CMV

[1459] Jiangang Liao, Alvaro Munoz, and Bernard Rosner. On Cornfield’s mean and variance for the hypergeometric distribution. Biometrics, 43(3):727, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532010.

Copas:1987:BRA

[1460] J. B. Copas. Book review: Applied Linear Statistical Models, 2nd Edition, by J. Neter, W. Wasserman, M. H. Kutner. Biometrics, 43(3):729–730, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532011.

Campbell:1987:BRM

[1461] M. J. Campbell. Book review: Mathematical Methods in Medicine,Part 1: Statistical and Analytical Techniques;Part 2: Applications in Clinical Specialities, by D. Ingram, R. Bloch. Biometrics, 43(3):730, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532012.

Cormack:1987:BRM

[1462] R. M. Cormack. Book review: Multivariate Analysis of Ecological Communities, by P. G. N. Digby, R. A. Kempton. Biometrics, 43(3):730–731, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532013.

Holgate:1987:BRP

[1463] P. Holgate. Book review: The Predator–Prey Model: Do We Live in a Volterra World?, by M. Peschel, W. Mende. Biometrics, 43(3):731–732, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532014.

Berger:1987:BRP

[1464] J. Berger. Book review: Population Dynamics of Rabies in Wildlife, by P. J. Bacon. Biometrics, 43(3):732, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532015.

Cohen:1987:BRT

[1465] G. R. Cohen. Book reviews: Technical Demography, by R. Ramakumar; Deterministic Aspects of Mathematical Demography, by J. Impagliazzo. Biometrics, 43(3):732–733, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532016.

North:1987:BRS

[1466] P. M. North. Book review: Statistics for Ornithologists, by J. Fowler, L. Cohen. Biometrics, 43(3):733–734, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532017.

Anonymous:1987:BRGb

[1467] Anonymous. Book review: A Guide to Statistical Methods and to the Pertinent Literature: Literatur zur Angewandten Statistik, by L. A. Sachs. Biometrics, 43(3):734, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532019.

Clark:1987:BRP

[1468] G. M. Clark. Book review: Practical Statistics for Experimental Biologists, by A. C. Wardlaw. Biometrics, 43(3):734, September 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532018.

Anonymous:1987:BRSb

[1469] Anonymous. Book review: The Sequential Statistical Analysis of Hypothesis Testing, Point and Interval Estimation, and Decision Theory, by Z. Govindarajulu. Biometrics, 43(3):734–735, September 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532020.

Anonymous:1987:BRH

[1470] Anonymous. Book review: Handbook of Statistics, Vol. 4. Nonparametric Methods, by P. R. Krishnaiah, P. K. Sen. Biometrics, 43(3):735, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532021.

Anonymous:1987:BRO

[1471] Anonymous. Book review: Optimal Use of the Clinical Laboratory, by O. Zinder. Biometrics, 43(3):735, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532025.

Anonymous:1987:BRP

[1472] Anonymous. Book review: Principles and Practice of Statistics in Medicine, by R. N. Srivastava, B. L. Verma, G. D. Shukla. Biometrics, 43(3):735, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532022.

Anonymous:1987:BRQ

[1473] Anonymous. Book review: Quotes, Damned Quotes, and …, 2nd Edition, by J. Bibby. Biometrics, 43(3):735, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532026.

Anonymous:1987:BRSc

[1474] Anonymous. Book review: Statistische Analyse von Zeitereihen, by J. Andăl. Biometrics, 43(3):735, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532024.

Anonymous:1987:BRT

[1475] Anonymous. Book review: Les Théories Mathématiques des Populations, by A. Hillion. Biometrics, 43(3):735, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532023.

Anonymous:1987:CIS

[1476] Anonymous. Correction: An improved sequentially rejective Bonferroni test procedure. Biometrics, 43(3):737, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532027. See [1414].

Holland:1987:CIS

[1477] B. S. Holland and M. D. Copenhaver. Correction to: “An improved sequentially rejective Bonferroni test procedure”. Biometrics, 43(3):737, September 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1414].

Anonymous:1987:FMd

[1478] Anonymous. Front matter. Biometrics, 43(4):i–iv, December 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531528.

Anonymous:1987:VI

[1479] Anonymous. Volume information. Biometrics, 43(4):xiii–xxviii, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531527.

Brookmeyer:1987:BPC

[1480] Ron Brookmeyer and Mitchell H. Gail. Biases in prevalent cohorts. Biometrics, 43(4):739–749, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531529.

Wickramaratne:1987:CES

[1481] Priya J. Wickramaratne and Theodore R. Holford. Confounding in epidemiologic studies: The adequacy of the control group as a measure of confounding. Biometrics, 43(4):751–765, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531530. See correction [1843, 1847].

Catchpole:1987:MAI

[1482] E. A. Catchpole, B. J. T. Morgan, and D. J. O’Dowd. Modelling ant– insect interactions on flowerheads of Helichrysum bracteatum. Biometrics, 43(4):767–782, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531531.

Chao:1987:EPS

[1483] Anne Chao. Estimating the population size for capture–recapture data with unequal catchability. Biometrics, 43(4):783–791, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531532.

Raaijmakers:1987:SAM

[1484] Jeroen G. W. Raaijmakers. Statistical analysis of the Michaelis–Menten equation. Biometrics, 43(4):793–803, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531533.

Louis:1987:ETH

[1485] Edward J. Louis and Everett R. Dempster. An exact test for Hardy– Weinberg and multiple alleles. Biometrics, 43(4):805–811, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531534.

Tamura:1987:SME

[1486] Roy N. Tamura and S. Stanley Young. A stabilized moment estimator for the beta-binomial distribution. Biometrics, 43(4):813–824, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531535. See corrections [1906].

Sanathanan:1987:TLM

[1487] L. P. Sanathanan, E. T. Gade, and N. L. Shipkowitz. Trimmed logit method for estimating the ED50 in quantal bioassay. Biometrics, 43 (4):825–832, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531536.

Kunert:1987:VEC

[1488] Joachim Kunert. On variance estimation in crossover designs. Biometrics, 43(4):833–845, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531537.

Racine-Poon:1987:TSP

[1489] A. Racine-Poon, A. P. Grieve, H. Fluhler, and A. F. M. Smith. A twostage procedure for bioequivalence studies. Biometrics, 43(4):847–856, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531538.

Kim:1987:CIF

[1490] Kyungmann Kim and David L. DeMets. Confidence intervals following group sequential tests in clinical trials. Biometrics, 43(4):857–864, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531539.

Chang:1987:DGS

[1491] Myron N. Chang, Terry M. Therneau, Harry S. Wieand, and Stephan S. Cha. Designs for group sequential Phase II clinical trials. Biometrics, 43(4):865–874, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531540.

Gross:1987:SSD

[1492] Alan J. Gross, Hurshell H. Hunt, Alan B. Cantor, and B. Christine Clark. Sample size determination in clinical trials with an emphasis on exponentially distributed responses. Biometrics, 43(4):875–883, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531541.

Matts:1987:REZ

[1493] John Matts and Richard McHugh. Randomization and efficiency in Zelen’s single-consent design. Biometrics, 43(4):885–894, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531542.

Crager:1987:ACP

[1494] Michael R. Crager. Analysis of covariance in parallel-group clinical trials with pretreatment baselines. Biometrics, 43(4):895–901, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531543.

Morgan:1987:NED

[1495] Timothy M. Morgan. Nonparametric estimation of duration of accrual and total study length for clinical trials. Biometrics, 43(4):903–912, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531544.

Edwards:1987:ESB

[1496] Don Edwards and Jack J. Berry. The efficiency of simulation-based multiple comparisons. Biometrics, 43(4):913–928, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531545.

Carriquiry:1987:MMA

[1497] Alicia L. Carriquiry, Daniel Gianola, and Rohan L. Fernando. Mixedmodel analysis of a censored normal distribution with reference to animal breeding. Biometrics, 43(4):929–939, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531546. See correction [1800].

Beal:1987:ACI

[1498] S. L. Beal. Asymptotic confidence intervals for the difference between two binomial parameters for use with small samples. Biometrics, 43(4):941– 950, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531547.

Bonney:1987:LRD

[1499] George Ebow Bonney. Logistic regression for dependent binary observations. Biometrics, 43(4):951–973, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531548.

Grayson:1987:SDI

[1500] D. A. Grayson. Statistical diagnosis and the influence of diagnostic error. Biometrics, 43(4):975–984, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531549.

Bonett:1987:ESC

[1501] Douglas G. Bonett, J. Arthur Woodward, and P. M. Bentler. Estimating the size of closed populations using inverse multiple-recapture sampling. Biometrics, 43(4):985–991, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531550.

Bedrick:1987:FCI

[1502] Edward J. Bedrick. A family of confidence intervals for the ratio of two binomial proportions. Biometrics, 43(4):993–998, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531551.

Streit:1987:TEI

[1503] Franz Streit. Testing for effective intervention. Biometrics, 43(4):999– 1000, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531552.

Espeland:1987:GAA

[1504] Mark A. Espeland and Siu L. Hui. A general approach to analyzing epidemiologic data that contain misclassification errors. Biometrics, 43 (4):1001–1012, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531553.

Williams:1987:DRM

[1505] D. A. Williams and Kamta Rai. Dose-response models for teratological experiments. Biometrics, 43(4):1013–1016, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531554.

Brownie:1987:RMM

[1506] Cavell Brownie, Jean Clobert, and Jean-Dominique Lebreton. Recent models for mark–recapture and mark-resighting data. Biometrics, 43 (4):1017–1022, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531555.

James:1987:EF

[1507] William H. James. Estimating fecundability. Biometrics, 43(4):1023– 1024, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531556.

John:1987:BRB

[1508] J. A. John. Book review: Beyond ANOVA, Basics of Applied Statistics, by R. G. Miller. Biometrics, 43(4):1025–1026, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531557.

Lachenbruch:1987:BRM

[1509] P. A. Lachenbruch. Book review: Multivariate Statistical Methods: a Primer, by B. F. J. Manly. Biometrics, 43(4):1026, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531558.

Freeman:1987:BRS

[1510] G. H. Freeman. Book review: Statistical Design: Theory and Practice, by C. E. McCulloch, S. J. Schwager, G. Casella, S. R. Searle. Biometrics, 43(4):1026–1027, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531559.

Kemp:1987:BRH

[1511] A. W. Kemp. Book reviews: The History of Statistics: The Measurement of Uncertainty Before 1900, by S. M. Stigler; The Rise of Statistical Thinking 1820–1900, by T. M. Porter. Biometrics, 43(4):1027–1028, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531560.

Armitage:1987:BRD

[1512] P. Armitage. Book review: The Design and Analysis of Clinical Experiments, by J. L. Fleiss. Biometrics, 43(4):1028, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531561.

Faddy:1987:BRC

[1513] M. J. Faddy. Book review: Compartmental Analysis in Biology and Medicine, 2nd edition, by J. A. Jacquez. Biometrics, 43(4):1028–1029, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531562.

Fluhler:1987:BRB

[1514] H. Flühler. Book review: Biometrie in der Chemisch-Pharmazeutischen Industrie 2: Biometrische Analyse von Kombinationswirkungen, by J. Vollmar, H. D. Unkelbach. Biometrics, 43(4):1029, December 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531563.

Fluhler:1987:BRQ

[1515] H. Flühler. Book review: Qualitative Dosis–Wirkungs–Analysen, by H. D. Unkelbach, T. Wolf. Biometrics, 43(4):1030, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531564.

Rutherford:1987:BRA

[1516] A. A. Rutherford. Book review: Adaptive Management of Renewable Resources, by C. Walters. Biometrics, 43(4):1030, December 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531565.

Anonymous:1987:BRAc

[1517] Anonymous. Book review: Análisis Multivariado: Método de Componentes Principales, by L. E. Pla. Biometrics, 43(4):1031, December 1987.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531567.

Anonymous:1987:BRAd

[1518] Anonymous. Book reviews: Advances in the Statistical Sciences, Vol. 1. Applied Probability, Stochastic Processes, and Sampling Theory, by I. B. MacNeill, G. J. Umphrey; Advances in the Statistical Sciences, Vol. 5. Biostatistics, by I. B. MacNeill, G. J. Umphrey. Biometrics, 43(4):1031, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531566.

Anonymous:1987:BRBe

[1519] Anonymous. Book review: The Bellman Continuum, by R. S. Roth. Biometrics, 43(4):1031, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531569.

Anonymous:1987:BRE

[1520] Anonymous. Book review: Expert Systems in Statistics, by R. Haux. Biometrics, 43(4):1031, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531568.

Anonymous:1987:BRN

[1521] Anonymous. Book review: Notes Towards a History of Teaching Statistics, by J. Bibby. Biometrics, 43(4):1031–1032, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531570.

Anonymous:1987:BRAe

[1522] Anonymous. Book review: Alpha Mathematics Handbook, by L. Rade. Biometrics, 43(4):1032, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531571.

Anonymous:1987:BRMc

[1523] Anonymous. Book review: Mathematical Programming for Economic Analysis in Agriculture, by P. B. R. Hazell, R. D. Norton. Biometrics, 43(4):1032, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531573.

Anonymous:1987:BRSd

[1524] Anonymous. Book reviews: Selected Tables in Mathematical Statistics, Vol. 7, by Institute of Mathematical Statistics; Selected Tables in Mathematical Statistics, Vol. 10, by Institute of Mathematical Statistics. Biometrics, 43(4):1032, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531572.

Anonymous:1987:BRW

[1525] Anonymous. Book reviews: World Health Statistics Annual 1986, by World Health Organization; World Health Statistics Quarterly 1986, Vol. 39 (3), (4), 1987, Vol. 40 (1), (2), by World Health Organization. Biometrics, 43(4):1033, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531574.

Anonymous:1987:CCC

[1526] Anonymous. Corrections: A comparison of confidence interval methods for the intraclass correction coefficient. Biometrics, 43(4):1035, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531575. See [1240, 1351].

Anonymous:1987:CIE

[1527] Anonymous. Corrections: Interval estimation of the median lethal dose. Biometrics, 43(4):1035, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531576. See [1278].

Donner:1987:CCC

[1528] A. Donner and G. Wells. Correction to: “A comparison of confidence interval methods for the intraclass correlation coefficient” [Biometrics 42 (1986), no. 2, 401–412]. Biometrics, 43(4):1035, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1240].

Williams:1987:CIE

[1529] D. A. Williams. Correction to: “Interval estimation of the median lethal dose” [Biometrics 42 (1986), no. 3, 641–645]. Biometrics, 43(4):1035, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1278].

Anonymous:1987:BSF

[1530] Anonymous. The Biometric Society financial statements and auditors’ report January 31, 1987. Biometrics, 43(4):1037–1039, December 1987.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531577.

Anonymous:1988:FMa

[1531] Anonymous. Front matter. Biometrics, 44(1):i–iv, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531891.

Rice:1988:NPM

[1532] William R. Rice. A new probability model for determining exact P - values for 2×2 contingency tables when comparing binomial proportions. Biometrics, 44(1):1–22, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531892.

Krzanowski:1988:CDN

[1533] W. J. Krzanowski and Y. T. Lai. A criterion for determining the number of groups in a data set using sum-of-squares clustering. Biometrics, 44(1): 23–34, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531893.

Segal:1988:RTC

[1534] Mark Robert Segal. Regression trees for censored data. Biometrics, 44(1):35–47, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531894.

Green:1988:SBG

[1535] D. G. Green, R. E. Reichelt, and R. H. Bradbury. Statistical behaviour of the GMDH algorithm. Biometrics, 44(1):49–69, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531895.

Lee:1988:API

[1536] Andy H. Lee. Assessing partial influence in generalized linear models. Biometrics, 44(1):71–77, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531896.

Self:1988:PSS

[1537] Steven G. Self and Robert H. Mauritsen. Power/sample size calculations for generalized linear models. Biometrics, 44(1):79–86, March 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531897.

Buonaccorsi:1988:BIR

[1538] John P. Buonaccorsi and Constantine A. Gatsonis. Bayesian inference for ratios of coefficients in a linear model. Biometrics, 44(1):87–101, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531898.

Berger:1988:TCS

[1539] Roger L. Berger, Dennis D. Boos, and Frank M. Guess. Tests and confidence sets for comparing two mean residual life functions. Biometrics, 44(1):103–115, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531899.

Kremers:1988:ESR

[1540] Walter K. Kremers. Estimation of survival rates from a mark–recapture study with tag loss. Biometrics, 44(1):117–130, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531900.

Lyons:1988:DPN

[1541] N. I. Lyons and K. Hutcheson. Distributional properties of the number of moves index of diversity. Biometrics, 44(1):131–140, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531901.

Srivastava:1988:EII

[1542] M. S. Srivastava, K. J. Keen, and R. S. Katapa. Estimation of interclass and intraclass correlations in multivariate familial data. Biometrics, 44(1):141–150, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531902.

Gail:1988:EPA

[1543] Mitchell H. Gail. The effect of pooling across strata in perfectly balanced studies. Biometrics, 44(1):151–162, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531903.

Haber:1988:MSA

[1544] Michael Haber, Ira M. Longini, Jr., and George A. Cotsonis. Models for the statistical analysis of infectious disease data. Biometrics, 44(1):163– 173, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531904.

Wu:1988:ECC

[1545] Margaret C. Wu and Raymond J. Carroll. Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring process. Biometrics, 44(1):175–188, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531905. See corrections [1896, 1905].

Conover:1988:LMP

[1546] William J. Conover and David S. Salsburg. Locally most powerful tests for detecting treatment effects when only a subset of patients can be expected to “Respond” to treatment. Biometrics, 44(1):189–196, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531906.

Thall:1988:MPL

[1547] Peter F. Thall. Mixed Poisson likelihood regression models for longitudinal interval count data. Biometrics, 44(1):197–209, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531907. See correction [1844, 1846].

Willan:1988:UMT

[1548] Andrew R. Willan. Using the maximum test statistic in the two-period crossover clinical trial. Biometrics, 44(1):211–218, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531908.

Berry:1988:OSS

[1549] Donald A. Berry and Chih-Hsiang Ho. One-sided sequential stopping boundaries for clinical trials: a decision- theoretic approach. Biometrics, 44(1):219–227, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531909.

Lakatos:1988:SSB

[1550] Edward Lakatos. Sample sizes based on the log-rank statistic in complex clinical trials. Biometrics, 44(1):229–241, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531910. See corrections [1670, 1671].

Thomas:1988:ACV

[1551] D. Roland Thomas and J. N. K. Rao. On the analysis of categorical variables by linear models having singular covariance matrices. Biometrics, 44(1):243–248, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531911.

George:1988:NEU

[1552] V. T. George and R. C. Elston. A note on estimation using self-contained subsets. Biometrics, 44(1):249–252, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531912.

Dallal:1988:PBT

[1553] Gerard E. Dallal. Paired Bernoulli trials. Biometrics, 44(1):253–257, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531913.

Worsley:1988:EPP

[1554] K. J. Worsley. Exact percentage points of the likelihood-ratio test for a change-point hazard-rate model. Biometrics, 44(1):259–263, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531914.

Slud:1988:HDC

[1555] E. Slud and D. Byar. How dependent causes of death can make risk factors appear protective. Biometrics, 44(1):265–269, March 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531915.

Hall:1988:CIS

[1556] Peter Hall. On confidence intervals for spatial parameters estimated from nonreplicated data. Biometrics, 44(1):271–277, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531916.

Barker:1988:PGM

[1557] F. Barker, Y. C. Soh, and R. J. Evans. Properties of the geometric mean functional relationship. Biometrics, 44(1):279–281, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531917.

Glasbey:1988:ECS

[1558] C. A. Glasbey. Effects of conditioning, swath structure, and weather on the drying rate of mown grass. Biometrics, 44(1):283–288, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531918.

Gill:1988:RMS

[1559] J. L. Gill. Repeated measurement: Split-plot trend analysis versus analysis of first differences. Biometrics, 44(1):289–297, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531919.

Le:1988:TLT

[1560] Chap T. Le. Testing for linear trends in proportions using correlated otolaryngology or ophthalmology data. Biometrics, 44(1):299–303, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531920.

Williams:1988:EBU

[1561] D. A. Williams, Lawrence L. Kupper, Christopher Portier, and Michael D. Hogan. Estimation bias using the beta-binomial distribution in teratology. Biometrics, 44(1):305–309, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531921.

Edland:1988:SAV

[1562] Steven D. Edland and Ronaldo Iachan. Stratification by an auxiliary variable. Biometrics, 44(1):311–312, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531922.

Kemp:1988:BRK

[1563] C. D. Kemp. Book review: Kendall’s Advanced Theory of Statistics, Volume 1, Distribution Theory, by A. Stuart, J. K. Ord. Biometrics, 44(1): 313–314, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531923.

Shotter:1988:BRE

[1564] M. V. Shotter. Book review: Essentials of Biostatistics, by R. C. Elston, W. D. Johnson. Biometrics, 44(1):314–315, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531924.

Corsten:1988:BRS

[1565] L. C. A. Corsten. Book review: Statistische Methoden der Modellbildung III, by K. M. Humak. Biometrics, 44(1):315, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531925.

Upton:1988:BRE

[1566] G. J. G. Upton. Book review: The Elements of Graphing Data, by W. S. Cleveland. Biometrics, 44(1):315–316, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531926.

Spurr:1988:BRG

[1567] B. D. Spurr. Book review: Goodness-of-Fit Techniques, by R. B. D’Agostino, M. A. Stephens. Biometrics, 44(1):316, March 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531927.

Weir:1988:BRE

[1568] B. S. Weir. Book review: Evolutionary Processes and Theory, by S. Karlin, E. Nevo. Biometrics, 44(1):316–317, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531928.

Anonymous:1988:BRB

[1569] Anonymous. Book review: Biostatistics: a Foundation for Analysis in the Health Sciences, by W. W. Daniel. Biometrics, 44(1):317–318, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531929.

Anonymous:1988:BRAa

[1570] Anonymous. Book review: Applied Statistics: Analysis of Variance and Regression, by O. J. Dunn, V. A. Clark. Biometrics, 44(1):318, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531930.

Anonymous:1988:BRH

[1571] Anonymous. Book review: Handbook of Introductory Statistical Methods, by C. P. Cox. Biometrics, 44(1):318, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531931.

Anonymous:1988:BRMa

[1572] Anonymous. Book review: Modern Statistical Selection, Parts I and II by M. H. Rizvi. Biometrics, 44(1):318, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531932.

Anonymous:1988:BRTa

[1573] Anonymous. Book review: Theory and Analysis of Sample Survey Designs, by D. Singh, F. S. Chaudhary. Biometrics, 44(1):318–319, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531933.

Anonymous:1988:BRI

[1574] Anonymous. Book review: Italian Contributions to the Methodology of Statistics, by A. Naddeo. Biometrics, 44(1):319, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531936.

Anonymous:1988:BRSa

[1575] Anonymous. Book review: State and Metropolitan Area Data Book, by U. S. Department of Commerce. Biometrics, 44(1):319, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531937.

Anonymous:1988:BRSb

[1576] Anonymous. Book review: Statistical Methods for Environmental Pollution Monitoring, by R. O. Gilbert. Biometrics, 44(1):319, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531935.

Anonymous:1988:BRTb

[1577] Anonymous. Book review: Taxonomic Analysis in Biology, by L. A. Abbott, F. A. Bisby, D. J. Rogers. Biometrics, 44(1):319, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531934.

Anonymous:1988:CEH

[1578] Anonymous. Corrections: Estimation of heritability from experiments with inbred and related individuals. Biometrics, 44(1):321, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531938. See [1].

Anonymous:1988:CFN

[1579] Anonymous. Corrections: Fitting the negative binomial distribution. Biometrics, 44(1):321, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531940. See [1332].

Anonymous:1988:CMB

[1580] Anonymous. Corrections: Multivariate bioassay, combination of bioassays, and Fieller’s theorem. Biometrics, 44(1):321, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531939. See [1196].

Srivastava:1988:CMB

[1581] M. S. Srivastava. Correction to: “Multivariate bioassay, combination of bioassays, and Fieller’s theorem” [Biometrics 42 (1986), no. 1, 131–141; MR0845630 (87g:62161)]. Biometrics, 44(1):321, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1196].

Anonymous:1988:FMb

[1582] Anonymous. Front matter. Biometrics, 44(2):i–iv, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531847.

Gart:1988:AIE

[1583] John J. Gart and Jun mo Nam. Approximate interval estimation of the ratio of binomial parameters: a review and corrections for skewness. Biometrics, 44(2):323–338, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531848.

OSullivan:1988:DEH

[1584] Finbarr O’Sullivan and Janet O’Sullivan. Deconvolution of episodic hormone data: an analysis of the role of season on the onset of puberty in cows. Biometrics, 44(2):339–353, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531849.

Kuritz:1988:ARE

[1585] Stephen J. Kuritz and J. Richard Landis. Attributable risk estimation from matched case-control data. Biometrics, 44(2):355–367, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531850.

Donner:1988:ECO

[1586] Allan Donner and Walter Hauck. Estimation of a common odds ratio in case-control studies of familial aggregation. Biometrics, 44(2):369– 378, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531851.

Hauck:1988:ARE

[1587] Walter W. Hauck and Allan Donner. The asymptotic relative efficiency of the Mantel–Haenszel estimator in the increasing-number-of-strata case. Biometrics, 44(2):379–384, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531852.

Yandell:1988:MIN

[1588] Brian S. Yandell and David B. Hogg. Modelling insect natality using splines. Biometrics, 44(2):385–395, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531853.

Pollock:1988:DFN

[1589] K. H. Pollock and W. L. Cornelius. A distribution-free nest survival model. Biometrics, 44(2):397–404, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531854.

Dinse:1988:ETI

[1590] Gregg E. Dinse. Estimating tumor incidence rates in animal carcinogenicity experiments. Biometrics, 44(2):405–415, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531855.

Bailer:1988:ETI

[1591] A. John Bailer and Christopher J. Portier. Effects of treatment-induced mortality and tumor-induced mortality on tests for carcinogenicity in small samples. Biometrics, 44(2):417–431, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531856.

Yuh:1988:AR

[1592] Lianng Yuh and Robert V. Hogg. On adaptive M-regression. Biometrics, 44(2):433–445, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531857.

Heltshe:1988:JEM

[1593] James F. Heltshe. Jackknife estimate of the matching coefficient of similarity. Biometrics, 44(2):447–460, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2531858. See correction [1898].

Laska:1988:ESP

[1594] Eugene M. Laska, Morris Meisner, and Carole Siegel. Estimating the size of a population from a single sample. Biometrics, 44(2):461–472, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531859. See correction [1897].

Binns:1988:BCS

[1595] M. R. Binns and N. J. Bostanian. Binomial and censored sampling in estimation and decision making for the negative binomial distribution. Biometrics, 44(2):473–483, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531860.

Smith:1988:EED

[1596] Stephen J. Smith. Evaluating the efficiency of the Δ-distribution mean estimator. Biometrics, 44(2):485–493, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531861.

Papageorgiou:1988:EPB

[1597] H. Papageorgiou and S. Loukas. On estimating the parameters of a bivariate probability model applicable to traffic accidents. Biometrics, 44(2): 495–504, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531862.

Rosner:1988:STC

[1598] Bernard Rosner and Roy C. Milton. Significance testing for correlated binary outcome data. Biometrics, 44(2):505–512, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531863.

Lachin:1988:ETA

[1599] John M. Lachin and L. J. Wei. Estimators and tests in the analysis of multiple nonindependent 2×2 tables with partially missing observations. Biometrics, 44(2):513–528, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2531864. See corrections [1669].

Klein:1988:BNS

[1600] John P. Klein and M. L. Moeschberger. Bounds on net survival probabilities for dependent competing risks. Biometrics, 44(2):529–538, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531865.

Agresti:1988:MAB

[1601] Alan Agresti. A model for agreement between ratings on an ordinal scale. Biometrics, 44(2):539–548, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531866.

Starmer:1988:CAD

[1602] C. F. Starmer. Characterizing activity-dependent processes with a piecewise exponential model. Biometrics, 44(2):549–559, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531867.

Singh:1988:SGI

[1603] Murari Singh, B. Gilliver, and M. R. Rao. Stability of genotypes in intercropping. Biometrics, 44(2):561–570, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531868.

McLachlan:1988:FMM

[1604] G. J. McLachlan and P. N. Jones. Fitting mixture models to grouped and truncated data via the EM algorithm. Biometrics, 44(2):571–578, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531869.

Lan:1988:VTM

[1605] K. K. Gordon Lan and Janet Wittes. The B-value: a tool for monitoring data. Biometrics, 44(2):579–585, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531870.

Faddy:1988:FTD

[1606] M. J. Faddy and M. C. Jones. Fitting time-dependent multicompartment models: a case study. Biometrics, 44(2):587–593, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531871.

Piegorsch:1988:ESI

[1607] Walter W. Piegorsch, Clarice R. Weinberg, and Barry H. Margolin. Exploring simple independent action in multifactor tables of proportions.

Biometrics, 44(2):595–603, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531872.

Fuchs:1988:EAM

[1608] Camil Fuchs and Joel B. Greenhouse. The EM algorithm for maximum likelihood estimation in the mover-stayer model. Biometrics, 44(2):605– 613, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531873.

Brownie:1988:ECD

[1609] Cavell Brownie, Jeffrey S. Simonoff, Yosef Hochberg, and Benjamin Reiser. Estimating P r(X < Y) in categorized data using “ROC” analysis. Biometrics, 44(2):615–621, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531874.

Mantel:1988:ELR

[1610] Nathan Mantel. Exact limits on the ratio or difference of two independent binomial proportions? Biometrics, 44(2):623, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531875.

Bailey:1988:BRSa

[1611] R. A. Bailey. Book review: Statistical Design for Research, by L. Kish. Biometrics, 44(2):625–626, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531876.

Atkinson:1988:BRR

[1612] A. C. Atkinson. Book reviews: Robust Statistics, by R. F. Hampel, E. M. Ronchetti, P. J. Rousseeuw, W. A. Stahel; Robust Regression and Outlier Detection, by P. J. Rousseeuw, A. M. Leroy. Biometrics, 44(2): 626–627, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531877.

Kemp:1988:BRR

[1613] A. W. Kemp. Book review: Robust Inference, by M. L. Tiku, W. Y. Tan, N. Balakrishnan. Biometrics, 44(2):627–628, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531878.

Morgan:1988:BRS

[1614] B. J. T. Morgan. Book review: Stochastic Simulation, by B. D. Ripley. Biometrics, 44(2):628–629, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531879.

Jupp:1988:BRS

[1615] P. E. Jupp. Book review: Statistical Analysis of Spherical Data, by N. I. Fisher, T. Lewis, B. J. J. Embleton. Biometrics, 44(2):629–630, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531880.

Spiegelhalter:1988:BRAa

[1616] D. J. Spiegelhalter. Book reviews: Artificial Intelligence Memoranda 1958–1979: MIT AI Laboratory, by Comtex Scientific Corporation (Scientific Datalink).; Artificial Intelligence Technical Reports 1968–1983: SRI International Artificial Intelligence Center. Biometrics, 44(2):630, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531881.

Spiegelhalter:1988:BRAb

[1617] D. J. Spiegelhalter. Book review: Artificial Intelligence Technical Reports 1968–1983: SRI International Artificial Intelligence Center, by Comtex Scientific Corporation (Scientific Data Link). Biometrics, 44(2):630, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531882.

Anonymous:1988:BRF

[1618] Anonymous. Book review: Foundations of Statistics, by D. G. Rees. Biometrics, 44(2):631, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531884.

Anonymous:1988:BRSc

[1619] Anonymous. Book review: Small Area Statistics, by R. Platek, J. K. N. Rao, C. E. Sarndal, M. P. Singh. Biometrics, 44(2):631, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531885.

Anonymous:1988:BRSd

[1620] Anonymous. Book review: Statistics Explained: Basic Concepts and Methods, by R. Kapadia, G. Andersson. Biometrics, 44(2):631, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531883.

Anonymous:1988:BRU

[1621] Anonymous. Book review: Unimodal Models to Relate Species to Environment, by C. J. F. Braak. Biometrics, 44(2):631–632, June 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531886.

Anonymous:1988:BRMb

[1622] Anonymous. Book review: Mathematical Modeling and digital Simulation for Engineers and Scientists, by J. M. Smith. Biometrics, 44(2):632, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531888.

Anonymous:1988:BRSe

[1623] Anonymous. Book review: Über Statistische Methoden bei Punktprozessen, by F. Boker. Biometrics, 44(2):632, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531887.

Anonymous:1988:CAD

[1624] Anonymous. Correction: Analysis of data from incomplete block designs. Biometrics, 44(2):633, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531889. See [1244].

Giesbrecht:1988:CAD

[1625] F. G. Giesbrecht. Correction to: “Analysis of data from incomplete block designs” [Biometrics 42 (1986), no. 2, 437–448]. Biometrics, 44(2): 633, June 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1244].

Anonymous:1988:FMc

[1626] Anonymous. Front matter. Biometrics, 44(3):i–iv, September 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531579.

Nedelman:1988:PMG

[1627] Jerry Nedelman. The prevalence of malaria in Garki, Nigeria: double sampling with a fallible expert. Biometrics, 44(3):635–655, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531580.

Kitsos:1988:ODP

[1628] C. P. Kitsos, D. M. Titterington, and B. Torsney. An optimal design problem in rhythmometry. Biometrics, 44(3):657–671, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531581.

Mentre:1988:HEU

[1629] France Mentré, Alain Mallet, and Jean-Louis Steimer. Hyperparameter estimation using stochastic approximation with application to population pharmacokinetics. Biometrics, 44(3):673–683, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531582.

Ball:1988:UAV

[1630] Frank Ball and Peter Donnelly. A unified approach to variability in compartmental models. Biometrics, 44(3):685–694, September 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531583.

Cameron:1988:TCE

[1631] M. A. Cameron, G. K. Eagleson, M. E. Willcox, D. G. Laing, and H. Panhuber. Testing for correlated errors in the analysis of variance. Biometrics, 44(3):695–704, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531584.

Gauch:1988:MSV

[1632] Hugh G. Gauch, Jr. Model selection and validation for yield trials with interaction. Biometrics, 44(3):705–715, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531585.

Wette:1988:PML

[1633] Reimut Wette, Matthew K. McGue, Dabeeru C. Rao, and C. Robert Cloninge. On the properties of maximum likelihood estimators of familial correlations under variable sibship size. Biometrics, 44(3):717–725, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531586.

Donner:1988:CIC

[1634] Allan Donner and Michael Eliasziw. Confidence interval construction for parent-offspring correlations. Biometrics, 44(3):727–737, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531587.

Piegorsch:1988:CBL

[1635] Walter W. Piegorsch and George Casella. Confidence bands for logistic regression with restricted predictor variables. Biometrics, 44(3):739–750, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531588.

Ko:1988:DMU

[1636] Daijin Ko and Judith E. Zeh. Detection of migration using sound location. Biometrics, 44(3):751–763, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531589.

Schwarz:1988:PRS

[1637] Carl J. Schwarz, Kenneth P. Burnham, and A. Neil Arnason. Postrelease stratification in band-recovery models. Biometrics, 44(3):765– 785, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531590.

Coffey:1988:REM

[1638] Mary Coffey. A random effects model for binary data from dependent samples. Biometrics, 44(3):787–801, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www. jstor.org/stable/2531591. See corrections [1903, 1904].

Hirji:1988:EIM

[1639] Karim F. Hirji, Cyrus R. Mehta, and Nitin R. Patel. Exact inference for matched case-control studies. Biometrics, 44(3):803–814, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531592.

Kalish:1988:EBT

[1640] Leslie A. Kalish and David P. Harrington. Efficiency of balanced treatment allocation for survival analysis. Biometrics, 44(3):815–821, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531593.

Sylvester:1988:BAD

[1641] Richard J. Sylvester. A Bayesian approach to the design of Phase II clinical trials. Biometrics, 44(3):823–836, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531594.

DeLong:1988:CAU

[1642] Elizabeth R. DeLong, David M. DeLong, and Daniel L. Clarke-Pearson. Comparing the areas under two or more correlated receiver operating characteristic curves: a nonparametric approach. Biometrics, 44(3):837– 845, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531595.

Collings:1988:EPT

[1643] Bruce Jay Collings and Martin A. Hamilton. Estimating the power of the two-sample Wilcoxon test for location shift. Biometrics, 44(3):847–860, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531596.

Williams:1988:ECI

[1644] D. A. Williams. An exact confidence interval for the relative potency estimated from a multivariate bioassay. Biometrics, 44(3):861–867, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531597.

Gart:1988:ETT

[1645] John J. Gart and Jun mo Nam. The equivalence of two tests and models for HLA data with no observed double blanks. Biometrics, 44(3):869–873, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531598.

delaReguera:1988:MSC

[1646] P. A. Fernandez de la Reguera, F. H. C. Marriott, and J. Burley. Multiple-set canonical analysis: an application to forestry genetics. Biometrics, 44(3):875–880, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531599.

Perry:1988:FDR

[1647] Joe N. Perry, L. R. Taylor, and A. W. Kemp. Families of distributions for repeated samples of animal counts. Biometrics, 44(3):881–890, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531600.

Breslow:1988:LRS

[1648] N. E. Breslow, L. P. Zhao, Thomas R. Fears, and Charles C. Brown. Logistic regression for stratified case-control studies. Biometrics, 44(3):891– 899, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531601.

Shih:1988:TCC

[1649] W. J. Shih and Jack Cuzick. Testing in case-control studies with ordinal exposure variables. Biometrics, 44(3):901–905, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531602.

Johnson:1988:VT

[1650] Roger Johnson, C. W. Dunnett, and M. Gent. P -values in 2 × 2 tables. Biometrics, 44(3):907–910, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531603.

Weiss:1988:OPI

[1651] Gary B. Weiss and James A. Hokanson. Optimal Phase II trials. Biometrics, 44(3):910–911, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531604.

Peace:1988:STO

[1652] Karl E. Peace. Some thoughts on one-tailed tests. Biometrics, 44(3):911– 912, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531605.

Derr:1988:BRS

[1653] J. A. Derr. Book review: Statistical Analysis with Missing Data, by R. J. A. Little, D. B. Rubin. Biometrics, 44(3):913–914, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531606.

Spurr:1988:BRD

[1654] B. D. Spurr. Book review: Density Estimation for Statistics and Data Analysis, by B. W. Silverman. Biometrics, 44(3):914, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531607.

Kemp:1988:BRD

[1655] A. W. Kemp. Book review: Design, Data and Analysis, by C. L. Mallows. Biometrics, 44(3):914–915, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531608.

Brown:1988:BRB

[1656] R. A. Brown. Book review: Biostatistics in Clinical Medicine, by J. A. Ingelfinger, F. Mosteller, L. A. Thibodeau, J. H. Ware. Biometrics, 44(3): 915, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531609.

Whitehead:1988:BRC

[1657] J. R. Whitehead. Book review: Clinimetrics, by A. R. Feinstein. Biometrics, 44(3):915–916, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531610.

Bailey:1988:BRSb

[1658] N. T. J. Bailey. Book review: Spatial Aspects of Influenza Epidemics, by A. D. Cliff, P. Haggett, J. K. Ord. Biometrics, 44(3):916–917, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531611.

McDonald:1988:BRL

[1659] J. W. McDonald. Book review: Life Table Techniques and Their Applications, by K. Namboodiri, C. M. Suchindran. Biometrics, 44(3):917–918, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531612.

Reyment:1988:BRS

[1660] R. A. Reyment. Book review: Statistics and Data Analysis in Geology, by J. C. Davis. Biometrics, 44(3):918, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531613.

Lachenbruch:1988:BRA

[1661] P. A. Lachenbruch. Book review: Applied Statistics: a Handbook of BMDP Analyses, by E. J. Snell. Biometrics, 44(3):918–919, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531614.

Morgan:1988:BRN

[1662] B. J. T. Morgan. Book review: Non-Uniform Random Variate Generation, by L. Devroye. Biometrics, 44(3):919, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531615.

Anonymous:1988:BRAb

[1663] Anonymous. Book review: Applied Multivariate Statistical Analysis, by A. R. Johnson, D. W. Wichern. Biometrics, 44(3):920, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531616.

Anonymous:1988:BRC

[1664] Anonymous. Book review: Contributions to the Theory and Application of Statistics: a Volume in Honor of Herbert Solomon, by A. E. Gelfand. Biometrics, 44(3):920, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531618.

Anonymous:1988:BRL

[1665] Anonymous. Book review: Lecture Notes on Medical Statistics, by A. Petrie. Biometrics, 44(3):920, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531617.

Anonymous:1988:BRP

[1666] Anonymous. Book review: Proceedings of the Second International Tampere Conference in Statistics, by T. Pukkila, S. Puntanen. Biometrics, 44(3):920–921, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531619.

Anonymous:1988:BRE

[1667] Anonymous. Book review: Einführung in das Relationale Datenbanksystem SIR/DBMS, by U. Heitfeld. Biometrics, 44(3):921, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531621.

Anonymous:1988:STM

[1668] Anonymous. Selected tables in mathematical statistics, vol. 11 by institute of mathematical statistics. Biometrics, 44(3):921, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531620.

Anonymous:1988:CET

[1669] Anonymous. Corrections: Estimators and tests in the analysis of multiple nonindependent 2×2 tables with partially missing observations. Biometrics, 44(3):923, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531623. See [1669].

Anonymous:1988:CSS

[1670] Anonymous. Corrections: Sample sizes based on the log-rank statistic in complex clinical trials. Biometrics, 44(3):923, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531622. See [1550].

Lakatos:1988:CSS

[1671] E. Lakatos. Correction to: “Sample sizes based on the log-rank statistic in complex clinical trials” [Biometrics 44 (1988), no. 1, 229–241]. Biometrics, 44(3):923, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1550].

Anonymous:1988:BS

[1672] Anonymous. The Biometric Society. Biometrics, 44(3):925–927, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531624.

Anonymous:1988:FMd

[1673] Anonymous. Front matter. Biometrics, 44(4):i–iv, December 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531723.

Anonymous:1988:VI

[1674] Anonymous. Volume information. Biometrics, 44(4):xv–xxxii, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531722.

Visser:1988:KFA

[1675] H. Visser and J. Molenaar. Kalman filter analysis in dendroclimatology. Biometrics, 44(4):929–940, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531724.

Green:1988:HHD

[1676] Jack Green and Mamdouh Montasser. HLA haplotype discordance. Biometrics, 44(4):941–950, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531725.

Shute:1988:PPL

[1677] Nereda C. E. Shute. Power properties of the likelihood ratio test in an analysis of genetic diseases. Biometrics, 44(4):951–958, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531726.

Diggle:1988:AAR

[1678] Peter J. Diggle. An approach to the analysis of repeated measurements. Biometrics, 44(4):959–971, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531727.

Stanek:1988:GCM

[1679] Edward J. Stanek III and Scott R. Diehl. Growth curve models of repeated binary response. Biometrics, 44(4):973–983, December 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531728.

Brown:1988:NCG

[1680] G. H. Brown and J. B. Donnelly. A note on comparing growth rates of animals between groups. Biometrics, 44(4):985–993, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531729.

Jolicoeur:1988:LAG

[1681] Pierre Jolicoeur, Jacques Pontier, Marie-Odile Pernin, and Michel Sempé. A lifetime asymptotic growth curve for human height. Biometrics, 44(4):995–1003, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531730.

Davis:1988:NMA

[1682] Charles S. Davis and L. J. Wei. Nonparametric methods for analyzing incomplete nondecreasing repeated measurements. Biometrics, 44 (4):1005–1018, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531731.

Zeger:1988:MRM

[1683] Scott L. Zeger and Bahjat Qaqish. Markov regression models for time series: a quasi-likelihood approach. Biometrics, 44(4):1019–1031, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531732.

Prentice:1988:CBR

[1684] Ross L. Prentice. Correlated binary regression with covariates specific to each binary observation. Biometrics, 44(4):1033–1048, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531733.

Zeger:1988:MLD

[1685] Scott L. Zeger, Kung-Yee Liang, and Paul S. Albert. Models for longitudinal data: a generalized estimating equation approach. Biometrics, 44(4):1049–1060, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531734. See correction [1754, 1755].

Knuiman:1988:IPI

[1686] M. W. Knuiman and T. P. Speed. Incorporating prior information into the analysis of contingency tables. Biometrics, 44(4):1061–1071, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531735.

vanHouwelingen:1988:UAV

[1687] J. C. van Houwelingen. Use and abuse of variance models in regression. Biometrics, 44(4):1073–1081, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531736.

Morris:1988:SSC

[1688] Max D. Morris. Small-sample confidence limits for parameters under inequality constraints with application to quantal bioassay. Biometrics, 44(4):1083–1092, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531737.

Does:1988:UJM

[1689] Ronald J. M. M. Does, Leo W. G. Strijbosch, and Willem Albers. Using jackknife methods for estimating the parameter in dilution series. Biometrics, 44(4):1093–1102, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531738.

Staniswalis:1988:KED

[1690] Joan G. Staniswalis and Vanessa Cooper. Kernel estimates of dose response. Biometrics, 44(4):1103–1119, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531739.

Raubertas:1988:STA

[1691] Richard F. Raubertas. Spatial and temporal analysis of disease occurrence for detection of clustering. Biometrics, 44(4):1121–1129, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531740.

Jones:1988:SCA

[1692] Richard H. Jones, Paul M. Ford, and Richard F. Hamman. Seasonality comparisons among groups using incidence data. Biometrics, 44(4):1131– 1144, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531741.

Liang:1988:UCP

[1693] Kung-Yee Liang and Scott L. Zeger. On the use of concordant pairs in matched case-control studies. Biometrics, 44(4):1145–1156, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531742.

Dupont:1988:PCM

[1694] William D. Dupont. Power calculations for matched case-control studies. Biometrics, 44(4):1157–1168, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531743.

Mullen:1988:MDT

[1695] Kenneth Mullen, Daniel M. Ennis, Elise de Doncker, and John A. Kapenga. Models for the duo-trio and triangular methods. Biometrics, 44(4):1169–1175, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531744.

Smith:1988:BMM

[1696] Philip J. Smith. Bayesian methods for multiple capture–recapture surveys. Biometrics, 44(4):1177–1189, December 1988. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531745.

Farewell:1988:UMM

[1697] V. T. Farewell and D. A. Sprott. The use of a mixture model in the analysis of count data. Biometrics, 44(4):1191–1194, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531746.

Thode:1988:SPP

[1698] Henry C. Thode, Jr., Stephen J. Finch, and Nancy R. Mendell. Simulated percentage points for the null distribution of the likelihood ratio test for a mixture of two normals. Biometrics, 44(4):1195–1201, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531747.

Slud:1988:DCR

[1699] Eric V. Slud, David P. Byar, Arthur Schatzkin, Ross Prentice, and Jack Kalbfleisch. Dependent competing risks and the latent-failure model. Biometrics, 44(4):1203–1205, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531748.

Newton:1988:BRS

[1700] J. Newton and C. D. Sinclair. Book review: StatLab Microcomputer-Based Practical Classes and Demonstrations for Teaching Statistical Concepts, by W. D. Stirling. Biometrics, 44(4):1207–1208, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531749.

Kemp:1988:BRP

[1701] A. W. Kemp. Book review: Pascal’s Arithmetical Triangle, by A. W. F. Edwards. Biometrics, 44(4):1208–1209, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531750.

Anonymous:1988:BRAc

[1702] Anonymous. Book review: Applied Statistics, by M. L. Berenson, D. M. Levine, D. Rindskopf. Biometrics, 44(4):1209, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531751.

Anonymous:1988:BRAd

[1703] Anonymous. Book review: Automation in Survey Processing, by Netherlands Central Bureau of Statistics. Biometrics, 44(4):1209, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531753.

Anonymous:1988:BRQ

[1704] Anonymous. Book review: Quantitative Risk Assessment, by J. M. Humber, R. F. Almeder. Biometrics, 44(4):1209, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531752.

Anonymous:1988:BRMc

[1705] Anonymous. Book review: Matrix Differential Calculus with Applications in Statistics and Econometrics, by J. R. Magnus, H. Neudecker. Biometrics, 44(4):1209–1210, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531754.

Anonymous:1988:BRW

[1706] Anonymous. Book reviews: World Health Statistics Annual 1987, by World Health Organization; World Health Statistics Quarterly 1987, by World Health Organization. Biometrics, 44(4):1210, December 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531755.

Anonymous:1989:FMa

[1707] Anonymous. Front matter. Biometrics, 45(1):i–iv, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532029.

DeGruttola:1989:ADC

[1708] Victor De Gruttola and Stephen W. Lagakos. Analysis of doublycensored survival data, with application to AIDS. Biometrics, 45(1): 1–11, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532030.

Whitehead:1989:SMM

[1709] John Whitehead. Sequential methods for monitoring declining quality, with application to the long-term storage of seeds. Biometrics, 45(1):13– 22, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532031.

Walters:1989:CES

[1710] David K. Walters, Timothy G. Gregoire, and Harold E. Burkhart. Consistent estimation of site index curves fitted to temporary plot data. Biometrics, 45(1):23–33, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532032.

Gentleman:1989:DFO

[1711] Jane F. Gentleman and Ronald C. Mullin. The distribution of the frequency of occurrence of nucleotide subsequences, based on their overlap capability. Biometrics, 45(1):35–52, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532033.

Hernandez:1989:DCA

[1712] J. L. Hernández and B. S. Weir. A disequilibrium coefficient approach to Hardy–Weinberg testing. Biometrics, 45(1):53–70, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532034.

Hoenig:1989:EIP

[1713] John M. Hoenig and Victor R. Restrepo. Estimating the intermolt periods in asynchronously molting crustacean populations. Biometrics, 45(1):71–82, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532035.

Robson:1989:TBA

[1714] Douglas Robson and Cynthia M. Jones. The theoretical basis of an access site angler survey design. Biometrics, 45(1):83–98, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532036.

Pollock:1989:EAS

[1715] Kenneth H. Pollock, Scott R. Winterstein, and Michael J. Conroy. Estimation and analysis of survival distributions for radio-tagged animals. Biometrics, 45(1):99–109, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532037.

Routledge:1989:RME

[1716] R. D. Routledge. The removal method for estimating natural populations: Incorporating auxiliary information. Biometrics, 45(1):111–121, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532038.

Clifford:1989:ASC

[1717] Peter Clifford, Sylvia Richardson, and Denis Hémon. Assessing the significance of the correlation between two spatial processes. Biometrics, 45(1): 123–134, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532039.

OQuigley:1989:STH

[1718] John O’Quigley and Fabienne Pessione. Score tests for homogeneity of regression effect in the proportional hazards model. Biometrics, 45(1): 135–144, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532040.

Huster:1989:MPS

[1719] William J. Huster, Ron Brookmeyer, and Steven G. Self. Modelling paired survival data with covariates. Biometrics, 45(1):145–156, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532041.

Jones:1989:GCN

[1720] Michael P. Jones and John Crowley. A general class of nonparametric tests for survival analysis. Biometrics, 45(1):157–170, March 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532042.

Jones:1989:MCI

[1721] Michael P. Jones, Thomas W. O’Gorman, Jon H. Lemke, and Robert F. Woolson. A Monte Carlo investigation of homogeneity tests of the odds ratio under various sample size configurations. Biometrics, 45(1):171– 181, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532043.

Greenland:1989:GMH

[1722] Sander Greenland. Generalized Mantel–Haenszel estimators for K 2 × J tables. Biometrics, 45(1):183–191, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532044.

Rochon:1989:AGM

[1723] James Rochon. The application of the GSK method to the determination of minimum sample sizes. Biometrics, 45(1):193–205, March 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532045.

Rochon:1989:MLE

[1724] James Rochon and Ronald W. Helms. Maximum likelihood estimation for incomplete repeated-measures experiments under an ARMA covariance structure. Biometrics, 45(1):207–218, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532046.

Finkelstein:1989:AMT

[1725] Dianne M. Finkelstein and David A. Schoenfeld. Analysis of multiple tumor data from a rodent carcinogenicity experiment. Biometrics, 45(1): 219–230, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532047.

Paul:1989:TDB

[1726] S. R. Paul, K. Y. Liang, and S. G. Self. On testing departure from the binomial and multinomial assumptions. Biometrics, 45(1):231–236, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532048.

McGilchrist:1989:SR

[1727] C. A. McGilchrist, S. Lianto, and D. M. Byron. Signed residuals. Biometrics, 45(1):237–246, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532049.

Chang:1989:CIN

[1728] Myron N. Chang. Confidence intervals for a normal mean following a group sequential test. Biometrics, 45(1):247–254, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532050.

Lin:1989:CCC

[1729] Lawrence I-Kuei Lin. A concordance correlation coefficient to evaluate reproducibility. Biometrics, 45(1):255–268, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532051.

Bloch:1989:KCM

[1730] Daniel A. Bloch and Helena Chmura Kraemer. 2 × 2 kappa coefficients: Measures of agreement or association. Biometrics, 45(1):269–287, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532052.

Snapinn:1989:EER

[1731] Steven M. Snapinn and James D. Knoke. Estimation of error rates in discriminant analysis with selection of variables. Biometrics, 45(1):289– 299, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532053.

Edmondson:1989:GDR

[1732] R. N. Edmondson. Glasshouse design for repeatedly harvested crops. Biometrics, 45(1):301–307, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532054.

Clark:1989:ENB

[1733] Suzanne J. Clark and Joe N. Perry. Estimation of the negative binomial parameter κ by maximum quasi-likelihood. Biometrics, 45(1):309–316, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532055.

Choi:1989:MCT

[1734] S. C. Choi and P. A. Pepple. Monitoring clinical trials based on predictive probability of significance. Biometrics, 45(1):317–323, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532056.

Brookmeyer:1989:CEA

[1735] Ron Brookmeyer and James J. Goedert. Censoring in an epidemic with an application to hemophilia-associated AIDS. Biometrics, 45(1):325– 335, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532057.

Hoekstra:1989:EE

[1736] Janneke A. Hoekstra. Estimation of the ED50. Biometrics, 45(1):337– 338, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532058.

Kemp:1989:BREa

[1737] C. D. Kemp. Book review: Encyclopedia of Statistical Sciences, Vol. 7. Plackett Family of Distributions-Regression, Wrong, by S. Kotz, N. L. Johnson, C. B. Read. Biometrics, 45(1):339, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532059.

Blom:1989:BRM

[1738] G. Blom. Book review: Model Selection, by H. Linhart, W. Zucchini. Biometrics, 45(1):340, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532060.

Pearce:1989:BRD

[1739] S. C. Pearce. Book review: Discovering Causal Structure, by C. Glymour, R. Scheines, P. Spirtes, K. Kelly. Biometrics, 45(1):340–341, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532061.

Berg:1989:BRG

[1740] S. Berg. Book review: Grundriss der Biologischen Statistik, 9 Aufl, by E. Weber. Biometrics, 45(1):341, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532062.

Berg:1989:BRS

[1741] S. Berg. Book review: Sequentielle statistische Verfahren, by P. Bauer, V. Scheiber, F. X. Wohlzogen. Biometrics, 45(1):341–342, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532063.

Hasted:1989:BRR

[1742] A. M. Hasted. Book review: Regression Analysis with Applications, by G. B. Wetherill. Biometrics, 45(1):342, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532064.

Everitt:1989:BRM

[1743] B. S. Everitt. Book review: Multivariate Exploratory Data Analysis: a Perspective on Exploratory Factor Analysis, by A. Yates. Biometrics, 45(1):342–343, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532065.

Gordon:1989:BRC

[1744] A. D. Gordon. Book review: Component and Correspondence Analysis: Dimension Reduction, by Functional Approximation, by J. L. A. van Rijckevorsel, J. Leeuw. Biometrics, 45(1):343, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532066.

Clarke:1989:BRL

[1745] G. M. Clarke. Book review: Linear Estimation and Design of Experiments, by D. D. Joshi. Biometrics, 45(1):343–344, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532067.

Fletcher:1989:BRC

[1746] D. J. Fletcher. Book review: Cyclic Designs, by J. A. John. Biometrics, 45(1):344, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532068.

Kershaw:1989:BRM

[1747] C. D. Kershaw. Book review: Multiple Imputation for Nonresponse in Surveys, by D. B. Rubin. Biometrics, 45(1):345, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532070.

Papageorgiou:1989:BRS

[1748] H. Papageorgiou. Book review: The Statistical Analysis of Compositional Data, by J. Aitchison. Biometrics, 45(1):345, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532069.

Anonymous:1989:BRC

[1749] Anonymous. Book review: The Collected Works of John W. Tukey, Volume V, Graphics 1965–1985, by W. S. Cleveland. Biometrics, 45(1): 346, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532071.

Anonymous:1989:BRL

[1750] Anonymous. Book review: Linear Structures, by J. R. Magnus. Biometrics, 45(1):346, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532073.

Anonymous:1989:BRM

[1751] Anonymous. Book review: Multivariate Statistical Simulation, by M. E. Johnson. Biometrics, 45(1):346, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532072.

Anonymous:1989:BRTa

[1752] Anonymous. Book review: Tables for Selecting the Best Population, by K. L. Butler, D. G. Butler. Biometrics, 45(1):346, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532074.

Anonymous:1989:BRTb

[1753] Anonymous. Book review: TIMESLAB: a Times Series Analysis Laboratory, by H. J. Newton. Biometrics, 45(1):346, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532075.

Anonymous:1989:CML

[1754] Anonymous. Correction: Models for longitudinal data: a generalized estimating equation approach. Biometrics, 45(1):347, March 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532076. See [1685].

Zeger:1989:CML

[1755] S. L. Zeger, K.-Y. Liang, and P. S. Albert. Correction to: “Models for longitudinal data: a generalized estimating equation approach” [Biometrics 44 (1988), no. 4, 1049–1060]. Biometrics, 45(1):347, March 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1685].

Anonymous:1989:FMb

[1756] Anonymous. Front matter. Biometrics, 45(2):i–iv, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531481.

Rudemo:1989:REM

[1757] M. Rudemo, D. Ruppert, and J. C. Streibig. Random-effect models in nonlinear regression with applications to bioassay. Biometrics, 45(2): 349–362, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531482.

Roeder:1989:AML

[1758] Kathryn Roeder, B. Devlin, and Bruce G. Lindsay. Application of maximum likelihood methods to population genetic data for the estimation of individual fertilities. Biometrics, 45(2):363–379, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531483.

Moulton:1989:ARM

[1759] Lawrence H. Moulton and Scott L. Zeger. Analyzing repeated measures on generalized linear models via the bootstrap. Biometrics, 45(2):381– 394, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531484.

Cormack:1989:LLM

[1760] R. M. Cormack. Log-linear models for capture–recapture. Biometrics, 45(2):395–413, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531485.

Steinhorst:1989:SAM

[1761] R. Kirk Steinhorst and Michael D. Samuel. Sightability adjustment methods for aerial surveys of wildlife populations. Biometrics, 45(2):415–425, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531486.

Chao:1989:EPS

[1762] Anne Chao. Estimating population size for sparse data in capture– recapture experiments. Biometrics, 45(2):427–438, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531487.

Heimbuch:1989:CRE

[1763] Douglas G. Heimbuch and John M. Hoenig. Change-in-ratio estimators for habitat usage and relative population size. Biometrics, 45(2):439– 451, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531488.

Zahl:1989:LTS

[1764] Samuel Zahl. Line transect sampling with unknown probability of detection along the transect. Biometrics, 45(2):453–470, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531489.

Yip:1989:IPC

[1765] Paul Yip. An inference procedure for a capture and recapture experiment with time- dependent capture probabilities. Biometrics, 45(2):471–479, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531490.

Laird:1989:EBC

[1766] Nan M. Laird and Thomas A. Louis. Empirical Bayes confidence intervals for a series of related experiments. Biometrics, 45(2):481–495, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531491.

Pepe:1989:WKM

[1767] Margaret Sullivan Pepe and Thomas R. Fleming. Weighted Kaplan– Meier statistics: a class of distance tests for censored survival data. Biometrics, 45(2):497–507, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531492.

Halperin:1989:DFC

[1768] Max Halperin, Mohamed I. Hamdy, and Peter F. Thall. Distribution-free confidence intervals for a parameter of Wilcoxon–Mann–Whitney type for ordered categories and progressive censoring. Biometrics, 45(2):509– 521, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531493.

Andersen:1989:SPN

[1769] Per Kragh Andersen and Michael Væth. Simple parametric and nonparametric models for excess and relative mortality. Biometrics, 45(2): 523–535, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531494.

Thall:1989:TSD

[1770] Peter F. Thall, Richard Simon, and Susan S. Ellenberg. A two-stage design for choosing among several experimental treatments and a control in clinical trials. Biometrics, 45(2):537–547, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531495.

Shrout:1989:DTP

[1771] Patrick E. Shrout and Stephen C. Newman. Design of two-phase prevalence surveys of rare disorders. Biometrics, 45(2):549–555, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531496.

Rosenbaum:1989:RKE

[1772] Paul R. Rosenbaum. The role of known effects in observational studies. Biometrics, 45(2):557–569, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531497.

Kianifard:1989:URR

[1773] Farid Kianifard and William H. Swallow. Using recursive residuals, calculated on adaptively-ordered observations, to identify outliers in linear regression. Biometrics, 45(2):571–585, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531498.

Espeland:1989:ULC

[1774] Mark A. Espeland and Stanley L. Handelman. Using latent class models to characterize and assess relative error in discrete measurements. Biometrics, 45(2):587–599, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531499.

Scott:1989:ASH

[1775] David J. Scott and Andrew H. Sinclair. Analysis of in-situ hybridization data for unique genes using GLIM. Biometrics, 45(2):601–604, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531500.

Donner:1989:SMO

[1776] Allan Donner. Statistical methods in ophthalmology: an adjusted chi-square approach. Biometrics, 45(2):605–611, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531501.

Kim:1989:PEF

[1777] Kyungmann Kim. Point estimation following group sequential tests. Biometrics, 45(2):613–617, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531502.

Permutt:1989:SEE

[1778] Thomas Permutt and J. Richard Hebel. Simultaneous-equation estimation in a clinical trial of the effect of smoking on birth weight. Biometrics, 45(2):619–622, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531503.

Shih:1989:PAS

[1779] W. J. Shih. Prediction approaches to sequentially searching for an optimal dose. Biometrics, 45(2):623–628, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531504.

Feuer:1989:TSS

[1780] Eric J. Feuer and Larry G. Kessler. Test statistic and sample size for a two-sample McNemar test. Biometrics, 45(2):629–636, June 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531505. See correction [1845].

Ruppert:1989:TWM

[1781] David Ruppert, Noel Cressie, and Raymond J. Carroll. A transformation/weighting model for estimating Michaelis–Menten parameters. Biometrics, 45(2):637–656, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531506.

Johnson:1989:LSE

[1782] Douglas H. Johnson. Least squares estimation of avian molt rates. Biometrics, 45(2):657–661, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531507.

Elswick:1989:NAA

[1783] R. K. Elswick, Jr. and Vincent A. Uthoff. A nonparametric approach to the analysis of the two-treatment, two-period, four-sequence crossover model. Biometrics, 45(2):663–667, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531508.

Chiang:1989:IDP

[1784] Yu-Kun Chiang, Robert J. Hardy, C. Morton Hawkins, and Asha S. Kapadia. An illness-death process with time-dependent covariates. Biometrics, 45(2):669–681, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531509.

Hansen:1989:LME

[1785] Poul Einer Hansen. Leslie matrix elements. Biometrics, 45(2):683–686, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531510.

Hall:1989:OPA

[1786] Peter G. Hall. Obituary: Patrick Alfred Pierce Moran 1917–1988. Biometrics, 45(2):687–692, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531511.

Kemp:1989:BREb

[1787] A. W. Kemp. Book reviews: Encyclopedia of Statistical Sciences, Vol. 8. Regressograms-St. Petersburg Paradox, by S. Kotz, N. L. Johnson, C. B. Read; Encyclopedia of Statistical Sciences, Vol. 9. Strata Chart– Zyskind–Martin Models–Index, by S. Kotz, N. L. Johnson, C. B. Read. Biometrics, 45(2):693–694, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531512.

Ford:1989:BRN

[1788] I. Ford. Book review: Nonlinear Statistical Models, by A. R. Gallant. Biometrics, 45(2):694–695, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531513.

Harris:1989:BRT

[1789] R. R. Harris. Book review: The Teaching of Practical Statistics, by C. W. Anderson, R. W. Loynes. Biometrics, 45(2):695–696, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531514.

Williamson:1989:BRQ

[1790] M. H. Williamson. Book review: Quantitative Ecological Theory — an Introduction to Basic Models, by M. R. Rose. Biometrics, 45(2):696, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531515.

Stene:1989:BRT

[1791] J. M. Stene. Book review: Theoretical Studies on Sex Ratio Evolution, by S. Karlin, S. Lessard. Biometrics, 45(2):696–697, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531516.

Wilson:1989:BRM

[1792] S. R. Wilson. Book review: Mapping and Sequencing the Human Genome, by National Research Council. Biometrics, 45(2):697, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531517.

Kempson:1989:BRB

[1793] R. E. Kempson. Book review: A Biologist’s Advanced Mathematics, by D. R. Causton. Biometrics, 45(2):697–698, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531518.

Bissell:1989:BRQ

[1794] A. F. Bissell. Book review: Quality Assurance of Chemical Measurements, by J. K. Taylor. Biometrics, 45(2):698, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531519.

Anonymous:1989:BRG

[1795] Anonymous. Book review: Graphs as Structural Models: The Application of Graphs and Multigraphs in Cluster Analysis, by E. Godehardt. Biometrics, 45(2):699, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531522.

Anonymous:1989:BRI

[1796] Anonymous. Book review: Introduction to Optimization Methods and Their Application in Statistics, by B. S. Everitt. Biometrics, 45(2):699, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531520.

Anonymous:1989:BRNa

[1797] Anonymous. Book review: The New S Language, by R. A. Becker, J. M. Chambers, A. R. Wilks. Biometrics, 45(2):699, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531523.

Anonymous:1989:BRSa

[1798] Anonymous. Book review: Stochastic Geometry and Its Applications, by D. Stoyan, W. S. Kendall, J. Mecke. Biometrics, 45(2):699, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531521.

Anonymous:1989:CAE

[1799] Anonymous. Correction: Alternative estimation procedures for P r(X < Y) in categorized data. Biometrics, 45(2):701, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531524. See [1320].

Anonymous:1989:CMM

[1800] Anonymous. Correction: Mixed-model analysis of a censored normal distribution with reference to animal breeding. Biometrics, 45(2):701, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531525. See [1497].

Simonoff:1989:CAE

[1801] J. S. Simonoff, Y. Hochberg, and B. Reiser. Correction to: “Alternative estimation procedures for Pr(X < Y) in categorized data” [Biometrics 42 (1986), no. 4, 895–907; MR0872965 (88c:62045)]. Biometrics, 45(2): 701, June 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1320].

Anonymous:1989:FMc

[1802] Anonymous. Front matter. Biometrics, 45(3):i–iv, September 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531677.

Matis:1989:SGS

[1803] James H. Matis, Thomas E. Wehrly, and William C. Ellis. Some generalized stochastic compartment models for digesta flow. Biometrics, 45(3):703–720, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531678.

Lwin:1989:PM

[1804] T. Lwin and P. J. Martin. Probits of mixtures. Biometrics, 45(3):721– 732, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531679.

Laurence:1989:OSM

[1805] A. F. Laurence and B. J. T. Morgan. Observations on a stochastic model for quantal assay data. Biometrics, 45(3):733–744, September 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531680.

Raz:1989:SSP

[1806] Jonathan Raz, Bruce Turetsky, and George Fein. Selecting the smoothing parameter for estimation of slowly changing evoked potential signals. Biometrics, 45(3):745–762, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531681.

VanDeusen:1989:MBA

[1807] Paul C. Van Deusen. A model-based approach to tree ring analysis. Biometrics, 45(3):763–779, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531682.

Gelber:1989:QLO

[1808] Richard D. Gelber, Rebecca S. Gelman, and Aron Goldhirsch. A qualityof-life-oriented endpoint for comparing therapies. Biometrics, 45(3):781– 795, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531683.

Garber:1989:DTM

[1809] Alan M. Garber. A discrete-time model of the acquistin of antibioticresistant infections in hospitalized patients. Biometrics, 45(3):797–816, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531684.

Wallenstein:1989:TPS

[1810] Sylvan Wallenstein, Clarice R. Weinberg, and Madelyn Gould. Testing for a pulse in seasonal event data. Biometrics, 45(3):817–830, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531685.

Izenman:1989:CLS

[1811] Alan J. Izenman and J. S. Williams. A class of linear spectral models and analyses for the study of longitudinal data. Biometrics, 45(3):831–849, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531686.

Raz:1989:ARM

[1812] Jonathan Raz. Analysis of repeated measurements using nonparametric smoothers and randomization tests. Biometrics, 45(3):851–871, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531687.

Smith:1989:ALS

[1813] Woollcott Smith. ANOVA-like similarity analysis using expected species shared. Biometrics, 45(3):873–881, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531688.

Tarone:1989:STC

[1814] Robert E. Tarone and John J. Gart. Significance tests for cancer screening trials. Biometrics, 45(3):883–890, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531689.

Lau:1989:RST

[1815] Tai Shing Lau. On repeated screening tests. Biometrics, 45(3):891–898, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531690.

Gray:1989:LRT

[1816] Robert J. Gray and Anastasios A. Tsiatis. A linear rank test for use when the main interest is in differences in cure rates. Biometrics, 45(3):899– 904, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531691.

Emerson:1989:SGS

[1817] Scott S. Emerson and Thomas R. Fleming. Symmetric group sequential test designs. Biometrics, 45(3):905–923, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531692.

Storer:1989:DAP

[1818] Barry E. Storer. Design and analysis of Phase I clinical trials. Biometrics, 45(3):925–937, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531693.

Wu:1989:ECC

[1819] Margaret C. Wu and Kent R. Bailey. Estimation and comparison of changes in the presence of informative right censoring: Conditional linear model. Biometrics, 45(3):939–955, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www. jstor.org/stable/2531694. See corrections [1901, 1902].

Kupper:1989:AIA

[1820] Lawrence L. Kupper and Kerry b. Hafner. On assessing interrater agreement for multiple attribute responses. Biometrics, 45(3):957–967, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531695.

Beal:1989:SSD

[1821] S. L. Beal. Sample size determination for confidence intervals on the population mean and on the difference between two population means. Biometrics, 45(3):969–977, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531696.

Choi:1989:RFG

[1822] Jai Won Choi and Richard B. McHugh. A reduction factor in goodnessof-fit and independence tests for clustered and weighted observations. Biometrics, 45(3):979–996, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531697.

Huhn:1989:TSN

[1823] M. Hühn and R. Nassar. On tests of significance for nonparametric measures of phenotypic stability. Biometrics, 45(3):997–1000, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531698.

Bockenholt:1989:AOE

[1824] Ulf Böckenholt. Analyzing optima in the exploration of multiple response surfaces. Biometrics, 45(3):1001–1008, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531699.

Holland:1989:NCM

[1825] Paul W. Holland. A note on the covariance of the Mantel–Haenszel log-odds-ratio estimator and the sample marginal rates. Biometrics, 45 (3):1009–1016, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531700.

Lan:1989:CFI

[1826] K. K. Gordon Lan and David L. DeMets. Changing frequency of interim analysis in sequential monitoring. Biometrics, 45(3):1017–1020, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531701.

Singh:1989:SOP

[1827] Murari Singh. A stochastic optimum proportion under a shared system of sole corps. Biometrics, 45(3):1021–1026, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531702.

Senn:1989:COM

[1828] Stephen Senn. Combining outcome measures: Statistical power is irrelevant. Biometrics, 45(3):1027–1028, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531703.

Morgan:1989:EAM

[1829] Byron J. T. Morgan, D. M. Titterington, Camil Fuchs, and Joel B. Greenhouse. The EM algorithm for the mover-stayer model. Biometrics, 45(3):1029–1030, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531704.

Lui:1989:NER

[1830] Kung-Jong Lui and Nicholas P. Jewell. A note on the estimate of the relative risk when sample size are small. Biometrics, 45(3):1030–1032, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531705.

Elandt-Johnson:1989:CCR

[1831] Regina C. Elandt-Johnson. Correction: Competing risk models. Biometrics, 45(3):1032, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531706.

Kemp:1989:BRC

[1832] C. D. Kemp, A. W. Kemp, and R. O. Kuehl. Book review: Comparative Experiments with Field Crops, by G. V. Dyke. Biometrics, 45(3):1033– 1034, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531707.

Barath:1989:BRM

[1833] E. Baráth. Book review: Multidimensional Data Analysis, by J. de Leeuw, W. Heiser, J. Meulman, F. Critchley. Biometrics, 45(3):1034, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531708.

Ripley:1989:BRI

[1834] B. D. Ripley. Book review: Introduction to the Theory of Coverage Processes, by P. Hall. Biometrics, 45(3):1034–1035, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531709.

Williams:1989:BRS

[1835] D. A. Williams. Book review: Statistical Techniques in Bioassay, by Z. Govindarajulu. Biometrics, 45(3):1035, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531710.

Pearce:1989:BRC

[1836] S. C. Pearce. Book review: Crimes of Official Science: a Casebook, by I. D. Bross. Biometrics, 45(3):1035–1036, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531711.

Anonymous:1989:BRB

[1837] Anonymous. Book review: Biometrie in der Chemisch-Pharmazeutischen Industrie 3: Auswertungs- und Darstellungsmethoden–Risikoextrapolation zur Kanzerogenität, by J. Vollmar. Biometrics, 45(3):1036, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531712.

Anonymous:1989:BRDa

[1838] Anonymous. Book review: A Dictionary of Epidemiology, by J. M. Last. Biometrics, 45(3):1036, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531713.

Anonymous:1989:BRNb

[1839] Anonymous. Book review: Numerical Methods in Quaternary Pollen Analysis, by H. J. B. Birks, A. D. Gordon. Biometrics, 45(3):1036–1037, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531714.

Anonymous:1989:BRDb

[1840] Anonymous. Book review: Dynamical Systems and Environmental Models, by H. G. Bothe, W. Ebeling, A. B. Kurzanski, M. Peschel. Biometrics, 45(3):1037, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531715.

Anonymous:1989:FSS

[1841] Anonymous. Fortschritte der Statistik-Software 1, by F. Faulbaum, H.-M. Uehlinger. Biometrics, 45(3):1037, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531717.

Anonymous:1989:BRSb

[1842] Anonymous. Book review: Search Theory, by D. V. Chudnovsky, G. V. Chudnovsky. Biometrics, 45(3):1037, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531716.

Anonymous:1989:CCE

[1843] Anonymous. Correction: Confounding in epidemiologic studies: The adequacy of the control group as a measure of confounding. Biometrics, 45(3):1039, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531718. See [1481].

Anonymous:1989:CMP

[1844] Anonymous. Correction: Mixed Poisson likelihood regression models for longitudinal interval count data. Biometrics, 45(3):1039, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531719. See [1547].

Anonymous:1989:CTS

[1845] Anonymous. Correction: Test statistic and sample size for a twosample McNemar test. Biometrics, 45(3):1039, September 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531720. See [1780].

Thall:1989:CMP

[1846] Peter F. Thall. Correction to: “Mixed Poisson likelihood regression models for longitudinal interval count data” [Biometrics 44 (1988), no. 1, 197–209]. Biometrics, 45(3):1039, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1547].

Wickramaratne:1989:CCE

[1847] Priya J. Wickramaratne and Theodore R. Holford. Correction to: “Confounding in epidemiologic studies: the adequacy of the control group as a measure of confounding” [Biometrics 43 (1987), no. 4, 751–765]. Biometrics, 45(3):1039, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1481].

Anonymous:1989:FMd

[1848] Anonymous. Front matter. Biometrics, 45(4):i–iv, December 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531758.

Anonymous:1989:VI

[1849] Anonymous. Volume information. Biometrics, 45(4):xi–xxviii, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531757.

Hougaard:1989:KFS

[1850] Philip Hougaard, Anne Plum, and Ulla Ribel. Kernel function smoothing of insulin absorption kinetics. Biometrics, 45(4):1041–1052, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531759.

Graubard:1989:ECS

[1851] Barry I. Graubard, Thomas R. Fears, and Mitchell H. Gail. Effects of cluster sampling on epidemiologic analysis in population-based casecontrol studies. Biometrics, 45(4):1053–1071, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531760.

Klein:1989:SMO

[1852] John P. Klein, Niels Keiding, and Claus Kamby. Semiparametric Marshall–Olkin models applied to the occurrence of metastases at multiple sites after breast cancer. Biometrics, 45(4):1073–1086, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531761.

Morgan:1989:MFY

[1853] Byron J. T. Morgan and Stephen N. Freeman. A model with first-year variation for ring-recovery data. Biometrics, 45(4):1087–1101, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531762.

Robins:1989:DSC

[1854] James M. Robins, Ross L. Prentice, and Donald Blevins. Designs for synthetic case-control studies in open cohorts. Biometrics, 45(4):1103– 1116, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531763.

Schell:1989:IFR

[1855] Michael J. Schell and Frederick W. Leysieffer. An increasing failure rate approach to low-dose extrapolation. Biometrics, 45(4):1117–1123, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531764.

Robins:1989:PCU

[1856] James Robins and Sander Greenland. The probability of causation under a stochastic model for individual risk. Biometrics, 45(4):1125–1138, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531765.

Laska:1989:TWI

[1857] Eugene M. Laska and Morris J. Meisner. Testing whether an identified treatment is best. Biometrics, 45(4):1139–1151, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531766.

Pettitt:1989:ODF

[1858] A. N. Pettitt. One degree of freedom for nonadditivity: Applications with generalized linear models and link functions. Biometrics, 45(4):1153– 1162, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531767.

Krzanowski:1989:NCT

[1859] W. J. Krzanowski and D. Radley. Nonparametric confidence and tolerance regions in canonical variate analysis. Biometrics, 45(4):1163–1173, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531768.

Norwood:1989:MCP

[1860] Paula K. Norwood, Allan R. Sampson, Kathleen McCarroll, and Roger Staum. A multiple comparisons procedure for use in conjunction with the benard-van elteren test. Biometrics, 45(4):1175–1182, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531769.

Geary:1989:MCS

[1861] D. N. Geary. Modelling the covariance structure of repeated measurements. Biometrics, 45(4):1183–1195, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531770.

Berry:1989:MAD

[1862] Donald A. Berry. Monitoring accumulating data in a clinical trial. Biometrics, 45(4):1197–1211, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531771.

Marshall:1989:PVS

[1863] R. J. Marshall. The predictive value of simple rules for combining two diagnostic tests. Biometrics, 45(4):1213–1222, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531772.

McDonald:1989:IES

[1864] J. F. McDonald and D. Palanacki. Interval estimation of the size of a small population from a mark–recapture experiment. Biometrics, 45 (4):1223–1231, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531773.

Jensen:1989:CIN

[1865] A. L. Jensen. Confidence intervals for nearly unbiased estimators in single-mark and single-recapture experiments. Biometrics, 45(4):1233– 1237, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531774.

Leitnaker:1989:EDT

[1866] M. G. Leitnaker. Estimation of delay times in stochastic compartmental models. Biometrics, 45(4):1239–1247, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531775.

Ekbohm:1989:SFI

[1867] Gunnar Ekbohm and Hans Melander. The subject-by-formulation interaction as a criterion of interchangeability of drugs. Biometrics, 45 (4):1249–1254, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531776.

Diggle:1989:TRD

[1868] Peter J. Diggle. Testing for random dropouts in repeated measurement data. Biometrics, 45(4):1255–1258, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531777.

Kopp:1989:NAC

[1869] Annette Kopp and Christopher J. Portier. A note on approximating the cumulative distribution function of the time to tumor onset in multistage models. Biometrics, 45(4):1259–1263, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531778.

Hosmer:1989:BSL

[1870] David W. Hosmer, Borko Jovanovic, and Stanley Lemeshow. Best subsets logistic regression. Biometrics, 45(4):1265–1270, December 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531779. See corrections [1900].

Kampert:1989:GFA

[1871] James B. Kampert and Abraham Silvers. Goodness of fit in armitage-doll Poisson models. Biometrics, 45(4):1271–1280, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531780.

McKean:1989:RTS

[1872] Joseph W. McKean, Thomas J. Vidmar, and Gerald L. Sievers. A robust two-stage multiple comparison procedure with application to a random drug screen. Biometrics, 45(4):1281–1297, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531781.

Petkau:1989:MQR

[1873] A. John Petkau, Randy R. Sitter, Edward M. Carter, and John J. Hubert. Models for quantal response experiments over time. Biometrics, 45(4):1299–1308, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531782.

Greenland:1989:CES

[1874] Sander Greenland, Paul W. Holland, Nathan Mantel, Paul J. Wickramaratne, and Theodore R. Holford. Confounding in epidemiologic studies.

Biometrics, 45(4):1309–1322, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531783.

Sato:1989:VEM

[1875] Tosiya Sato, Sander Greenland, and James M. Robins. On the variance estimator for the Mantel–Haenszel risk difference. Biometrics, 45 (4):1323–1324, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531784.

Ulm:1989:ETV

[1876] K. Ulm and Christopher Cox. On the estimation of threshold values. Biometrics, 45(4):1324–1328, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531785.

Stonehouse:1989:NUO

[1877] John Stonehouse. A note on the use of the O’Brien–Dyck runs test with a blocking variable. Biometrics, 45(4):1328–1329, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531786.

Mielke:1989:CK

[1878] Paul W. Mielke, Jr., Daniel A. Bloch, and Helena C. Kraemer. Cohen’s kappa. Biometrics, 45(4):1329–1330, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531787.

Royston:1989:TMT

[1879] Patrick Royston, Henry C. Thode, Stephen J. Finch, and Nancy R. Mendell. Test for mixture of two normals. Biometrics, 45(4):1330–1332, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531788.

Searle:1989:OCR

[1880] Shayle R. Searle. Obituary: Charles Roy Henderson 1911–1989. Biometrics, 45(4):1333–1335, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531789.

Read:1989:BRM

[1881] C. B. Read. Book review: Multiple Comparison Procedures, by Y. Hochberg, A. C. Tamhane. Biometrics, 45(4):1337–1338, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531790.

Kocherlakota:1989:BRP

[1882] K. Kocherlakota and S. Kocherlakota. Book review: Principles of Multivariate Analysis, A User’s Perspective, by W. J. Krzanowski. Biometrics, 45(4):1338, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531791.

Kocherlakota:1989:BRC

[1883] S. Kocherlakota and K. Kocherlakota. Book review: Common Principal Components and Related Multivariate Models, by B. Flury. Biometrics, 45(4):1338–1339, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531792.

Laycock:1989:BRC

[1884] P. J. Laycock. Book review: Characterizations and Analysis of Block Designs, by A. K. Nigam, P. D. Puri, V. K. Gupta. Biometrics, 45(4): 1339, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531793.

Gates:1989:BRM

[1885] C. E. Gates. Book review: Manual of Crop Experimentation, by S. C. Pearce, G. M. Clarke, G. V. Dyke, R. E. Kempson. Biometrics, 45 (4):1339–1340, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531794.

Raab:1989:BRM

[1886] G. M. Raab. Book review: Multivariate Interpretation of Clinical Laboratory Data, by A. Albert, E. K. Harris. Biometrics, 45(4):1340–1341, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531795.

Dunnett:1989:BRM

[1887] C. W. Dunnett. Book review: Modern Statistical Methods in Chronic Disease Epidemiology, by S. H. Moolgavkar, R. L. Prentice. Biometrics, 45(4):1341, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531796.

Darby:1989:BRH

[1888] S. C. Darby. Book review: Health Risks of Radon and Other Internally Deposited Alpha-Emitters-BEIR IV, by National Research Council. Biometrics, 45(4):1341–1342, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531797.

Jeffers:1989:BRE

[1889] J. N. R. Jeffers. Book review: Ecological Simulation Primer, by G. L. Swartzman, S. P. Kaluzny. Biometrics, 45(4):1342, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531798.

Pope:1989:BRD

[1890] J. A. Pope. Book review: Design and Analysis Methods for Fish Survival Experiments Based on Release- Recapture, by K. P. Burnham, D. R. Anderson, G. C. White, C. Brownie, K. H. Pollock. Biometrics, 45(4): 1343, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531799.

Jamieson:1989:BRP

[1891] A. Jamieson. Book review: Population Genetics and Fishery Management, by N. Ryman, F. Utter. Biometrics, 45(4):1343–1344, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531800.

Conolly:1989:BRR

[1892] B. W. Conolly. Book review: Reaction-Diffusion Equations and Their Applications to Biology, by N. F. Britton. Biometrics, 45(4):1344, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531801.

Anonymous:1989:BRQ

[1893] Anonymous. Book review: Quality Control, Robust Design, and the Taguchi Method, by K. Dehnad. Biometrics, 45(4):1345, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531803.

Anonymous:1989:BRR

[1894] Anonymous. Book review: Readings in Decision Analysis, by S. French. Biometrics, 45(4):1345, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531802.

Anonymous:1989:BRW

[1895] Anonymous. Book reviews: World Health Statistics Annual 1988, by World Health Organization; World Health Statistics Quarterly 1988, Vol. 41 (3/4), by World Health Organization; World Health Statistics Quarterly 1989, Vol. 42 (1), by World Health Organization. Biometrics, 45(4): 1345, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531804.

Anonymous:1989:CEC

[1896] Anonymous. Correction: Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring by modeling the censoring process. Biometrics, 45(4):1347, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531805. See [1819].

Anonymous:1989:CES

[1897] Anonymous. Correction: Estimating the size of a population from a single sample. Biometrics, 45(4):1347, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531807. See [1594].

Anonymous:1989:CJE

[1898] Anonymous. Correction: Jackknife estimate of the matching coefficient of similarity. Biometrics, 45(4):1347, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531806. See [1593].

Anonymous:1989:BSF

[1899] Anonymous. The Biometric Society financial statements and independent auditors’ report January 31, 1989. Biometrics, 45(4):1351–1353, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531808.

Anonymous:1990:CBS

[1900] Anonymous. Corrections: Best subsets logistic regression. Biometrics, 46(3):889, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2532127. See [1870].

Anonymous:1990:CEC

[1901] Anonymous. Corrections: Estimation and comparison of changes in the presence of informative right censoring: Conditional linear model. Biometrics, 46(3):889, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2532126. See [1819].

Wu:1990:CEC

[1902] M. C. Wu and K. R. Bailey. Correction to: “Estimation and comparison of changes in the presence of informative right censoring: conditional linear model” [Biometrics 45 (1989), no. 3, 939–955; MR1029611 (90k:62151)]. Biometrics, 46(3):889, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1819].

Anonymous:1990:CRE

[1903] Anonymous. Correction: A random effects model for binary data from dependent samples. Biometrics, 46(4):1247, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532486. See [1638].

Coffey:1990:CRE

[1904] M. Coffey. Correction to: “A random effects model for binary data from dependent samples” [Biometrics 44 (1988), no. 3, 787–801; MR0963914 (89h:62166)]. Biometrics, 46(4):1247, December 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1638].

Wu:1991:CEC

[1905] M. C. Wu and R. J. Carroll. Correction to: “Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring process” [Biometrics 44 (1988), no. 1, 175–188; MR0931633 (89d:62109)]. Biometrics, 47(1):357, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1545].

Anonymous:1994:CSM

[1906] Anonymous. Corrections: A stabilized moment estimator for the betabinomial distribution. Biometrics, 50(1):321, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533245. See [1486].